A Complete Bibliography of Publications in the Proceedings of the American Philosophical

Society (1850–1899)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.01

Title word cross-reference

2 [Smi88]. _{3} [Smi88]. a [Smi78a]. m [Smi88].

&c [Ano89r, Bri77b, Cla84, Woo71b].

-Monochlordinitrophenol [Smi78a]. -Oxybenzoic [Smi88].

1 [Ano50j, Ano50p, Ano50t, Ano52-27, Ano53c, Ano54f, Ano56h, Ano57t, Ano58k]. 12th [Kir78]. 13 [Ano54r]. 13th [Kir78]. 14 [Ano52c, Cha68a]. 14th [Kir85a]. 15 [Ano50k, Ano50q, Ano50u, Ano52f, Ano52-28, Ano53d, Ano53l, Ano54g, Ano54v, Ano55n, Ano56e, Ano56i, Ano57u, Ano58l]. 15th [Ano61t, Ano71-39, Ano71-44, Ano74q, Ano74r, Ano75u, Ano76k, Ano77n, Ano78o]. 16 [Ano50e, Ano50v, Ano51l, Ano52d, Ano52n, Ano52p, Ano53f, Ano53v, Ano54m, Ano55i, Ano55o, Ano55s, Ano56p, Ano57n, Ano57x, Ano58d]. 16th [Ano63u, Ano67e, Ano69w, Ano73q, Ano74k, Ano74v, Ano75m, Ano77i, Ano77o]. 17 [Ano50r, Ano51g, Ano51b, Ano52h, Ano52-29, Ano52-31, Ano53m, Ano54h, Ano54n, Ano54s, Ano55f, Ano56t, Ano57g, Ano57p]. 1757 [Hay98a]. 17th [Ano68z, Ano69u, Ano71-33, Ano73k, Ano73u, Ano74d, Ano74m, Ano75g, Ano75w, Ano76v, Ano77f, Ano78p, Ano81p]. 18 [Ano50l, Ano50x, Ano51c, Ano52q, Ano52r, Ano53h, Ano53n, Ano53r, Ano54e, Ano55q, Ano56c, Ano56j, Ano56l, Ano57j, Ano57z, Ano58n]. 1852 [Ano52o]. 1858 [Ano58f, Ano58h, Ano58g, Ano58m, Ano58t, Ano58s, Ano58u, Ano58v, Ano58w]. 1859 [Ano59f, Ano59g, Ano59h, Ano59j, Ano59i, Ano59l, Ano59k, Ano59n, Ano59m, Ano59p, Ano59o, Ano59r, Ano59q, Ano59t, Ano59s, Ano59v, Ano59u, Ano59x, Ano59w, Ano59y]. 1860 [Ano60f, Ano60e, Ano60g, Ano60i, Ano60h, Ano60k, Ano60j, Ano60m, Ano60l, Ano60o, Ano60n, Ano60q, Ano60p, Ano60s, Ano60r, Ano60u, Ano60t, Ano60w, Ano60v, Ano60x]. 1861 [Ano61f, Ano61e, Ano61g, Ano61i, Ano61h, Ano61j, Ano61k, Ano61m, Ano61l, Ano61o, Ano61n, Ano61p, Ano61q, Ano61s, Ano61r, Ano61u, Ano61t, Ano61w, Ano61v, Ano61x]. 1862 [Ano62g, Ano62f, Ano62h, Ano62j, Ano62i, Ano62l, Ano62k, Ano62n, Ano62m, Ano62o, Ano62q, Ano62p, Ano62s, Ano62r, Ano62u, Ano62t, Ano62w, Ano62v, Ano62x]. 1863 [Ano63r, Ano63q, Ano63t, Ano63x, Ano63v, Ano63-28, Ano63z, Ano63-32, Ano63-31, Ano63-36, Ano63-34, Ano63-40, Ano63-38, Ano63-41, Ano63-42, Ano63-48, Ano63-46, Ano63-50, Ano63-49, Ano63-54]. 1864 [Ano63o, Ano63p, Ano63s, Ano63w, Ano63u, Ano63-27, Ano63y, Ano63-29, Ano63-30, Ano63-35, Ano63-33, Ano63-39, Ano63-37, Ano63-44, Ano63-43, Ano63-47, Ano63-45, Ano63-52, Ano63-51, Ano63-53]. 1865 [Ano65f, Ano65e, Ano65g, Ano65h, Ano65i, Ano65k, Ano65j, Ano65m, Ano65l, Ano65o, Ano65n, Ano65q, Ano65p, Ano65s, Ano65r, Ano65u, Ano65t, Ano65w, Ano65v, Ano65x]. 1866 [Ano66d, Ano66c, Ano66e, Ano66g, Ano66f, Ano66i, Ano66h, Ano66k, Ano66j, Ano66m, Ano66l, Ano66o, Ano66n, Ano66q, Ano66p, Ano66s, Ano66r, Ano66u, Ano66t, Ano66v]. 1867 [Ano67d, Ano67c, Ano67e, Ano67f, Ano67h, Ano67g, Ano67b, Ano67j, Ano67i, Ano67l, Ano67k, Ano67m, Ano67n, Ano67p, Ano67o, Ano67q, Ano67r, Ano67t, Ano67s, Ano67u, Ano68v]. 1868 [Ano68h, Ano68r, Ano68q, Ano68s, Ano68u, Ano68t, Ano68w, Ano68y, Ano68x, Ano68-27, Ano68z, Ano68-29, Ano68-28, Ano68-30, Ano68-31, Ano68-33, Ano68-32, Ano68-35, Ano68-34, Ano68-36, Kir68]. 1869 [Ano69r, Ano69p, Ano69t, Ano69v, Ano69u, Ano69-28, Ano69z, Ano69-29, Ano69-30, Ano69-36, Ano69-34, Ano69-40, Ano69-38, Ano69-44, Ano69-42, Ano69-48, Ano69-45, Ano69-50, Ano69-49, Ano69-53, May69]. 1870 [Ano69q, Ano69o, Ano69s, Ano69x, Ano69n, Ano69w, Ano69-27, Ano69y, Ano69-32, Ano69-31, Ano69-35, Ano69-33, Ano69-39, Ano69-37, Ano69-43, Ano69-41, Ano69-47, Ano69-46, Ano69-52, Ano69-51, Ano69-54, Ano69-55].

1871 [Ano71j, Ano71h, Ano71k, Ano71m, Ano71n, Ano71t, Ano71r, Ano71v, Ano71-27, Ano71z, Ano71-32, Ano71-31, Ano71-36, Ano71-34, Ano71-37, Ano71-40, Ano71-43, Ano71-44, Cop71m]. 1872 [Ano71i, Ano71g, Ano71l, Ano71p, Ano71o, Ano71s, Ano71q, Ano71w, Ano71u, Ano71-28, Ano71y, Ano71-29, Ano71-30, Ano71-35, Ano71-33, Ano71-38, Ano71-39, Ano71-42, Ano71-41, Ano71-45, Fra71]. 1873 [Ano73f, Ano73e, Ano73h, Ano73g, Ano73j, Ano73i, Ano73l, Ano73k, Ano73n, Ano73m, Ano73p, Ano73o, Ano73r, Ano73q, Ano73t, Ano73s, Ano73v, Ano73u, Ano73w, Wil75]. 1874 [Ano74a, Ano74e, Ano74d, Ano74f, Ano74h, Ano74g, Ano74j, Ano74i, Ano74l, Ano74k, Ano74n, Ano74m, Ano74p, Ano74o, Ano74s, Ano74q, Ano74r, Ano74u, Ano74t, Ano74w, Ano74v, Ano74x]. 1875 [Ano75e, Ano75d, Ano75f, Ano75h, Ano75g, Ano75j, Ano75i, Ano75k, Ano75l, Ano75n, Ano75m, Ano75p, Ano75o, Ano75r, Ano75q, Ano75t, Ano75s, Ano75v, Ano75u, Ano75w]. 1876 [Ano76h, Ano76g, Ano76i, Ano76j, Ano76k, Ano76m, Ano76l, Ano76o, Ano76n, Ano76q, Ano76p, Ano76s, Ano76r, Ano76u, Ano76t, Ano76w, Ano76v, Ano76y, Ano76x, Ano76z, Kir77a, Str76]. 1877 [Ano77e, Ano77d, Ano77f, Ano77h, Ano77g, Ano77j, Ano77i, Ano77l, Ano77k, Ano77n, Ano77m, Ano77p, Ano77o, Ano77r, Ano77q, Ano77t, Ano77s, Ano77v, Ano77u, Ano77w, Kir77a]. 1878 [Ano78e, Ano78d, Ano78f, Ano78h, Ano78g, Ano78i, Ano78j, Ano78l, Ano78k, Ano78m, Ano78n, Ano78o, Ano78q, Ano78p, Ano78s, Ano78r, Ano78u, Ano78t, Ano78v, Bar78, Phi79b, Wil78]. 1879 [Ano79d, Ano79c, Ano79e, Ano79g, Ano79f, Ano79i, Ano79h, Ano79k, Ano79j, Ano79m, Ano79l, Ano79o, Ano79n, Ano79q, Ano79p, Ano79s, Ano79r, Ano79u, Ano79t, Ano79v]. 1880 [Ano80f, Ano80e, Ano80g, Ano80i, Ano80h, Ano80k, Ano80j, Ano80m, Ano80l, Ano80o, Ano80n, Ano80q, Ano80p, Ano80s, Ano80r, Ano80t, Ano80u, Ano80v, Ano80d, Ano80w, Ano80x, Les80a]. 1881 [Ano81g, Ano81h, Ano81j, Ano81i, Ano81l, Ano81k, Ano81n, Ano81m, Ano81p, Ano81o, Ano81r, Ano81q, Ano81t, Ano81s, Ano81v, Ano81u, Ano81x, Ano81w, Ano81y]. 1882 [Ano82d, Ano82c, Ano82e, Ano82f, Ano82g, Ano82i, Ano82h, Ano82k, Ano82j, Ano82m, Ano82l, Ano82o, Ano82n, Ano82r, Ano82p, Ano82t, Ano82s, Ano82v, Ano82u, Ano82w, Lew82a]. 1883 [Ano83f, Ano83e, Ano83g, Ano83i, Ano83h, Ano83k, Ano83j, Ano83m, Ano83l, Ano83p, Ano83o, Ano83r, Ano83q, Ano83t, Ano83s, Ano83u, Ano83v, Ano83x, Ano83w, Ano83y]. 1884 [Ano83n, Ano84d, Ano84c, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84k, Ano84j, Ano84m, Ano84l, Ano84o, Ano84n, Ano84q, Ano84p, Ano84s, Ano84r, Ano84t, Ano85-74]. 1885 [Ano85-73, Ano85-76, Ano85-75, Ano85-78, Ano85-77, Ano85-80, Ano85-79, Ano85-82, Ano85-81, Ano85-83, Ano85-84, Ano86g, Ano86j, Ano86h, Ano86v, Ano86t, Ano86z, Ano86y, Ano86-27, Fra86c]. 1886 [Ano86f, Ano86e, Ano86i, Ano86l, Ano86k, Ano86m, Ano86n, Ano86p, Ano86o, Ano86r, Ano86q, Ano86u, Ano86s, Ano86w, Ano86x, Ano86-28, Ano87i]. 1887 [Ano87f, Ano87g, Ano87j, Ano87h, Ano87l, Ano87k, Ano87n, Ano87m, Ano87p, Ano87o, Ano87r, Ano87q, Ano87t, Ano87s, Ano87v, Ano87u, Ano87x, Ano87w].

1888 [Ano88d, Ano88c, Ano88f, Ano88e, Ano88h, Ano88g, Ano88j, Ano88i, Ano88l, Ano88k, Ano88n, Ano88m, Ano88p, Ano88o, Ano88r, Ano88q, Ano88t, Ano88s]. 1889 [Ano89g, Ano89s, Ano89u, Ano89t, Ano89v, Ano89w, Ano89y, Ano89x, Ano89z, Ano89-27, Ano89-29, Ano89-28, Ano89-30, Ano89-31, Ano89-33, Ano89-32, Ano89-35, Ano89-34]. 1890 [Ano90f, Ano90h, Ano90g, Ano90j, Ano90i, Ano90l, Ano90k, Ano90n, Ano90m, Ano90p, Ano90o, Ano90r, Ano90q, Ano90s]. 1891 [Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano91k, Ano91j, Ano91m, Ano91l, Ano91n, Ano91o, Ano91q, Ano91p, Ano91s, Ano91r, Ano91u, Ano91t]. 1892 [Ano92e, Ano92g, Ano92f, Ano92i, Ano92h, Ano92j, Ano92k, Ano92m, Ano92l, Ano92o, Ano92n, Ano92q, Ano92p, Ano92r, Ano92t, Ano92s]. 1893 [Ano93g, Ano93f, Ano93h, Ano93i, Ano93k, Ano93j, Ano93m, Ano93l, Ano93o, Ano93n, Ano93q, Ano93p, Ano93s, Ano93r, Ano93u, Ano93t, Ano93v, Ano93w, Ano94f, Pot95]. 1894 [Ano94k, Ano94j, Ano94m, Ano94l, Ano94o, Ano94n, Ano94q, Ano94p, Ano94s, Ano94r, Ano94u, Ano94t, Ano94w, Ano94v, Ano94y, Ano94x, Ano94-27, Ano94z, Ros95b]. 1895 [Ano95f, Ano95e, Ano95h, Ano95g, Ano95i, Ano95j, Ano95l, Ano95k, Ano95m, Ano95n, Ano95p, Ano95o, Ano95q, Ano95r, Ano95t, Ano95s, Ano95v, Ano95u]. 1896 [Ano96e, Ano96f, Ano96g, Ano96h, Ano96j, Ano96i, Ano96l, Ano96k, Ano96n, Ano96m, Ano96p, Ano96o, Ano96q, Ano96r, Ano96t, Ano96s, Ano96v, Ano96u, Ano96x, Ano96w]. 1897 [Ano97d, Ano97e, Ano97k, Ano97h, Ano97m, Ano97l, Ano97o, Ano97n, Ano97p, Ano97q, Ano97s, Ano97r, Ano97u, Ano97t, Ano97w, Ano97v, Ano97x, Ano97i, Ano97y, Ano97j, Ano97-27, Ano97z]. 1898 [Ano98c, Ano98d, Ano98e, Ano98f, Ano98h, Ano98g, Ano98j, Ano98i, Ano98l, Ano98k, Ano98n, Ano98m, Ano98p, Ano98o, Ano98r, Ano98q, Ano98t, Ano98s]. 1899 [Ano99g, Ano99f, Ano99h, Ano99i, Ano99k, Ano99j, Ano99m, Ano99l, Ano99o, Ano99n, Ano99q, Ano99p, Ano99s, Ano99r, Ano99u, Ano99t, Hau99]. 18th [Ano63v, Ano68-36, Ano69y, Ano71k, Ano71-41, Ano73e, Ano73m, Ano74g, Ano74x, Ano75n, Ano76l]. 19 [Ano50c, Ano50n, Ano50y, Ano52i, Ano52s, Ano52w, Ano52-32, Ano53e, Ano54p, Ano55k, Ano55u, Ano56f, Ano56v, Ano57q, Ano58i, Ano58o]. 19th [Ano67k, Ano69-45, Ano71g, Ano73g, Ano73w, Ano74n, Ano75s, Ano77k, Ano77u, Ano78d]. 1st [Ano61u, Ano67h, Ano69q, Ano74s, Ano75v, Ano76j].

2 [Ano50w, Ano51m, Ano52e, Ano53g, Ano55j, Ano55p, Ano55t, Ano56q, Ano57o, Ano57y, Ano58e, Hay69a]. 20 [Ano50f, Ano50-28, Ano51i, Ano52g, Ano52l, Ano53p, Ano54j, Ano54t, Ano55d, Ano55m, Ano56m, Ano57l, Ano57r, Ano57v]. 20th [Ano68-32, Ano71o, Ano73n, Ano74i, Ano74o, Ano74t, Ano75f, Ano76x, Ano77d, Ano77m, Ano77q, Ano82u]. 21 [Ano50g, Ano50o, Ano51e, Ano51j, Ano52u, Ano52x, Ano53j, Ano53t, Ano54c, Ano54l, Ano55g, Ano55w, Ano56n, Ano56r, Ano57i, Ano58q]. 21st [Ano68s, Ano69-46, Ano71h, Ano73i, Ano73o, Ano73s, Ano74f, Ano77w, Ano78m]. 22d [Ano94f]. 23 [Ano58c]. 26th [Ano94f]. 27 [Ano57d]. 29 [Ano57f]. 29th [Bar78]. 2d [Ano69x, Ano71s, Ano73r, Ano73v, Ano74l, Ano74w, Ano81j]. 2nd [Ano77j, Ano77p, Kir73].

3 [Ano50s, Ano51h, Ano52j, Ano52-30, Ano54i, Ano54o, Ano56u, Ano57h, Ano57e]. 30 [Kir87a]. 3d [Ano69v, Ano71-35, Ano71-37, Ano73l, Ano74e, Ano75h, Ano78q]. 3rd [Ano76w].

4 [Ano50m, Ano50z, Ano51d, Ano53i, Ano53o, Ano53s, Ano55r, Ano56d, Ano56k, Ano57k]. 4th [Ano63x, Ano69-27, Ano69-47, Ano71-42, Ano73f, Ano74h, Ano76m, Ano77r].

5 [Ano50d, Ano50-27, Ano52k, Ano52t, Ano52y, Ano54q, Ano55l, Ano55v, Ano56g, Ano58j, Ano58p]. 5th [Ano71w, Ano73h, Ano75t, Ano77l, Ano77v, Ano78e].

6 [Ano50h, Ano52m, Ano53q, Ano54k, Ano54u, Ano55e, Ano57m, Ano57s, Ano57w, Ano71x]. 6th [Ano60f, Ano68-33, Ano71p, Ano71-43, Ano74j, Ano74p, Ano74u, Ano82v, Ano87r].

7 [Ano50i, Ano51f, Ano51k, Ano52v, Ano52z, Ano53k, Ano53u, Ano54d, Ano55h, Ano56o, Ano56s, Ano58r]. 7th [Ano60i, Ano69-32, Ano69-52, Ano71j, Ano73j, Ano73p, Ano73t, Ano76o, Ano76y].

9 [Mor96]. 9th [Ano69n].

A. [Ano69-55]. A.M [Bri81a, Wil88]. Aachen [Phi79b]. Abdominal [God96]. Aboriginal [Jor88]. Aborigines [Mat97b, Mat98b]. Above [Pat75, Abb98, Bar77]. Absorption [Dor96]. Abstract [Ano74a, May69]. Accessibility [Ano50a, Ano50b, Ano51a, Ano52a, Ano52b, Ano53a, Ano53b, Ano54a, Ano54b, Ano55a, Ano55b, Ano56a, Ano56b, Ano57a, Ano57b, Ano58a, Ano58b, Ano59a, Ano59b, Ano60a, Ano60b, Ano61a, Ano61b, Ano62a, Ano62b, Ano63a, Ano63b, Ano65a, Ano65b, Ano66a, Ano66b, Ano67a, Ano68a, Ano68b, Ano69a, Ano71a, Ano73a, Ano74b, Ano74c, Ano75a, Ano75b, Ano76a, Ano76b, Ano76c, Ano77a, Ano77b, Ano78a, Ano78b, Ano79a, Ano79b, Ano80a, Ano80b, Ano80c, Ano81a, Ano81b, Ano81c, Ano82a, Ano82b, Ano83a, Ano83b, Ano84a, Ano84b, Ano85-65, Ano85-66, Ano85-67, Ano85-68, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano88a, Ano88b, Ano89a, Ano89b, Ano89c, Ano90a, Ano90b, Ano90c, Ano91a, Ano91b, Ano92a, Ano92b, Ano92c, Ano93a, Ano93b, Ano93c, Ano94a, Ano94b, Ano94c, Ano94d]. Accessibility [Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano98a, Ano98b, Ano99a, Ano99b]. Accessory [Sto88a]. Account [Cop71y, Cop71-27, Gro79, Phi79b, Phi80a, Phi80b, Phi83a, Phi89a, Wil81]. Acetobenzoic [Gre80b]. Achaz [Sac95]. Acid [Dav83b, Gre80b, Sad75, Smi77b, SKHE87, Smi88]. Acids [Mar78]. Acoustic [P.73]. Acquired [Hya94]. across [Ano63k, Ano69g, Lew85]. Act [BP89]. Action [Cha76a, Fra95a, Gre79a, Gre80b, Hal75a, Hau89, Smi88]. Activity [Woo69a]. Adams [Fra75b, Fra75a]. Adaptive [Ryd94]. Addison [Ley60]. Addition [Bau92a, Cop69m, Cop83a, Cop83f, Cop93b]. Additional [Gra58a, LeC78, Phi79a]. Address [Kea78]. Adjacent [Ste79]. Adjourned [Ano91c, Ano96e, Ano97d, Ano97e, Ano98c, Ano98d]. Adocidæ [Cop69i, Cop77f]. Adocus [Cop69f]. Adulterations [Car85b]. Aerial [Tha83]. Aerolites [Kir87c]. Aerolitic [Kir78]. Æthereal [Cha71b, Cha71c]. Affected [Ker73]. Affinities [Bri89a, Bri92e, Bri97a, Cop86e, Wak90]. Affinity [Cha80f]. Africa [Phi80c]. After [Oli86]. After-Color [Oli86]. Age [Hal80, Lym94a, Lym94c, Mor62, Win69a, Win69b]. Agency [Hof86c]. Ages [Ano77c, Kir77b]. Aggregation [Cha77b]. Aggregatus [Hub94]. Agricultural [Cop71o]. Agriochærus [Sco94]. Aix [Phi79b]. Alaguilac [Bri87b]. Alaska [Cop73b]. Alaskaite [Koe85, Kön81]. Albert [All86, Les90a]. Albirupean [Uhl88]. Alexander [Ley60]. Alfred [Bai91]. Algae [Woo69b]. Algeria [Pep96]. Alleged [Ste74]. Alleghany [Les77d]. Allegheny [Ash77a]. Allen [Les76]. Allied [Gro95a]. Allotropic [Hou74]. Almanacs [Phi81a]. Almighty [Les86b]. along [Hal75a, Les74, SS78, Ste81d, Les73a]. Alphabet [Ano63d, Ano63e, Phi83b]. Alphabetical [Phi89b]. Alpine [Hei92]. al’tä [Hof94]. Alterations [Gen73]. Alternating [Mil97]. alticoronatus [Ano69e]. Altitude [Abb98, Hag82]. Altitudes [Dav66]. Alto [Ano63j]. Alum [Ano68k]. Aluminium [Gre79a]. Amazon [Cop78b, Cop86b]. Amazonas [Der79a]. Amblypoda [Cop88b]. Amblyrhiza [Ano69i, Ano69j]. Amended [Ano89q]. America [Cop71j, Ant76, Bri81b, Bri85a, Bri89b, Bri93a, Bri95b, Bri98b, Cop69b, Cop69p, Cop77p, Cop79a, Cop85d, Cop85f, Cop86j, Cop87b, Cop88a, Cop96b, Dav83a, LH76, LeC77, Pea61, Phi80b, Woo69a, Woo69b]. American [Ano59e, Ano62y, Ano75c, Ano89g, Ano89q, Ano99c, Cop86i, FGG^{+}80, Hay98b, Phi89b, Wil88, Ano74a, Ano85-72, Ano89i, Ano89j, Ano89l, Ano89k, Ano89m, Ano89r, Ano94f, Ano94i, Ano99c, Ano99d, Bri85c, Bri86a, Bri86d, Bri86c, Bri92f, Bri94b, Bri77b, Cha71d, Cha71i, Cha80d, Cop89b, Cou93, Gen74a, Gro74, Gro83, Les71c, Les77b, Mab97, Mar73, Mon78, Phi83a, Str76, Wil82, vHB85]. Americana [All89b]. among [Cop77a, Cop85e, Hof75, Mer97b]. Amphiumidæ [Cop86e]. Analogues [Cha60, Cha61]. Analyses [BC75, Les78a, Sad78]. Analysis [Bri76b, Cha81a, Kön78, Les77a, Out74, Smi79]. Analytical [Cop86a]. Anatomy [Pac94, Wil81]. Ancient [Ano63k, Bri87a, Cus96, Ano63e, Bri89a]. Andagueda [Bri95d]. André [MS89]. Andrew [Car85a]. Aneroid [Fra83]. Angustatus [Ano69f]. Anhydride [Gre80b]. Aniline [Smi78a]. Animal [Ryd89c]. Anisodactylus [Hor80b]. Anniversary [Ano89g, Ano90f, Ano94f, Ano94i, FGG^{+}80]. Annual [Sel99]. Anoki [Les91a]. Anoplonassa [Ano69j]. Anse [Cop95b]. Antagonism [Hay77]. Anthracite [Cha81a, She69, Woo76]. Antilocapra [All89b]. Antique [Pea63]. Antonii [Dug86]. ANX [Les91a]. Apatela [Gro95b]. Apollonius [Smy98]. Appalachian [Ano69g]. Apparatus [Cop86f, Woo76]. Apparent [Hag82, Les71c, Cha79b]. Application [Hag83, Tay87]. Applied [Fra85, Fra86a, Hau79]. Appointed [Ano89q]. Approaches [Fra77b]. Approximate [Cha79a]. Approximations [Cha71j, Cha79b, Cha80a, Cha80b, Cha80c]. April [Ano50d, Ano50c, Ano51d, Ano51c, Ano52e, Ano52d, Ano53c, Ano53d, Ano54d, Ano54c, Ano55e, Ano55d, Ano56d, Ano56c, Ano57h, Ano57g, Ano58e, Ano58d, Ano59f, Ano59g, Ano60f, Ano60e, Ano61f, Ano61e, Ano62g, Ano62f, Ano63r, Ano63q, Ano63o, Ano63p, Ano65f, Ano65e, Ano66d, Ano66c, Ano67d, Ano67c, Ano68r, Ano68q, Ano69r, Ano69p, Ano69q, Ano69o, Ano71j, Ano71h, Ano71i, Ano71g, Ano73f, Ano73e, Ano74e, Ano74d, Ano75e, Ano75d, Ano76h, Ano76g, Ano77e, Ano77d, Ano78e, Ano78d, Ano79d, Ano79c, Ano80f, Ano80e, Ano81g, Ano81h, Ano82d, Ano82c, Ano83f, Ano83e, Ano84d, Ano84c, Ano85-73, Ano86f, Ano86e, Ano87f, Ano87g, Ano88d, Ano88c, Ano89s, Ano90f, Ano90h, Ano90g, Ano91e, Ano91d, Ano92e]. April [Ano93g, Ano93f, Ano94k, Ano94j, Ano95f, Ano95e, Ano96g, Ano96h, Ano97k, Ano97h, Ano98e, Ano98f, Ano99g, Ano99f, Fra71, Lew82a, Pot95]. Aransas [Hau99]. Arbitration [Edm97]. Arch [Sla95]. Archæan [Fra84]. Archaeology [Bri86d, Phi83a]. Archer [Rot89]. Archetypes [Ano63e]. Arctic [Rin85]. Area [Ash97]. Arfvedsonite [Kön77a]. Arkansas [Ash97, Smi96]. Armature [Pac93b]. Armstrong [Les71e]. Army [Car85a]. Arrangement [Cop81c, Cro73]. Art [Mer97b, Phi79a, Phi79c]. Artesian [LC91]. Arthur [Ano99d]. Article [Gat90]. Artifex [Sac95]. Artiodactyla [Cop87a, Cop85d]. Artiodactyls [Sco98]. Artionyx [Sco94]. Aruba [Gat85a]. Arval [Bri92a]. Ascertain [Cre74]. Ashburner [Les90a]. Ashes [Woo71b]. Asia [Bri95c, Phi80c]. Asiatic [Wak90]. Asphalt [Day98]. Asphaltum [Pec68]. Assay [Out74]. Assaying [Out74]. Asserted [Hay77]. Assist [Ano89q]. Assistant [Ano76f, Les76]. Associated [Gen73, MKAP71]. Association [Bai94, Str76]. Assyrian [Dol93]. Astaci [Cop69k]. Asteroidal [Cha73j]. Asteroids [Kir69b, Kir83, Kir89, Kir92]. Astrological [Cha73k]. Astronomical [Cha73h, Cha79b, Cha80a, Cha80b, Cha80c]. Astrophyllite [Kön77a]. Atalapha [All91]. Athinson [Car85a]. Atlantic [Jor88]. Atmosphere [Blo73]. Atmospheres [Tro76]. Atmospheric [Ano94e, Blo74, Cha71s]. Atomic [PS94, SD94]. Atrophy [Oli94]. Atrypa [Gin78]. Attachment [Key88]. Attaci [Ano75c, Gro74]. Attractive [Cha74f]. Auditory [Cop86f]. Aufgezeichnet [Kra94]. Aug. [Ano71k, Ano84e]. August [Hay98a, Phi79b, Ano50e, Ano52f, Ano52g, Ano53e, Ano54e, Ano55f, Ano56e, Ano57i, Ano58f, Ano59h, Ano60g, Ano61g, Ano62h, Ano63t, Ano63s, Ano65g, Ano66e, Ano67e, Ano68s, Ano69t, Ano69s, Ano71l, Ano74f, Ano75f, Ano76i, Ano77f, Ano78f, Ano79e, Ano80g, Ano82e, Ano83g, Ano86g, May69]. Augustus [G.86]. Auriferous [Cha81b]. Aurora [Abb98, Fra71, Lew82a]. Auroral [Cha71h]. Auroras [Cha71k, Cha71m, Cha71p]. Australian [Ano97f, Mat97a, Mat97b, Mat98a, Mat98c, Mat99]. Austroriparian [Cop77h]. Axis [Dav66]. Azimuth [Hag82]. Azimuth-Observations [Hag82].

B [Har80, Hau77, Les81c, Rob75, Vau90]. B. [Ano68e]. B.C. [Ano63d]. Back [Ano55c, Ano57c, Ano59c, Ano62c, Ano73b, Ano76d, Ano81d, Ano85-69, Ano89d, Ano89e, Ano89f, Ano90d, Ano93d]. Bad [Rot83]. Bahamas [Rot91]. Baltimore [Ano71e, Hau77]. Bancker [Cad69]. Bank [Rei71, Ano63j, Ano63k]. Banks [Gen74b, Les74, Woo76, Ano63j]. Bar [Str76]. Barbadoes [Cha74d]. Baric [Cha85c]. Barite [SKHE87]. Barrett [Hau77]. Bars [Hau89]. Bart [Fie71]. Barton [Rus93]. Based [Cou93]. Basin [Cop71b, Cop71s, She69, Spe81, Ano63g, Jon80]. Basins [New82]. Bathmodon [Cop71t]. Batrachia [Cop77a, Cop85e, Cop71j, Cop78a, Cop86i, Cop87c, Cop93b]. Batrachians [Cop86b, Gar87a]. Battle [Phi78]. Bawerk [Bai91]. Be [Woo76, Cha73k, Fra85]. Beadle [Agn85]. Bear [Woo71a, Hay69b]. Bearing [Les81b, Les81a, Les79]. Beaver [Ano63g, Les77d]. became [Kra94]. Bed [Car91, Cop77m, Les77d, Ste81d]. Bedford [Les73b]. Beds [Cla83c, Cop69h, Cop71v, Cop77g, Cop81b, Cop86g, Les71c, Les77a, Ste74, Ful74]. Beegerite [Koe85]. before [Kön77a, Str76]. Beginnings [Cha75a]. Behavior [Bro94]. Being [Hei92]. Bellicosa [Cop71e]. Belonging [Hay69b]. Belt [Ano69g, Les71d, Les73d]. Bend [Les71e]. Benjamin [Ano90f]. Benneville [Bri94a]. Benzol [Gre79a]. Beothuk [Gat85b, Gat86, Gat90]. Berlin [Fra86c]. Beryllium [Smi78a]. Bethune [Ano62d]. Betoya [Bri92c]. between [Cha78a, Hay77, Kir69b, Kir71a, Kir92, Phi81a, Pri95, Rot83, SS78, Wil75]. beyond [Dav66]. Biblical [Gro79]. Biddle [Cou93]. Biela [Kir87a]. Bielids [Kir87b]. Big [Mer97a]. Binney [Sti69, Str76]. Biographical [Agn85, Ano60c, Da 85, Lei87, Pot95, Rog80, Rot82, Rot86, Rot89, Rus86, Vau85, Wil88]. Birds [vD99]. Bismuth [Gen74a, Kön81]. Bitter [Cop71p, Cop71z]. Bitumen [Ano63f]. Bitumens [Pec68, Pec98]. Bituminous [Cre77]. Black [Hau77]. Blair [Gen74b, Les73c, PS78]. Bland [Ste87a]. Blarina [Ryd90a]. Blast [Bri77b]. Blood [Ryd94]. Blood-Corpuscles [Ryd94]. Blooded [Cop77o]. Bloomington [Kir68]. Blow [Kön77a]. Blow-Pipe [Kön77a]. Blowpipe [Kön78]. Blue [Hal75a]. Board [Bar77]. Bog [Lew82b]. Böhm [Bai91]. Böhm-Bawerk [Bai91]. Boiler [Cre74]. Boilers [Woo76]. Bois [Ack69]. Bombycine [Pac94]. Bone [Cop71i, Cop77m, Cop80a, Mer97a]. Book [But99, Ros95b]. Books [Cra83]. Booth [DuB88]. Borate [Smi78a]. Boric [SKHE87]. Boring [Dub71b, Les91c, She69]. Borneo [Fur96, Hil96]. Bosnien [Kra86].

Bostrichidæ [Hor78a]. Botanical [Bai96, Day83]. Botanique [Les81c]. Botanist [MS89]. Botocúdus [Bra89]. Boxing [Bac94a]. Bradford [Lil83, Ash80, Lil86]. Brady [Les71e]. Brain [Cop77l, Cop77k, Cop86f, Wil81]. Brains [Cop83c]. Branch [Kön78, Les79]. Brazil [Cop86c, Cop87c, Der79b]. Breakwater [Hau99]. Brethren [Bri92a]. Breton [Daw63, Les62]. Brevet [Car85a]. brevicauda [All90a]. Bridge [Ash84]. Bridger [Cop71c, Cop71y, Cop71-27, Cop73d]. Bridges [Rus84]. Brief [Bla76, Hay68a, Hil96, Mee69, Phi79b, Phi83a]. Brigadier [Car85a, Ano99c]. Brigadier-General [Car85a, Ano99c]. Brightley [Ano71c]. Brinton [Sto85]. British [Boa87]. Broad [Ano63g]. Broadside [Sac98]. Brook [Lil83, Lil86]. Brothers [Rus86]. Brown [Les74]. Brownian [Bac94b]. Brownstone [Lym94a, Lym94c]. Brush [Les73c]. Bufo [Cop86i]. Building [Hau89, Ano85-72]. Built [Pri82]. Bundles [Bra84]. Bureau [Bla87]. Burette [Koe74]. Burettes [Bri76b]. Burned [Woo76]. Butterflies [Gro99a].

C [Ano68e, Bau92a, Bau92b, Fra77a, Hau77]. C. [Cop77j]. Cachryx [Ano69d]. Cadmium [Smi78b]. Cairo [Wil75]. Cakchiquel [Bri84, Sto85]. Cal. [Hof86c]. Calcification [Ryd89b]. Calculation [Sad78]. Calculus [Smi79]. Caledonia [Ano63j, Ano63k]. Calendar [Ano99c, Ano99d, Bri93a]. California [Mee69, vD99]. Called [Bri87b, Lym95a, Oli94]. Calospasta [Hor91]. Cambala [Pac83b]. Cambrian [Les77a]. Cambridge [Gar87b]. Cambro [Ash77a]. Camden [Pot95]. Can [Woo76]. Canada [MS89]. Cannelton [Les77d]. Cañon [Ste81b]. Cantabrigensis [Mei86]. Cape [Daw63, Les62]. Capital [Gra58a]. Capt. [Cop69j]. Carbon [Hal75a]. Carboniferous [Les77b, Les78b, Pac86]. Carey [Bai91]. Carll [Ano76f]. Carnivora [Cop73c, Cop82c]. Carolina [Cha81b, Cop69l, Cop69g, Cop71q, Les71d]. Carollia [All90a]. Cartography [Bri98a]. Carved [Rei71]. Carvings [Ano97f, Mat97a, Mat97b]. Case [All84, MKAP71]. Cases [Tha78]. Caspar [Bid89]. Cat [Sto82, Sto86c, Sto87b, Sto88a, Sto88b, Sto88c, Wil81]. Catalogue [Ano94e, Cop71a, Cop86b]. Cataract [Wil75]. Catskill [Cla83b, SS78, She78]. Cause [Blo74, Fra77b, Loc83, Low69a]. Causes [Bla77, Cre74, Dut71, Hau88]. Cave [Mer97a, Cop71i, Mer95]. Cavities [Phi97b]. Cayman [Gar87a]. Ceasing [Woo71a]. Celebrated [Ano99c]. Celebration [FGG^{+}80]. Centennial [Ano89g]. Central [Cha76b, Cha77a, Ash77a, Bri81b, Bri85a, Bri93a, Bri95b]. Centre [Les74, Gen74b]. Centres [Cha77b]. Ceratocampadae [Ano75c, Gro74]. Ceratocampidæ [Pac93b]. Ceremonies [Mat98c]. Certain [Ano77c, Bla71, Cop71q, Kir69d, Kir77b, Kir92, Pac93a, Pac93b, Phi81a, Sad76, Sad77]. Château [Ros94]. Chêne [Ack69]. Chabas [Cha71a]. Chaco [Bri98a]. Chairman [Sel99]. Chalfont [Lym95a]. Chalk [Cop71d]. Channel [Ste80b, Kea78]. Channel-Way [Ste80b]. Chapelle [Phi79b]. Character [Pla79, Str76]. Characteristic [Hya94]. Characteristics [Bri86c, Fra78a].

Characters [Ano68g, Ano99c]. Charcoal [Dav83b]. Charge [Ano76f, Les76]. Charles [Muh83, Ano65d, Bel73, Les76, Les90a, Rob75]. Chase [Ano63d, Ano63e, Cha85a, Gar87c]. Chemical [Bro94, Cha80f, Fra78a, Les77a, Pec98, Sad97, Gen87]. Chemung [Cla83c, Lil83, Lil86]. Chevalier [Kea78, Rob80]. Chicago [Gra58b]. Chief [Car85a, Pri74]. Chilognath [Pac83b]. Chilopoda [Mei86]. Chinantec [Bri92d]. Chinese [Ano63d, Ano63e, Cha61]. Chloride [Gre79b, Gre79a]. Chlorine [Dav83b, Gre80b, Smi77a]. Choco [Bri95d, Bri96]. Choctaw [BB69]. Christianity [Cha79d]. Christophorus [Sac95]. Chromic [Smi77c]. Chromometry [Kön78]. Circle [Cha79a]. Circulation [Ant76, Blo74]. Circumstances [HT78]. Cities [Tha78]. City [Ano62y, Bac91b, Hay69b, Ste81b]. Civil [Bac92, Rob71]. Civilized [Bri85a]. Clairsville [Les73b]. Clara [vD99]. Clarence [Cop69j]. Clarion [Hau77]. Clark [Cou93]. Classification [Bau93, Cop82a, Cop87a, Hun88, Pea61]. Claudius [Ano69f]. Clawed [Cop73c]. Cleeman [Pri94]. Cliff [Ano68d]. Climatological [Pep96]. Climatology [Ano94e]. Clinton [Cha78a, Cla84]. Co [Ano71e, Cha78a, Cha75b, GvR85, Kön77a, Ste81c]. Co-Ordinations [Cha75b]. Co. [Cla83e]. Coahuila [Kan77]. Coal [Ano63g, Ano71e, Ash77a, BC75, Cha78a, Cha81a, Cop71w, Cop77a, Cop85e, Cop86h, Daw63, Dra97, Ful74, Jon80, Les62, Les71c, Les73d, Les77d, Lym94b, Smi96, St.69, Ste74, Ste81b, Ste81c, Ste81d, Woo76]. Coal-Field [Ste81b]. Coal-Measures [Daw63, Les62]. Coal-Waste [Woo76]. Coals [Pla79]. Coast [Bri91, Les62, Cus96]. Cobitidæ [Cop71g]. Cochliopodidæ [Pac93a]. Codex [Phi84a]. Coin [Dub71b]. Coinage [Out74]. Coins [Phi79a, Phi79c]. Coke [Ano71e]. Cold [Blo74, Cop77o]. Coleoptera [Cro73, HSL78, Sch78]. Collation [Les76]. Collected [Mee69, SO87]. Collection [Alg85, Ano71f, Cop86b, Les76, Net77, Phi79a, Phi79c]. Collections [Ano76f, Phi83a]. College [Cop71o, Sha83]. Colliery [Ano63g]. Color [Fra74b, Oli86, Sac96]. Colorado [Cop74a, Cop77g, Cop77n, Hay68b, Hay69a, Kön77a, Mee69, Scu94, Ste81b]. Colored [Dor96]. Colorimetric [Kön77a]. Columbia [Boa87]. Columella [Mon78]. Colwell [Car71]. Colydiidæ [Hor78b]. Combination [Woo76]. Comet [Kir85a, Kir87a]. Comets [Kir69a]. Coming [Les71c]. Commemorate [Phi78]. Commemoration [Ano89g, Ano90f, Ano94f]. Comment [Les83d]. Commission [Ano89q, Hay99]. Committee [Ano89q, Hei92]. Commodities [Dud86]. Communicated [Cha80d]. Communication [Ano69-55, Mar71, Pea63]. Comparative [Bri90b, Bro94, Dav83a, Gar87b]. Comparison [Cha69b]. Compass [Fra85, Hau77]. Complementary [Oli86]. Composing [Cla84]. Composite [Fra86a, Tay85]. Composition [Mab97, Sad76, Sad77]. Computation [Hau71]. Conception [Bri86a]. Conclusion [Ano74r, Les81a, Taf62]. Condensed [Bri77b]. Condition [Del74, Del75]. Conditions [Eme71]. Conductor [HK97]. Confirmation [Cha73h]. Confirmations [Cha79c]. Congenital [All84]. Conglomerate [Ano68d, Whi80]. Congo [Phi89a].

Congress [Fra86c, Sac98]. Connection [Bla75, Mar74]. Connotations [Cha68b]. Consequences [Kir85b]. Conservation [Bac93]. Constant [Cha78b]. Constitution [Ash80]. Construction [Fra83]. Consumption [Dud86]. Contained [Cop86b]. Continuation [Cop77a, Cop85e]. Contractions [Hou90]. Contribution [Cop69a, Cop69b, Cop69c, Cop69p, Cop73b, Cop77p, Cop79a, Cop79b, Cop80c, Cop82e, Cop83b, Cop85a, Cop85c, Cop85f, Cop86c, Cop86j, Cop87b, Cop92a, Cop95a, Cop96a, Cop96b, Day83, Der79a, Fra78a, Hal77, Oli94, Phi88, Phi92, Wil82]. Contributions [Bar75, Bri69, Cop82b, Gen73, Gen74a, Gen77, Gen82, GvR85, Gen86, Gen87, Gra58b, Gra58a, Hor73a, Koe74, Kön77a, Kön78, Kön81, Les76, Mar78, Sad75, Sad76, Sad77, Smi77a, Smi77b, Smi77c, Smi78a, Smi78b, Gen82, Gen86, Gen87]. Convenient [Fra85]. Conversion [Dav83b]. Coördination [Hau79]. Cope [Ano69f, Ano69d, Ano69i, Ano69e, Ano69-55, Ano74a, Cop92f]. Copied [Bar77]. Copland [Ano69l]. Coppée [Ros95a]. Copper [Mor62]. Cordaites [Les78b, Les79]. Corpuscles [Ryd94]. Correct [Phi83b]. Correction [Cla83g]. Corrections [Ano97f]. Correlations [Cha71e, Cha71u]. Correspondence [Ano99c, Ano99d]. Corundum [ST82, Gen73]. Coryphodon [Cop77l]. Co’s [Gen74b]. Cosalite [Koe85]. Cosmical [Cha69c, Cha71e, Cha73a, Cha74b, Cha74c, Cha80f, Low69a, Low69b]. Cosmogony [Kir79a]. Cosmography [Phi80a]. Costa [Bri97a, Cop93b, Gab75]. Cotylosauria [Cop95c, Cop96a]. Counties [Ash78, Fra75a, Gen74b, Hal75a, Les74, Pri77b, Ste80a, Ste81a, Ste85a, Ste87a, Fra75b, Les71a]. Country [Ano68d]. County [Ano83d, Ash81, Cla83g, Cop75b, Cre77, Fra86b, Ful74, Gen79, Hau77, Les71e, Les73a, Les73c, Les74, Les77a, Les78a, Les77d, Lil83, Lil86, SS78, Ste81d, Wal84, vD99, Bar77, Cla83d, Net77, PS78, Pla79]. Cours [Les81c]. Course [Bra84]. Courses [Hau77]. Court [Ano62y]. Created [Ano89q]. Creation [Cop71k, Gro79]. Creative [Cha75f]. Creek [Ano68k, Cop71p, Cop71z, Gen74b, She78, Les73b, Les74, SS78]. Cremastochilus [Hor79a]. Cremation [Hof74, Hof75]. Creodonta [Cop80b]. Cresson [Fra77a]. Cretaceous [Ano69m, Cop69f, Cop69m, Cop69h, Cop71a, Cop71f, Cop71o, Cop77i, Cop94c, Gab61]. Crinoid [Wil83]. crispata [Pac94]. Criteria [Cha78c]. Critical [Bri87c, Hor80a]. Critics [Bai91]. Cross [Bri89b, Vau89]. Crucial [Cha78d]. Crural [Gin78]. Crustacean [Cla83b]. Crystalline [Hal80]. Crystallography [Fra77c]. Cumberland [Ano63k, Cop75b, Jon80, Les73a, Les77a, Les78a]. Cumulative [Ryd90b]. Curculionidæ [Hor73a]. Curious [Cha73k]. Current [HK97]. Currents [Cha71p, Cha71s]. Curtis [DuB88]. Customs [Moo88, Moo89]. Cycle [Cha71v]. Cyclical [Cha71g, Cha71f, Cha71n, Cha74d, Cha74e]. Cymatogaster [Hub94]. Cypher [Cha73k].

D [Ano69-55, Bel73, Da 85]. D. [Sel73, Sto85]. D.D [Agn85, Ano63n, Mer90].

D.D. [Ley60]. Daily [Cha71h]. Dakota [Cop77g, Cop77n]. Damp [Cha81a]. Daniel [Mer90]. Danish [Rin85]. Dark [Cha78a]. Darlington [Ano63m, Les77d]. Datames [Han88]. Date [Mor62]. Day [Cha79b, Cop88a]. Dead [But99]. Death [Hou90]. Debituminisation [Les71c]. Dec. [Ano69v, Ano69u, Ano76j, Ano80h, Ano82f, Ano85-74]. Decaying [Woo71a]. Decease [Ano90f]. December [Ano50i, Ano50g, Ano50h, Ano50f, Ano52k, Ano52i, Ano52j, Ano52h, Ano53g, Ano53f, Ano54f, Ano54g, Ano55h, Ano55g, Ano56g, Ano56f, Ano57k, Ano57j, Ano58h, Ano58g, Ano59j, Ano59i, Ano60i, Ano60h, Ano61i, Ano61h, Ano62j, Ano62i, Ano63x, Ano63v, Ano63w, Ano63u, Ano65h, Ano65i, Ano66g, Ano66f, Ano67f, Ano68u, Ano68t, Ano69x, Ano69n, Ano69w, Ano71m, Ano71n, Ano71p, Ano71o, Ano73h, Ano73g, Ano74h, Ano74g, Ano75h, Ano75g, Ano76k, Ano77h, Ano77g, Ano78h, Ano78g, Ano79g, Ano79f, Ano80i, Ano81j, Ano81i, Ano82g, Ano83i, Ano83h, Ano84f, Ano86j, Ano86h, Ano86i, Ano87i, Ano87j, Ano87h, Ano88f, Ano88e, Ano89u, Ano89t, Ano90j, Ano90i, Ano91g, Ano91f, Ano92g, Ano92f, Ano93h, Ano93i]. December [Ano94m, Ano94l, Ano95h, Ano95g, Ano96j, Ano96i, Ano97m, Ano97l, Ano98h, Ano98g, Ano99h, Ano99i]. Declaration [Hay98b]. Declination [Fra77b]. Decomposition [Smi77c]. Deer [Smi79]. Defects [Hau88]. Defensor [Ano69d]. Definite [Cla87]. Degenerated [Bla89]. Degree [Mil97]. Delesse [Fra76]. Delivered [Str76]. Demonstrate [Hau99]. Demonstration [Ano96f]. Density [Cha71b]. Dentition [Cop71u, Cop88b]. Department [Out74]. Deposition [Dav83b]. Deposits [Cop86h, Hal75b]. Derivative [Smi78a]. Derivatives [Mar78, Smi77a, SKHE87]. Dermestidæ [Jay82]. Description [All89a, All90a, All90c, All90b, Ano73c, Ash77b, Cop71n, Cou93, Han88, MKAP71, Pea61, Rei71, Ric98]. Descriptions [Cop71b, Cop71c, Cop77b, Cop77c, Cop78a, Cop86i, LeC78, Mee69, St.69, Win69c]. Descriptive [Phi78]. Desor [Les83c, Les82]. Destruction [Cre75]. Destructive [SM79]. Detailed [Cha78a, Lil83]. Detection [Car85b, Fra94b]. Determination [SKHE87]. Determine [Bri77b]. Determining [Rit93]. deux [Dug88]. Development [Cha75a]. Device [Fra85, Sel73]. Devonian [Cha78a, She78]. Dialect [Bri86b, Bri93b, Bri95d, Bri96, Hof86c, Hof89b]. Dialects [Bri92c]. Diamantiferous [Der79b]. Diameter [Cha79b]. Dibenzyl [Gre79a]. Dichlor [Mar78]. Dichlor-Salicylic [Mar78]. Dichlorsalicylic [Smi77b]. Dicotylinæ [Cop88a]. Didymodus [Cop84a]. Died [Pot95]. Diego [Bri87c]. Digger [Hof75]. Dihalogen [SKHE87]. Dinner [Ano89o]. Dinosauria [Cop71v]. Dinosaurian [Cop77d, Cop92f]. Dionysos [Dol93]. Dioxyethyl [Gre79b]. Dioxyethyl-Methylene [Gre79b]. Direction [Cla87]. Dirt [Woo76]. Dirt-Banks [Woo76]. Disaster [Bla77]. Discourse [Sel99, Str76]. Discovered [Cop69j, Cop71i, Les77c, Les77d, Ste96]. Discoveries [Phi80c]. Discovery [Dra77, Gin78, Mar71, Pac86, Spe81]. Discussion [Bla76, Con96, Hau89, Net77]. Diseases [Loc83]. Disputed [Har76]. Dissemination [Ano61d]. Dissociation [Cha77b]. Distance [Cha71j, Cha73b]. Distillation [Day98, SM79]. Distinctions [Rot83].

Distribution [Bla71, Blo73, Cop71q, Cop83d, Dav83a]. District [Ano76f]. Division [Hau77, Hau77]. Divisions [Les78b, Mat98a, Mat98b, Mat99]. do [Cop94a]. Dodecahedron [Mor97]. Dolomitic [Les78a]. Domestic [Sto82, Sto86c, Sto87b, Sto88a, Sto88b, Sto88c]. Domestica [Sto82, Wil81, Sto86c]. Dominating [Bai94]. Dopplerite [Lew82b]. D’Orbiney [Les68a]. down [Ash77a, SS78]. Downward [Blo74]. Dr [Muh83]. Dr. [All86, Ano61c, Ano62d, Ano63m, Bri80, Bri95a, Gen74c, Mee69, Rus84, Rus91, Wil78]. Drainage [New82]. Draper [Bar87, Ham82, Bar83a]. Draught [Hay98b]. Drift [Wil98]. Dromatherium [Osb87]. Dry [Eme71]. DuBois [Pat82]. Duchy [Kra94]. Dudley [Pot95]. Duméril [Bau92a, Bau92b]. Dunning [Les73b]. during [Cop77o, Cop78b, Cop82b, Hay98a, Hay99, Kir79b, MS89, Wil75]. Dust [Wha94]. Dutch [Hof94]. Dweller [Cus96]. Dyke [Lew85]. Dykes [Fra84]. Dynamic [Cha75b, Hau89]. Dynamics [Bac94a]. Dynamo [HT78]. Dynamo-Electric [HT78].

E. [Ano69-55]. Earle [Gar87c]. Early [Jam95, Les83a, Law88]. Earth [Abb98, Ker73]. Earthquake [Phi79b]. East [Ano63k, Ano68d, Ano89i, Ano89j, Les71c, Ano89i, Ano89j, Les71b, Les73d]. Eastern [Hau77, Cop77m, Hal80, Les91c, Woo69b]. Eckfeldt [Dub71a]. Eclipse [Bar78, May69]. Ecuador [Bro84]. Edited [Sto85]. Editions [Bri87c]. Edouard [Les82, Les83c]. Edward [Ano63n, Fra95b, Har71a, Rob71]. Effect [Cre75, Hau71]. Effects [Key88, Ryd89c]. Efficiency [HT78]. effodiens [Ano69j]. Efforts [Hau99]. Egg [Ryd93b]. Egypt [Del74, Ste96]. Egyptian [Ano63d, Les81b, Les81a, Les91a]. Egyptology [McC81]. Eight [Kir77a]. Eighth [Cop69b]. Elasmobranch [Cop84a]. Elastic [Cha73b]. Elbow [Cop92b]. Electric [HT78]. Electrical [Out74]. Electricity [Hou90, Sel99]. Electro [Cha85c]. Electrolysis [SKHE87, War91]. Electrolytic [Smi78b]. Electromotive [Bar83b]. Element [Ano63c, Les81b, Les81a]. Elements [Bri92a]. Elevations [Bar77, Dut71]. Eleventh [Cop79a]. Eli [Rot86, Day83]. Elias [Agn85]. Elk [Ash81, Ash78]. Ellipsoid [Cha71o]. Ellis [Kön77b]. Embreeville [Les71b]. Embryological [Con96]. Embryology [Pac85]. Emerson [Ame82, Rus91]. Emydidæ [Cop69f]. Emys [Bau92a, Bau92b]. Ending [Kir79b]. Energies [Cha80f]. Energy [Ryd93a, Sel99]. Engineer [Rob71]. Engineers [Car85a]. England [Cha74a, Hay99]. English [Cha60, Fra78c, Kea78, Law88, Phi83b, Taf61, Taf62]. Enoch [Les91a]. Enos [Les91a]. Entrances [Hau88]. Environment [Cla87]. Eocene [Cop71b, Cop71-27, Cop69c, Cop71c, Cop71r, Cop71l, Cop71y, Cop73c, Cop73d, Cop73e, Cop75a, Cop81b, Cop82b, Cop82d, Cop83c, Cop83a, Mar73]. Epitaph [Kin88]. Epoch [Kir78, Cop82d, Cop77n, Cop83f, Cop86f]. Epochs [Pri76a]. Equations [Hag83]. Equilibrium [Cha73b]. Equivalent [Ano83d]. Equivalents [Cha69b, Win69a, Win69b]. Erected [Ano85-72]. Erie [Spe81]. erläutert [Kra94]. Eroded [Ste80b]. Erosion [Les73c]. Errata [Ano59d, Ano59e, Ano60d, Ano69b, Ano75c, Ano77c, Ano78c, Ano85-70, G.86]. Erratum [Ano81e, Ano87c, Rat80]. Errors [Hag82]. Eruption [Wha94]. espèces [Dug88]. Esq [Goo69]. Esq. [Bri88b]. Essay [Jam95]. Essays [Ano95d]. Estimation [Cha73b, Cha80b, Kön77a, Smi78b]. etc [Mar74, Pac93b]. etc. [Bri81a]. Etheostomine [Cop69g]. Ethics [Mor94]. Ethnic [Bri97a]. Ethnography [Bri95c]. Ethnologic [Bri85b, Bri89a]. Ethnology [Boa87]. Ethylene [Gre79a]. Etruscan [Bri90b]. Etruscans [Bri89a]. Etrusco [Bri92a]. Etymologies [Les91a]. Eucalyptus [Pep96]. Eugen [Bai91]. Euphoriæ [Hor79b]. Europe [Cha71w]. European [Cha61, Cha71i]. Evening [Str76]. Evidences [Fra95a]. Evolution [Bai96, Cha74b, Con96, Cop89a, Ryd93a]. Exactitude [Kön77b]. Examination [Cre74, Day83, Out74, Sel73]. Examinations [May69]. Exception [Edd82]. Exfoliation [Fra74a]. Exhibited [Les80a]. Exhibition [Phi79a, Phi79c, Ros95b]. Existence [Cop71v, Kir87b]. Exodus [Les81b, Les81a]. Expansion [Mar74]. Expedition [Cop94a, Cou93, Nan97, SO87]. Expeditions [Cop78b]. Experiment [Cha85b]. Experimental [Bri77b]. Experiments [Cre74, Pec68]. explained [Kra94]. Exploded [Cre74]. Exploration [Cop77o, Cus96]. Explorations [Rin85]. Explosion [Cre74]. Explosions [Cha81a]. Explosive [Cha71q]. Exposed [Lil83, Lil86]. Expressing [Hau79]. Extension [Bac91a]. Extinct [Bri86b, Cop69k, Cop71t, Cop71n, Cop71b, Cop71q, Cop71z, Cop71j, Cop71h, Cop73a, Cop77b, Cop78a, Cop85b, Cop85d]. Extremes [Blo74]. Eye [Ryd90a].

F [Ano76f, Hau77, LeC73, Ros95c]. F. [Cha71a]. Facial [Sto87b]. Factor [Ryd93a]. Factors [Bai96, Con96]. Fails [Cha71i]. Fairmount [Phi79a, Phi79c]. Fall [Cha75e]. Falls [Bla85]. False [Cop92b]. Families [Cop71f, Cro73, Pac93b]. Family [Ano63c, Hor78a, Pac83b, Pac93a]. Farenheit [Bro84]. Father [Rus86]. Fault [Les71b, Lym95a, Cla83g, Lym93, Lym95b]. Fauna [Cop69c, Cop79b, Cop83a, Cop85c, Cop86g, Cop95a, Cop96b]. Fayette [Wal84]. Features [Les68a]. Feb [Ano69-55]. Feb. [Ano78i, Ano78j, Ano83k, Ano83j]. February [Wil75, Ano50j, Ano50k, Ano51f, Ano51e, Ano52m, Ano52l, Ano53i, Ano53h, Ano54i, Ano54h, Ano55j, Ano55i, Ano56h, Ano56i, Ano57m, Ano57l, Ano57d, Ano58j, Ano58i, Ano58c, Ano59l, Ano59k, Ano60k, Ano60j, Ano61j, Ano61k, Ano62l, Ano62k, Ano63-28, Ano63z, Ano63-27, Ano63y, Ano65k, Ano65j, Ano66i, Ano66h, Ano67h, Ano67g, Ano67b, Ano68w, Ano68v, Ano69-28, Ano69z, Ano69-27, Ano69y, Ano71t, Ano71r, Ano71s, Ano71q, Ano73j, Ano73i, Ano74j, Ano74i, Ano75j, Ano75i, Ano76m, Ano76l, Ano77j, Ano77i, Ano79i, Ano79h, Ano80k, Ano80j, Ano81l, Ano81k, Ano82i, Ano82h, Ano84g, Ano84h, Ano85-76, Ano85-75, Ano86l, Ano86k, Ano87l, Ano87k, Ano88h, Ano88g, Ano89v, Ano89w, Ano90l, Ano90k, Ano91i, Ano91h]. February [Ano92i, Ano92h, Ano93k, Ano93j, Ano94o, Ano94n, Ano95i, Ano95j, Ano96e, Ano96f, Ano96l, Ano96k, Ano97o, Ano97n, Ano98j, Ano98i, Ano99k, Ano99j, Hau99, Kir77a, Wil78]. Feet [Cop85d, Les91c, Pac86]. Feldspar [Car91]. Felis [Sto82, Sto86c, Wil81]. Ferdinand [Les88]. Ferns [Dav83a]. Few [God96, Mee69]. Fiat [Cha75f]. Fibro [Bra84]. Fibro-Vascular [Bra84]. Field [Ano73c, Cou93, Ste81b, Les73d]. Field-Note [Ano73c]. Fifteen [Ant76]. Fifth [Cla84, Cop85c]. Fiftieth [Ano94f, Ano94i]. Fig [Ano63j, Ano63k, Ano68j, Ano68n, Ano68o, Ano68p, Ano69g]. Figures [Dol93]. Finding [Mer97a]. Fine [Les71b]. Fire [Cha81a]. Fire-Damp [Cha81a]. Fireballs [Kir77a, Kir79b, Kir87b]. First [Cop83a, Hay99, Net77, Phi88, Cla84, Wil75]. Fish [Cla83c, Day98]. Fish-Plate [Cla83c]. Fishes [Cop69j, Cop69e, Cop69l, Cop69m, Cop71f, Cop77h, Cop77i, Cop78b, Cop94a, St.69]. Fissipedia [Cop82c]. Fixed [Ano77c, Kir77b]. Flaccus [Kin88]. Flash [Hou86]. Flat [Cop73c]. Flat-Clawed [Cop73c]. Flinders [Ste96]. Flinders-Petrie [Ste96]. Flora [Cla83d, Hei91, Les77b]. Floras [Hei92]. Florida [Cus96, Sch78]. Florissant [Scu94]. Flow [Bri76a, Bri77a]. Fluids [Sel73]. Folk [Bri94b, Hof89a, Phi88, Phi92]. Folk-Lore [Bri94b, Phi88, Phi92]. Folk-Medicine [Hof89a]. Following [Hou90]. Fond [Fra76]. Foot [Cha71a, Cop81a]. Foot-Mark [Cha71a]. Footed [Cop73e]. Footprint [Bri87a]. Foramina [Cop80a]. Force [Bar83b, Cha71c, Cha71e, Cha74f, Cha75c, Cha76b, Cha77a]. Forces [Cha73b, Cha73f, Cre75, Low69b]. forcipata [Ano69j]. Forecast [Cha85b]. Foreigners [Ste96]. Forestry [Rot94b]. Forests [Rot94a]. Forgery [Fra94b]. Fork [Cop83d, Kön77b]. Form [Cop71n, Key88, Ryd93b]. Formamations [Cop77b]. Formates [War91]. Formation [Ano69m, Cop71f, Cop78a, Cop80c, Cop82e, Cop83b, Cop85c, Gab61, Gre79a, Kir71b, Les78b, Pec68, Uhl88, Cop83d, SO87, Sco98]. Formations [Cop77c, Win69c]. Formerly [Bri86b]. Forms [Cop71k, Cop71h, Fra79a, Ryd94]. Forschern [Rol94]. Forsyth [Pep83]. Fort [Hay98a]. Forth [Bri85c]. Fossil [Cop71g, Fra79a, Les73b, Mer97a, St.69]. Fossile [Les81c]. Fossils [Cop71w, Mar71, Mee69, Phi97b, SO87, Smi96, Win69c]. Foulke [Les68b]. Found [Cop71a, Cop77j, Dub71b, Mor97, Smi79]. Foundation [Ano94f, Ano94i]. Four [Gra58a, Ric98]. Fourteen [Les71e]. Fourth [Cop69c, Cop83b, Cop95a]. Fowl [Ryd93b]. Fox [Hau77]. Fragment [Bac91a, But99, Mor96]. France [Cha71a]. Francisco [Cha71f]. Franklin [Ano69k, Ano90f, Vau90]. Frazer [LeC73]. Frederick [Fie71]. Free [Wil94]. Freeport [Ste81d]. French [Ant76, Dug86, Dug88, Kön77b]. Fresh [Cop69j, Cop69l, Cop69k, Sto86a, Sto87a, Sto90, Woo69b, Woo71b]. Fresh-Water [Cop69k, Sto86a, Sto87a, Sto90]. Frieze [Ang90]. Front [Ano65c, Ano69c, Ano71b, Ano73d, Ano76e, Ano85-71, Ano89h, Ano90e, Ano92d, Ano93e, Ano94g, Ano94h, Ano97g, Ano99e]. Fruit [Les79, Woo71a]. Fuegian [Bri92b]. fun [Hof94]. Function [Bri97b]. Fund [Cha74g, Str76]. Fundamental [Cha76b]. Funeral [Moo88]. Fungus [Les77d]. Furnace [Bri77b, Ano63j, Les71b]. Furnaces [Woo76]. Further [Bri92c, Bri92b, Cha71j, Cha75b, Cha75c, Cha77a, Cha79c, Lym94c].

G [Sto85]. Galera [Ano69h]. Galvanic [Cre75]. Galvanometer [Bar75]. Gamuts [Cha73c]. Gas [Les91d, Phi97a, Sad76, Sad77, Sad78, Sad97, Smi88]. Gas-Analyses [Sad78]. Gas-Wells [Les91d]. Gen. [Bar77]. Genealogical [Gro99a]. Genera [Cop80b, Cop86a, Hor73b, Hor73c]. General [Car85a, Cha68b, Cha71k, Fra86b, Rin85, Rit93, Ano99c]. Generic [Les78b]. Genesis [Pec98, Phi97a, Ryd90b, Ryd93b, Sad97, Mor96]. Genus [All89a, Bau92a, Bau92b, Cla83e, Cop69f, Cop71t, Cop71n, Cop71r, Cop71s, Cop84a, Cop92g, Gin78, Pac83b, Ric98, Scu97, Cop93a]. Geographical [Gab61, Les84]. Geography [Bla71, Gra58b, Gra58a, Les76]. Geological [Ash81, Cla83a, Cla83g, Dra97, Fra86c, Hal76, Hal77, Les71a, Les76, Les78a, Les83b, Les77b, Ste75b, Ste81a, Ste85a, Ste87a, Win69a, Win69b, Ano76f]. Geologist [Ano76f]. Geology [Ash97, Bla71, Der79a, Der79b, Fra86b, Hay68b, Hay69a, Pec98, Ste75a, Ste79, Ste80a, Ste87b, Whi81, Whi82]. George [Ano62d, Ano62e, Ano99c, FGG^{+}80, Sel86, Bri94a, Har80]. Georgia [Cop77f]. German [Hof89b, Kra88, Kra94, Rol94]. Germanized [Kra94]. Germans [Hof89a, Mer97b]. Gettysburg [Fra74a]. geworden [Kra94]. Giant [Hei92]. Gibbons [Hub94]. Gigantic [Mar73]. Gilboa [SS78]. Gildas [Law88]. Giles [Ste87a]. Gilsonite [Day98]. Glacial [Hal75b, Hal75a, Pri76a]. Glaciation [Bra86]. Glands [Ryd89a, Ryd90a]. Glimpses [Fur96]. Glosso [Sto88b]. Glosso-Pharyngeal [Sto88b]. Glows [Bla85]. Glyceric [Sad75]. Gneiss [Car91]. Gods [Bri81b]. Gold [Ano61d, Dav83b, Pri95]. Good [Rot83]. Goodwin [Mer90]. Gorge [Lil83, Lil86]. Gouverneur [Rus91]. Government [Ano63c, Hay99]. Gowen [Vau90]. Graduation [Les59]. Grammar [Bri69, Bri84, Bri85c, BB69, Sto85]. Grammatic [Hof89b]. Gramophone [Ano87e, Hou88]. Grand [Gar87a]. Grande [Cop94a]. Grapeville [Les91d]. Gravels [Cha81b]. Gravitating [Cha71p, Cha73b, Cha75c, Cha75d]. Gravitation [Cha78b]. Gravity [Cha69c, Cha73i]. Gray [Ryd90a]. Great [Lew85, New82, Spe81, Lym93]. Gredients [Hau71]. Green [Cop71b, Cop69e, Cop71s]. Greenland [Rin85]. Greensboro [Les71d]. Griffin [Les84]. Gross [Wil81, Da 85]. Grosso [Cop87c]. Grote [G.86]. Group [Cla83b, Cla84, Cop71c, Mil97, Mil98, Cla83e, Hay69b, Ste81c, Ste82, Win69a, Win69b, Win69c]. Groups [Gro95a, Mil97, Ste75b]. Grown [Woo69a]. Growth [Bra84, Rot94b, Woo71b]. Gshicht [Hof94]. Guatemala [Bri84, Bri85b, Bri87b, Sto85]. Guetares [Bri97a]. Gulf [Cus96, Lil83, Lil86]. Guslar [Kra94]. Guslarenlied [Kra88, Kra94]. gusle [Kra88].

H [All86, Bar77, Bri88b, Pot95]. H. [Cop87c]. Habitual [Ryd89c]. Haida [Boa91]. Haldeman [Bri81a]. Half [Gen74b, Les74]. Hall [Ano89g, Phi79a, Phi79c, Ano89p, Str76]. Hamilton [Cla83e]. Hammer [Pea63]. Hancock [Han88]. Hand [Fra85, Pea63]. Hand-Hammer [Pea63].

Hands [Fra95a]. Handwriting [Fra86a]. Harbor [Hau88]. Harden [Les80b]. Harlan [Cop69d]. Harmonies [Cha73a, Cha78d]. Harrisburg [Ano63k, Les73a]. Harrison [Sel75]. Hartranft [Les78a]. Hartshorne [Les78a]. Haupt [Bar77]. Hauxwell [Cop86b]. Haverford [Sha83]. Hayden [Les88, Mee69]. Hays [Bri79]. haysii [Ano69h]. Hayti [Ack69]. Health [Bla75]. Heat [Blo73, Edd82, Mar74]. Hebrew [Les81b, Les81a, Les86b, Les91a, Mor91]. Held [Ano91c, Ano94f, Fra86c]. Helderberg [SS78]. Heller [Ano71c]. Hematite [Les74]. Hemicaulodon [Ano69j]. Hemileucidæ [Pac93b]. Hemileucini [Ano75c, Gro74]. Hemp [Woo69a]. Henry [Ano99d, Hay98a, Ang90, Ano95d, Bar83a, Bar87, Rob83, Rog80, Ros95a, Vau85]. Hepiali [Ano75c, Gro74]. Hercegovina [Kra86, Kra88]. Herderite [Gen84]. Hereditary [Oli94]. Herpetological [vD98]. Herpetology [Cop69b, Cop69p, Cop77p, Cop79a, Cop85a, Cop85f, Cop86j]. Herschel [Fie71]. Herschell [Cha71l]. Herzogtum [Kra94]. Hidalgo [Cop86h]. Hieroglyphic [Ano68l, Ano68m, Cha71a]. Hieroglyphs [Ano63d]. Hill [Ano68l, Ano68m, Fra82]. Him [Bri85c]. Hippotherium [Cop89b]. Histeridæ [Hor73d]. Historic [Bri87d]. Historical [Hau99, Les68a]. Histories [Pac93b]. History [Bri94b, Cop69c, Cop69m, Cop80c, Cop82b, Cop82e, Cop83b, Cop87b, Cop96a, Cou93, Hay98b, Law88, Les81b, Les81a, Pac93a]. Hitchcock [Ano63n]. HNO [Smi88]. Holiday [Moo89]. Holman [Sel73]. Home [Ano63j]. Hon. [Pot95]. Honorable [Rot86]. Horace [Sti69, Str76]. Horizon [Fra82, Hag82]. Horizons [Lym94d]. Horologium [Sac95]. Horses [Cop86g, Far96]. Hotel [FGG^{+}80]. Hudson [Hal80]. Human [All84, Bri87a]. Humboldt [Bri85c]. Humphreys [Car85a]. Hundred [Ano94f, Ano94i]. Hundredth [Ano90f, FGG^{+}80]. Hunt [Gen74c]. Huntingdon [Gen74b, Les74]. Huron [Bla77]. Hydrobiini [Hor73b]. Hydrochloric [Dav83b, Gre80b]. Hydromice [Hal80]. Hygienic [Pep96]. Hyner [Cha78a]. Hypenoid [Gro95a]. Hyperostosis [MKAP71]. Hypoglossal [Sto88c]. Hypothesis [Cha78c]. Hypotrichous [Sto86b]. Hypsometrical [Fra83]. Hyrachyus [Cop73a]. Hyrtl [Ano71f].

Ibrâhim [Kra88]. Ichthyology [Cop69a, Cop73b, Cop74b]. Idaho [Cop69j, Cop69k, Cop71g]. Idea [Rot94b]. Ideal [Ano63k]. Identification [Dor96]. Iguaninæ [Cop86d]. II [Cop87c, Ano63e, Ano69g, Cha73a, Cha80a, Cha81d, Gro99c, Koe74, Smi78a, Ste75a]. III [Ano63f, Ano69h, Cha80a, Cha81e, Gen74a, Pac85]. Ikonomatic [Bri86d]. Illinois [Cop77m, Cop97, Dub71b, Gra58b]. Illuminating [Fra80]. Illuminative [Mer97b]. Illustrate [Ano68d]. Illustration [Ano63d, Ano63e, Ano63f, Ano63g, Ano63h, Ano63i, Ano63j, Ano63k, Ano63l, Ano68c, Ano68d, Ano68e, Ano68f, Ano68j, Ano68i, Ano68k, Ano68l, Ano68m, Ano68n, Ano68o, Ano68p, Ano68g, Ano68h, Ano69f, Ano69d, Ano69g, Ano69h, Ano69i, Ano69e, Ano69j, Ano71d, Ano71e, Ano85-72, Ano89i, Ano89j, Ano89l, Ano89k].

Illustrations [Cha71q, Cha75f, Cha77a]. Illustrative [Str76]. im [Kra94]. Implement [Ano69j]. Implements [Pea61, Phi80c]. Important [Cla83g, Les91c, Phi83a]. Improved [Bri76b]. Improvement [Fra83, Hau88, Koe74]. Inclination [Dav66, Hag82]. Inclinations [Kir89]. Included [Cha78a]. Including [Gab61]. Incorporation [Bri86c, FGG^{+}80]. incrassatus [Cop92f]. Indefinite [Cla87]. Independence [Hay98b]. Independent [Phi89a]. Index [Ano81f]. India [Bla71]. Indian [Ano68l, Ano68m, Ano68g, Cop85c, Dra97, Gab75, Gar87b, Hof86a, Wal84]. Indiana [Kir68, Bar77]. Indians [Bri85b, Gat85b, Gat86, Gat90, Gat91, Hof75, Mer95, Pea61]. Indo [Cha61]. Indo-European [Cha61]. Industrial [Phi79a, Phi79c]. Inflorescence [Dol93]. Influence [Cha71m, Eme71, Woo71b]. Influencing [HT78]. Information [Ano53w, Ano54w, Ano58x, Ano59z, Ano60y, Ano61y, Ano62z, Ano63-55, Ano65y, Ano68-37, Ano74y, Ano75x, Ano76-27, Ano77x, Ano77y, Ano78w, Ano78x, Ano80y, Ano80z, Ano81z, Ano81-27, Ano83z, Ano83-27, Ano84u, Ano85-86, Ano85-87, Ano86-29, Ano86-30, Ano87y, Ano87z, Ano88u, Ano88v, Ano89-36, Ano89-37, Ano90t, Ano90u, Ano91v, Ano91w, Ano92u, Ano92v, Ano93x, Ano93y, Ano94-28, Ano94-29, Ano95w, Ano95x, Ano96y, Ano96z, Ano97-28, Ano97-29, Ano98u, Ano98v, Ano99v, Ano99w]. Infusoria [Sto86a, Sto86b, Sto87a, Sto90, Sto94]. Ingersoll [Sha68]. Inhabiting [Hor80b]. Initiation [Mat98c]. Inks [Dor96, Sha95]. Inscription [Phi84b]. Insensible [Les59]. Institute [Wil94, Cha71a]. Instruction [Hau79]. Instruments [Dav66]. Insulating [HK97]. Insurance [Cha74g]. Insurrection [Tha78]. Integration [Ryd90b]. Interconversion [Fra78c]. Interesting [But99]. Interferences [Cha78f]. International [Edm97, Fra86c]. Intra [Cha73j]. Intra-Asteroidal [Cha73j]. Introduction [Gro83]. inundata [Ano69i, Ano69j]. Invertebrates [Mey88]. Investigation [Gen74b, SM79]. Investigations [Taf61, Taf62]. Iodyrite [GvR85]. Ireland [Moo87, Moo88, Moo89]. Iron [Ano71e, Cre75, Gen74b, Les71d, Les73a, Les73b, Smi77c]. Isaac [Bri79, Lei87]. Islands [Bla71]. Isleta [Gat91]. Isomorphic [Mil97]. IV [Ano63g, Ano68f, Ano69i, Cha80b, Cha81f]. IX [Ano63h, Ano68j, Cha84, Smi77a].

J [Ano76f, Hau77, Kön77b]. J. [Ano63h, Ano63j, Ano63k, Ano63l, Ano68i, Ano68h]. Jacob [Dub71a]. Jamaica [Rot91]. James [Ano69l, Cop77o, DuB88, Har95, Les86a, Rot82, Vau87]. Jan. [Ano78l, Ano78k, Ano83m, Ano83l, Str76]. January [Ano74a, Ano89q, Ano50m, Ano50l, Ano51h, Ano51g, Ano52o, Ano52n, Ano53k, Ano53j, Ano54k, Ano54j, Ano55l, Ano55k, Ano56k, Ano56j, Ano57o, Ano57n, Ano58k, Ano58l, Ano59n, Ano59m, Ano60m, Ano60l, Ano61m, Ano61l, Ano62n, Ano62m, Ano63-32, Ano63-31, Ano63-29, Ano63-30, Ano65m, Ano65l, Ano66k, Ano66j, Ano67j, Ano67i, Ano68y, Ano68x, Ano69-29, Ano69-30, Ano69-32, Ano69-31, Ano71x, Ano71v, Ano71w, Ano71u, Ano73l, Ano73k, Ano74l, Ano74k, Ano75k, Ano75l, Ano76o, Ano76n, Ano77l, Ano77k, Ano79k, Ano79j, Ano80m, Ano80l, Ano81n, Ano81m, Ano82k, Ano82j, Ano83n, Ano84i, Ano85-78, Ano85-77, Ano86m, Ano86n, Ano87n, Ano87m, Ano88j, Ano88i, Ano89y, Ano89x, Ano90n, Ano90m, Ano91k, Ano91j, Ano92j, Ano92k]. January [Ano93m, Ano93l, Ano94q, Ano94p, Ano95l, Ano95k, Ano96n, Ano96m, Ano97p, Ano97q, Ano98l, Ano98k, Ano99m, Ano99l, Kir73, Wil75]. Jargon [Gat85a]. Jasper [Mer95]. Jefferson [Hay98b]. Jersey [Cop75b, Lym93]. John [Bri80, Fie71, Les80b, Pri74, Cop86b, Cop88a, Dol94, Fra77a, Ham82, Hor83, LeC73, Pep83]. Joint [Fra95a]. Joints [Cop92b]. Joly [Sac96]. Joseph [Rog80, Sel75, Ley60, Oli93, Rus92, Sha68]. Journal [Hay98a, MS89]. Journals [Cou93]. Journey [Hil96]. Jr [Hau77]. Jr. [Sel75, Sti69]. Jugal [Sla95]. July [Ano50n, Ano52c, Ano52q, Ano52p, Ano53l, Ano54l, Ano55m, Ano56l, Ano57e, Ano57p, Ano58m, Ano59o, Ano60n, Ano61n, Ano63-34, Ano63-33, Ano65n, Ano66l, Ano67k, Ano68z, Ano69-34, Ano69-33, Ano71z, Ano71y, Ano73m, Ano74m, Ano75m, Ano76p, Ano77m, Ano79l, Ano80n, Ano81o, Ano82l, Ano83o, Ano84j, Ano85-79, Bar78, Kir77a]. June [Ano50o, Ano51i, Ano52r, Ano53m, Ano54m, Ano55n, Ano56m, Ano57q, Ano58n, Ano59p, Ano60o, Ano61o, Ano62o, Ano63-36, Ano63-35, Ano65o, Ano66m, Ano67l, Ano68-27, Ano69-36, Ano69-35, Ano71-27, Ano71-28, Ano73n, Ano74n, Ano75n, Ano76q, Ano77n, Ano78m, Ano79m, Ano80o, Ano81p, Ano82m, Ano83p, Ano84k, Ano85-80]. Jupiter [Cha74e, Kir69b, Kir71a]. Jupiter-Cyclical [Cha74e]. Justice [Pri74].

Kansan [Wil98]. Kansas [Cop71o, Ano68k, Ano68l, Ano68m, Cop69h, Cop71a, Cop71f, Cop71d, Cop71x, Cop77i]. Kauterskill [She78]. Kawi’a [Hof86c]. Keim [Bri94a]. Kennedy [Cop71i]. Kentucky [Les73d]. Kept [Hay98a]. Key [Cus96]. Keynote [Bai94]. Ki [Bri89b]. Kiche [Bri81b]. Kinetogenesis [Cop92d]. King [Cop69j]. Kings [Les81b, Les81a]. Kingsmill [Cla83f, Les83d]. Kirkbride [Cur85]. Kittatinny [Hal75a]. Kneass [Gra84]. Knowledge [Ano94f, Cop79b, Cop83f, Cop85c, Cop93b, Cop96b, Hor73a]. Known [Cop77h, Cop77i, Cop92c]. Köprülü [Kra94]. Krakatoa [Wha94]. Krauth [Muh83]. Kwakiutl [Boa93]. Ky. [Ano68d].

L [Hor83]. Laboratory [Bar75, Day83, Gen87, Gen73, Gen74a, Gen77, Gen82, GvR85, Gen86, Koe74, Kön77a, Les78a, Mar78, Sad75, Sad76, Sad77, Smi77a, Smi77b, Smi77c, Smi78a, Smi78b]. Lacertilia [Cop92h]. Lachneides [Ano75c, Gro74]. Lachrymal [Ryd90a]. Lackawanna [Bra86]. Lælaps [Cop92f]. Laemanctus [Ano69e]. Laemanctus-alticoronatus [Ano69e]. laevidens [Ano69h]. Lagoa [Pac94]. Lake [GvR85, Gra58a, Spe81]. Lakes [New82, Spe81]. Land [Coo79, Les77c]. Land-Shells [Coo79]. Landa [Bri87c]. Lands [Ano71e]. Language [Boa93, Bri69, Bri73, Bri84, Bri85b, Bri87b, Bri88a, Bri90a, Bri92e, Bri92d, BB69, Gat76, Gat77, Gat78, Gat80, Gat85a, Hal83, Hof86b, Hof86c, Sto85, Wak90]. Languages [Boa91, Bri85c, Bri86a, Bri86c, Bri92b, Bri92f, Gab75, Hay68a, vHB85]. Lantern [Bar75]. Laplace [Kir79a]. Laramie [Ste82]. Large [Cla83b, Cla83c, Kir87a]. Larvæ [Pac93b]. Last [Phi83b]. Late [Bri79, Wil88, Ano60c, Ano69l, Les80b, Rot86, Rot89, Sel86]. Latent [Mar74]. Latitude [Doo97, Gra58b, Gra58a, Sha83]. Laurel [Ste81d]. Laurentian [Car91]. Law [Bri88b, Edd82, Les71c, Str76]. Laws [Ano89m, Taf61, Taf62]. Lea [Lei87]. Lead [SKHE87]. Lec [Hor91]. LeConte [Hor83]. Lectures [Rot77]. Lee [Ano99d, Ste80a, Ste81a]. Legacy [Hei92]. Legislature [Ano89q]. Lehigh [Pri77b]. Leidy [Rus92, Sco94]. Leo [Les90b, Les83e]. Lepidopterous [Gro98, Gro99b, Gro99c]. LeRoy [Lil86]. Lesley [Ano68h, Ano63h, Ano63j, Ano63k, Ano63l, Ano68f, Ano68i, Ano73c, Daw63]. Lesquereux [Les90b, Les83e]. Letter [Les83e, Pat75, Phi83b, Wil78]. Letters [Ano63e]. Level [Dav66, Hau77]. Levels [Car77, Les76]. Lewis [Cou93]. Lewisburg [Les74]. Libellulidæ [God96]. Library [Ano99d, Ano99c, Hay98b]. Libyan [Bri90b, Bri92a]. Lick [Ano68d]. Life [Cha69e, Cha74g, Fie97, Pac93a, Pac93b, Rus86, Rus91, Rus92, Str76]. Life-Insurance [Cha74g]. Light [Cha69c, Cha73c, Cha73i, Cha80b]. Lightning [Hou86, Ran81]. Lignitic [Ste75b]. Limbs [Cop92d]. Limestone [Les77a, Les78a, Ash77a]. Limestones [Gen74b, Hal80]. Limiting [Cha78b]. Limits [Kir85b]. Limonites [Fra75b]. Limulus [Pac85]. Line [Les73a, Les74, Bar77, Hau77]. Lineal [Bri85a]. Lines [Hau77]. Linguistic [Bri95b, Bri98a]. Linguistics [Bri93a]. Lisbon [Cha71g]. List [Bar77, Cha80d, Cop86i, Gro74, Mee69, Phi78, Phi89b, Ano75c]. Lithologie [Fra76]. Lithology [Fra78a]. Little [Cop77h, Cop77i, Cop92c, Ano63j]. LL.D [Wil88, DuB88]. LL.D. [Agn85, Ang90, Ano63n, Ham82, Har80, Har95, Lei87, Les88, Mer90, Ros95a, Rot86, Rus86, Rus92]. Locomotive [Cre74, Woo76]. longidens [Ano69j]. Longitude [Gra58b, Gra58a]. Looking [Ano89i, Ano89j, Ano89l, Ano89k]. Lore [Bri94b, Phi88, Phi92]. Louis [Ant76, Ant76]. Loup [Cop83d]. Love [Bri86a]. Lower [Cha78a, Ano63j, Ash77a, Cop82b, Der79a, Fra75b, Fra79a, Les77d, Spe81]. Lowest [Cop81b, LC91]. Loxomylus [Ano69j]. Ludlow [Vau87]. Luminosity [Mar74]. Luminous [Cha75c, Cha80f]. Lunar [Cha71n, Cha75e, Eme71]. Lunar-Cyclical [Cha71n]. Lunar-Monthly [Cha75e]. Lungs [Cop94b]. Luxor [Ano68h]. Lyman [Fra94a]. Lyon [Gen74b]. Lysiopetalidæ [Pac83b].

M [Ano95d, Cop77j, Dol94, Kea78, Les81c, Vau85]. M. [Fra76, Kin88]. M.D [Ham82, Har80, Har95, Les88, Rus86, Rus92]. M.D. [Ano65d, Ano69l, Bel73, Bid89, Bri79, Cur85, Har71a, Hor83, Pep83, Rot89, Sto85]. Macauley [Fra95b]. Macfarlane [Les86a]. Machines [HT78]. Macrotus [All90b]. Made [Ack69, Ano96f, Car77, Cop82b, Cop86b, Hau99, Les78a, Out74, She78].

Madison [Gra58a]. Magna [Han88]. Magnetic [Cha71s, Cha75c, Cre75, Fra77b]. Magnetism [Cha68b]. Magnetizable [Ano87d]. Maintaining [Low69b]. Maisch [Dol94]. Major [Car85a]. Major-General [Car85a]. Making [Pea61]. Malay [Wak90]. Malformation [All84]. Mallophaga [Pac87]. Mammalia [Cop80a, Cop81b, Cop82a, Cop83c, Cop83e, Cop85b, Cop95b]. Mammals [Mar73, Osb87]. Man [Bai94, Bri88a]. Mangue [Bri86b]. Manual [McC81]. Manufactured [Dud86]. Manuscript [Cha73k, Cou93, Hay98b]. Map [Ano68d, Ano71e, Ano83c, Cla83g, Les73c]. Maps [Hau78, Phi80b]. Marañon [Cop69a]. Marcasite [Bro94]. March [FGG^{+}80, Kir79b, Ano50p, Ano50q, Ano51k, Ano51b, Ano51j, Ano52t, Ano52s, Ano53o, Ano53n, Ano54o, Ano54n, Ano55p, Ano55o, Ano56o, Ano56n, Ano57s, Ano57r, Ano58p, Ano58o, Ano59r, Ano59q, Ano60q, Ano60p, Ano61p, Ano61q, Ano62q, Ano62p, Ano63-40, Ano63-38, Ano63-39, Ano63-37, Ano65q, Ano65p, Ano66o, Ano66n, Ano67m, Ano67n, Ano68-29, Ano68-28, Ano69-40, Ano69-38, Ano69-39, Ano69-37, Ano71-32, Ano71-31, Ano71-29, Ano71-30, Ano73p, Ano73o, Ano74p, Ano74o, Ano75p, Ano75o, Ano76s, Ano76r, Ano77p, Ano77o, Ano78n, Ano78o, Ano79o, Ano79n, Ano80q, Ano80p, Ano81r, Ano81q, Ano82o, Ano82n, Ano83r, Ano83q, Ano84m, Ano84l, Ano85-82, Ano85-81, Ano86p, Ano86o, Ano87p, Ano87o, Ano88l, Ano88k, Ano89z, Ano89-27, Ano90p]. March [Ano90o, Ano91m, Ano91l, Ano92m, Ano92l, Ano93o, Ano93n, Ano94s, Ano94r, Ano95m, Ano95n, Ano96p, Ano96o, Ano97s, Ano97r, Ano98n, Ano98m, Ano99o, Ano99n]. Marietta [Mor97]. Marine [Cop95a, Cop96b, Smi96]. Mark [Cha71a, Gen74b, Les74]. Marks [Fra95a]. Mars [Kir69b]. Marshall [Bai91, Win69a, Win69b, Win69c]. Mary [Ash81]. Maryland [Ste79, Uhl88]. Mass [Cha73b, Kir69b, PS94, SD94, Mei86]. Mass. [Gar87b]. Matagalpan [Bri95b]. Material [Cha71c, Cre77]. Mato [Cop87c]. Matter [Ano55c, Ano57c, Ano59c, Ano62c, Ano65c, Ano69c, Ano71b, Ano73b, Ano73d, Ano76d, Ano76e, Ano81d, Ano85-69, Ano85-71, Ano89d, Ano89e, Ano89f, Ano89h, Ano90d, Ano90e, Ano92d, Ano93d, Ano93e, Ano94g, Ano94h, Ano97g, Ano99e, DuB95]. Maxwell [Cha85a]. May [Ano82q, Ano91c, Ano94f, Ano50s, Ano50r, Ano51m, Ano51l, Ano52v, Ano52u, Ano53q, Ano53p, Ano54q, Ano54p, Ano55r, Ano55q, Ano56q, Ano56p, Ano57t, Ano57u, Ano57f, Ano58r, Ano58q, Ano59t, Ano59s, Ano60s, Ano60r, Ano61s, Ano61r, Ano62s, Ano62r, Ano63-41, Ano63-42, Ano63-44, Ano63-43, Ano65s, Ano65r, Ano66q, Ano66p, Ano67p, Ano67o, Ano68-30, Ano68-31, Ano69-44, Ano69-42, Ano69-43, Ano69-41, Ano71-36, Ano71-34, Ano71-35, Ano71-33, Ano73r, Ano73q, Ano74s, Ano74q, Ano74r, Ano75r, Ano75q, Ano76u, Ano76t, Ano77r, Ano77q, Ano78q, Ano78p, Ano79q, Ano79p, Ano80s, Ano80r, Ano81t, Ano81s, Ano82r, Ano82p, Ano83t, Ano83s, Ano84o, Ano84n, Ano85-83, Ano85-84, Ano86r, Ano86q, Ano87r, Ano87q, Ano88n, Ano88m, Ano89-29]. May [Ano89-28, Ano90r, Ano90q, Ano91n, Ano91o, Ano92o, Ano92n, Ano93q, Ano93p, Ano94u, Ano94t, Ano95p, Ano95o, Ano96q, Ano96r, Ano97d, Ano97u, Ano97t, Ano98c, Ano98p, Ano98o, Ano99q, Ano99p]. Mazatec [Bri92e]. McCall [Phi81b]. McKean [Ash78]. Md. [Hau77]. Meadow [Ano63g]. Mean [Bar77, Kir71a]. Meaning [Les84]. Means [Cha71h, Out74]. Measure [Lym94b]. Measured [Ash77a]. Measurement [Bar83b, Bri97b, Dav66]. Measurements [May69]. Measures [Bri85a, Cha78a, Cop77a, Cop85e, Daw63, Les62, Les77d, Smi96, Tay87, St.69]. Mechanical [Cha69b, Cop88b, Ryd93b]. Medals [Phi78, Phi79a, Phi79c]. Medical [Moo87, Woo69a]. Medicine [Hof89a]. Medieval [Cra83, Mer97b]. Medium [HK97]. Meeting [Ano50d, Ano50c, Ano50e, Ano50i, Ano50g, Ano50h, Ano50f, Ano50j, Ano50k, Ano50m, Ano50l, Ano50o, Ano50n, Ano50p, Ano50q, Ano50s, Ano50r, Ano50w, Ano50v, Ano50t, Ano50u, Ano50-27, Ano50y, Ano50z, Ano50x, Ano50-28, Ano51d, Ano51c, Ano51f, Ano51e, Ano51h, Ano51g, Ano51i, Ano51k, Ano51b, Ano51j, Ano51m, Ano51l, Ano52f, Ano52e, Ano52d, Ano52g, Ano52k, Ano52i, Ano52j, Ano52h, Ano52m, Ano52l, Ano52c, Ano52q, Ano52o, Ano52n, Ano52r, Ano52p, Ano52t, Ano52s, Ano52v, Ano52u, Ano52z, Ano52x, Ano52y, Ano52w, Ano52-30, Ano52-29, Ano52-27, Ano52-28, Ano52-32, Ano52-31, Ano53c, Ano53d, Ano53e, Ano53g, Ano53f, Ano53i, Ano53h, Ano53k, Ano53j, Ano53m, Ano53l, Ano53o, Ano53n, Ano53q, Ano53p, Ano53s]. Meeting [Ano53r, Ano53u, Ano53t, Ano53v, Ano54d, Ano54c, Ano54e, Ano54f, Ano54g, Ano54i, Ano54h, Ano54k, Ano54j, Ano54m, Ano54l, Ano54o, Ano54n, Ano54q, Ano54p, Ano54r, Ano54s, Ano54u, Ano54t, Ano54v, Ano55e, Ano55d, Ano55f, Ano55h, Ano55g, Ano55j, Ano55i, Ano55l, Ano55k, Ano55n, Ano55m, Ano55p, Ano55o, Ano55r, Ano55q, Ano55t, Ano55s, Ano55v, Ano55u, Ano55w, Ano56d, Ano56c, Ano56e, Ano56g, Ano56f, Ano56h, Ano56i, Ano56k, Ano56j, Ano56m, Ano56l, Ano56o, Ano56n, Ano56q, Ano56p, Ano56s, Ano56r, Ano56u, Ano56t, Ano56v, Ano57h, Ano57g, Ano57i, Ano57k, Ano57j, Ano57m, Ano57l, Ano57d, Ano57o, Ano57n, Ano57q, Ano57e, Ano57p, Ano57s, Ano57r, Ano57t, Ano57u, Ano57f, Ano57w, Ano57v]. Meeting [Ano57y, Ano57x, Ano57z, Ano58e, Ano58d, Ano58f, Ano58h, Ano58g, Ano58j, Ano58i, Ano58c, Ano58k, Ano58l, Ano58n, Ano58m, Ano58p, Ano58o, Ano58r, Ano58q, Ano58t, Ano58s, Ano58u, Ano58v, Ano58w, Ano59f, Ano59g, Ano59h, Ano59j, Ano59i, Ano59l, Ano59k, Ano59n, Ano59m, Ano59p, Ano59o, Ano59r, Ano59q, Ano59t, Ano59s, Ano59v, Ano59u, Ano59x, Ano59w, Ano59y, Ano60f, Ano60e, Ano60g, Ano60i, Ano60h, Ano60k, Ano60j, Ano60m, Ano60l, Ano60o, Ano60n, Ano60q, Ano60p, Ano60s, Ano60r, Ano60u, Ano60t, Ano60w, Ano60v, Ano60x, Ano61f, Ano61e, Ano61g, Ano61i, Ano61h, Ano61j, Ano61k, Ano61m, Ano61l, Ano61o, Ano61n, Ano61p, Ano61q, Ano61s, Ano61r, Ano61u, Ano61t, Ano61w, Ano61v, Ano61x]. Meeting [Ano62g, Ano62f, Ano62h, Ano62j, Ano62i, Ano62l, Ano62k, Ano62n, Ano62m, Ano62o, Ano62q, Ano62p, Ano62s, Ano62r, Ano62u, Ano62t, Ano62w, Ano62v, Ano62x, Ano63r, Ano63q, Ano63t, Ano63o, Ano63p, Ano63s, Ano63x, Ano63v, Ano63w, Ano63u, Ano63-28, Ano63z, Ano63-27, Ano63y, Ano63-32, Ano63-31, Ano63-36, Ano63-34, Ano63-29, Ano63-30, Ano63-35, Ano63-33, Ano63-40, Ano63-38, Ano63-41, Ano63-42, Ano63-39, Ano63-37, Ano63-44, Ano63-43, Ano63-48, Ano63-46, Ano63-47, Ano63-45, Ano63-50, Ano63-49, Ano63-52, Ano63-51, Ano63-54, Ano63-53, Ano65f, Ano65e, Ano65g, Ano65h, Ano65i, Ano65k, Ano65j, Ano65m, Ano65l, Ano65o, Ano65n, Ano65q, Ano65p, Ano65s, Ano65r, Ano65u, Ano65t, Ano65w, Ano65v, Ano65x, Ano66d, Ano66c, Ano66e, Ano66g, Ano66f]. Meeting [Ano66i, Ano66h, Ano66k, Ano66j, Ano66m, Ano66l, Ano66o, Ano66n, Ano66q, Ano66p, Ano66s, Ano66r, Ano66u, Ano66t, Ano66v, Ano67d, Ano67c, Ano67e, Ano67f, Ano67h, Ano67g, Ano67b, Ano67j, Ano67i, Ano67l, Ano67k, Ano67m, Ano67n, Ano67p, Ano67o, Ano67q, Ano67r, Ano67t, Ano67s, Ano67u, Ano68r, Ano68q, Ano68s, Ano68u, Ano68t, Ano68w, Ano68v, Ano68y, Ano68x, Ano68-27, Ano68z, Ano68-29, Ano68-28, Ano68-30, Ano68-31, Ano68-33, Ano68-32, Ano68-35, Ano68-34, Ano68-36, Ano69r, Ano69p, Ano69t, Ano69q, Ano69o, Ano69s, Ano69v, Ano69u, Ano69x, Ano69n, Ano69w, Ano69-28, Ano69z, Ano69-27, Ano69y, Ano69-29, Ano69-30, Ano69-36, Ano69-34, Ano69-32, Ano69-31, Ano69-35, Ano69-33, Ano69-40, Ano69-38, Ano69-44, Ano69-42, Ano69-39, Ano69-37]. Meeting [Ano69-43, Ano69-41, Ano69-48, Ano69-45, Ano69-47, Ano69-46, Ano69-50, Ano69-49, Ano69-52, Ano69-51, Ano69-53, Ano69-54, Ano69-55, Ano71j, Ano71h, Ano71k, Ano71i, Ano71g, Ano71l, Ano71m, Ano71n, Ano71p, Ano71o, Ano71t, Ano71r, Ano71s, Ano71q, Ano71x, Ano71v, Ano71-27, Ano71z, Ano71w, Ano71u, Ano71-28, Ano71y, Ano71-32, Ano71-31, Ano71-36, Ano71-34, Ano71-29, Ano71-30, Ano71-35, Ano71-33, Ano71-37, Ano71-40, Ano71-38, Ano71-39, Ano71-43, Ano71-42, Ano71-41, Ano71-44, Ano71-45, Ano73f, Ano73e, Ano73h, Ano73g, Ano73j, Ano73i, Ano73l, Ano73k, Ano73n, Ano73m, Ano73p, Ano73o, Ano73r, Ano73q, Ano73t, Ano73s, Ano73v, Ano73u, Ano73w, Ano74a, Ano74e, Ano74d, Ano74f, Ano74h, Ano74g, Ano74j, Ano74i, Ano74l, Ano74k, Ano74n, Ano74m, Ano74p]. Meeting [Ano74o, Ano74s, Ano74q, Ano74r, Ano74u, Ano74t, Ano74w, Ano74v, Ano74x, Ano75e, Ano75d, Ano75f, Ano75h, Ano75g, Ano75j, Ano75i, Ano75k, Ano75l, Ano75n, Ano75m, Ano75p, Ano75o, Ano75r, Ano75q, Ano75t, Ano75s, Ano75v, Ano75u, Ano75w, Ano76h, Ano76g, Ano76i, Ano76j, Ano76k, Ano76m, Ano76l, Ano76o, Ano76n, Ano76q, Ano76p, Ano76s, Ano76r, Ano76u, Ano76t, Ano76w, Ano76v, Ano76y, Ano76x, Ano76z, Ano77e, Ano77d, Ano77f, Ano77h, Ano77g, Ano77j, Ano77i, Ano77l, Ano77k, Ano77n, Ano77m, Ano77p, Ano77o, Ano77r, Ano77q, Ano77t, Ano77s, Ano77v, Ano77u, Ano77w, Ano78e, Ano78d, Ano78f, Ano78h, Ano78g, Ano78i, Ano78j, Ano78l, Ano78k, Ano78m, Ano78n, Ano78o, Ano78q, Ano78p, Ano78s]. Meeting [Ano78r, Ano78u, Ano78t, Ano78v, Ano79d, Ano79c, Ano79e, Ano79g, Ano79f, Ano79i, Ano79h, Ano79k, Ano79j, Ano79m, Ano79l, Ano79o, Ano79n, Ano79q, Ano79p, Ano79s, Ano79r, Ano79u, Ano79t, Ano79v, Ano80f, Ano80e, Ano80g, Ano80i, Ano80h, Ano80k, Ano80j, Ano80m, Ano80l, Ano80o, Ano80n, Ano80q, Ano80p, Ano80s, Ano80r, Ano80t, Ano80u, Ano80v, Ano80d, Ano80w, Ano80x, Ano81g, Ano81h, Ano81j, Ano81i, Ano81l, Ano81k, Ano81n, Ano81m, Ano81p, Ano81o, Ano81r, Ano81q, Ano81t, Ano81s, Ano81v, Ano81u, Ano81x, Ano81w, Ano81y, Ano82d, Ano82c, Ano82e, Ano82f, Ano82g, Ano82i, Ano82h, Ano82k, Ano82j, Ano82m, Ano82l, Ano82o, Ano82n, Ano82r, Ano82p, Ano82q, Ano82t, Ano82s, Ano82v, Ano82u]. Meeting [Ano82w, Ano83f, Ano83e, Ano83g, Ano83i, Ano83h, Ano83k, Ano83j, Ano83m, Ano83l, Ano83p, Ano83o, Ano83n, Ano83r, Ano83q, Ano83t, Ano83s, Ano83u, Ano83v, Ano83x, Ano83w, Ano83y, Ano84d, Ano84c, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84k, Ano84j, Ano84m, Ano84l, Ano84o, Ano84n, Ano84q, Ano84p, Ano84s, Ano84r, Ano84t, Ano85-73, Ano85-74, Ano85-76, Ano85-75, Ano85-78, Ano85-77, Ano85-80, Ano85-79, Ano85-82, Ano85-81, Ano85-83, Ano85-84, Ano86g, Ano86f, Ano86e, Ano86j, Ano86h, Ano86i, Ano86l, Ano86k, Ano86m, Ano86n, Ano86p, Ano86o, Ano86r, Ano86q, Ano86v, Ano86t, Ano86u, Ano86s, Ano86z, Ano86y, Ano86w, Ano86x, Ano86-27, Ano86-28, Ano87f, Ano87g, Ano87i, Ano87j, Ano87h, Ano87l, Ano87k, Ano87n]. Meeting [Ano87m, Ano87p, Ano87o, Ano87r, Ano87q, Ano87t, Ano87s, Ano87v, Ano87u, Ano87x, Ano87w, Ano88d, Ano88c, Ano88f, Ano88e, Ano88h, Ano88g, Ano88j, Ano88i, Ano88l, Ano88k, Ano88n, Ano88m, Ano88p, Ano88o, Ano88r, Ano88q, Ano88t, Ano88s, Ano89i, Ano89j, Ano89l, Ano89k, Ano89s, Ano89u, Ano89t, Ano89v, Ano89w, Ano89y, Ano89x, Ano89z, Ano89-27, Ano89-29, Ano89-28, Ano89-30, Ano89-31, Ano89-33, Ano89-32, Ano89-35, Ano89-34, Ano90h, Ano90g, Ano90j, Ano90i, Ano90l, Ano90k, Ano90n, Ano90m, Ano90p, Ano90o, Ano90r, Ano90q, Ano90s, Ano91c, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano91k, Ano91j, Ano91m, Ano91l, Ano91n, Ano91o, Ano91q, Ano91p, Ano91s, Ano91r, Ano91u, Ano91t, Ano92e, Ano92g]. Meeting [Ano92f, Ano92i, Ano92h, Ano92j, Ano92k, Ano92m, Ano92l, Ano92o, Ano92n, Ano92q, Ano92p, Ano92r, Ano92t, Ano92s, Ano93g, Ano93f, Ano93h, Ano93i, Ano93k, Ano93j, Ano93m, Ano93l, Ano93o, Ano93n, Ano93q, Ano93p, Ano93s, Ano93r, Ano93u, Ano93t, Ano93v, Ano93w, Ano94i, Ano94k, Ano94j, Ano94m, Ano94l, Ano94o, Ano94n, Ano94q, Ano94p, Ano94s, Ano94r, Ano94u, Ano94t, Ano94w, Ano94v, Ano94y, Ano94x, Ano94-27, Ano94z, Ano95f, Ano95e, Ano95h, Ano95g, Ano95i, Ano95j, Ano95l, Ano95k, Ano95m, Ano95n, Ano95p, Ano95o, Ano95q, Ano95r, Ano95t, Ano95s, Ano95v, Ano95u, Ano96e, Ano96f, Ano96g, Ano96h, Ano96j, Ano96i, Ano96l, Ano96k, Ano96n, Ano96m, Ano96p, Ano96o, Ano96q, Ano96r, Ano96t]. Meeting [Ano96s, Ano96v, Ano96u, Ano96x, Ano96w, Ano97d, Ano97e, Ano97k, Ano97h, Ano97m, Ano97l, Ano97o, Ano97n, Ano97p, Ano97q, Ano97s, Ano97r, Ano97u, Ano97t, Ano97w, Ano97v, Ano97x, Ano97i, Ano97y, Ano97j, Ano97-27, Ano97z, Ano98c, Ano98d, Ano98e, Ano98f, Ano98h, Ano98g, Ano98j, Ano98i, Ano98l, Ano98k, Ano98n, Ano98m, Ano98p, Ano98o, Ano98r, Ano98q, Ano98t, Ano98s, Ano99g, Ano99f, Ano99h, Ano99i, Ano99k, Ano99j, Ano99m, Ano99l, Ano99o, Ano99n, Ano99q, Ano99p, Ano99s, Ano99r, Ano99u, Ano99t, Cop71m, FGG^{+}80, Les80a].

Megaptera [Cop71e]. Meigs [Bel73, Pep83]. Melanoplus [Scu97]. Meloidæ [Hor73c]. Member [Kön81, Wil88]. Memoir [Bri81a, Bri85c, Gar87c, Har80, Har95, Hor83, Ros95c, Rot86]. Memorial [Bar87, Phi79a, Phi79c]. Memory [Fra95b]. Mercer [Ste81c]. Meredith [Pri74]. Merrick [Goo69]. Mers [Fra76]. Mesozoic [Lym93, Fra75a, Fra77d]. Messrs. [Gen74b, Les78a]. Metallic [War91]. Metalline [Hun88]. Metalophodon [Cop71u]. Metals [Out74]. Metamorphism [Ste85b]. Meteoric [Cha71h, Cha71m, Kir69d, Kir77a, Kir79b]. Meteorites [Kir87b]. Meteorological [Ack69, Cha74a, Wil75]. Meteorology [Bla75, Bla76]. Meteors [Cha68a, Kir68, Kir69a, Kir73, Kir85a, Kir87a, Mar74]. Method [Bri86d, Cop71k, Rit93, Smi77c]. Methods [Bla87, Fra94b, Hau79, Hau88, Hou88, Pea61]. Methylene [Gre79b]. Metric [Fra78c]. Mexico [Bri85a, Bri92e, Bri92d, Bri93a, Cop69h, Cop75a, Cop81b, Cop82b, Cop83d, Cop85a, Cop85b, Cop86h, Dug88, GvR85, Hei92, LH76, Mee69, Ste82]. Mexique [Dug88]. Michael [Kea78]. Michaux [Hei92, MS89, Pri76b]. Michel [Rob80]. Michigan [Gra58a, HSL78]. Microconodon [Osb87]. Micrometer [Ano73c]. Microscope [Fra80]. Microscopic [Day83, Rot83, Sel73]. Microscopical [Fra78b]. Middle [Ano83d, Hau77, Jor88]. Middleburg [SS78]. Miles [Ano63k, Ano69g]. Military [Bac92]. Miller [Rob71]. Mineralogy [Gen82, Gen86, Gen87]. Minerals [Gen73, Gen74a, Gen77, Hun88]. Mines [Cha81a, Woo76]. Minor [Rit93]. Mint [Out74]. Minutes [Ano94i]. Miocene [Cop69c, Cop75b, Cop79b, Cop86g, Cop88a, Cop95a, Cop96b, Mey88]. Mirror [Fra80]. Mixophagus [Ano69h]. Mixture [Day98]. Models [Les80a]. Modern [Pri71a, Pri71b]. Modification [Hou74, Ryd89c]. Modifying [Key88]. Mohammed [Kra94, Kra94]. Mohammedan [Kra88]. Mohammedaner [Kra94]. Mohammedanischen [Kra88]. Mohammedans [Kra94]. Molar [Cop83e]. Mole [Ryd90a]. Molecular [Cha71e, Cha73a, Sad75]. Mollusca [Bla71, Gab61]. Monday [Phi79b]. Mono [Mar78]. Mono- [Mar78]. Monochlordinitrophenol [Smi78a]. Monograph [Wil82]. Monographic [Hor79a]. Monongahela [Rei71]. Monroe [Hal75a]. Mont [Ano63j]. Montgomery [Ste87a]. Monthly [Cha69d, Cha71r, Cha75e]. Monuments [Dol93]. Moon [Gen74b, Les74, Fra74b]. Moore [Ano61c]. Moraine [Ano83c]. Morning [Cha68a]. Morphology [Pac83a]. Moth [Pac94]. Moths [Gro95a, Pac93a, Pac93b]. Motion [Low69b, Ryd94]. Motions [Kir71a]. Mountain [Ano63j, BC75, Hal75a, Mar71, Les73a, Les73c, She78]. Mountains [P.73]. Mounted [Bri76b]. Mouth [Hau77]. Movable [Wil83]. Movements [Bac94b, Tha78]. Mr. [Ano76f, Cad69, Daw63, Dub71a, Fra94a, Kön77b, Pat75, Sel75, Ste96]. Multiplex [Hou83]. Muscles [Ryd90a]. Muscular [Hou90]. Musei [Mei86]. Museum [Hal76, Phi79a, Phi79c, Cop86i, Gar87b, Hal77, Ste96]. Mushroom [Ano68k]. Music [Cha73d]. Musical [Str76]. Muskokee [Bri69]. Musquito [Bri91]. Mutter [Pri94]. Mutual [Kir85b, Kir92]. Myriapoda [Mei86]. Myriopoda [Pac83a, Pac83b]. Mythic [Gat91]. Mythological [Les68a]. Mythology [Moo87]. Myths [Bri81b].

N.J. [Pot95]. N.Y. [SS78]. Nagualism [Bri94b]. Name [Phi83b]. Names [Bri81b, Bri90b, Hof86a, Les81b, Les81a]. Nanticoke [Bri93b]. Natchez [Bri73]. Nathaniel [Rot89]. National [Cop86i]. Nationality [Bri87d]. Nations [Bri85a]. Native [Bri92f, Bri93a, Cla83d, Bri94b]. Natural [Ano61d, Ort96, Phi97a, Sad76, Sad77, Sad97, Ash84]. Naturalist [Cop94a]. Nature [Pec97, Pri78]. near [Ack69, Cop75a, Cop86h, Fra74a, Hau77, Hay69b, Les71d, Lym94b, Mor97, Ste81b]. Nearest [Uhl88]. Nebraska [St.69]. Nebular [Cha76a, Cha78c, Cha79b]. Nebulous [Kir85b]. Need [Bai94]. Neill [Bri80]. Neotropical [Cop71h]. Neptune [Kir71a]. Nerve [Oli94, Sto82, Sto86c, Sto87b, Sto88a, Sto88b, Sto88c]. Nevada [Cop71w]. Newark [Lym94a, Lym94c]. Nicaragua [Bri86b, Bri87a]. Nicotin [Hay77]. Nile [Wil75]. No [Ano59e, Ano77c, Cop77i, Gra58a, Kön81, Hay68b]. No. [Bar75, Cha73a, Gen73, Gen74a, Gen77, Gen82, GvR85, Gen86, Gen87, Gra58b, Hay69a, Koe74, Kön77a, Kön78, Mar78, Sad75, Sad76, Sad77, Smi77a, Smi77b, Smi77c, Smi78a, Smi78b, Ste75a]. Noanama [Bri96]. Noctuidæ [Gro83]. Nodal [Cha80b]. Nodes [Low69a]. Nomenclature [Hun88]. Non [Ano87d, Blo73]. Non-Magnetizable [Ano87d]. Non-Periodic [Blo73]. Normal [Cha71o, Kön77b, Low69a]. Norristown [LC91]. North [Ano75c, Ano89k, Cop71j, LH76, Les71d, Woo69b, Car77, Cha81b, Cop69l, Cop69g, Cop71q, Cop75a, Cop85d, Cop86i, Cop87b, Cop88a, Cop89b, Cop96b, Dav83a, Gro74, Gro83, Les77b, Mat99, Mon78, Pea61, Wil82, Woo69a]. North-West [Ano89k]. North-Western [Cop75a]. Northampton [Hal75a, Pri77b]. Northern [Cha71n, Cop71s, Cla83c, Hei91]. Norwegian [Nan97]. Note [Ano73c, Bau92a, Bri90a, Car75, Cha71a, Cha73e, Cla83e, Cla83b, Cla83c, Cla83d, Cop69d, Cop71o, Cop81a, Daw63, Fra74b, Fra76, Ful74, Hay98b, Hay99, Kir87b, Koe85, Kön77b, Les71c, Les71b, Les71d, Les83b, Les84, Lew82a, Oli94, Phi83b, Ran81, Sco98, Ste82, Cla83g]. Notebooks [Cou93]. Notes [Ash84, Bar77, Bau93, Bla71, Boa87, Bra89, Bri86b, Bri92c, Bri92b, Cha71d, Cha81c, Cha81d, Cha81e, Cha81f, Cha82a, Cha82b, Cha83a, Cha83b, Cha84, Cla83a, Coo79, Cop86g, Far96, Fra86b, Gil84, Gro74, Hal75a, Hau77, Hay68a, Hay68b, Hay69a, Hof89b, Hor80a, Hor91, Jon80, Les68a, Les78a, Les80a, Les81b, Les91a, Mor91, Pac93a, Pac93b, Pec68, Phi79a, Phi79c, Phi84a, Sco94, Sha95, Spe81, Ste75a, Ste80a, Ste80b, Ste81b, Ste81c, Ste85a, Ste85b, Ste87b, Whi80, Whi81, Whi82, Wil98, Win69c, vD98, vD99, Ano75c, Les81a]. Notice [All86, Ame82, Ang90, Ano60c, Ano61c, Ano62d, Ano62e, Ano63m, Ano63n, Ano65d, Ano69k, Ano69l, Bel73, Bid89, Bri79, Bri80, Bri88b, Bri94a, Bri95a, Cad69, Car71, Cop71l, Cop71z, Cur85, Dol94, Dub71a, Fie71, Fra77a, Fra79b, Gin78, Goo69, Gra84, Ham82, Har71a, Kön78, Law85, LeC73, Lei87, Les68b, Les80b, Les82, Les83c, Les86a, Les88, Les90a, Les90b, Les91b, Ley60, Mer90, Muh83, Oli93, Pac83b, Pat82, Phi81b, Pri74, Pri94, Rob71, Rob75, Rob82, Rob80, Rob83, Rog80, Ros95a, Rus84, Rus93, SM79, Sel75, Sel86, Sha68, Sti69, Vau85, Vau90, Wil88]. Notices [Cop71p, Phi89b, Rus86, Sto86a, Sto87a, Sto90, Sto94, Win69c]. nouvelles [Dug88]. Nov. [Ano60t, Ano69-48, Ano69-45, Ano69-47, Ano69-46, Ano71-37, Ano78r, Ano80t, Ano82s, Ano83u]. Nova [Phi84b]. Novaculite [Ash97]. November [Ano89g, Ano97e, Kir68, Ano50w, Ano50v, Ano50t, Ano50u, Ano52z, Ano52x, Ano52y, Ano52w, Ano53s, Ano53r, Ano54r, Ano54s, Ano55t, Ano55s, Ano56s, Ano56r, Ano57w, Ano57v, Ano58t, Ano58s, Ano59v, Ano59u, Ano60u, Ano61u, Ano61t, Ano62u, Ano62t, Ano63-48, Ano63-46, Ano63-47, Ano63-45, Ano65u, Ano65t, Ano66s, Ano66r, Ano67q, Ano67r, Ano68-33, Ano68-32, Ano71-40, Ano71-38, Ano71-39, Ano73t, Ano73s, Ano74u, Ano74t, Ano75t, Ano75s, Ano76w, Ano76v, Ano77t, Ano77s, Ano78s, Ano79s, Ano79r, Ano80u, Ano81v, Ano81u, Ano82t, Ano83v, Ano84q, Ano84p, Ano86v, Ano86t, Ano86u, Ano86s, Ano87t, Ano87s, Ano88p, Ano88o, Ano89-30, Ano89-31, Ano90s, Ano91q, Ano91p, Ano92q, Ano92p, Ano93s, Ano93r, Ano94w, Ano94v]. November [Ano95q, Ano95r, Ano96t, Ano96s, Ano97e, Ano97w, Ano97v, Ano98d, Ano98r, Ano98q, Ano99s, Ano99r, Cha68a, Kir68, Kir78, Kir85a, Kir87a, Les80a]. Nucleus [Hub94]. Nukic [Kra88, Kra88]. Numeration [Tay87]. Numerical [Hag83]. Nuttall [Ano60c]. Nyctinomus [All89a]. Nymphalidæ [Gro98].

O [Sco94, Smi88]. O. [Hau77]. Obituary [All86, Ame82, Ang90, Ano61c, Ano62d, Ano62e, Ano63m, Ano63n, Ano65d, Ano69k, Ano69l, Bel73, Bid89, Bri79, Bri80, Bri88b, Bri94a, Bri95a, Cad69, Car71, Cur85, Dol94, Dub71a, Fie71, Fra77a, Fra79b, Goo69, Gra84, Ham82, Har71a, Law85, LeC73, Les68b, Les80b, Les82, Les86a, Les88, Les90a, Les90b, Les91b, Ley60, Mer90, Muh83, Oli93, Pat82, Pep83, Phi81b, Phi89b, Pri74, Pri94, Rob71, Rob75, Rob82, Rob80, Rob83, Ros95a, Rus84, Sel75, Sel86, Sha68, Sti69, Vau90, Les83c]. Objectionable [Bac91a]. Objects [Fra80]. Observation [Fra71]. Observations [Ack69, Bri92d, Bro84, Cha68a, Cop69e, Cop71q, Dwi93, Fra78b, Hag82, Hei91, Law88, Rot91, Wil75]. Observatory [Sha83]. Observed [Cop75a, Dav83b, Kir68]. Obtained [Cop78b, Cop87c, Cop94a]. Occidentalis [Cop77k]. Occupation [Ano89g]. Occurrence [Cop71g, Phi97b, Sad75]. Occurring [Fra78a]. Ocean [Hau89]. Oceanic [Cha71s]. Oct [Fra86c]. Oct. [Ano69-49, Ano71-43, Ano71-42, Ano71-41, Ano82v, Ano82u, Cop71m]. October [Ano50-27, Ano50y, Ano50z, Ano50x, Ano52-30, Ano52-29, Ano52-27, Ano52-28, Ano53u, Ano53t, Ano54u, Ano54t, Ano55v, Ano55u, Ano56u, Ano56t, Ano57y, Ano57x, Ano58u, Ano58v, Ano59x, Ano59w, Ano60w, Ano60v, Ano61w, Ano61v, Ano62w, Ano62v, Ano63-50, Ano63-49, Ano63-52, Ano63-51, Ano65w, Ano65v, Ano66u, Ano66t, Ano67t, Ano67s, Ano68-35, Ano68-34, Ano69-50, Ano69-52, Ano69-51, Ano73v, Ano73u, Ano74w, Ano74v, Ano75v, Ano75u, Ano76y, Ano76x, Ano77v, Ano77u, Ano78u, Ano78t, Ano79u, Ano79t, Ano80v, Ano80d, Ano80w, Ano81x, Ano81w, Ano83x, Ano83w, Ano84s, Ano84r, Ano86z, Ano86y, Ano86w, Ano86x, Ano87v, Ano87u, Ano88r, Ano88q, Ano89-33, Ano89-32, Ano91s, Ano91r, Ano92r, Ano93u, Ano93t, Ano94y, Ano94x, Ano95t]. October [Ano95s, Ano96v, Ano96u, Ano97x, Ano97i, Ano97y, Ano97j, Ano98t, Ano98s, Ano99u, Ano99t]. Octonary [Tay87]. Ocular [Ryd90a]. Ohio [Cop77a, Cop85e, Cop97, Mor97]. Oil [Ano76f, Ash78, Ash80, Car77, Cha78a, Hau77, Les71e, Net77, Ran81, Wyc86]. Oils [Car85b]. Old [Phi80a, Sac98]. olden [Hof94]. Omaha [Hay68a]. Omission [Ano81f]. One [Ano90f, Ano94f, Ano94i, Blo74]. Ontario [Spe81]. Opaque [Fra80]. Opening [Bri76a, Bri77a]. Ophidia [Cop94b]. Ophidians [Dug88]. Ophidiens [Dug88]. Opinions [Bla76]. Optic [Oli94]. Optic-Nerve [Oli94]. Optics [Har76]. Orbits [Kir92]. Order [Ano89r, Cop81c, Cop95c]. Ordinary [Dav66, Woo76]. Ordinations [Cha75b]. Ore [Ano63j, Ano63k, Gen74b, Les71d, Les73b, Les74]. Oregon [Cop79b]. Oreodontidæ [Cop84b]. Ores [Fra77d, Gen74b, Les73a]. Organic [Bai96, Cla87, Con96, Cop71k, Har71b, Ryd93a]. Organisms [Ryd89c]. Origin [Ano63d, Ano63e, Cha74f, Cha79b, Cop88b, Day97, Fra75b, Kir80, New82, Pec68, Pec97, Ryd90b, Spe81]. Original [Cou93]. Ornaments [Bra89]. Ornithosaurians [Cop71x]. Orthografy [Ano89q]. Orthography [Taf61, Taf62]. Orton [Cop77o, Cop78b]. Oscillation [Cha71c, Cha71q]. Oscillatory [Cha73f]. Osino [Cop71w]. Osmazome [Bac93]. Osteology [Cop73a, Cop92h, Far96, Sco94]. Osteoporosis [MKAP71]. Other [Cla84, Les76, Win69c]. Our [Spe81]. Outcome [Cla87]. Outcrop [Les73d]. Outlet [Spe81]. Outline [Cop89a]. Oxybenzoic [Smi88]. Oxygen [Dra77].

P [Ano63h, Ano63j, Ano63k, Ano63l, Ano68i, Ano68h, Ano77c]. P. [Ano63d, Ano63e, Les91b, Ros95c]. PA [Gen74b]. Pa. [Ano68g, Hau77, Kön78, Les78a, Rei71, Wal84]. Pacific [Coo79]. Pah [Hof74]. Pah-Ute [Hof74]. Paint [Ano68d]. Paintings [Mat97b]. Palæocrinoids [Key88]. Palenville [She78]. Paleolithic [Bri88a]. Paleontological [Hal76]. Paleontology [Cop86c, Cop92a, Hal77]. Paleozoic [Ash77a, Ash97, Cop92c, Cop97, PS78, Pri77b, Whi80]. Palm [Dol93]. Palms [Bra84]. panolius [Ano69h]. Paper [Daw63, Fra77e, G.86, Kön77b, Fra94a]. Papers [Cha80d]. Papiamento [Gat85a]. Papilionidæ [Gro99b, Gro99c]. Papyrus [Les68a]. Parallelism [Ste74]. Paraná [Der79b]. Parasitic [Loc83]. Paris [Ros95b]. Park [Phi79a, Phi79c]. Parker [Les68b]. Parnassi [Gro99b, Gro99c]. Parnassi-Papilionidæ [Gro99b, Gro99c]. Part [Cop71s, Cop80a, Mei86]. Partial [Cop69l]. Parts [Bra86, Ste81a]. Parvus [Dwi93]. Paso [Kön77a]. Pass [Hau99]. Passamaquoddy [Alg85, Pri97, Pri99]. Patapsco [Hau77]. Path [HK97]. Patterson [Pat75]. Pawnee [Hay68a]. Peale [Ano69k]. Peat [Bri77b, Lew82b]. Pebas [Cop86b]. Peculiarities [Cha74a]. Peculiarity [Oli94]. Pelasgic [Ano63d]. Penn [Hay99]. Penn’a [Ano71e]. Pennsylvania [Ano76f, Ano83d, Ano89q, Ash77a, Ash78, Ash81, Bar75, Bar77, Car77, Cha78a, Cla83b, Cla83c, Cla83d, Cla83g, Cla84, Cop77j, Cop97, Day83, Day97, Fra75a, Fra78a, Fra82, Fra84, Fra86b, Ful74, Gen73, Gen74a, Gen77, Gen79, Gen82, GvR85, Gen86, Gen87, Hal75a, Hal76, Hal77, Hal80, Hay99, Hof89a, Hof89b, Hof94, Koe74, Kön77a, Kön81, Les71e, Les73a, Les73c, Les74, Les76, Les77a, Les78a, Les83b, Les91c, Les77b, Les77d, Lew85, Lil83, Lil86, Mar78, Mer97b, Net77, Phi79a, Phi79c, Pri82, Pri77b, Rot94a, Rot94b, Sad75, Sad76, Sad77, Smi77a, Smi77b, Smi77c, Smi78a, Smi78b, Ste79, Ste96, Vau84, Wil98]. Pensilfâni [Hof94]. Pensylvania [Ano69g]. Pentagonal [Mor97]. Penumbra [Wil78]. Perch [Cop69g]. perdicida [Ano69h]. Perforating [Cop80a]. Pericles [Smy98]. Period [Cha73g, Cop95a]. Periodic [Blo73]. Periodicity [Kir69c]. Periods [Cop69c, Kir69d]. Periptychus [Cop83c]. Perissodactyla [Cop81c]. Perissodactyles [Cop73d]. Permian [Cop77b, Cop78a, Cop80c, Cop82e, Cop83b, Cop85c, Cop86f]. Perry [Ano83d, Cla83e, Cla83d, Cla83g]. Perturbations [Rit93]. Peru [Bri90a, Cop77o]. Peruvian [Cop78b]. Peter [Phi81b]. Petite [Cop95b]. Petrie [Ste96]. Petrocene [SM79]. Petroleum [Day97, Mab97, Pec97, Phi97a, Phi97b, SM79, Sad97]. Peytona [Lym94b]. Ph.D [Bri81a, DuB88, Les88, Wil88]. Pharyngeal [Sto88b]. Phase [Pri71a]. Phases [Pri71b]. Phenacodus [Cop83c]. Phenomena [Hau88, Leo96]. Phenomenon [P.73]. Phil [Ano69-55]. Philadelphia [Ano62y, Ano94f, Cha68c, Cha69d, Cha69e, Cou93, Hal75b, Out74, Phi79a, Phi79c, Phi81a, Phi88, Phi92, Str76, Wil94]. Philip [Bri88b]. Phillips [Ano95d, Vau85]. Philosophic [Bri85c]. Philosophical [Ano59e, Ano62y, Ano74a, Ano85-72, Ano89g, Ano89i, Ano89j, Ano89l, Ano89k, Ano89m, Ano89q, Ano89r, Ano94f, Ano94i, Ano99c, Ano99d, Cha80d, Cou93, FGG^{+}80, Hay98b, Phi89b, Str76, Wil88]. Philosophy [Cha79d, Cop89a, Gro79, Pri71a, Pri71b]. Phonetic [Bri86d]. Phonetics [Mor91]. Phonograph [Fra78b]. Phosphorus [Hou74]. Photo [Cha85c]. Photodynamic [Cha81c, Cha81d, Cha81e, Cha81f, Cha82a, Cha82b, Cha83a, Cha83b, Cha84]. Photodynamics [Cha80e]. Photogrammetry [Bac92]. Photographic [Sha95]. Photographs [May69, Bar87]. Photography [Dra77, Fra86a, Hou86, Sac96, Tay85]. Phrases [Alg85]. Phyllocaridan [Pac86]. Phylogeny [Cop87a, Cop92e, Hya94, Ryd89a]. Physical [Bla71, Fra78a, Hau88, Leo96, Les76, Bar75]. Physics [Ano94e, Har71b]. Physiological [Har76, Ryd89b]. Physostomi [Cop71h]. Picture [Wal84]. Pierced [Bri76a, Bri77a]. Pieri [Gro98]. Pieri-Nymphalidæ [Gro98]. Pipe [Bar77, Hau77, Kön77a]. Pittsburg [Ano71e]. Pl [Ano68f, Ano68g, Ano68h, Ano69f, Ano69d, Ano69e]. Pl. [Ano68e, Ano69g]. Place [Whi80]. Plagopterinæ [Cop74b]. Plains [Cop74a]. Planetary [Cha71q, Cha71u, Cha71t, Cha73g, Cha75f]. Planeto [Cha73e].

Planeto-Taxis [Cha73e]. Planets [Cha73j, Kir80, Kir85b, Rit93, Tro76]. Plant [Woo69a]. Plants [Les77c]. Plate [Ano63f, Ano63g, Ano63h, Ano63i, Ano63j, Ano63k, Ano63l, Ano68j, Ano68i, Ano68k, Ano68l, Ano68m, Ano68n, Ano68o, Ano68p, Ano69h, Ano69i, Ano69j, Ano85-72, Bri76a, Bri77a, Cla83c, Ano63d, Ano63e]. Plateau [Cop75a]. Platyceras [Key88]. Platypterices [Ano75c, Gro74]. Plesiosaurian [Cop94c]. Pleurodira [Cop71r]. Pliny [Gar87c]. Plistocene [Cop95b]. Plotting [Ano73c]. Poetic [Mor96]. Point [Pep96, Fra78a]. Points [Cop92d, Har76]. Polar [Nan97]. Polarity [Cha71b]. polyphemus [Pac85]. Polysynthesis [Bri86c]. Pond [Ano63j]. Pool [Cop74a]. Population [Cop75a]. Port [Ack69, Cop71i]. Port-au-Prince [Ack69]. Portable [Dav66]. Portage [Ano83d]. Porter [Ano68e]. Porterfield [Muh83]. Portions [MS89, Ste79, Ste87a]. Position [Bri85b, Cha71o, Gin78, Pac87]. Positions [Gra58a]. Possession [Cou93]. Possible [Bac91b, Hou88, Kir87b, Les84]. Posterior [Cop80a, Cop81a]. Postulate [Cha71l]. Potatoes [Woo71b]. Potomac [Les73a, Jon80]. Pottery [Jor88]. Potts [Rot82]. Practicability [Hau99, Out74]. Prediction [Cha73h, Cha79c]. Preface [Phi85]. Preglacial [Spe81]. Preliminary [Cop71i, Mee69, SM79, SO87, Sco98, Kön78, Wil81]. Prematurely [Woo71a]. Preparation [Gre79b, Hou88, Les77b]. Presence [Gre79a]. Present [Ano89g, Hau88]. President [Kea78]. Presidents [Ano89n]. Preston [Ste81d]. Presumably [Sto94]. Price [Day83, Rot86]. Primary [Loc83]. Prime [Fra77e]. Primitive [Kir71b]. Primordial [Cha71c]. Prince [Ack69]. Princeton [SO87]. Priority [DuB95]. Prison [Vau84]. Prize [Ano95d]. Proboscians [Cop71l]. Proboscidian [Cop77e]. Procamelus [Cop77k]. Proceedings [Ano89o, Ano89p, Phi89b, Ano77c, Ano59e]. Process [Sac96]. Processes [Gin78]. Procured [Cop77o]. Prodromus [Woo69b]. Product [SM79]. Production [Day98]. Productions [Del74, Del75]. Productive [Cha78a, Les77d]. Prof [Ano63n]. Prof. [Ano62e, Ano74a, Cop77o, Les81c]. Professor [Cop78b, Da 85, Fra77e, Rog80]. Profitably [Woo76]. Progress [Hal76, Les83b, Les77b]. Progressive [Les71c]. Projecting [Fra80]. Promotion [Ano94f]. Pronghorn [All89b]. Pronunciation [Taf61, Taf62]. Proofs [Ryd89c]. Proportional [Jam95]. Propositions [Cha76b]. Prospect [Ano63g]. Protection [Ran81]. Protohistoric [Bri95c]. Protophytes [Loc83]. Protostega [Cop71n]. Province [Cop87c, Der79b]. Psoas [Dwi93]. Pteropus [All90c]. Public [Hau79, Phi83a]. Published [Phi80b, Phi81a, Phi89b]. Puddling [Bri77b]. Puerco [Cop82d, Cop83a, Cop83f]. Pulaski [Cre77, Ste87a]. Puquina [Bri90a]. Purpose [Car77]. Pyramidalis [Dwi93]. Pyrite [Bro94]. Pyrophyllite [Gen79]. Pythonomorpha [Cop69h, Cop71a].

Quadrature [Cha79a]. Quantitative [Kön78]. Quarried [Mer95]. Quartzose [Fra79a]. Quasi [Dub71b]. Quaternion [Mil98]. Queensland [Mat98b]. Question [Oli94]. Quinnimont [Ste81c]. Quito [Bro84].

R [Agn85, Dub71a, G.86, Vau87]. Radiant [Edd82]. Radiation [Cha78e]. Rail [Car77]. Railroad [Les76, Les73b, Les74]. Railway [Kea78, Les73a]. Rain [Cha71i, Cha75e]. Rain-Fails [Cha71i]. Rain-Fall [Cha75e]. Rainfall [Cha68c, Cha69d, Cha69a, Cha71f, Cha71k, Cha71n, Cha71r, Cha74d, Cha74e, Cha75g]. Rainfalls [Cha71g]. Ralph [Ame82]. Ramirez [Phi84a]. Rana [Cop86i]. Randolph [Rot89]. Range [Gab61]. Ratio [Cha85a, Pri95]. Raton [Ste82]. Ray [Ano96f, Leo96]. Raynes [Mer90]. Re [Ste80b]. Re-Eroded [Ste80b]. Reach [Dav66]. Reaction [Hau99]. Read [Wil78, Pri74]. Real [HK97]. Recently [Les77c, Les77d]. Recession [Low69a]. Reciprocity [Dud86]. Reconnaissance [Dra97, Ste81a, Ste87a]. Record [Fra78b, Les71e]. Recorded [Kra94]. Records [Ano76f, Ash78, Cha79e, Hou88, Les83a, Pri97, She69]. Rectifying [Car77]. Red [Lym94d, Ryd94]. Reed [Sha68]. Reference [Bri86d, Sac98, Scu94]. Referred [Pat75]. Reforesting [Pri78]. Refraction [Man77]. Regard [Day83]. Regarding [Fra77d]. Region [Ash97, Cop71h, Der79b, Bri98a, Cop77h]. Regional [Dut71]. Register [Ack69]. Regular [Les71c]. Rejang [Hil96]. Related [Les78b, Pec98]. Relating [Bla71]. Relation [Cha71k, Cha71p, Kir87c, Mor97, Ryd90b]. Relations [Cha69c, Cha73g, Cha75c, Cha80f, Cop82c, Hal80, Kir71a, Kir85b, Kir92, Sad97, Ste75b]. Relative [Ano77c, Cha73i, Kir77b]. Relatives [Uhl88]. Relic [Cla83d]. Remains [Cop75a, Cop77g, Cus96, Mer97a, Ste96]. Remarkable [Bla85, Kir71a]. Remarks [All89b, All90a, Ano74a, Ano96f, Bla75, Bri87c, Cop69o, Fra77e, Fra94a, Gat76, Hof86a, Sad75, Sto85, Ano71f]. Remedies [Hau88]. Renault [Les81c]. Renovo [Cha78a]. Rensselaeria [Cla83e]. Reply [Gen74c, Kön77b]. Report [Ano89q, Cop71i, Cop86h, Hei92, Hil96, SO87]. Reported [Dub71b]. Representation [Jam95]. Reptile [Cop86f]. Reptiles [Cop71b, Cop77h, Cop77i, Cop86b, Gar87a, Gar87b]. Reptilia [Ano69m, Cop71j, Cop78a, Cop87c, Cop93b, Cop94c]. Reptilian [Cop77g, Cop95c]. Researchers [Rol94]. Researches [Cop77a, Cop85e]. Resemblance [Cha71s]. Resembling [Day98, Lew82b]. Resolution [Ano89q]. Resources [Del74, Del75]. Respecting [Phi83b, Ste80b, Ste85b]. Respectively [Phi80b]. Resulting [Kir85b]. Results [Bla87, Car77, Cha78f, Cre74, Hau99, May69, Nan97, Sad78, Out74]. Résumé [Fra86c, Hau99]. Rev [Muh83]. Rev. [Agn85, Ano63n]. Reversion [Hag83]. Review [Cop89b, Hor80b]. Revision [Hor73b, Hor73c, Hor78a, Hor79a, Jay82, Lil86, Pac83b]. Revival [Woo71a]. Revolution [Ano99c]. Rhinocheilus [Dug86]. Rhoads [Har71a, Har95]. Rhynchophora [LH76, LeC77]. Rica [Bri97a, Cop93b, Gab75]. Richard [Ano99d]. Ridge [Ste81d]. Ring [Kir83]. Rings [Kir69d]. Rio [Cop94a]. Riot [Tha78]. Rising [Hag82]. River [Ano68l, Ano68m, Ano68g, Ash77a, Cop69e, Cop71b, Hay69b, Les77d, Rei71, Cop71s, Far96, Hal80, Hau77, Hil96]. Road [Car77]. Roasting [Bac93]. Robert [Rus84, Fra79b, Rus86]. Roberts [Rob82]. Rochambeau [Ros94]. Rocinela [Ric98]. Rock [Ano68k, Ano68l, Ano68m, Ano97f, Lym95a, Mat97a, Mat97b, Rei71]. Rockery [Pri82]. Rocks [Ano68g, Ash77a, Cha78a, Fra74a, Fra79a, Fra82, Fra84, Hal80, Les78a, Les77c, Lil83, Lil86, PS78, Pri77b, She78, Wal84]. Rocky [BC75, Mar71, P.73]. Rogers [Rus86, Cla84, Rus93]. Role [Loc83]. Röntgen [Ano96f]. Room [Ano89i, Ano89j, Ano89l, Ano89k]. Roots [Cha61]. Rotation [Cha71t, Cha73j, Cha78e, Ker73]. Rothermel [Ros95c]. Roy [Lil83]. Rules [Ano89r]. Run [Hau77, Gen74b, Hau77, Les74]. Runic [Phi84b]. Ruschenberger [Bri95a]. Russell [Les71a, Ste85a].

S [Ano77c, Bri81a, Car85a, Cur85, Law85, Out74, Sel73]. S. [Bri81a]. Salicylic [Mar78, SKHE87]. Saligenin [Gre80a]. Salishan [Boa95]. Same [Phi84a, Rot83]. Samuel [Ano61c, Bri95a, Da 85, Goo69]. San [Cha71f]. Sand [Ash80]. Sandstone [Cla83f, Fra75a, Les83d]. Sanscrit [Cha60]. Santa [vD99]. Saturn [Kir71a, Kir83]. Saurians [Cop77j]. Saurocephalus [Cop69d]. Saurodontidæ [Cop69n]. Saving [Cha74g]. Saving-Fund [Cha74g]. Scales [Hau78]. Schissler [Sac95]. Schists [Hal80]. Schoharie [SS78]. School [Hau79, Phi79a, Phi79c]. Schuylkill [Gen79]. Science [Bai94, Rin85, Wil94]. Sciurus [Ano69h]. Scotia [Phi84b]. Scott [Ste81a, Ste80a]. Scranton [Lew82b]. Script [Cha71a]. Sculpture [Fra77c]. Sea [Wyc86, Bar77]. Seaboard [Hau77]. Search [Low69a]. Second [Cop69m, Cop71y, Cop71z, Cop79b, Cop80c, Cop83f, Cop85e, Cop93b, Cop96a, Edd82, God96, Hal77, Les76, Les78a, Phi92, Hal76, Les83b, Les77b]. Secondary [Loc83]. Secret [Bac94b]. Section [Ano63k, Ash77a, Ash81, Cha78a, Les62, Lil83, Lil86, PS78, SS78, She78, Ano63j, Ano69g]. Sections [Hay69b, Lym94b]. Seen [Kir77a, Kir79b]. Segment [God96]. Selected [Ano76f]. Selection [Ort96]. Selenodont [Sco98]. Selenophorus [Hor80a]. Selish [Hof86b]. Semi [Bri85a, Cha79b]. Semi-Civilized [Bri85a]. Semi-Diameter [Cha79b]. Separation [Ort96]. Sept [Fra86c]. Sept. [Ano52-31, Ano69-53, Ano71-44, Ano71-45, Ano84t]. September [Ano50-28, Ano52-32, Ano53v, Ano54v, Ano55w, Ano56v, Ano57z, Ano58w, Ano59y, Ano60x, Ano61x, Ano62x, Ano63-54, Ano63-53, Ano65x, Ano66v, Ano67u, Ano68-36, Ano69-54, Ano73w, Ano74x, Ano75w, Ano76z, Ano77w, Ano78v, Ano79v, Ano80x, Ano81y, Ano82w, Ano83y, Ano86-27, Ano86-28, Ano87x, Ano87w, Ano88t, Ano88s, Ano89-35, Ano89-34, Ano91u, Ano91t, Ano92t, Ano92s, Ano93v, Ano93w, Ano94-27, Ano94z, Ano95v, Ano95u, Ano96x, Ano96w, Ano97-27, Ano97z]. Series [Hag83, Kön81, Les77a, Les78a, Whi80, Ash77a, Les77d]. Sermon [Cra83]. Sermon-Books [Cra83]. Service [Bla87, Bla89]. Set [Bri85c, Les84]. Seventh [Cop69p]. Several [Hor73c]. Sex [Ryd90b]. Sexuality [Ryd90b]. Seybert [Rob83]. Shaddai [Les86b]. Shadows [Wil78]. Shakespeare [Smy98]. Shales [Cla84, Cop69e, Les77d]. Shamanism [Mor97]. Sharon [Whi80]. ShDI [Les86b]. Sheafer [Les91b]. Shell [Key88, Ano69j]. Shells [Coo79]. Ships [Tha83]. Shooting [Kir87c]. Shorb [Gen74b]. Short [Ano65d, Cop73e, Oli94]. Shower [Kir68]. Showers [Cha71m]. Showing [Dav83a]. Shrew [Ryd90a]. Shrew-Mole [Ryd90a]. Siege [Hay98a]. Sierra [GvR85]. Signal [Bla89, Bla87]. Signature [Fra95a]. Significance [Rin85, Sla95]. Silurian [Ash77a, Fra75b, Hal80, Les77c]. Siluro [Les77a]. Silver [Ant76, Pri95]. Silvyculture [Rot77]. Simmons [Ang90]. Simply [Mil97]. Sir [Fie71]. Six [Mil97]. Sixth [Cop96b]. Skeleton [Ryd89b]. Sketch [Agn85, Da 85, Pot95, Rei71, Rot82, Rot89, Rus86, Rus91, Rus92, Fra86b, Hof94]. Skull [Cop84a, Cop92f, Cop94c]. Slates [Cha78a]. Slaven [Kra88]. Slaves [Kra88]. Slavic [Kra94]. slavischen [Kra94]. Slope [Coo79]. Sloth [Mer97a]. Smith [All86, Cop87c]. Smoky [Ano68l, Ano68m]. Smyth [Ste85a]. Snakes [Cop86a]. So-Called [Bri87b, Oli94]. Soc [Ano69-55]. Social [Bai94, Del74, Del75]. Society [Ano74a, Ano89g, Ano89q, Ano94f, Cou93, FGG^{+}80, Les83a, Ano62y, Ano85-72, Ano89i, Ano89j, Ano89l, Ano89k, Ano89m, Ano89r, Ano94i, Ano99c, Ano99d, Cha80d, Hay98b, Str76, Wil88, Ano59e, Phi89b]. Solar [Bar78, Cha71t, Cha73b, Cha73f, Cha76a, Cha79e, Dra77, Kir69c, Kir71b, May69, Rei63]. Solomon [Mor94]. Solution [Hag83]. Solutions [Dav83b, SKHE87]. Some [Ano69m, Bla75, Bla76, Bri86a, Bri95d, Cha68b, Cha76b, Coo79, Cop69g, Cop69h, Cop71c, Cop71o, Cop77h, Cop77i, Cop77j, Cop81b, Cop86i, Cop92c, Cop92d, Cop95b, Dav83a, Fra77d, Fra78b, Fra78c, Gen77, Har76, Hou88, Kir71a, Les68a, Low69b, Lym94b, Lym94d, Mar78, May69, Mon78, Nan97, Phi80c, Pla79, Pri71b, Pri99, Rot83, Rot91, Sad75, Smi77a, Ste85b, Sto86b, Vau89, Wal84, Wil94, vD99]. Somerset [Ano71e, Ful74]. Song [Bri92a, Kra94, Kra88]. Sound [Cha73c]. Sources [Bri92c]. Sous [Ant76]. South [Ano89j, Ano89l, Ash97, Bri92f, Bri98b, Fra82, Les73a, Ste79, Ste87b]. South-East [Ano89j]. South-West [Ano89l, Ste79, Ste87b]. Southeastern [Fra84, Lew85]. Southern [Cop71l]. Spain [Del75]. Span [Fie97]. Special [Ano51b, Ano52c, Ano57d, Ano57e, Ano57f, Ano58c, Ano67b, Ano69n, Ano80d, Ano90g, Ano97h, Ano97i, Ano97j, Bri86d, FGG^{+}80, Scu94]. Specializations [Gro98, Gro99b, Gro99c]. Species [All89a, All90a, All90c, All90b, All91, Cop69h, Cop77f, Cop84b, Cop86b, Cop86d, Cop86i, Cop86g, Cop89b, Cop94c, Hor73b, Hor73c, Hor78a, Hor79a, Hor80a, Hor80b, LeC78, Les77d, Mee69, Ric98, Rot83, Ryd90b, Scu97, Dug88, Cop86i]. Specimens [Cop86i, MKAP71]. Spectra [Bar87, Dor96, Out74]. Spectroscopic [Fra71]. Spectrum [Dra77, Out74]. spelaeus [Ano69h]. Spheres [Cha73d]. Spines [Pac93a, Wil83]. Spoken [Bri86b]. Spot [Cha71v, Cha73g]. Spots [Kir69c, Rei63]. Spouting [Ash77b]. Spruce [Gen74b, Les74]. Squamosal [Cop80a]. St. [Ash81, FGG^{+}80, Les73b]. Stability [Kir85b]. Stalagmite [Mer95]. Standpoint [Bai96, Con96]. Stapes [Mon78]. Stars [Kir77b, Kir87c, Ano77c]. State [Gra58a, Smi95, Cop71o, Cop86h, Phi89a]. Stated [Ano50d, Ano50c, Ano50e, Ano50i, Ano50g, Ano50h, Ano50f, Ano50j, Ano50k, Ano50m, Ano50l, Ano50o, Ano50n, Ano50p, Ano50q, Ano50s, Ano50r, Ano50w, Ano50v, Ano50t, Ano50u, Ano50-27, Ano50y, Ano50z, Ano50x, Ano50-28, Ano51d, Ano51c, Ano51f, Ano51e, Ano51h, Ano51g, Ano51i, Ano51k, Ano51j, Ano51m, Ano51l, Ano52f, Ano52e, Ano52d, Ano52g, Ano52k, Ano52i, Ano52j, Ano52h, Ano52m, Ano52l, Ano52q, Ano52o, Ano52n, Ano52r, Ano52p, Ano52t, Ano52s, Ano52v, Ano52u, Ano52z, Ano52x, Ano52y, Ano52w, Ano52-30, Ano52-29, Ano52-27, Ano52-28, Ano52-32, Ano52-31, Ano53c, Ano53d, Ano53e, Ano53g, Ano53f, Ano53i, Ano53h, Ano53k, Ano53j, Ano53m, Ano53l, Ano53o, Ano53n, Ano53q, Ano53p, Ano53s, Ano53r, Ano53u]. Stated [Ano53t, Ano53v, Ano54d, Ano54c, Ano54e, Ano54f, Ano54g, Ano54i, Ano54h, Ano54k, Ano54j, Ano54m, Ano54l, Ano54o, Ano54n, Ano54q, Ano54p, Ano54r, Ano54s, Ano54u, Ano54t, Ano54v, Ano55e, Ano55d, Ano55f, Ano55h, Ano55g, Ano55j, Ano55i, Ano55l, Ano55k, Ano55n, Ano55m, Ano55p, Ano55o, Ano55r, Ano55q, Ano55t, Ano55s, Ano55v, Ano55u, Ano55w, Ano56d, Ano56c, Ano56e, Ano56g, Ano56f, Ano56h, Ano56i, Ano56k, Ano56j, Ano56m, Ano56l, Ano56o, Ano56n, Ano56q, Ano56p, Ano56s, Ano56r, Ano56u, Ano56t, Ano56v, Ano57h, Ano57g, Ano57i, Ano57k, Ano57j, Ano57m, Ano57l, Ano57o, Ano57n, Ano57q, Ano57p, Ano57s, Ano57r, Ano57t, Ano57u, Ano57w, Ano57v, Ano57y, Ano57x, Ano57z, Ano58e, Ano58d]. Stated [Ano58f, Ano58h, Ano58g, Ano58j, Ano58i, Ano58k, Ano58l, Ano58n, Ano58m, Ano58p, Ano58o, Ano58r, Ano58q, Ano58t, Ano58s, Ano58u, Ano58v, Ano58w, Ano59f, Ano59g, Ano59h, Ano59j, Ano59i, Ano59l, Ano59k, Ano59n, Ano59m, Ano59p, Ano59o, Ano59r, Ano59q, Ano59t, Ano59s, Ano59v, Ano59u, Ano59x, Ano59w, Ano59y, Ano60f, Ano60e, Ano60g, Ano60i, Ano60h, Ano60k, Ano60j, Ano60m, Ano60l, Ano60o, Ano60n, Ano60q, Ano60p, Ano60s, Ano60r, Ano60u, Ano60t, Ano60w, Ano60v, Ano60x, Ano61f, Ano61e, Ano61g, Ano61i, Ano61h, Ano61j, Ano61k, Ano61m, Ano61l, Ano61o, Ano61n, Ano61p, Ano61q, Ano61s, Ano61r, Ano61u, Ano61t, Ano61w, Ano61v, Ano61x, Ano62g, Ano62f, Ano62h, Ano62j, Ano62i, Ano62l]. Stated [Ano62k, Ano62n, Ano62m, Ano62o, Ano62q, Ano62p, Ano62s, Ano62r, Ano62u, Ano62t, Ano62w, Ano62v, Ano62x, Ano63r, Ano63q, Ano63t, Ano63o, Ano63p, Ano63s, Ano63x, Ano63v, Ano63w, Ano63u, Ano63-28, Ano63z, Ano63-27, Ano63y, Ano63-32, Ano63-31, Ano63-36, Ano63-34, Ano63-29, Ano63-30, Ano63-35, Ano63-33, Ano63-40, Ano63-38, Ano63-41, Ano63-42, Ano63-39, Ano63-37, Ano63-44, Ano63-43, Ano63-48, Ano63-46, Ano63-47, Ano63-45, Ano63-50, Ano63-49, Ano63-52, Ano63-51, Ano63-54, Ano63-53, Ano65f, Ano65e, Ano65g, Ano65h, Ano65i, Ano65k, Ano65j, Ano65m, Ano65l, Ano65o, Ano65n, Ano65q, Ano65p, Ano65s, Ano65r, Ano65u, Ano65t, Ano65w, Ano65v, Ano65x, Ano66d, Ano66c, Ano66e, Ano66g, Ano66f, Ano66i, Ano66h, Ano66k, Ano66j, Ano66m, Ano66l]. Stated [Ano66o, Ano66n, Ano66q, Ano66p, Ano66s, Ano66r, Ano66u, Ano66t, Ano66v, Ano67d, Ano67c, Ano67e, Ano67f, Ano67h, Ano67g, Ano67j, Ano67i, Ano67l, Ano67k, Ano67m, Ano67n, Ano67p, Ano67o, Ano67q, Ano67r, Ano67t, Ano67s, Ano67u, Ano68r, Ano68q, Ano68s, Ano68u, Ano68t, Ano68w, Ano68v, Ano68y, Ano68x, Ano68-27, Ano68z, Ano68-29, Ano68-28, Ano68-30, Ano68-31, Ano68-33, Ano68-32, Ano68-35, Ano68-34, Ano68-36, Ano69r, Ano69p, Ano69t, Ano69q, Ano69o, Ano69s, Ano69v, Ano69u, Ano69x, Ano69w, Ano69-28, Ano69z, Ano69-27, Ano69y, Ano69-29, Ano69-30, Ano69-36, Ano69-34, Ano69-32, Ano69-31, Ano69-35, Ano69-33, Ano69-40, Ano69-38, Ano69-44, Ano69-42, Ano69-39, Ano69-37, Ano69-43, Ano69-41, Ano69-48, Ano69-45, Ano69-47, Ano69-46, Ano69-50, Ano69-49]. Stated [Ano69-52, Ano69-51, Ano69-53, Ano69-54, Ano71j, Ano71h, Ano71k, Ano71i, Ano71g, Ano71l, Ano71m, Ano71n, Ano71p, Ano71o, Ano71t, Ano71r, Ano71s, Ano71q, Ano71x, Ano71v, Ano71-27, Ano71z, Ano71w, Ano71u, Ano71-28, Ano71y, Ano71-32, Ano71-31, Ano71-36, Ano71-34, Ano71-29, Ano71-30, Ano71-35, Ano71-33, Ano71-37, Ano71-40, Ano71-38, Ano71-39, Ano71-43, Ano71-42, Ano71-41, Ano71-44, Ano71-45, Ano73f, Ano73e, Ano73h, Ano73g, Ano73j, Ano73i, Ano73l, Ano73k, Ano73n, Ano73m, Ano73p, Ano73o, Ano73r, Ano73q, Ano73t, Ano73s, Ano73v, Ano73u, Ano73w, Ano74e, Ano74d, Ano74f, Ano74h, Ano74g, Ano74j, Ano74i, Ano74l, Ano74k, Ano74n, Ano74m, Ano74p, Ano74o, Ano74s, Ano74q, Ano74u, Ano74t, Ano74w, Ano74v, Ano74x, Ano75e, Ano75d]. Stated [Ano75f, Ano75h, Ano75g, Ano75j, Ano75i, Ano75k, Ano75l, Ano75n, Ano75m, Ano75p, Ano75o, Ano75r, Ano75q, Ano75t, Ano75s, Ano75v, Ano75u, Ano75w, Ano76h, Ano76g, Ano76i, Ano76j, Ano76k, Ano76m, Ano76l, Ano76o, Ano76n, Ano76q, Ano76p, Ano76s, Ano76r, Ano76u, Ano76t, Ano76w, Ano76v, Ano76y, Ano76x, Ano76z, Ano77e, Ano77d, Ano77f, Ano77h, Ano77g, Ano77j, Ano77i, Ano77l, Ano77k, Ano77n, Ano77m, Ano77p, Ano77o, Ano77r, Ano77q, Ano77t, Ano77s, Ano77v, Ano77u, Ano77w, Ano78e, Ano78d, Ano78f, Ano78h, Ano78g, Ano78i, Ano78j, Ano78l, Ano78k, Ano78m, Ano78n, Ano78o, Ano78q, Ano78p, Ano78s, Ano78r, Ano78u, Ano78t, Ano78v, Ano79d, Ano79c, Ano79e, Ano79g, Ano79f, Ano79i, Ano79h]. Stated [Ano79k, Ano79j, Ano79m, Ano79l, Ano79o, Ano79n, Ano79q, Ano79p, Ano79s, Ano79r, Ano79u, Ano79t, Ano79v, Ano80f, Ano80e, Ano80g, Ano80i, Ano80h, Ano80k, Ano80j, Ano80m, Ano80l, Ano80o, Ano80n, Ano80q, Ano80p, Ano80s, Ano80r, Ano80t, Ano80u, Ano80v, Ano80w, Ano80x, Ano81g, Ano81h, Ano81j, Ano81i, Ano81l, Ano81k, Ano81n, Ano81m, Ano81p, Ano81o, Ano81r, Ano81q, Ano81t, Ano81s, Ano81v, Ano81u, Ano81x, Ano81w, Ano81y, Ano82d, Ano82c, Ano82e, Ano82f, Ano82g, Ano82i, Ano82h, Ano82k, Ano82j, Ano82m, Ano82l, Ano82o, Ano82n, Ano82r, Ano82p, Ano82t, Ano82s, Ano82v, Ano82u, Ano82w, Ano83f, Ano83e, Ano83g, Ano83i, Ano83h, Ano83k, Ano83j, Ano83m, Ano83l, Ano83p, Ano83o, Ano83n]. Stated [Ano83r, Ano83q, Ano83t, Ano83s, Ano83u, Ano83v, Ano83x, Ano83w, Ano83y, Ano84d, Ano84c, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84k, Ano84j, Ano84m, Ano84l, Ano84o, Ano84n, Ano84q, Ano84p, Ano84s, Ano84r, Ano84t, Ano85-73, Ano85-74, Ano85-76, Ano85-75, Ano85-78, Ano85-77, Ano85-80, Ano85-79, Ano85-82, Ano85-81, Ano85-83, Ano85-84, Ano86g, Ano86f, Ano86e, Ano86j, Ano86h, Ano86i, Ano86l, Ano86k, Ano86m, Ano86n, Ano86p, Ano86o, Ano86r, Ano86q, Ano86v, Ano86t, Ano86u, Ano86s, Ano86z, Ano86y, Ano86w, Ano86x, Ano86-27, Ano86-28, Ano87f, Ano87g, Ano87i, Ano87j, Ano87h, Ano87l, Ano87k, Ano87n, Ano87m, Ano87p, Ano87o, Ano87r, Ano87q, Ano87t, Ano87s, Ano87v, Ano87u, Ano87x, Ano87w, Ano88d, Ano88c]. Stated [Ano88f, Ano88e, Ano88h, Ano88g, Ano88j, Ano88i, Ano88l, Ano88k, Ano88n, Ano88m, Ano88p, Ano88o, Ano88r, Ano88q, Ano88t, Ano88s, Ano89s, Ano89u, Ano89t, Ano89v, Ano89w, Ano89y, Ano89x, Ano89z, Ano89-27, Ano89-29, Ano89-28, Ano89-30, Ano89-31, Ano89-33, Ano89-32, Ano89-35, Ano89-34, Ano90h, Ano90j, Ano90i, Ano90l, Ano90k, Ano90n, Ano90m, Ano90p, Ano90o, Ano90r, Ano90q, Ano90s, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano91k, Ano91j, Ano91m, Ano91l, Ano91n, Ano91o, Ano91q, Ano91p, Ano91s, Ano91r, Ano91u, Ano91t, Ano92e, Ano92g, Ano92f, Ano92i, Ano92h, Ano92j, Ano92k, Ano92m, Ano92l, Ano92o, Ano92n, Ano92q, Ano92p, Ano92r, Ano92t, Ano92s, Ano93g, Ano93f, Ano93h, Ano93i, Ano93k]. Stated [Ano93j, Ano93m, Ano93l, Ano93o, Ano93n, Ano93q, Ano93p, Ano93s, Ano93r, Ano93u, Ano93t, Ano93v, Ano93w, Ano94k, Ano94j, Ano94m, Ano94l, Ano94o, Ano94n, Ano94q, Ano94p, Ano94s, Ano94r, Ano94u, Ano94t, Ano94w, Ano94v, Ano94y, Ano94x, Ano94-27, Ano94z, Ano95f, Ano95e, Ano95h, Ano95g, Ano95i, Ano95j, Ano95l, Ano95k, Ano95m, Ano95n, Ano95p, Ano95o, Ano95q, Ano95r, Ano95t, Ano95s, Ano95v, Ano95u, Ano96f, Ano96g, Ano96h, Ano96j, Ano96i, Ano96l, Ano96k, Ano96n, Ano96m, Ano96p, Ano96o, Ano96q, Ano96r, Ano96t, Ano96s, Ano96v, Ano96u, Ano96x, Ano96w, Ano97d, Ano97e, Ano97k, Ano97m, Ano97l, Ano97o, Ano97n, Ano97p, Ano97q, Ano97s, Ano97r, Ano97u, Ano97t, Ano97w, Ano97v, Ano97x]. Stated [Ano97y, Ano97-27, Ano97z, Ano98e, Ano98f, Ano98h, Ano98g, Ano98j, Ano98i, Ano98l, Ano98k, Ano98n, Ano98m, Ano98p, Ano98o, Ano98r, Ano98q, Ano98t, Ano98s, Ano99g, Ano99f, Ano99h, Ano99i, Ano99k, Ano99j, Ano99m, Ano99l, Ano99o, Ano99n, Ano99q, Ano99p, Ano99s, Ano99r, Ano99u, Ano99t, Cop71m, Ano74r, Ano82q]. States [Win69c, Ano69m, Cha71r, Cha71x, Cha75e, Cha75g, Cop69c, Cop69m, Cop77b, Cop95a, Dav83a, Hor73a, Hor73c, Hor73d, Hor78a, Hor78b, Hor79a, Hor79b, Hor80a, Hor80b, Jay82, Jor88, Kir77a, Kir79b, Les76, Les77c, Les78b, MS89, Mor62, Phi83a]. Station [Cha78a]. Stationary [Woo76]. Steel [Cre75]. Stellar [Bar87, Cha71u]. Stephen [Car71]. Stephenson [Cha71l]. Stereodeetes [Ano69h]. Sterilization [Bac91b]. Sterry [Gen74c]. Stock [Bri95b, Bri95d, Bri96]. Stone [Pea61, Pea63, Phi80c]. Stories [Cra83]. Storms [Wyc86]. Strata [Cop71a, Hay69b]. Stratigraphical [Gab61]. Stream [Kir87b]. Strength [Cre75, Day83]. Strickland [Gra84]. Striding [Dav66]. Stroke [Ran81]. Stromateidæ [Gil84]. Struck [Ant76, Phi78]. Structural [Oli94]. Structure [Cop81a, Cop84a, Cop85d, Cop86e, Cop86f, Cop94c, Kir71b, Les71a, Les73c, Sad75, Ste85a]. Structures [Cre75]. Strychnia [Hay77]. Students [McC81]. Studies [Bri92f]. Study [Bri90b, Bri93a, Bri94b, Bro94, Dwi93, Gro83, Les73c, Mar78, Pac94, Phi92, Woo69b]. Sub [Hor78a]. Sub-Family [Hor78a]. Subjective [Oli86]. Subsidences [Dut71]. Substance [Lew82b, Ryd94]. Substitution [Mil97]. Successfully [Woo76]. Suggestion [Day97]. Suggestions [Low69b]. Sul [Cop94a]. Sullivan [Pla79]. Sulphosalts [Kön81]. Sun [Ano77c, Bla85, Cha71v, Cha73g, Cha71j, Cha73j, Cha79b, Dra77, Kir77b, Tro76, Vau89]. Sun-Glows [Bla85]. Sun-Spot [Cha71v, Cha73g]. Supplement [Cop71j]. Supplementary [Sto85]. Supposed [Cha73k, Hou74, Phi84b]. Supreme [Ano62y]. Surface [Abb98, Ano63k, Ste79, Ste87b]. Surrounding [HK97]. Survey [Ano76f, Bar77, Cla84, Hal76, Hal77, Les76, Les78a, Les83b, Les77b]. Surveys [Car77]. Survival [Mer97b]. Surviving [Cla83d]. Susquehanna [Fra79a]. Susquehannah [Ano68g]. Svastika [Bri89b]. Sweat [Ryd89a]. Sweat-Glands [Ryd89a]. Syllabus [Rot77]. Sylviculture [Pri77a]. Symbolism [Bri93a]. Symbols [Vau89]. Symmetric [Mil97]. Synchronous [Hou83]. Synchronous-Multiplex [Hou83]. Synonymic [Cop86i]. Synonymy [Gab61]. Synopsis [Cop69l, Cop71j, Cop75b, Cop77o, Cop78b, Cop82d, Cop84b, Cop87c, Gab61, Hor73d, Hor78b, Hor79b, LeC77, Ric98]. Synthesis [Gre80a]. Syrphidæ [Wil82]. System [Car77, Cha73f, Cha76a, Hau79, Kir71b, Tay87, Vau84, Ash77a]. Systematic [Cop81c, Cop82c, Net77, Pac87].

T [Ano68e, Gen74c]. Table [Cop86a]. Tables [Cha69e, Dav83a, Fra78c, Man77]. Tabular [LeC77]. Taken [Alg85, Wil75]. Tale [Gat91]. Tales [Pri99]. talpoides [Ryd90a]. Tamias [Ano69h]. Tanks [Ran81]. Tapiras [Ano69h]. Tapiroid [Cop73a]. Tartronic [Sad75]. Taxis [Cha73e]. Taxonomy [Bau92a, Bau92b, Bau93]. Tazewell [Les71a, Ste81c, Ste85a]. Teaching [Bac91a]. Teiidæ [Cop92g]. Telegraph [Hou83]. Telephone [Hou88]. Tellurium [Gen74a, Gen77]. Temperate [Cha71n, Hei92]. Temporary [Cop74a]. Tennessee [Cop69g, Les71b, Mer97a]. Tenth [Cop77p]. Teredines [Ano75c, Gro74]. Terminal [Ano83c]. Terrestrial [Bla71, Cha79b, Doo97]. Territories [Hay68b, Hay69a]. Territory [Cop69k, Cop69e, Cop71p, Cop71-27, Cop85c, Dra97, Hay69b]. Tertiary [Cop69j, Cop69e, Cop71s, Cop71w, Cop77c, Scu94, Cop77f]. Testing [Sha95]. Tests [Bri77b]. Testudinata [Bau93, Cop71n]. Testudinate [Cop71d]. Tex. [Hau99]. Texas [Cop78a, Cop80c, Cop82e, Cop83b, Cop85c, Cop92a]. Texts [Boa95]. Their [Bra89, Dor96, Kir85b, Les78b, Pac93a, Rin85, Hau88]. Them [Pea61]. Theory [Dra77, Fra77b, Ryd89b, Smi95]. There [Dud86]. Thereof [Hag82]. Thermo [Cha85c]. Thermo-Electro-Photo-Baric [Cha85c]. Thermodynamics [Cha74c, Edd82]. Thermometrical [Bro84]. Theromorphous [Cop86f]. Third [Cop71-27, Cop82e, Gat90]. Thirteenth [Cop86j]. Thomas [Ano60c, Pot95, Cur85, Pri94, Rot82, Wil88]. Thoracic [Pac86]. Those [Scu94]. Thought [Bri97b, Hau79]. Thoughts [Vau89]. Three [Cop69k, Cop86g, Fra94b]. Three-Toed [Cop86g]. Throne [Sac98]. throughout [Car77]. Thyrsos [Dol93]. Tiaporus [Cop92g]. Ticholeptus [Cop86g]. Tidal [Cha68c, Cha69a, Cha71o]. Tide [Bar77]. Timber [Day83, Rot83]. times [Hof94]. Timucua [Gat77, Gat78, Gat80]. Tipulidæ [Scu94]. Titaniferous [Les71d]. Titanium [Kön77a]. Tlingit [Boa91]. Toed [Cop86g]. Together [Pec68]. Toltecs [Bri87d]. Toluol [Smi77a]. Tomiopsis [Cop93a]. Tongue [Alg85]. Tónkawa [Gat76]. Tooth [Cop83e]. Top [Ano63g, Ash77a]. Topography [Ker73]. tortus [Ano69h]. Total [Bar78, May69]. Toxodon [Cop81a]. Trade [Dud86]. Transactions [Phi89b]. Transcript [Cha73k]. Transformations [Pac94]. Transit [Ano71c, Dav66, Hag82]. Transition [Cop71v]. Transitive [Mil97]. Translated [Les86b]. Translation [Bri85c, Phi84a]. Transmission [Oli94, Sel99]. Trap [Fra78a, Fra84, Lew85]. Traps [Fra75a]. Travels [MS89]. Trees [Gro99a, Pri76b, Woo71a]. Trego [Rob75]. Trenton [Ash77a]. Trias [Cop77d, Cop87b, LC91, Les91c]. Triassic [Cop77b, Cop77j, Osb87]. Tribal [Hof86a]. Tribe [Hal83, Hor73b]. Tribes [Gab75, Mat98a, Mat98c, Mat99]. Trigeminus [Sto86c]. Trituberculate [Cop83e]. Troops [Tha78]. Tropical [Cop69b, Cop69p, Cop77p, Cop79a, Cop85f, Cop86j]. True [Hag82]. tsai’tä [Hof94]. Tsimshian [Boa91]. Tubercles [Pac93a]. Tucker [Ano62e]. Tule [Hof86c]. Tungsten [PS94, SD94]. Tunisia [Pep96]. Turtles [Mon78]. Tutelo [Hal83]. Twelfth [Cop85f]. Two [All89a, Bai91, Bri98b, Cop71h, Cop71x, Cop86g, Cop94c, Fra95a, Phi80b, Dug88]. Type [Cop83e]. Tyre [Smy98]. Tyrone [Les74].

U [Car85a, Out74]. U.S. [Cre77]. U.S.A. [Ano99c, Cop71o]. U.S.N. [Fra95b]. Uinta [SO87, Sco98]. Uncertainties [Law88]. Unclassified [Bri98b]. Undescribed [Sto94]. Ungulata [Cop73e]. Ungulate [Cop82a]. Ungulates [Cop71t]. Unit [Cha85c]. United [Ano69m, Cha71r, Cha71x, Cha75e, Cha75g, Cop69c, Cop69m, Cop77b, Cop95a, Dav83a, Hor73a, Hor73c, Hor73d, Hor78a, Hor78b, Hor79a, Hor79b, Hor80a, Hor80b, Jay82, Kir77a, Kir79b, Les76, Les77c, Les78b, MS89, Mor62, Phi83a]. Units [Fra78c]. Universal [MKAP71]. University [Bac91a, Bar75, Day83, Gen73, Gen74a, Gen77, Gen82, GvR85, Gen86, Gen87, Koe74, Kön77a, Mar78, Sad75, Sad76, Sad77, Smi77a, Smi77b, Smi77c, Smi78a, Smi78b, Ste96, Kön78, Kön81, Pri82]. University-Extension [Bac91a]. Unpublished [Bri85c, Bri92c]. upon [Cre75, Eme71, Gat76, Oli94, Pec68, Phi79a, Phi79c, Phi84a, Smi88, Smi77a]. Upper [Cop71p, Cop77c, Cop86b, Ano63j, Ano68d, Cla83c, Cop71b, Cop86g, Cop94c, SS78, St.69, Ste81d]. Upthrow [Les71b]. Uranus [Kir71a]. Ursina [Ano71e]. Use [Bri77b, McC81, Ryd89c, Wyc86]. Used [Out74]. Useful [Ano94f]. Ustick [Wil88]. Utah [Cop74b, Cop77d]. Ute [Hof74].

V [Ano63i, Bai91, Ano62y, Ano69j, Cha80c, Cha82a, Sad75]. Va. [Ano63f]. Vagus [Sto82]. Valley [Fra82, GvR85, Ano63k, Cop85b]. Valleys [Bra86]. Value [Bri77b]. Valve [Koe74]. Vanadates [GvR85]. Vanadium [Gen77]. Vandevere [Les88]. Variation [Cla87, Doo97, Dwi93]. Variations [Cha69d]. Various [Hau79]. Varying [Pri95]. Vascular [Bra84]. Vaughan [Goo69].

Vaux [Law85]. Vein [Ano63f]. Velocities [Cha73i]. Velocity [Cha80b]. Venango [Net77]. Verb [Bri85c, vHB85]. Verbal [Ano69-55]. verdeutscht [Kra94]. Verrius [Kin88]. Vertebrata [Cop71c, Cop71p, Cop71o, Cop71q, Cop71i, Cop71y, Cop71-27, Cop75b, Cop77b, Cop77c, Cop77o, Cop77m, Cop77n, Cop80c, Cop82b, Cop82d, Cop82e, Cop87b, Cop92e, Cop97]. Vertebrate [Cop71s, Cop86c, Cop92a, SO87]. Vertebrates [Cop71z, Cop92c, Cop92d, Ryd94]. Vertical [Bar75]. Vertical-Lantern [Bar75]. Vezier [Kra94, Kra94]. VI [Ano59e, Ano68k, Cha80c, Cha82b, Sad76]. Vicinity [Cha78a, Phi88, Phi92, Ste82]. View [Ano63k, Pep96]. VII [Ano59e, Ano68l, Cha83a, Kön77a]. VIII [Cha78a, Ano63j, Ano68m, Cha83b, Sad77]. Violation [Les71c]. Virginia [Cre77, Ash84, Les71a, Lym94b, Mey88, Ste75a, Ste79, Ste80a, Ste81a, Ste81c, Ste81d, Ste85a, Ste87a, Ste87b, Whi81, Whi82]. Virtue [Ano89q]. Visible [Hal75a]. Visit [Hay99]. Vocabularies [Boa91, Bri91, Bri98b]. Vocabulary [Boa93, Bri93b, Bri96, Hof86b, Hof86c, Hof89b]. Vol [Ano59e, Ano77c]. Vol. [Ano81f]. Volcanoes [Hei92]. Volume [Ano53w, Ano54w, Ano58x, Ano59z, Ano60y, Ano61y, Ano62z, Ano63-55, Ano65y, Ano68-37, Ano74y, Ano75x, Ano76-27, Ano77x, Ano77y, Ano78w, Ano78x, Ano80y, Ano80z, Ano81z, Ano81-27, Ano83z, Ano83-27, Ano84u, Ano85-85, Ano85-86, Ano85-87, Ano86-29, Ano86-30, Ano87y, Ano87z, Ano88u, Ano88v, Ano89-36, Ano89-37, Ano90t, Ano90u, Ano91v, Ano91w, Ano92u, Ano92v, Ano93x, Ano93y, Ano94-28, Ano94-29, Ano95w, Ano95x, Ano96y, Ano96z, Ano97-28, Ano97-29, Ano98u, Ano98v, Ano99v, Ano99w]. Volumetric [Bri76b]. Vortex [Ryd94]. Vortex-Motion [Ryd94].

W [Ano62d, Ano63f, Ham82, Les80b, Les91b, Sco94]. Wagner [Wil94]. Waithman [Bri95a]. Waitshum’ni [Hof86c]. Waldo [Ame82]. Walter [Wil88]. Wampum [Pri97]. Warren [Hau77]. Warrior [Gen74b, Les74]. Was [Ano91c, Cou93]. Washington [Ste81a, Ste85a]. Waste [Woo76]. Watch [Ano87d]. Water [Ash77b, Bac91b, Bri76a, Bri77a, Cop69j, Cop69k, Sto86a, Sto87a, Sto90, Woo69b]. Water-Well [Ash77b]. Waterloo [Phi78]. Waters [Ano68d, Cop69l, Cop71p]. Wave [Cha78f]. Wavellite [ST82]. Waves [Cha75d]. Way [Ste80b]. Weather [Cha71d, Cha85b, Eme71]. Weedon [Ano99c]. Weights [Tay87]. Well [Ano76f, Ash77b, Ash78, Car77, Cha78a, LC91]. Wells [Les71e, Les91d, Net77, Sad76, Sad77]. Were [Bri87d]. West [Ano89l, Ano89k, Car77, Cop77c, Lym94b, Ste81c, Ste81d, Ano89l, Ano89k, Bla71, Gar87b, Hal75b, Ste75a, Ste79, Ste87b, Whi81, Whi82]. Western [Hau77, Bri95c, Cop75a, Net77, Sad76, Sad77, Win69c]. Wet [Eme71]. Wheat [Woo71b]. Wheatley [Cop77j]. Which [Cou93, Woo76]. White [Cla83f, Les83d, Far96]. Whitney [Sel86]. Who [Pot95]. Wie [Kra94]. Wilcocks [Wil78]. Wilcox [Ash77b]. Wilhelm [Bri85c]. Wilkins [Ano65d]. William [Ano63m, Bri95a, Fie71, Hay98a, Hay99, Law85, Pat82, Rob82, Rus93]. Williamson [Fra78a]. Winds [Cha71w, Cha71x]. Wing [Gro98, Gro99b, Gro99c]. Winged [Dol93]. Winnebago [Hay68a]. Wisconsin [Gra58a]. Wise [Ste81a, Les71a, Ste80a, Ste85a]. Wister [Bid89]. Witchcraft [Pri99]. without [Wil78]. Wm. [Les68b]. Wood [Day98, Har80]. Wood-Ashes [Woo71b]. Word [Les86b]. Words [Alg85, Bri95d, Les59]. Work [Cha73k, Fra86c, Hal76, Phi80a, Wil94]. Works [Ano94e]. Writing [Bri86d, Mer97b]. Writings [Bri87c]. Written [MS89]. Wyandotte [Mer95]. Wyoming [Cop69e, Cop71p, Cop71-27, Hay69b, Bra86, Cop71b, Cop71v, Cop71r, Cop71l, Cop71z, Cop73c, Cop73e, Cop82b, Hay68b, Hay69a]. Wythe [Ste87a].

X [Ano63k, Ano68n, Leo96, Smi77b]. X-Ray [Leo96]. XI [Ano63l, Ano68o, Gen77]. XI.1 [Mor96]. XII [Ano68p, Smi77c]. XIII [Cha78a, Ano68g, Mar78]. Xinca [Bri85b]. XIV [Ano68h, Ant76, Smi78a]. XIX [Ano81f]. XV [Kön78]. XVI [Ano77c, Smi78b]. XX [Gen82]. XXIII [GvR85]. XXIV [Gen86]. XXIX [Gen87].

Yardley [Lym95b]. Yarmouth [Phi84b]. Year [Hal76, Kir79b]. Yearly [Cha75g]. Years [Cha71v, Phi80b]. Yolk [Hub94]. York [Ano83d, Fra75b, Fra75a, Fra86b, Sad77, She78]. Yorke [Fra95b]. Yucatan [Hei91].

Zacualtipan [Cop86h]. Zentmayer [Oli93]. Zircon [Kön77a]. Zone [Cha71n, Kir83]. Zoology [Cop74a, Gar87b]. Zymotic [Loc83].

References

Abbe:1898:AAA

[Abb98] Cleveland Abbe. The altitude of the aurora above the Earth’s surface. Proceedings of the American Philosophical Society, 37(157):4–7, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

         (electronic). URL https://www.jstor.org/stable/983685.

Ackerman:1869:RMO

[Ack69] A. Ackerman. A register of meteorological observations made at Bois Chêne, near Port-au-Prince, Hayti. Proceedings of the American Philosophical Society, 11(81):499–512, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

         https://www.jstor.org/stable/981502.

Agnew:1885:BSR

[Agn85] D. Hayes Agnew. Biographical sketch of the Rev. Elias R. Beadle, D.D., LL.D. Proceedings of the American Philosophical Society, 22(120):227–240, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982981.

Alger:1885:CWP

[Alg85] Abby Langdon Alger. A collection of words and phrases taken from the Passamaquoddy tongue. Proceedings of the American Philosophical Society, 22(120):240–255, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982982.

Allen:1884:CHC

[All84] Harrison Allen. On a case of human congenital malformation. Proceedings of the American Philosophical Society, 21(115):413–418, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982730.

Allen:1886:OND

[All86] Harrison Allen. Obituary notice of Dr. Albert H. Smith. Proceedings of the American Philosophical Society, 23(124):606–611, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983340.

Allen:1889:GND

[All89a] Harrison Allen. On the genus Nyctinomus and description of two new species. Proceedings of the American Philosophical Society, 26 (130):558–563, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983197.

Allen:1889:RPA

[All89b] Harrison Allen. Remarks on the pronghorn (Antilocapra Americana). Proceedings of the American Philosophical Society, 26(129): 366–367, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982959.

Allen:1890:DNSa

[All90a] Harrison Allen. Description of a new species of Carollia and remarks on Carollia brevicauda. Proceedings of the American Philosophical Society, 28(132):19–26, January/June 1890. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983303.

Allen:1890:DNSc

[All90b] Harrison Allen. Description of a new species of Macrotus. Proceedings of the American Philosophical Society, 28(132):72–74, January/ June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983314.

Allen:1890:DNSb

[All90c] Harrison Allen. Description of a new species of Pteropus. Proceedings of the American Philosophical Society, 28(132):70–72, January/ June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983313.

Allen:1891:NSA

[All91] Harrison Allen. On a new species of Atalapha. Proceedings of the American Philosophical Society, 29(135):5–7, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983011.

Ames:1882:ONR

[Ame82] Charles G. Ames. Obituary notice of Ralph Waldo Emerson. Proceedings of the American Philosophical Society, 20(112):498–503, June/ December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982699.

Angell:1890:ONH

[Ang90] James B. Angell. Obituary notice of Henry Simmons Frieze, LL.D. Proceedings of the American Philosophical Society, 28(132):59–61, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
983310.

Anonymous:1850:Aa

[Ano50a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(??):??, ???? 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1850:Ab

[Ano50b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(45):??, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1850:SMAb

[Ano50c] Anonymous. Stated meeting, April 19. Proceedings of the American Philosophical Society, 5(45):143–144, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143373.

Anonymous:1850:SMAa

[Ano50d] Anonymous. Stated meeting, April 5. Proceedings of the American Philosophical Society, 5(45):141–143, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143372.

Anonymous:1850:SMAc

[Ano50e] Anonymous. Stated meeting, August 16. Proceedings of the American Philosophical Society, 5(45):156–157, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143378.

Anonymous:1850:SMDd

[Ano50f] Anonymous. Stated meeting, December 20. Proceedings of the American Philosophical Society, 5(45):171–174, April/December 1850. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143385.

Anonymous:1850:SMDb

[Ano50g] Anonymous. Stated meeting, December 21. Proceedings of the American Philosophical Society, 5(??):113–119, ???? 1850. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143489.

Anonymous:1850:SMDc

[Ano50h] Anonymous. Stated meeting, December 6. Proceedings of the American Philosophical Society, 5(45):169–171, April/December 1850. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143384.

Anonymous:1850:SMDa

[Ano50i] Anonymous. Stated meeting, December 7. Proceedings of the American Philosophical Society, 5(??):111–113, ???? 1850. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143488.

Anonymous:1850:SMFa

[Ano50j] Anonymous. Stated meeting, February 1. Proceedings of the American Philosophical Society, 5(??):123–134, ???? 1850. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143492.

Anonymous:1850:SMFb

[Ano50k] Anonymous. Stated meeting, February 15. Proceedings of the American Philosophical Society, 5(??):134–135, ???? 1850. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143493.

Anonymous:1850:SMJb

[Ano50l] Anonymous. Stated meeting, January 18. Proceedings of the American Philosophical Society, 5(??):121–123, ???? 1850. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143491.

Anonymous:1850:SMJa

[Ano50m] Anonymous. Stated meeting, January 4. Proceedings of the American Philosophical Society, 5(??):119–121, ???? 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143490.

Anonymous:1850:SMJd

[Ano50n] Anonymous. Stated meeting, July 19. Proceedings of the American Philosophical Society, 5(45):154–155, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143377.

Anonymous:1850:SMJc

[Ano50o] Anonymous. Stated meeting, June 21. Proceedings of the American Philosophical Society, 5(45):151–153, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143376.

Anonymous:1850:SMMa

[Ano50p] Anonymous. Stated meeting, March 1. Proceedings of the American Philosophical Society, 5(??):136–138, ???? 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143494.

Anonymous:1850:SMMb

[Ano50q] Anonymous. Stated meeting, March 15. Proceedings of the American Philosophical Society, 5(??):138–139, ???? 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143495.

Anonymous:1850:SMMd

[Ano50r] Anonymous. Stated meeting, May 17. Proceedings of the American Philosophical Society, 5(45):147–150, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143375.

Anonymous:1850:SMMc

[Ano50s] Anonymous. Stated meeting, May 3. Proceedings of the American Philosophical Society, 5(45):145–147, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143374.

Anonymous:1850:SMNc

[Ano50t] Anonymous. Stated meeting, November 1. Proceedings of the American Philosophical Society, 5(45):165–167, April/December 1850. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143382.

Anonymous:1850:SMNd

[Ano50u] Anonymous. Stated meeting, November 15. Proceedings of the American Philosophical Society, 5(45):168–169, April/December 1850. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143383.

Anonymous:1850:SMNb

[Ano50v] Anonymous. Stated meeting, November 16. Proceedings of the American Philosophical Society, 5(??):110–111, ???? 1850. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143487.

Anonymous:1850:SMNa

[Ano50w] Anonymous. Stated meeting, November 2. Proceedings of the American Philosophical Society, 5(??):109–110, ???? 1850. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143486.

Anonymous:1850:SMOd

[Ano50x] Anonymous. Stated meeting, October 18. Proceedings of the American Philosophical Society, 5(45):164–165, April/December 1850. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143381.

Anonymous:1850:SMOb

[Ano50y] Anonymous. Stated meeting, October 19. Proceedings of the American Philosophical Society, 5(??):107–109, ???? 1850. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143485.

Anonymous:1850:SMOc

[Ano50z] Anonymous. Stated meeting, October 4. Proceedings of the American Philosophical Society, 5(45):163, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143380.

Anonymous:1850:SMOa

[Ano50-27] Anonymous. Stated meeting, October 5. Proceedings of the American Philosophical Society, 5(??):105–107, ???? 1850. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143484.

Anonymous:1850:SMS

[Ano50-28] Anonymous. Stated meeting, September 20. Proceedings of the American Philosophical Society, 5(45):157–163, April/December 1850. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/3143379.

Anonymous:1851:A

[Ano51a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(46):??, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1851:SMMb

[Ano51b] Anonymous. Special meeting, March 17. Proceedings of the American Philosophical Society, 5(46):185, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143168.

Anonymous:1851:SMAb

[Ano51c] Anonymous. Stated meeting, April 18. Proceedings of the American Philosophical Society, 5(46):193–196, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143171.

Anonymous:1851:SMAa

[Ano51d] Anonymous. Stated meeting, April 4. Proceedings of the American Philosophical Society, 5(46):190–193, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143170.

Anonymous:1851:SMFb

[Ano51e] Anonymous. Stated meeting, February 21. Proceedings of the American Philosophical Society, 5(46):183–184, January/July 1851. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143166.

Anonymous:1851:SMFa

[Ano51f] Anonymous. Stated meeting, February 7. Proceedings of the American Philosophical Society, 5(46):181–183, January/July 1851. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143165.

Anonymous:1851:SMJb

[Ano51g] Anonymous. Stated meeting, January 17. Proceedings of the American Philosophical Society, 5(46):178–181, January/July 1851. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143164.

Anonymous:1851:SMJa

[Ano51h] Anonymous. Stated meeting, January 3. Proceedings of the American Philosophical Society, 5(46):175–178, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143163.

Anonymous:1851:SMJc

[Ano51i] Anonymous. Stated meeting, June 20. Proceedings of the American Philosophical Society, 5(46):202–209, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143174.

Anonymous:1851:SMMc

[Ano51j] Anonymous. Stated meeting, March 21. Proceedings of the American Philosophical Society, 5(46):186–189, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143169.

Anonymous:1851:SMMa

[Ano51k] Anonymous. Stated meeting, March 7. Proceedings of the American Philosophical Society, 5(46):184–185, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143167.

Anonymous:1851:SMMe

[Ano51l] Anonymous. Stated meeting, May 16. Proceedings of the American Philosophical Society, 5(46):200–202, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143173.

Anonymous:1851:SMMd

[Ano51m] Anonymous. Stated meeting, May 2. Proceedings of the American Philosophical Society, 5(46):196–200, January/July 1851. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143172.

Anonymous:1852:Aa

[Ano52a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(??):??, ???? 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1852:Ab

[Ano52b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(48):??, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1852:SMJa

[Ano52c] Anonymous. Special meeting, July 14. Proceedings of the American Philosophical Society, 5(??):211, ???? 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143454.

Anonymous:1852:SMAc

[Ano52d] Anonymous. Stated meeting, April 16. Proceedings of the American Philosophical Society, 5(48):263–265, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143311.

Anonymous:1852:SMAb

[Ano52e] Anonymous. Stated meeting, April 2. Proceedings of the American Philosophical Society, 5(48):255–262, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143310.

Anonymous:1852:SMAa

[Ano52f] Anonymous. Stated meeting, August 15. Proceedings of the American Philosophical Society, 5(??):214–216, ???? 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143456.

Anonymous:1852:SMAd

[Ano52g] Anonymous. Stated meeting, August 20. Proceedings of the American Philosophical Society, 5(48):280–282, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143316.

Anonymous:1852:SMDd

[Ano52h] Anonymous. Stated meeting, December 17. Proceedings of the American Philosophical Society, 5(48):297–300, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143323.

Anonymous:1852:SMDb

[Ano52i] Anonymous. Stated meeting, December 19. Proceedings of the American Philosophical Society, 5(??):234–236, ???? 1852. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143463.

Anonymous:1852:SMDc

[Ano52j] Anonymous. Stated meeting, December 3. Proceedings of the American Philosophical Society, 5(48):295–297, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143322.

Anonymous:1852:SMDa

[Ano52k] Anonymous. Stated meeting, December 5. Proceedings of the American Philosophical Society, 5(??):231–233, ???? 1852. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143462.

Anonymous:1852:SMFb

[Ano52l] Anonymous. Stated meeting, February 20. Proceedings of the American Philosophical Society, 5(48):247–250, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143307.

Anonymous:1852:SMFa

[Ano52m] Anonymous. Stated meeting, February 6. Proceedings of the American Philosophical Society, 5(48):245–247, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143306.

Anonymous:1852:SMJd

[Ano52n] Anonymous. Stated meeting, January 16. Proceedings of the American Philosophical Society, 5(??):239–244, ???? 1852. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143465.

Anonymous:1852:SMJc

[Ano52o] Anonymous. Stated meeting, January 2, 1852. Proceedings of the American Philosophical Society, 5(??):236–238, ???? 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143464.

Anonymous:1852:SMJf

[Ano52p] Anonymous. Stated meeting, July 16. Proceedings of the American Philosophical Society, 5(48):276–280, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143315.

Anonymous:1852:SMJb

[Ano52q] Anonymous. Stated meeting, July 18. Proceedings of the American Philosophical Society, 5(??):211–214, ???? 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143455.

Anonymous:1852:SMJe

[Ano52r] Anonymous. Stated meeting, June 18. Proceedings of the American Philosophical Society, 5(48):270–275, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143314.

Anonymous:1852:SMMb

[Ano52s] Anonymous. Stated meeting, March 19. Proceedings of the American Philosophical Society, 5(48):253–255, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143309.

Anonymous:1852:SMMa

[Ano52t] Anonymous. Stated meeting, March 5. Proceedings of the American Philosophical Society, 5(48):250–253, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143308.

Anonymous:1852:SMMd

[Ano52u] Anonymous. Stated meeting, May 21. Proceedings of the American Philosophical Society, 5(48):268–270, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143313.

Anonymous:1852:SMMc

[Ano52v] Anonymous. Stated meeting, May 7. Proceedings of the American Philosophical Society, 5(48):265–268, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143312.

Anonymous:1852:SMNd

[Ano52w] Anonymous. Stated meeting, November 19. Proceedings of the American Philosophical Society, 5(48):292–295, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143321.

Anonymous:1852:SMNb

[Ano52x] Anonymous. Stated meeting, November 21. Proceedings of the American Philosophical Society, 5(??):229–230, ???? 1852. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143461.

Anonymous:1852:SMNc

[Ano52y] Anonymous. Stated meeting, November 5. Proceedings of the American Philosophical Society, 5(48):287–291, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143320.

Anonymous:1852:SMNa

[Ano52z] Anonymous. Stated meeting, November 7. Proceedings of the American Philosophical Society, 5(??):225–228, ???? 1852. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143460.

Anonymous:1852:SMOc

[Ano52-27] Anonymous. Stated meeting, October 1. Proceedings of the American Philosophical Society, 5(48):285, February/December 1852. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143318.

Anonymous:1852:SMOd

[Ano52-28] Anonymous. Stated meeting, October 15. Proceedings of the American Philosophical Society, 5(48):286–287, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143319.

Anonymous:1852:SMOb

[Ano52-29] Anonymous. Stated meeting, October 17. Proceedings of the American Philosophical Society, 5(??):223–224, ???? 1852. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143459.

Anonymous:1852:SMOa

[Ano52-30] Anonymous. Stated meeting, October 3. Proceedings of the American Philosophical Society, 5(??):222, ???? 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/3143458.

Anonymous:1852:SMSb

[Ano52-31] Anonymous. Stated meeting, Sept. 17. Proceedings of the American Philosophical Society, 5(48):283–284, February/December 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143317.

Anonymous:1852:SMSa

[Ano52-32] Anonymous. Stated meeting, September 19. Proceedings of the American Philosophical Society, 5(??):216–221, ???? 1852. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143457.

Anonymous:1853:Aa

[Ano53a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(49):??, January/June 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1853:Ab

[Ano53b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 5(50):??, July/December 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1853:SMAa

[Ano53c] Anonymous. Stated meeting, April 1. Proceedings of the American Philosophical Society, 5(49):322–324, January/June 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143225.

Anonymous:1853:SMAb

[Ano53d] Anonymous. Stated meeting, April 15. Proceedings of the American Philosophical Society, 5(49):324–327, January/June 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143226.

Anonymous:1853:SMAc

[Ano53e] Anonymous. Stated meeting, August 19. Proceedings of the American Philosophical Society, 5(50):338–340, July/December 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143084.

Anonymous:1853:SMDb

[Ano53f] Anonymous. Stated meeting, December 16. Proceedings of the American Philosophical Society, 5(50):358–360, July/December 1853. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143091.

Anonymous:1853:SMDa

[Ano53g] Anonymous. Stated meeting, December 2. Proceedings of the American Philosophical Society, 5(50):356–358, July/December 1853. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143090.

Anonymous:1853:SMFb

[Ano53h] Anonymous. Stated meeting, February 18. Proceedings of the American Philosophical Society, 5(49):314–317, January/June 1853. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143222.

Anonymous:1853:SMFa

[Ano53i] Anonymous. Stated meeting, February 4. Proceedings of the American Philosophical Society, 5(49):309–314, January/June 1853. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143221.

Anonymous:1853:SMJb

[Ano53j] Anonymous. Stated meeting, January 21. Proceedings of the American Philosophical Society, 5(49):307–309, January/June 1853. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143220.

Anonymous:1853:SMJa

[Ano53k] Anonymous. Stated meeting, January 7. Proceedings of the American Philosophical Society, 5(49):301–306, January/June 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143219.

Anonymous:1853:SMJd

[Ano53l] Anonymous. Stated meeting, July 15. Proceedings of the American Philosophical Society, 5(50):335–337, July/December 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143083.

Anonymous:1853:SMJc

[Ano53m] Anonymous. Stated meeting, June 17. Proceedings of the American Philosophical Society, 5(49):332–334, January/June 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143229.

Anonymous:1853:SMMb

[Ano53n] Anonymous. Stated meeting, March 18. Proceedings of the American Philosophical Society, 5(49):319–322, January/June 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143224.

Anonymous:1853:SMMa

[Ano53o] Anonymous. Stated meeting, March 4. Proceedings of the American Philosophical Society, 5(49):317–319, January/June 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143223.

Anonymous:1853:SMMd

[Ano53p] Anonymous. Stated meeting, May 20. Proceedings of the American Philosophical Society, 5(49):330–332, January/June 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143228.

Anonymous:1853:SMMc

[Ano53q] Anonymous. Stated meeting, May 6. Proceedings of the American Philosophical Society, 5(49):327–329, January/June 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143227.

Anonymous:1853:SMNb

[Ano53r] Anonymous. Stated meeting, November 18. Proceedings of the American Philosophical Society, 5(50):352–356, July/December 1853. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143089.

Anonymous:1853:SMNa

[Ano53s] Anonymous. Stated meeting, November 4. Proceedings of the American Philosophical Society, 5(50):347–351, July/December 1853. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143088.

Anonymous:1853:SMOb

[Ano53t] Anonymous. Stated meeting, October 21. Proceedings of the American Philosophical Society, 5(50):345–347, July/December 1853. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143087.

Anonymous:1853:SMOa

[Ano53u] Anonymous. Stated meeting, October 7. Proceedings of the American Philosophical Society, 5(50):343–345, July/December 1853. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/3143086.

Anonymous:1853:SMS

[Ano53v] Anonymous. Stated meeting, September 16. Proceedings of the American Philosophical Society, 5(50):340–342, July/December 1853. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/3143085.

Anonymous:1853:VI

[Ano53w] Anonymous. Volume information. Proceedings of the American Philosophical Society, 5(50):361–367, July/December 1853. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/3143082.

Anonymous:1854:Aa

[Ano54a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(51):??, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1854:Ab

[Ano54b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(52):??, July/December 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1854:SMAb

[Ano54c] Anonymous. Stated meeting, April 21. Proceedings of the American Philosophical Society, 6(51):22–26, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981428.

Anonymous:1854:SMAa

[Ano54d] Anonymous. Stated meeting, April 7. Proceedings of the American Philosophical Society, 6(51):21–22, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981427.

Anonymous:1854:SMAc

[Ano54e] Anonymous. Stated meeting, August 18. Proceedings of the American Philosophical Society, 6(52):45–47, July/December 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981386.

Anonymous:1854:SMDa

[Ano54f] Anonymous. Stated meeting, December 1. Proceedings of the American Philosophical Society, 6(52):58–65, July/December 1854. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981392.

Anonymous:1854:SMDb

[Ano54g] Anonymous. Stated meeting, December 15. Proceedings of the American Philosophical Society, 6(52):65–68, July/December 1854. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981393.

Anonymous:1854:SMFb

[Ano54h] Anonymous. Stated meeting, February 17. Proceedings of the American Philosophical Society, 6(51):9–16, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981424.

Anonymous:1854:SMFa

[Ano54i] Anonymous. Stated meeting, February 3. Proceedings of the American Philosophical Society, 6(51):6–9, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981423.

Anonymous:1854:SMJb

[Ano54j] Anonymous. Stated meeting, January 20. Proceedings of the American Philosophical Society, 6(51):4–6, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981422.

Anonymous:1854:SMJa

[Ano54k] Anonymous. Stated meeting, January 6. Proceedings of the American Philosophical Society, 6(51):1–4, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981421.

Anonymous:1854:SMJd

[Ano54l] Anonymous. Stated meeting, July 21. Proceedings of the American Philosophical Society, 6(52):43–44, July/December 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981385.

Anonymous:1854:SMJc

[Ano54m] Anonymous. Stated meeting, June 16. Proceedings of the American Philosophical Society, 6(51):35–41, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981431.

Anonymous:1854:SMMb

[Ano54n] Anonymous. Stated meeting, March 17. Proceedings of the American Philosophical Society, 6(51):19–20, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981426.

Anonymous:1854:SMMa

[Ano54o] Anonymous. Stated meeting, March 3. Proceedings of the American Philosophical Society, 6(51):16–19, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981425.

Anonymous:1854:SMMd

[Ano54p] Anonymous. Stated meeting, May 19. Proceedings of the American Philosophical Society, 6(51):33–35, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981430.

Anonymous:1854:SMMc

[Ano54q] Anonymous. Stated meeting, May 5. Proceedings of the American Philosophical Society, 6(51):27–33, January/June 1854. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981429.

Anonymous:1854:SMNa

[Ano54r] Anonymous. Stated meeting, November 13. Proceedings of the American Philosophical Society, 6(52):54–56, July/December 1854. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981390.

Anonymous:1854:SMNb

[Ano54s] Anonymous. Stated meeting, November 17. Proceedings of the American Philosophical Society, 6(52):57–58, July/December 1854. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981391.

Anonymous:1854:SMOb

[Ano54t] Anonymous. Stated meeting, October 20. Proceedings of the American Philosophical Society, 6(52):51–54, July/December 1854. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981389.

Anonymous:1854:SMOa

[Ano54u] Anonymous. Stated meeting, October 6. Proceedings of the American Philosophical Society, 6(52):49–51, July/December 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981388.

Anonymous:1854:SMS

[Ano54v] Anonymous. Stated meeting, September 15. Proceedings of the American Philosophical Society, 6(52):47–49, July/December 1854. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981387.

Anonymous:1854:VI

[Ano54w] Anonymous. Volume information. Proceedings of the American Philosophical Society, 6(51):??, January/June 1854. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981420.

Anonymous:1855:Aa

[Ano55a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(53):??, January/April 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1855:Ab

[Ano55b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(54):??, May/December 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1855:BM

[Ano55c] Anonymous. Back matter. Proceedings of the American Philosophical Society, 6(54):166, May/December 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981860.

Anonymous:1855:SMAb

[Ano55d] Anonymous. Stated meeting, April 20. Proceedings of the American Philosophical Society, 6(53):119–121, January/April 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981559.

Anonymous:1855:SMAa

[Ano55e] Anonymous. Stated meeting, April 6. Proceedings of the American Philosophical Society, 6(53):113–119, January/April 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981558.

Anonymous:1855:SMAc

[Ano55f] Anonymous. Stated meeting, August 17. Proceedings of the American Philosophical Society, 6(54):138–139, May/December 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981852.

Anonymous:1855:SMDb

[Ano55g] Anonymous. Stated meeting, December 21. Proceedings of the American Philosophical Society, 6(54):164–165, May/December 1855. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981859.

Anonymous:1855:SMDa

[Ano55h] Anonymous. Stated meeting, December 7. Proceedings of the American Philosophical Society, 6(54):156–164, May/December 1855. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981858.

Anonymous:1855:SMFb

[Ano55i] Anonymous. Stated meeting, February 16. Proceedings of the American Philosophical Society, 6(53):85–100, January/April 1855. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981555.

Anonymous:1855:SMFa

[Ano55j] Anonymous. Stated meeting, February 2. Proceedings of the American Philosophical Society, 6(53):82–85, January/April 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981554.

Anonymous:1855:SMJb

[Ano55k] Anonymous. Stated meeting, January 19. Proceedings of the American Philosophical Society, 6(53):76–81, January/April 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981553.

Anonymous:1855:SMJa

[Ano55l] Anonymous. Stated meeting, January 5. Proceedings of the American Philosophical Society, 6(53):69–76, January/April 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981552.

Anonymous:1855:SMJd

[Ano55m] Anonymous. Stated meeting, July 20. Proceedings of the American Philosophical Society, 6(54):132–138, May/December 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981851.

Anonymous:1855:SMJc

[Ano55n] Anonymous. Stated meeting, June 15. Proceedings of the American Philosophical Society, 6(54):129–132, May/December 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981850.

Anonymous:1855:SMMb

[Ano55o] Anonymous. Stated meeting, March 16. Proceedings of the American Philosophical Society, 6(53):105–113, January/April 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981557.

Anonymous:1855:SMMa

[Ano55p] Anonymous. Stated meeting, March 2. Proceedings of the American Philosophical Society, 6(53):100–105, January/April 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981556.

Anonymous:1855:SMMd

[Ano55q] Anonymous. Stated meeting, May 18. Proceedings of the American Philosophical Society, 6(54):125–128, May/December 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981849.

Anonymous:1855:SMMc

[Ano55r] Anonymous. Stated meeting, May 4. Proceedings of the American Philosophical Society, 6(54):123–124, May/December 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981848.

Anonymous:1855:SMNb

[Ano55s] Anonymous. Stated meeting, November 16. Proceedings of the American Philosophical Society, 6(54):153–155, May/December 1855. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981857.

Anonymous:1855:SMNa

[Ano55t] Anonymous. Stated meeting, November 2. Proceedings of the American Philosophical Society, 6(54):151–153, May/December 1855. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981856.

Anonymous:1855:SMOb

[Ano55u] Anonymous. Stated meeting, October 19. Proceedings of the American Philosophical Society, 6(54):145–151, May/December 1855. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981855.

Anonymous:1855:SMOa

[Ano55v] Anonymous. Stated meeting, October 5. Proceedings of the American Philosophical Society, 6(54):143–145, May/December 1855. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981854.

Anonymous:1855:SMS

[Ano55w] Anonymous. Stated meeting, September 21. Proceedings of the American Philosophical Society, 6(54):140–143, May/December 1855. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981853.

Anonymous:1856:Aa

[Ano56a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(55):??, January/June 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1856:Ab

[Ano56b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(56):??, July/December 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1856:SMAb

[Ano56c] Anonymous. Stated meeting, April 18. Proceedings of the American Philosophical Society, 6(55):185–188, January/June 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981967.

Anonymous:1856:SMAa

[Ano56d] Anonymous. Stated meeting, April 4. Proceedings of the American Philosophical Society, 6(55):182–184, January/June 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981966.

Anonymous:1856:SMAc

[Ano56e] Anonymous. Stated meeting, August 15. Proceedings of the American Philosophical Society, 6(56):206–208, July/December 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981544.

Anonymous:1856:SMDb

[Ano56f] Anonymous. Stated meeting, December 19. Proceedings of the American Philosophical Society, 6(56):228–230, July/December 1856. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981551.

Anonymous:1856:SMDa

[Ano56g] Anonymous. Stated meeting, December 5. Proceedings of the American Philosophical Society, 6(56):221–228, July/December 1856. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981550.

Anonymous:1856:SMFa

[Ano56h] Anonymous. Stated meeting, February 1. Proceedings of the American Philosophical Society, 6(55):173–175, January/June 1856. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981962.

Anonymous:1856:SMFb

[Ano56i] Anonymous. Stated meeting, February 15. Proceedings of the American Philosophical Society, 6(55):176–177, January/June 1856. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981963.

Anonymous:1856:SMJb

[Ano56j] Anonymous. Stated meeting, January 18. Proceedings of the American Philosophical Society, 6(55):168–173, January/June 1856. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981961.

Anonymous:1856:SMJa

[Ano56k] Anonymous. Stated meeting, January 4. Proceedings of the American Philosophical Society, 6(55):167–168, January/June 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981960.

Anonymous:1856:SMJd

[Ano56l] Anonymous. Stated meeting, July 18. Proceedings of the American Philosophical Society, 6(56):203–206, July/December 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981543.

Anonymous:1856:SMJc

[Ano56m] Anonymous. Stated meeting, June 20. Proceedings of the American Philosophical Society, 6(55):197–202, January/June 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981970.

Anonymous:1856:SMMb

[Ano56n] Anonymous. Stated meeting, March 21. Proceedings of the American Philosophical Society, 6(55):180–182, January/June 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981965.

Anonymous:1856:SMMa

[Ano56o] Anonymous. Stated meeting, March 7. Proceedings of the American Philosophical Society, 6(55):178–180, January/June 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981964.

Anonymous:1856:SMMd

[Ano56p] Anonymous. Stated meeting, May 16. Proceedings of the American Philosophical Society, 6(55):191–197, January/June 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981969.

Anonymous:1856:SMMc

[Ano56q] Anonymous. Stated meeting, May 2. Proceedings of the American Philosophical Society, 6(55):188–191, January/June 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981968.

Anonymous:1856:SMNb

[Ano56r] Anonymous. Stated meeting, November 21. Proceedings of the American Philosophical Society, 6(56):220–221, July/December 1856. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981549.

Anonymous:1856:SMNa

[Ano56s] Anonymous. Stated meeting, November 7. Proceedings of the American Philosophical Society, 6(56):216–220, July/December 1856. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981548.

Anonymous:1856:SMOb

[Ano56t] Anonymous. Stated meeting, October, 17. Proceedings of the American Philosophical Society, 6(56):213–216, July/December 1856. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981547.

Anonymous:1856:SMOa

[Ano56u] Anonymous. Stated meeting, October 3. Proceedings of the American Philosophical Society, 6(56):212–213, July/December 1856. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981546.

Anonymous:1856:SMS

[Ano56v] Anonymous. Stated meeting, September 19. Proceedings of the American Philosophical Society, 6(56):208–212, July/December 1856. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981545.

Anonymous:1857:Aa

[Ano57a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(57):??, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1857:Ab

[Ano57b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(58):??, July/December 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1857:BM

[Ano57c] Anonymous. Back matter. Proceedings of the American Philosophical Society, 6(57):??, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981919.

Anonymous:1857:SMFc

[Ano57d] Anonymous. Special meeting, February 27. Proceedings of the American Philosophical Society, 6(57):240–241, January/June 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981910.

Anonymous:1857:SMJd

[Ano57e] Anonymous. Special meeting, July 3. Proceedings of the American Philosophical Society, 6(58):261, July/December 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981978.

Anonymous:1857:SMMe

[Ano57f] Anonymous. Special meeting, May 29. Proceedings of the American Philosophical Society, 6(57):257–258, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981917.

Anonymous:1857:SMAb

[Ano57g] Anonymous. Stated meeting, April 17. Proceedings of the American Philosophical Society, 6(57):249–252, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981914.

Anonymous:1857:SMAa

[Ano57h] Anonymous. Stated meeting, April 3. Proceedings of the American Philosophical Society, 6(57):247–248, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981913.

Anonymous:1857:SMAc

[Ano57i] Anonymous. Stated meeting, August 21. Proceedings of the American Philosophical Society, 6(58):263–267, July/December 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981980.

Anonymous:1857:SMDb

[Ano57j] Anonymous. Stated meeting, December 18. Proceedings of the American Philosophical Society, 6(58):279–280, July/December 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981987.

Anonymous:1857:SMDa

[Ano57k] Anonymous. Stated meeting, December 4. Proceedings of the American Philosophical Society, 6(58):278–279, July/December 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981986.

Anonymous:1857:SMFb

[Ano57l] Anonymous. Stated meeting, February 20. Proceedings of the American Philosophical Society, 6(57):237–240, January/June 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981909.

Anonymous:1857:SMFa

[Ano57m] Anonymous. Stated meeting, February 6. Proceedings of the American Philosophical Society, 6(57):235–237, January/June 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981908.

Anonymous:1857:SMJb

[Ano57n] Anonymous. Stated meeting, January 16. Proceedings of the American Philosophical Society, 6(57):233–234, January/June 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981907.

Anonymous:1857:SMJa

[Ano57o] Anonymous. Stated meeting, January 2. Proceedings of the American Philosophical Society, 6(57):231–232, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981906.

Anonymous:1857:SMJe

[Ano57p] Anonymous. Stated meeting, July 17. Proceedings of the American Philosophical Society, 6(58):261–263, July/December 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981979.

Anonymous:1857:SMJc

[Ano57q] Anonymous. Stated meeting, June 19. Proceedings of the American Philosophical Society, 6(57):258–260, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981918.

Anonymous:1857:SMMb

[Ano57r] Anonymous. Stated meeting, March 20. Proceedings of the American Philosophical Society, 6(57):244–247, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981912.

Anonymous:1857:SMMa

[Ano57s] Anonymous. Stated meeting, March 6. Proceedings of the American Philosophical Society, 6(57):241–244, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981911.

Anonymous:1857:SMMc

[Ano57t] Anonymous. Stated meeting, May 1. Proceedings of the American Philosophical Society, 6(57):252–254, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981915.

Anonymous:1857:SMMd

[Ano57u] Anonymous. Stated meeting, May 15. Proceedings of the American Philosophical Society, 6(57):254–257, January/June 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981916.

Anonymous:1857:SMNb

[Ano57v] Anonymous. Stated meeting, November 20. Proceedings of the American Philosophical Society, 6(58):276–278, July/December 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981985.

Anonymous:1857:SMNa

[Ano57w] Anonymous. Stated meeting, November 6. Proceedings of the American Philosophical Society, 6(58):273–276, July/December 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981984.

Anonymous:1857:SMOb

[Ano57x] Anonymous. Stated meeting, October 16. Proceedings of the American Philosophical Society, 6(58):271–273, July/December 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981983.

Anonymous:1857:SMOa

[Ano57y] Anonymous. Stated meeting, October 2. Proceedings of the American Philosophical Society, 6(58):270, July/December 1857. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981982.

Anonymous:1857:SMS

[Ano57z] Anonymous. Stated meeting, September 18. Proceedings of the American Philosophical Society, 6(58):267–269, July/December 1857. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981981.

Anonymous:1858:Aa

[Ano58a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(59):??, January/June 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1858:Ab

[Ano58b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 6(60):??, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1858:SMFc

[Ano58c] Anonymous. Special meeting, February 23. Proceedings of the American Philosophical Society, 6(59):289–290, January/June 1858. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982192.

Anonymous:1858:SMAb

[Ano58d] Anonymous. Stated meeting, April 16. Proceedings of the American Philosophical Society, 6(59):304–308, January/June 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982196.

Anonymous:1858:SMAa

[Ano58e] Anonymous. Stated meeting, April 2. Proceedings of the American Philosophical Society, 6(59):302–304, January/June 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982195.

Anonymous:1858:SMAc

[Ano58f] Anonymous. Stated meeting, August 20, 1858. Proceedings of the American Philosophical Society, 6(60):322–323, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981990.

Anonymous:1858:SMDb

[Ano58g] Anonymous. Stated meeting, December 17, 1858. Proceedings of the American Philosophical Society, 6(60):345–346, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981997.

Anonymous:1858:SMDa

[Ano58h] Anonymous. Stated meeting, December 3, 1858. Proceedings of the American Philosophical Society, 6(60):343–345, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981996.

Anonymous:1858:SMFb

[Ano58i] Anonymous. Stated meeting, February 19. Proceedings of the American Philosophical Society, 6(59):287–288, January/June 1858. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982191.

Anonymous:1858:SMFa

[Ano58j] Anonymous. Stated meeting, February 5. Proceedings of the American Philosophical Society, 6(59):284–287, January/June 1858. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982190.

Anonymous:1858:SMJa

[Ano58k] Anonymous. Stated meeting, January 1. Proceedings of the American Philosophical Society, 6(59):281–282, January/June 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982188.

Anonymous:1858:SMJb

[Ano58l] Anonymous. Stated meeting, January 15. Proceedings of the American Philosophical Society, 6(59):283–284, January/June 1858. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982189.

Anonymous:1858:SMJd

[Ano58m] Anonymous. Stated meeting, July 16, 1858. Proceedings of the American Philosophical Society, 6(60):321, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981989.

Anonymous:1858:SMJc

[Ano58n] Anonymous. Stated meeting, June 18. Proceedings of the American Philosophical Society, 6(59):317–320, January/June 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982199.

Anonymous:1858:SMMb

[Ano58o] Anonymous. Stated meeting, March 19. Proceedings of the American Philosophical Society, 6(59):298–302, January/June 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982194.

Anonymous:1858:SMMa

[Ano58p] Anonymous. Stated meeting, March 5. Proceedings of the American Philosophical Society, 6(59):290–298, January/June 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982193.

Anonymous:1858:SMMd

[Ano58q] Anonymous. Stated meeting, May 21. Proceedings of the American Philosophical Society, 6(59):316–317, January/June 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982198.

Anonymous:1858:SMMc

[Ano58r] Anonymous. Stated meeting, May 7. Proceedings of the American Philosophical Society, 6(59):308–316, January/June 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982197.

Anonymous:1858:SMNb

[Ano58s] Anonymous. Stated meeting, November 19, 1858. Proceedings of the American Philosophical Society, 6(60):332–342, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981995.

Anonymous:1858:SMNa

[Ano58t] Anonymous. Stated meeting, November 5, 1858. Proceedings of the American Philosophical Society, 6(60):330–332, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981994.

Anonymous:1858:SMOa

[Ano58u] Anonymous. Stated meeting, October 1, 1858. Proceedings of the American Philosophical Society, 6(60):325–329, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981992.

Anonymous:1858:SMOb

[Ano58v] Anonymous. Stated meeting, October 15, 1858. Proceedings of the American Philosophical Society, 6(60):329–330, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981993.

Anonymous:1858:SMS

[Ano58w] Anonymous. Stated meeting, September 17, 1858. Proceedings of the American Philosophical Society, 6(60):324–325, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981991.

Anonymous:1858:VI

[Ano58x] Anonymous. Volume information. Proceedings of the American Philosophical Society, 6(60):389–396, July/December 1858. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981988.

Anonymous:1859:Aa

[Ano59a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 7(61):??, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1859:Ab

[Ano59b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 7(62):??, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1859:BM

[Ano59c] Anonymous. Back matter. Proceedings of the American Philosophical Society, 7(62):166, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981827.

Anonymous:1859:E

[Ano59d] Anonymous. Errata. Proceedings of the American Philosophical Society, 7(61):??, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981933.

Anonymous:1859:ENV

[Ano59e] Anonymous. Errata in no. 60 (vol. VI.), and no. 61 (vol. VII.), of the Proceedings of the American Philosophical Society. Proceedings of the American Philosophical Society, 7(62):167–168, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981826.

Anonymous:1859:SMAa

[Ano59f] Anonymous. Stated meeting, April 1, 1859. Proceedings of the American Philosophical Society, 7(61):14–15, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981927.

Anonymous:1859:SMAb

[Ano59g] Anonymous. Stated meeting, April 15, 1859. Proceedings of the American Philosophical Society, 7(61):15–17, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981928.

Anonymous:1859:SMAc

[Ano59h] Anonymous. Stated meeting, August 19, 1859. Proceedings of the American Philosophical Society, 7(62):121–122, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981817.

Anonymous:1859:SMDb

[Ano59i] Anonymous. Stated meeting, December 16, 1859. Proceedings of the American Philosophical Society, 7(62):164–165, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981825.

Anonymous:1859:SMDa

[Ano59j] Anonymous. Stated meeting, December 2, 1859. Proceedings of the American Philosophical Society, 7(62):162–164, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981824.

Anonymous:1859:SMFb

[Ano59k] Anonymous. Stated meeting, February 18, 1859. Proceedings of the American Philosophical Society, 7(61):7–9, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981924.

Anonymous:1859:SMFa

[Ano59l] Anonymous. Stated meeting, February 4, 1859. Proceedings of the American Philosophical Society, 7(61):6–7, January/June 1859. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981923.

Anonymous:1859:SMJb

[Ano59m] Anonymous. Stated meeting, January 21, 1859. Proceedings of the American Philosophical Society, 7(61):3–5, January/June 1859. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981922.

Anonymous:1859:SMJa

[Ano59n] Anonymous. Stated meeting, January 7, 1859. Proceedings of the American Philosophical Society, 7(61):1–3, January/June 1859. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981921.

Anonymous:1859:SMJd

[Ano59o] Anonymous. Stated meeting, July 15, 1859. Proceedings of the American Philosophical Society, 7(61):23–119, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981932.

Anonymous:1859:SMJc

[Ano59p] Anonymous. Stated meeting, June 17, 1859. Proceedings of the American Philosophical Society, 7(61):21–22, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981931.

Anonymous:1859:SMMb

[Ano59q] Anonymous. Stated meeting, March 18, 1859. Proceedings of the American Philosophical Society, 7(61):12–14, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981926.

Anonymous:1859:SMMa

[Ano59r] Anonymous. Stated meeting, March 4, 1859. Proceedings of the American Philosophical Society, 7(61):9–12, January/June 1859. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981925.

Anonymous:1859:SMMd

[Ano59s] Anonymous. Stated meeting, May 20, 1859. Proceedings of the American Philosophical Society, 7(61):19–21, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981930.

Anonymous:1859:SMMc

[Ano59t] Anonymous. Stated meeting, May 6, 1859. Proceedings of the American Philosophical Society, 7(61):17–19, January/June 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981929.

Anonymous:1859:SMNb

[Ano59u] Anonymous. Stated meeting, November 18, 1859. Proceedings of the American Philosophical Society, 7(62):155–162, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981823.

Anonymous:1859:SMNa

[Ano59v] Anonymous. Stated meeting, November 4, 1859. Proceedings of the American Philosophical Society, 7(62):127–128, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981821.

Anonymous:1859:SMOb

[Ano59w] Anonymous. Stated meeting, October 21, 1859. Proceedings of the American Philosophical Society, 7(62):126–127, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981820.

Anonymous:1859:SMOa

[Ano59x] Anonymous. Stated meeting, October 7, 1859. Proceedings of the American Philosophical Society, 7(62):123–125, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981819.

Anonymous:1859:SMS

[Ano59y] Anonymous. Stated meeting, September 16, 1859. Proceedings of the American Philosophical Society, 7(62):122–123, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981818.

Anonymous:1859:VI

[Ano59z] Anonymous. Volume information. Proceedings of the American Philosophical Society, 7(61):??, January/June 1859. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981920.

Anonymous:1860:Aa

[Ano60a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 7(63):??, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1860:Ab

[Ano60b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 7(64):??, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1860:BNL

[Ano60c] Anonymous. Biographical notice of the Late Thomas Nuttall. Proceedings of the American Philosophical Society, 7(63):297–315, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981892.

Anonymous:1860:E

[Ano60d] Anonymous. Errata. Proceedings of the American Philosophical Society, 7(63):291, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981896.

Anonymous:1860:SMAb

[Ano60e] Anonymous. Stated meeting, April 20, 1860. Proceedings of the American Philosophical Society, 7(63):326–328, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981895.

Anonymous:1860:SMAa

[Ano60f] Anonymous. Stated meeting, April 6th, 1860. Proceedings of the American Philosophical Society, 7(63):319–320, 325–326, January/ June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981893.

Anonymous:1860:SMAc

[Ano60g] Anonymous. Stated meeting, August 17, 1860. Proceedings of the American Philosophical Society, 7(64):345–346, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982180.

Anonymous:1860:SMDb

[Ano60h] Anonymous. Stated meeting, December 21, 1860. Proceedings of the American Philosophical Society, 7(64):409–416, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982187.

Anonymous:1860:SMDa

[Ano60i] Anonymous. Stated meeting, December 7th, 1860. Proceedings of the American Philosophical Society, 7(64):396–408, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982186.

Anonymous:1860:SMFb

[Ano60j] Anonymous. Stated meeting, February 17, 1860. Proceedings of the American Philosophical Society, 7(63):176, 292, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981888.

Anonymous:1860:SMFa

[Ano60k] Anonymous. Stated meeting, February 3, 1860. Proceedings of the American Philosophical Society, 7(63):173–175, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981887.

Anonymous:1860:SMJb

[Ano60l] Anonymous. Stated meeting, January 20, 1860. Proceedings of the American Philosophical Society, 7(63):172–173, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981886.

Anonymous:1860:SMJa

[Ano60m] Anonymous. Stated meeting, January 6, 1860. Proceedings of the American Philosophical Society, 7(63):169–172, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981885.

Anonymous:1860:SMJd

[Ano60n] Anonymous. Stated meeting, July 20, 1860. Proceedings of the American Philosophical Society, 7(64):340–344, May/December 1860. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982179.

Anonymous:1860:SMJc

[Ano60o] Anonymous. Stated meeting, June 15, 1860. Proceedings of the American Philosophical Society, 7(64):337–340, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982178.

Anonymous:1860:SMMb

[Ano60p] Anonymous. Stated meeting, March 16, 1860. Proceedings of the American Philosophical Society, 7(63):294–296, 316–318, January/ June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981891.

Anonymous:1860:SMMa

[Ano60q] Anonymous. Stated meeting, March 2, 1860. Proceedings of the American Philosophical Society, 7(63):293–294, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981890.

Anonymous:1860:SMMd

[Ano60r] Anonymous. Stated meeting, May 18, 1860. Proceedings of the American Philosophical Society, 7(64):330–337, May/December 1860. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982177.

Anonymous:1860:SMMc

[Ano60s] Anonymous. Stated meeting, May 4, 1860. Proceedings of the American Philosophical Society, 7(64):329–330, May/December 1860. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982176.

Anonymous:1860:SMNb

[Ano60t] Anonymous. Stated meeting, Nov. 16, 1860. Proceedings of the American Philosophical Society, 7(64):394–396, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982185.

Anonymous:1860:SMNa

[Ano60u] Anonymous. Stated meeting November 2, 1860. Proceedings of the American Philosophical Society, 7(64):391–394, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982184.

Anonymous:1860:SMOb

[Ano60v] Anonymous. Stated meeting October 19, 1860. Proceedings of the American Philosophical Society, 7(64):390–391, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982183.

Anonymous:1860:SMOa

[Ano60w] Anonymous. Stated meeting, October 5, 1860. Proceedings of the American Philosophical Society, 7(64):385–390, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982182.

Anonymous:1860:SMS

[Ano60x] Anonymous. Stated meeting, September 21, 1860. Proceedings of the American Philosophical Society, 7(64):346–385, May/December 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982181.

Anonymous:1860:VI

[Ano60y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 7(64):417–421, May/December 1860. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982175.

Anonymous:1861:Aa

[Ano61a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 8(65):??, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1861:Ab

[Ano61b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 8(66):??, July/December 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1861:OND

[Ano61c] Anonymous. Obituary notice of Dr. Samuel Moore. Proceedings of the American Philosophical Society, 8(65):53–56, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982031.

Anonymous:1861:NDG

[Ano61d] Anonymous. On the natural dissemination of gold. Proceedings of the American Philosophical Society, 8(65):273–276, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982039.

Anonymous:1861:SMAb

[Ano61e] Anonymous. Stated meeting, April 19, 1861. Proceedings of the American Philosophical Society, 8(65):259–260, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982034.

Anonymous:1861:SMAa

[Ano61f] Anonymous. Stated meeting, April 5, 1861. Proceedings of the American Philosophical Society, 8(65):258–259, January/June 1861. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982033.

Anonymous:1861:SMAc

[Ano61g] Anonymous. Stated meeting, August 16, 1861. Proceedings of the American Philosophical Society, 8(66):279–280, July/December 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981776.

Anonymous:1861:SMDb

[Ano61h] Anonymous. Stated meeting, December 20, 1861. Proceedings of the American Philosophical Society, 8(66):394–405, July/December 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981784.

Anonymous:1861:SMDa

[Ano61i] Anonymous. Stated meeting, December 6, 1861. Proceedings of the American Philosophical Society, 8(66):381–394, July/December 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981783.

Anonymous:1861:SMFa

[Ano61j] Anonymous. Stated meeting, February 1, 1861. Proceedings of the American Philosophical Society, 8(65):49, January/June 1861. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982027.

Anonymous:1861:SMFb

[Ano61k] Anonymous. Stated meeting, February 15, 1861. Proceedings of the American Philosophical Society, 8(65):49–50, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982028.

Anonymous:1861:SMJb

[Ano61l] Anonymous. Stated meeting, January 18, 1861. Proceedings of the American Philosophical Society, 8(65):3–4, January/June 1861. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982025.

Anonymous:1861:SMJa

[Ano61m] Anonymous. Stated meeting, January 4, 1861. Proceedings of the American Philosophical Society, 8(65):1–2, January/June 1861. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982024.

Anonymous:1861:SMJd

[Ano61n] Anonymous. Stated meeting, July 19, 1861. Proceedings of the American Philosophical Society, 8(66):277–279, July/December 1861. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981775.

Anonymous:1861:SMJc

[Ano61o] Anonymous. Stated meeting, June 21, 1861. Proceedings of the American Philosophical Society, 8(65):263–264, 272, 276, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982037.

Anonymous:1861:SMMa

[Ano61p] Anonymous. Stated meeting, March 1, 1861. Proceedings of the American Philosophical Society, 8(65):50–51, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982029.

Anonymous:1861:SMMb

[Ano61q] Anonymous. Stated meeting, March 15, 1861. Proceedings of the American Philosophical Society, 8(65):51–52, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982030.

Anonymous:1861:SMMd

[Ano61r] Anonymous. Stated meeting, May 17, 1861. Proceedings of the American Philosophical Society, 8(65):261–262, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982036.

Anonymous:1861:SMMc

[Ano61s] Anonymous. Stated meeting, May 3, 1861. Proceedings of the American Philosophical Society, 8(65):260–261, January/June 1861. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982035.

Anonymous:1861:SMNb

[Ano61t] Anonymous. Stated meeting, November 15th, 1861. Proceedings of the American Philosophical Society, 8(66):377–381, July/December 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981782.

Anonymous:1861:SMNa

[Ano61u] Anonymous. Stated meeting, November 1st, 1861. Proceedings of the American Philosophical Society, 8(66):377, July/December 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981781.

Anonymous:1861:SMOb

[Ano61v] Anonymous. Stated meeting, October 18, 1861. Proceedings of the American Philosophical Society, 8(66):284, July/December 1861. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981779.

Anonymous:1861:SMOa

[Ano61w] Anonymous. Stated meeting, October 4, 1861. Proceedings of the American Philosophical Society, 8(66):283, July/December 1861. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981778.

Anonymous:1861:SMS

[Ano61x] Anonymous. Stated meeting, September 20, 1861. Proceedings of the American Philosophical Society, 8(66):280–283, July/December 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981777.

Anonymous:1861:VI

[Ano61y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 8(66):407–408, July/December 1861. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981774.

Anonymous:1862:Aa

[Ano62a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 9(67):??, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1862:Ab

[Ano62b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 9(68):??, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1862:BM

[Ano62c] Anonymous. Back matter. Proceedings of the American Philosophical Society, 9(68):122–124, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982053.

Anonymous:1862:OND

[Ano62d] Anonymous. Obituary notice of Dr. George W. Bethune. Proceedings of the American Philosophical Society, 9(68):70–85, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982045.

Anonymous:1862:ONP

[Ano62e] Anonymous. Obituary notice of Prof. George Tucker. Proceedings of the American Philosophical Society, 9(68):64–70, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982044.

Anonymous:1862:SMAb

[Ano62f] Anonymous. Stated meeting, April 18, 1862. Proceedings of the American Philosophical Society, 9(67):24–27, January 1862. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981539.

Anonymous:1862:SMAa

[Ano62g] Anonymous. Stated meeting, April 4, 1862. Proceedings of the American Philosophical Society, 9(67):13–14, 24, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981537.

Anonymous:1862:SMAc

[Ano62h] Anonymous. Stated meeting, August 15, 1862. Proceedings of the American Philosophical Society, 9(68):57–58, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982041.

Anonymous:1862:SMDb

[Ano62i] Anonymous. Stated meeting, December 19, 1862. Proceedings of the American Philosophical Society, 9(68):115–121, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982052.

Anonymous:1862:SMDa

[Ano62j] Anonymous. Stated meeting, December 5, 1862. Proceedings of the American Philosophical Society, 9(68):114–115, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982051.

Anonymous:1862:SMFb

[Ano62k] Anonymous. Stated meeting, February 21, 1862. Proceedings of the American Philosophical Society, 9(67):10–11, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981534.

Anonymous:1862:SMFa

[Ano62l] Anonymous. Stated meeting, February 7, 1862. Proceedings of the American Philosophical Society, 9(67):8–10, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981533.

Anonymous:1862:SMJb

[Ano62m] Anonymous. Stated meeting, January 17, 1862. Proceedings of the American Philosophical Society, 9(67):5–8, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981532.

Anonymous:1862:SMJa

[Ano62n] Anonymous. Stated meeting, January 3, 1862. Proceedings of the American Philosophical Society, 9(67):1–5, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981531.

Anonymous:1862:SMJc

[Ano62o] Anonymous. Stated meeting, June 20, 1862. Proceedings of the American Philosophical Society, 9(68):55–57, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982040.

Anonymous:1862:SMMb

[Ano62p] Anonymous. Stated meeting, March 21, 1862. Proceedings of the American Philosophical Society, 9(67):12–13, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981536.

Anonymous:1862:SMMa

[Ano62q] Anonymous. Stated meeting, March 7, 1862. Proceedings of the American Philosophical Society, 9(67):11–12, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981535.

Anonymous:1862:SMMd

[Ano62r] Anonymous. Stated meeting, May 16, 1862. Proceedings of the American Philosophical Society, 9(67):28–38, January 1862. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981541.

Anonymous:1862:SMMc

[Ano62s] Anonymous. Stated meeting, May 2, 1862. Proceedings of the American Philosophical Society, 9(67):27–28, January 1862. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981540.

Anonymous:1862:SMNb

[Ano62t] Anonymous. Stated meeting, November 21, 1862. Proceedings of the American Philosophical Society, 9(68):109–110, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982049.

Anonymous:1862:SMNa

[Ano62u] Anonymous. Stated meeting, November 7, 1862. Proceedings of the American Philosophical Society, 9(68):92, 109, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982047.

Anonymous:1862:SMOb

[Ano62v] Anonymous. Stated meeting, October 17, 1862. Proceedings of the American Philosophical Society, 9(68):88–91, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982046.

Anonymous:1862:SMOa

[Ano62w] Anonymous. Stated meeting, October 3, 1862. Proceedings of the American Philosophical Society, 9(68):64, 86–87, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982043.

Anonymous:1862:SMS

[Ano62x] Anonymous. Stated meeting, September 19, 1862. Proceedings of the American Philosophical Society, 9(68):58–63, June/December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982042.

Anonymous:1862:SCC

[Ano62y] Anonymous. Supreme Court. The City of Philadelphia v. the American Philosophical Society. Proceedings of the American Philosophical Society, 9(67):14–24, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981538.

Anonymous:1862:VI

[Ano62z] Anonymous. Volume information. Proceedings of the American Philosophical Society, 9(67):??, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981530.

Anonymous:1863:Aa

[Ano63a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 9(69):??, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1863:Ab

[Ano63b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 9(70):??, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1863:FEG

[Ano63c] Anonymous. The family as an element of government. Proceedings of the American Philosophical Society, 9(70):295–327, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982149.

Anonymous:1863:IOAa

[Ano63d] Anonymous. [Illustration]: Origin of alphabet by P. E. Chase. Plate I. Chinese and Egyptian hieroglyphs. Pelasgic alphabet (about 1400 B.C.). Proceedings of the American Philosophical Society, 9(69):??, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981415.

Anonymous:1863:IOAb

[Ano63e] Anonymous. [Illustration]: Origin of alphabet by P. E. Chase. Plate II. Chinese archetypes. Ancient letters. Proceedings of the American Philosophical Society, 9(69):??, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981416.

Anonymous:1863:IPIa

[Ano63f] Anonymous. [Illustration]: Plate III. Bitumen Vein, W. Va. Proceedings of the American Philosophical Society, 9(69):??, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981417.

Anonymous:1863:IPIb

[Ano63g] Anonymous. [Illustration]: Plate IV. Prospect Colliery, Broad Top. Beaver Meadow Coal Basin. Proceedings of the American Philosophical Society, 9(69):??, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981418.

Anonymous:1863:IPIc

[Ano63h] Anonymous. [Illustration]: Plate IX. J. P. Lesley. Proceedings of the American Philosophical Society, 9(70):??, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982172.

Anonymous:1863:IPVa

[Ano63i] Anonymous. [Illustration]: Plate v. Proceedings of the American Philosophical Society, 9(69):??, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981419.

Anonymous:1863:IPVb

[Ano63j] Anonymous. [Illustration]: Plate VIII. J. P. Lesley. Fig 1. Section at Mont Alto Furnace. Fig. 2. Home Bank. Fig. 3. Home Bank. Fig. 5. Caledonia Bank. Fig. 4. Little Mountain Ore Banks. Fig. 6. Lower Pond Bank. Fig. 7. Upper Pond Bank. Proceedings of the American Philosophical Society, 9(70):??, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982171.

Anonymous:1863:IPXa

[Ano63k] Anonymous. [Illustration]: Plate X. J. P. Lesley. Fig. 8. view of Caledonia Ore Bank. Fig. 9, section across the Cumberland Valley. Fig. 10, section 10 miles east of Harrisburg. Fig 11. Ideal section of the ancient surface. Proceedings of the American Philosophical Society, 9(70):??, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982173.

Anonymous:1863:IPXb

[Ano63l] Anonymous. [Illustration]: Plate XI. J. P. Lesley. Proceedings of the American Philosophical Society, 9(70):??, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982174.

Anonymous:1863:OND

[Ano63m] Anonymous. An obituary notice of Dr. William Darlington. Proceedings of the American Philosophical Society, 9(70):330–342, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982152.

Anonymous:1863:ONR

[Ano63n] Anonymous. Obituary notice of the Rev. Prof. Edward Hitchcock, D.D., LL.D. Proceedings of the American Philosophical Society, 9(70):443–449, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982168.

Anonymous:1863:SMAd

[Ano63o] Anonymous. Stated meeting, April 1, 1864. Proceedings of the American Philosophical Society, 9(70):355–366, June 1863. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982155.

Anonymous:1863:SMAe

[Ano63p] Anonymous. Stated meeting, April 15, 1864. Proceedings of the American Philosophical Society, 9(70):366–371, June 1863. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982156.

Anonymous:1863:SMAb

[Ano63q] Anonymous. Stated meeting, April 17, 1863. Proceedings of the American Philosophical Society, 9(69):205–207, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981412.

Anonymous:1863:SMAa

[Ano63r] Anonymous. Stated meeting, April 3, 1863. Proceedings of the American Philosophical Society, 9(69):204–205, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981411.

Anonymous:1863:SMAf

[Ano63s] Anonymous. Stated meeting, August 19, 1864. Proceedings of the American Philosophical Society, 9(70):412–414, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982162.

Anonymous:1863:SMAc

[Ano63t] Anonymous. Stated meeting, August 21, 1863. Proceedings of the American Philosophical Society, 9(70):240, June 1863. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982139.

Anonymous:1863:SMDd

[Ano63u] Anonymous. Stated meeting, December 16th, 1864. Proceedings of the American Philosophical Society, 9(70):461–496, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982170.

Anonymous:1863:SMDb

[Ano63v] Anonymous. Stated meeting, December 18th, 1863. Proceedings of the American Philosophical Society, 9(70):281–289, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982146.

Anonymous:1863:SMDc

[Ano63w] Anonymous. Stated meeting, December 2, 1864. Proceedings of the American Philosophical Society, 9(70):454–461, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982169.

Anonymous:1863:SMDa

[Ano63x] Anonymous. Stated meeting, December 4th, 1863. Proceedings of the American Philosophical Society, 9(70):278–280, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982145.

Anonymous:1863:SMFd

[Ano63y] Anonymous. Stated meeting, February 19, 1864. Proceedings of the American Philosophical Society, 9(70):330, 342–343, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982151.

Anonymous:1863:SMFb

[Ano63z] Anonymous. Stated meeting, February 20, 1863. Proceedings of the American Philosophical Society, 9(69):144–146, January 1863. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981407.

Anonymous:1863:SMFc

[Ano63-27] Anonymous. Stated meeting, February 5, 1864. Proceedings of the American Philosophical Society, 9(70):328–330, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982150.

Anonymous:1863:SMFa

[Ano63-28] Anonymous. Stated meeting, February 6, 1863. Proceedings of the American Philosophical Society, 9(69):130–143, January 1863. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981406.

Anonymous:1863:SMJe

[Ano63-29] Anonymous. Stated meeting, January 1, 1864. Proceedings of the American Philosophical Society, 9(70):290–294, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982147.

Anonymous:1863:SMJf

[Ano63-30] Anonymous. Stated meeting, January 15, 1864. Proceedings of the American Philosophical Society, 9(70):294–295, 327–328, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982148.

Anonymous:1863:SMJb

[Ano63-31] Anonymous. Stated meeting, January 16, 1863. Proceedings of the American Philosophical Society, 9(69):127–130, January 1863. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981405.

Anonymous:1863:SMJa

[Ano63-32] Anonymous. Stated meeting, January 2, 1863. Proceedings of the American Philosophical Society, 9(69):125–127, January 1863. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981404.

Anonymous:1863:SMJh

[Ano63-33] Anonymous. Stated meeting, July 15, 1864. Proceedings of the American Philosophical Society, 9(70):403–411, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982161.

Anonymous:1863:SMJd

[Ano63-34] Anonymous. Stated meeting, July 17, 1863. Proceedings of the American Philosophical Society, 9(70):237–239, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982138.

Anonymous:1863:SMJg

[Ano63-35] Anonymous. Stated meeting, June 17, 1864. Proceedings of the American Philosophical Society, 9(70):394–400, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982159.

Anonymous:1863:SMJc

[Ano63-36] Anonymous. Stated meeting, June 19, 1863. Proceedings of the American Philosophical Society, 9(70):233–234, 236–237, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982136.

Anonymous:1863:SMMf

[Ano63-37] Anonymous. Stated meeting, March 18, 1864. Proceedings of the American Philosophical Society, 9(70):349–355, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982154.

Anonymous:1863:SMMb

[Ano63-38] Anonymous. Stated meeting, March 20, 1863. Proceedings of the American Philosophical Society, 9(69):164–165, January 1863. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981409.

Anonymous:1863:SMMe

[Ano63-39] Anonymous. Stated meeting, March 4, 1864. Proceedings of the American Philosophical Society, 9(70):343–349, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982153.

Anonymous:1863:SMMa

[Ano63-40] Anonymous. Stated meeting, March 6, 1863. Proceedings of the American Philosophical Society, 9(69):146–164, January 1863. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981408.

Anonymous:1863:SMMc

[Ano63-41] Anonymous. Stated meeting, May 1, 1863. Proceedings of the American Philosophical Society, 9(69):207–226, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981413.

Anonymous:1863:SMMd

[Ano63-42] Anonymous. Stated meeting, May 15, 1863. Proceedings of the American Philosophical Society, 9(69):226–232, January 1863. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981414.

Anonymous:1863:SMMh

[Ano63-43] Anonymous. Stated meeting, May 20, 1864. Proceedings of the American Philosophical Society, 9(70):376–394, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982158.

Anonymous:1863:SMMg

[Ano63-44] Anonymous. Stated meeting, May 6, 1864. Proceedings of the American Philosophical Society, 9(70):372–376, June 1863. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982157.

Anonymous:1863:SMNd

[Ano63-45] Anonymous. Stated meeting, November 18, 1864. Proceedings of the American Philosophical Society, 9(70):442, 449–453, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982167.

Anonymous:1863:SMNb

[Ano63-46] Anonymous. Stated meeting, November 20, 1863. Proceedings of the American Philosophical Society, 9(70):277, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982144.

Anonymous:1863:SMNc

[Ano63-47] Anonymous. Stated meeting, November 4, 1864. Proceedings of the American Philosophical Society, 9(70):441–442, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982166.

Anonymous:1863:SMNa

[Ano63-48] Anonymous. Stated meeting, November 6, 1863. Proceedings of the American Philosophical Society, 9(70):275–276, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982143.

Anonymous:1863:SMOb

[Ano63-49] Anonymous. Stated meeting, October 16, 1863. Proceedings of the American Philosophical Society, 9(70):275, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982142.

Anonymous:1863:SMOa

[Ano63-50] Anonymous. Stated meeting, October 2, 1863. Proceedings of the American Philosophical Society, 9(70):259–274, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982141.

Anonymous:1863:SMOd

[Ano63-51] Anonymous. Stated meeting, October 21, 1864. Proceedings of the American Philosophical Society, 9(70):426–440, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982165.

Anonymous:1863:SMOc

[Ano63-52] Anonymous. Stated meeting, October 7, 1864. Proceedings of the American Philosophical Society, 9(70):424–426, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982164.

Anonymous:1863:SMSb

[Ano63-53] Anonymous. Stated meeting, September 16, 1864. Proceedings of the American Philosophical Society, 9(70):414–424, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982163.

Anonymous:1863:SMSa

[Ano63-54] Anonymous. Stated meeting, September 18, 1863. Proceedings of the American Philosophical Society, 9(70):241–259, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982140.

Anonymous:1863:VI

[Ano63-55] Anonymous. Volume information. Proceedings of the American Philosophical Society, 9(70):497–509, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982135.

Anonymous:1865:Aa

[Ano65a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 10(73):??, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1865:Ab

[Ano65b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 10(74):??, July 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1865:FM

[Ano65c] Anonymous. Front matter. Proceedings of the American Philosophical Society, 10(73):??, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982123.

Anonymous:1865:ONC

[Ano65d] Anonymous. Obituary notice of Charles Wilkins Short, M.D. Proceedings of the American Philosophical Society, 10(74):171–186, July 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981401.

Anonymous:1865:SMAb

[Ano65e] Anonymous. Stated meeting, April 21, 1865. Proceedings of the American Philosophical Society, 10(73):95–105, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982131.

Anonymous:1865:SMAa

[Ano65f] Anonymous. Stated meeting, April 7, 1865. Proceedings of the American Philosophical Society, 10(73):30–94, January 1865. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982130.

Anonymous:1865:SMAc

[Ano65g] Anonymous. Stated meeting, August 18, 1865. Proceedings of the American Philosophical Society, 10(74):148–149, July 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981395.

Anonymous:1865:SMDa

[Ano65h] Anonymous. Stated meeting, December 1, 1865. Proceedings of the American Philosophical Society, 10(74):187–192, July 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981402.

Anonymous:1865:SMDb

[Ano65i] Anonymous. Stated meeting, December 15, 1865. Proceedings of the American Philosophical Society, 10(74):192–194, July 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981403.

Anonymous:1865:SMFb

[Ano65j] Anonymous. Stated meeting, February 17, 1865. Proceedings of the American Philosophical Society, 10(73):8–23, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982127.

Anonymous:1865:SMFa

[Ano65k] Anonymous. Stated meeting, February 3, 1865. Proceedings of the American Philosophical Society, 10(73):7–8, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982126.

Anonymous:1865:SMJb

[Ano65l] Anonymous. Stated meeting, January 20, 1865. Proceedings of the American Philosophical Society, 10(73):3–7, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982125.

Anonymous:1865:SMJa

[Ano65m] Anonymous. Stated meeting, January 6, 1865. Proceedings of the American Philosophical Society, 10(73):1–3, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982124.

Anonymous:1865:SMJd

[Ano65n] Anonymous. Stated meeting, July 21, 1865. Proceedings of the American Philosophical Society, 10(74):147–148, July 1865. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981394.

Anonymous:1865:SMJc

[Ano65o] Anonymous. Stated meeting, June 16, 1865. Proceedings of the American Philosophical Society, 10(73):119–146, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982134.

Anonymous:1865:SMMb

[Ano65p] Anonymous. Stated meeting, March 17, 1865. Proceedings of the American Philosophical Society, 10(73):29, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982129.

Anonymous:1865:SMMa

[Ano65q] Anonymous. Stated meeting, March 3, 1865. Proceedings of the American Philosophical Society, 10(73):23–28, January 1865. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982128.

Anonymous:1865:SMMd

[Ano65r] Anonymous. Stated meeting, May 19, 1865. Proceedings of the American Philosophical Society, 10(73):107–118, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982133.

Anonymous:1865:SMMc

[Ano65s] Anonymous. Stated meeting, May 5, 1865. Proceedings of the American Philosophical Society, 10(73):105–106, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982132.

Anonymous:1865:SMNb

[Ano65t] Anonymous. Stated meeting, November 17, 1865. Proceedings of the American Philosophical Society, 10(74):169–171, July 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981400.

Anonymous:1865:SMNa

[Ano65u] Anonymous. Stated meeting, November 3, 1865. Proceedings of the American Philosophical Society, 10(74):167–169, July 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981399.

Anonymous:1865:SMOb

[Ano65v] Anonymous. Stated meeting, October 20, 1865. Proceedings of the American Philosophical Society, 10(74):167, July 1865. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981398.

Anonymous:1865:SMOa

[Ano65w] Anonymous. Stated meeting, October 6, 1865. Proceedings of the American Philosophical Society, 10(74):150–166, July 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981397.

Anonymous:1865:SMS

[Ano65x] Anonymous. Stated meeting, September 15, 1865. Proceedings of the American Philosophical Society, 10(74):149–150, July 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981396.

Anonymous:1865:VI

[Ano65y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 10(73):??, January 1865. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982122.

Anonymous:1866:Aa

[Ano66a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 10(75):??, January 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1866:Ab

[Ano66b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 10(76):??, June 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1866:SMAb

[Ano66c] Anonymous. Stated meeting, April 20, 1866. Proceedings of the American Philosophical Society, 10(75):226–242, January 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981567.

Anonymous:1866:SMAa

[Ano66d] Anonymous. Stated meeting, April 6, 1866. Proceedings of the American Philosophical Society, 10(75):208–226, January 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981566.

Anonymous:1866:SMAc

[Ano66e] Anonymous. Stated meeting, August 17, 1866. Proceedings of the American Philosophical Society, 10(76):259–260, June 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982212.

Anonymous:1866:SMDb

[Ano66f] Anonymous. Stated meeting, December 21, 1866. Proceedings of the American Philosophical Society, 10(76):291–296, June 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982220.

Anonymous:1866:SMDa

[Ano66g] Anonymous. Stated meeting, December 7, 1866. Proceedings of the American Philosophical Society, 10(76):278–288, 290, June 1866. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982218.

Anonymous:1866:SMFb

[Ano66h] Anonymous. Stated meeting, February 16, 1866. Proceedings of the American Philosophical Society, 10(75):202–205, January 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981563.

Anonymous:1866:SMFa

[Ano66i] Anonymous. Stated meeting, February 2, 1866. Proceedings of the American Philosophical Society, 10(75):201–202, January 1866. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981562.

Anonymous:1866:SMJb

[Ano66j] Anonymous. Stated meeting, January 19, 1866. Proceedings of the American Philosophical Society, 10(75):198–201, January 1866. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981561.

Anonymous:1866:SMJa

[Ano66k] Anonymous. Stated meeting, January 5, 1866. Proceedings of the American Philosophical Society, 10(75):195–198, January 1866. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981560.

Anonymous:1866:SMJd

[Ano66l] Anonymous. Stated meeting, July 20, 1866. Proceedings of the American Philosophical Society, 10(76):257–259, June 1866. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982211.

Anonymous:1866:SMJc

[Ano66m] Anonymous. Stated meeting, June 15, 1866. Proceedings of the American Philosophical Society, 10(76):255–257, June 1866. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982210.

Anonymous:1866:SMMb

[Ano66n] Anonymous. Stated meeting, March 16, 1866. Proceedings of the American Philosophical Society, 10(75):207–208, January 1866. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981565.

Anonymous:1866:SMMa

[Ano66o] Anonymous. Stated meeting, March 2, 1866. Proceedings of the American Philosophical Society, 10(75):205–207, January 1866. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981564.

Anonymous:1866:SMMd

[Ano66p] Anonymous. Stated meeting, May 18, 1866. Proceedings of the American Philosophical Society, 10(75):245–254, January 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981569.

Anonymous:1866:SMMc

[Ano66q] Anonymous. Stated meeting, May 4, 1866. Proceedings of the American Philosophical Society, 10(75):242–245, January 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981568.

Anonymous:1866:SMNb

[Ano66r] Anonymous. Stated meeting, November 16, 1866. Proceedings of the American Philosophical Society, 10(76):276–278, June 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982217.

Anonymous:1866:SMNa

[Ano66s] Anonymous. Stated meeting, November 2, 1866. Proceedings of the American Philosophical Society, 10(76):274–276, June 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982216.

Anonymous:1866:SMOb

[Ano66t] Anonymous. Stated meeting, October 19, 1866. Proceedings of the American Philosophical Society, 10(76):273, June 1866. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982215.

Anonymous:1866:SMOa

[Ano66u] Anonymous. Stated meeting, October 5, 1866. Proceedings of the American Philosophical Society, 10(76):271–272, June 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982214.

Anonymous:1866:SMS

[Ano66v] Anonymous. Stated meeting, September 21, 1866. Proceedings of the American Philosophical Society, 10(76):260–270, June 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982213.

Anonymous:1867:A

[Ano67a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 10(77):??, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1867:SMFc

[Ano67b] Anonymous. Special meeting, February 19, 1867. Proceedings of the American Philosophical Society, 10(77):326–327, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982225.

Anonymous:1867:SMAb

[Ano67c] Anonymous. Stated meeting, April 19, 1867. Proceedings of the American Philosophical Society, 10(77):335–337, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982229.

Anonymous:1867:SMAa

[Ano67d] Anonymous. Stated meeting, April 5, 1867. Proceedings of the American Philosophical Society, 10(77):332–335, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982228.

Anonymous:1867:SMAc

[Ano67e] Anonymous. Stated meeting, August 16th, 1867. Proceedings of the American Philosophical Society, 10(77):342, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982234.

Anonymous:1867:SMD

[Ano67f] Anonymous. Stated meeting, December 6, 1867. Proceedings of the American Philosophical Society, 10(77):359–360, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982240.

Anonymous:1867:SMFb

[Ano67g] Anonymous. Stated meeting, February 15, 1867. Proceedings of the American Philosophical Society, 10(77):320–326, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982224.

Anonymous:1867:SMFa

[Ano67h] Anonymous. Stated meeting, February 1st, 1867. Proceedings of the American Philosophical Society, 10(77):308–320, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982223.

Anonymous:1867:SMJb

[Ano67i] Anonymous. Stated meeting, January 18, 1867. Proceedings of the American Philosophical Society, 10(77):299–308, January 1867. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982222.

Anonymous:1867:SMJa

[Ano67j] Anonymous. Stated meeting, January 4, 1867. Proceedings of the American Philosophical Society, 10(77):297–298, January 1867. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982221.

Anonymous:1867:SMJd

[Ano67k] Anonymous. Stated meeting, July 19th, 1867. Proceedings of the American Philosophical Society, 10(77):341, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982233.

Anonymous:1867:SMJc

[Ano67l] Anonymous. Stated meeting, June 21, 1867. Proceedings of the American Philosophical Society, 10(77):340, January 1867. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982232.

Anonymous:1867:SMMa

[Ano67m] Anonymous. Stated meeting, March 1, 1867. Proceedings of the American Philosophical Society, 10(77):327–330, January 1867. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982226.

Anonymous:1867:SMMb

[Ano67n] Anonymous. Stated meeting, March 15, 1867. Proceedings of the American Philosophical Society, 10(77):330–332, January 1867. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982227.

Anonymous:1867:SMMd

[Ano67o] Anonymous. Stated meeting, May 17, 1867. Proceedings of the American Philosophical Society, 10(77):339, January 1867. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982231.

Anonymous:1867:SMMc

[Ano67p] Anonymous. Stated meeting, May 3, 1867. Proceedings of the American Philosophical Society, 10(77):337–338, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982230.

Anonymous:1867:SMNa

[Ano67q] Anonymous. Stated meeting, November 1, 1867. Proceedings of the American Philosophical Society, 10(77):354, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982238.

Anonymous:1867:SMNb

[Ano67r] Anonymous. Stated meeting, November 15, 1867. Proceedings of the American Philosophical Society, 10(77):355–359, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982239.

Anonymous:1867:SMOb

[Ano67s] Anonymous. Stated meeting, October 18, 1867. Proceedings of the American Philosophical Society, 10(77):351–353, January 1867. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982237.

Anonymous:1867:SMOa

[Ano67t] Anonymous. Stated meeting, October 4, 1867. Proceedings of the American Philosophical Society, 10(77):350–351, January 1867. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982236.

Anonymous:1867:SMS

[Ano67u] Anonymous. Stated meeting, September 20, 1867. Proceedings of the American Philosophical Society, 10(77):343–350, January 1867. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982235.

Anonymous:1868:Aa

[Ano68a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 10(78):??, January 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1868:Ab

[Ano68b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1868:I

[Ano68c] Anonymous. [Illustration]. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981807.

Anonymous:1868:IMI

[Ano68d] Anonymous. [Illustration]: Map to illustrate the conglomerate cliff country on the Upper Paint Lick Waters in East Ky. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981805.

Anonymous:1868:IPC

[Ano68e] Anonymous. [Illustration]: Pl. I. T. C. Porter. A. B. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981803.

Anonymous:1868:IPIa

[Ano68f] Anonymous. [Illustration]: Pl. IV. Lesley. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981806.

Anonymous:1868:IPXd

[Ano68g] Anonymous. [Illustration]: Pl. XIII. Indian characters on rocks in the Susquehannah River. Pa. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981815.

Anonymous:1868:IPXe

[Ano68h] Anonymous. [Illustration]: Pl. XIV. J. P. Lesley. Luxor. 1868. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981816.

Anonymous:1868:IPJ

[Ano68i] Anonymous. [Illustration]: Plate 2. J. P. Lesley. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981804.

Anonymous:1868:IPIb

[Ano68j] Anonymous. [Illustration]: Plate IX. Fig. 1. Fig. 2. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981811.

Anonymous:1868:IPVa

[Ano68k] Anonymous. [Illustration]: Plate VI. Mushroom rock on Alum Creek, Kansas. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981808.

Anonymous:1868:IPVb

[Ano68l] Anonymous. [Illustration]: Plate VII. Indian hieroglyphic rock on Smoky Hill River, Kansas. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981809.

Anonymous:1868:IPVc

[Ano68m] Anonymous. [Illustration]: Plate VIII. Indian hieroglyphic rock on Smoky Hill River, Kansas. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981810.

Anonymous:1868:IPXa

[Ano68n] Anonymous. [Illustration]: Plate X. Fig. 1. Fig. 2. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981812.

Anonymous:1868:IPXb

[Ano68o] Anonymous. [Illustration]: Plate XI. Fig. 1. Fig. 2. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981813.

Anonymous:1868:IPXc

[Ano68p] Anonymous. [Illustration]: Plate XII. Fig. 1. Fig. 2. Proceedings of the American Philosophical Society, 10(80):??, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981814.

Anonymous:1868:SMAb

[Ano68q] Anonymous. Stated meeting, April 17, 1868. Proceedings of the American Philosophical Society, 10(78):388–389, 421–424, January 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981591.

Anonymous:1868:SMAa

[Ano68r] Anonymous. Stated meeting, April 3, 1868. Proceedings of the American Philosophical Society, 10(78):386–388, January 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981590.

Anonymous:1868:SMAc

[Ano68s] Anonymous. Stated meeting, August 21st, 1868. Proceedings of the American Philosophical Society, 10(80):442, July 1868. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981787.

Anonymous:1868:SMDb

[Ano68t] Anonymous. Stated meeting, December 18, 1868. Proceedings of the American Philosophical Society, 10(80):540–541, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981800.

Anonymous:1868:SMDa

[Ano68u] Anonymous. Stated meeting, December 4, 1868. Proceedings of the American Philosophical Society, 10(80):512, 522, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981796.

Anonymous:1868:SMFb

[Ano68v] Anonymous. Stated meeting, February 21, 1867. Proceedings of the American Philosophical Society, 10(78):366–367, 380–381, January 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981586.

Anonymous:1868:SMFa

[Ano68w] Anonymous. Stated meeting, February 7, 1868. Proceedings of the American Philosophical Society, 10(78):365–366, January 1868. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981585.

Anonymous:1868:SMJb

[Ano68x] Anonymous. Stated meeting, January 17, 1868. Proceedings of the American Philosophical Society, 10(78):364–365, January 1868. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981584.

Anonymous:1868:SMJa

[Ano68y] Anonymous. Stated meeting, January 3, 1868. Proceedings of the American Philosophical Society, 10(78):361–364, January 1868. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981583.

Anonymous:1868:SMJd

[Ano68z] Anonymous. Stated meeting, July 17th, 1868. Proceedings of the American Philosophical Society, 10(80):441–442, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981786.

Anonymous:1868:SMJc

[Ano68-27] Anonymous. Stated meeting, June 19, 1868. Proceedings of the American Philosophical Society, 10(78):429–439, January 1868. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981595.

Anonymous:1868:SMMb

[Ano68-28] Anonymous. Stated meeting, March 20, 1868. Proceedings of the American Philosophical Society, 10(78):384–386, January 1868. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981589.

Anonymous:1868:SMMa

[Ano68-29] Anonymous. Stated meeting, March 6, 1868. Proceedings of the American Philosophical Society, 10(78):381–384, January 1868. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981588.

Anonymous:1868:SMMc

[Ano68-30] Anonymous. Stated meeting, May 1, 1868. Proceedings of the American Philosophical Society, 10(78):424–428, January 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981593.

Anonymous:1868:SMMd

[Ano68-31] Anonymous. Stated meeting, May 15, 1868. Proceedings of the American Philosophical Society, 10(78):428–429, January 1868. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981594.

Anonymous:1868:SMNb

[Ano68-32] Anonymous. Stated meeting, November 20th, 1868. Proceedings of the American Philosophical Society, 10(80):511, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981795.

Anonymous:1868:SMNa

[Ano68-33] Anonymous. Stated meeting, November 6th, 1868. Proceedings of the American Philosophical Society, 10(80):480, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981793.

Anonymous:1868:SMOb

[Ano68-34] Anonymous. Stated meeting, October 16, 1868. Proceedings of the American Philosophical Society, 10(80):479, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981792.

Anonymous:1868:SMOa

[Ano68-35] Anonymous. Stated meeting, October 2, 1868. Proceedings of the American Philosophical Society, 10(80):444, July 1868. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981789.

Anonymous:1868:SMS

[Ano68-36] Anonymous. Stated meeting, September 18th, 1868. Proceedings of the American Philosophical Society, 10(80):443, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981788.

Anonymous:1868:VI

[Ano68-37] Anonymous. Volume information. Proceedings of the American Philosophical Society, 10(80):583–586, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981785.

Anonymous:1869:A

[Ano69a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 11(81):??, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1869:E

[Ano69b] Anonymous. Errata. Proceedings of the American Philosophical Society, 11(81):??, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981529.

Anonymous:1869:FM

[Ano69c] Anonymous. Front matter. Proceedings of the American Philosophical Society, 11(81):611–614, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981432.

Anonymous:1869:IPCb

[Ano69d] Anonymous. [Illustration]: Pl. 10. Cachryx Defensor Cope. Proceedings of the American Philosophical Society, 11(81):??, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981527.

Anonymous:1869:IPL

[Ano69e] Anonymous. [Illustration]: Pl. 11. Laemanctus-alticoronatus Cope. Proceedings of the American Philosophical Society, 11(81):??, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981528.

Anonymous:1869:IPCa

[Ano69f] Anonymous. [Illustration]: Pl. 9. Claudius Angustatus Cope. Proceedings of the American Philosophical Society, 11(81):??, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981526.

Anonymous:1869:IPIa

[Ano69g] Anonymous. [Illustration]: Pl. II. Fig. 1. Section of 150 miles across the Appalachian Belt in Pensylvania. Fig. 2. Proceedings of the American Philosophical Society, 11(81):??, January 1869. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981522.

Anonymous:1869:IPIb

[Ano69h] Anonymous. [Illustration]: Plate III. 1. Galera perdicida. 2. Mixophagus spelaeus. 3. Stereodeetes tortus. 4. Tamias laevidens. 5. Sciurus panolius. 6. Tapiras haysii. Proceedings of the American Philosophical Society, 11(81):??, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981523.

Anonymous:1869:IPIc

[Ano69i] Anonymous. [Illustration]: Plate IV. Amblyrhiza inundata Cope. Proceedings of the American Philosophical Society, 11(81):??, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981524.

Anonymous:1869:IPV

[Ano69j] Anonymous. [Illustration]: Plate V. 1. Amblyrhiza inundata. 2–3 Loxomylus longidens. 4. Shell implement. 5. Anoplonassa forcipata. 6. Hemicaulodon effodiens. Proceedings of the American Philosophical Society, 11(81):??, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981525.

Anonymous:1869:ONF

[Ano69k] Anonymous. An obituary notice of Franklin Peale. Proceedings of the American Philosophical Society, 11(81):597–604, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981519.

Anonymous:1869:ONL

[Ano69l] Anonymous. Obituary notice of the Late James Copland, M.D. Proceedings of the American Philosophical Society, 11(81):525–528, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981508.

Anonymous:1869:SRC

[Ano69m] Anonymous. On some Reptilia of the Cretaceous formation of the United States. Proceedings of the American Philosophical Society, 11(81):271–274, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981471.

Anonymous:1869:SMDd

[Ano69n] Anonymous. Special meeting, December 9th, 1870. Proceedings of the American Philosophical Society, 11(81):572–573, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981515.

Anonymous:1869:SMAe

[Ano69o] Anonymous. Stated meeting, April 15, 1870. Proceedings of the American Philosophical Society, 11(81):367–368, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981487.

Anonymous:1869:SMAb

[Ano69p] Anonymous. Stated meeting, April 16, 1869. Proceedings of the American Philosophical Society, 11(81):110–112, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981447.

Anonymous:1869:SMAd

[Ano69q] Anonymous. Stated meeting, April 1st, 1870. Proceedings of the American Philosophical Society, 11(81):316, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981485.

Anonymous:1869:SMAa

[Ano69r] Anonymous. Stated meeting, April 2, 1869. Proceedings of the American Philosophical Society, 11(81):90–93, January 1869. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981443.

Anonymous:1869:SMAf

[Ano69s] Anonymous. Stated meeting, August 19, 1870. Proceedings of the American Philosophical Society, 11(81):496–498, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981500.

Anonymous:1869:SMAc

[Ano69t] Anonymous. Stated meeting, August 20, 1869. Proceedings of the American Philosophical Society, 11(81):192–193, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981454.

Anonymous:1869:SMDb

[Ano69u] Anonymous. Stated meeting, Dec. 17th, 1869. Proceedings of the American Philosophical Society, 11(81):234–239, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981467.

Anonymous:1869:SMDa

[Ano69v] Anonymous. Stated meeting, Dec. 3d, 1869. Proceedings of the American Philosophical Society, 11(81):233–234, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981466.

Anonymous:1869:SMDe

[Ano69w] Anonymous. Stated meeting, December 16th, 1870. Proceedings of the American Philosophical Society, 11(81):573–574, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981516.

Anonymous:1869:SMDc

[Ano69x] Anonymous. Stated meeting, December 2d, 1870. Proceedings of the American Philosophical Society, 11(81):571–572, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981514.

Anonymous:1869:SMFd

[Ano69y] Anonymous. Stated meeting, February 18th, 1870. Proceedings of the American Philosophical Society, 11(81):281–282, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981475.

Anonymous:1869:SMFb

[Ano69z] Anonymous. Stated meeting, February 19, 1869. Proceedings of the American Philosophical Society, 11(81):14–16, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981436.

Anonymous:1869:SMFc

[Ano69-27] Anonymous. Stated meeting, February 4th, 1870. Proceedings of the American Philosophical Society, 11(81):279–281, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981474.

Anonymous:1869:SMFa

[Ano69-28] Anonymous. Stated meeting, February 5, 1869. Proceedings of the American Philosophical Society, 11(81):8–13, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981435.

Anonymous:1869:SMJa

[Ano69-29] Anonymous. Stated meeting, January 1, 1869. Proceedings of the American Philosophical Society, 11(81):1–2, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981433.

Anonymous:1869:SMJb

[Ano69-30] Anonymous. Stated meeting, January 15, 1869. Proceedings of the American Philosophical Society, 11(81):3–7, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981434.

Anonymous:1869:SMJf

[Ano69-31] Anonymous. Stated meeting, January 21, 1870. Proceedings of the American Philosophical Society, 11(81):277–278, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981473.

Anonymous:1869:SMJe

[Ano69-32] Anonymous. Stated meeting, January 7th, 1870. Proceedings of the American Philosophical Society, 11(81):274–277, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981472.

Anonymous:1869:SMJh

[Ano69-33] Anonymous. Stated meeting, July 15, 1870. Proceedings of the American Philosophical Society, 11(81):446–447, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981498.

Anonymous:1869:SMJd

[Ano69-34] Anonymous. Stated meeting, July 16, 1869. Proceedings of the American Philosophical Society, 11(81):145–146, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981452.

Anonymous:1869:SMJg

[Ano69-35] Anonymous. Stated meeting, June 17, 1870. Proceedings of the American Philosophical Society, 11(81):442–446, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981497.

Anonymous:1869:SMJc

[Ano69-36] Anonymous. Stated meeting, June 18, 1869. Proceedings of the American Philosophical Society, 11(81):116–118, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981450.

Anonymous:1869:SMMf

[Ano69-37] Anonymous. Stated meeting, March 18, 1870. Proceedings of the American Philosophical Society, 11(81):310–313, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981482.

Anonymous:1869:SMMb

[Ano69-38] Anonymous. Stated meeting, March 19, 1869. Proceedings of the American Philosophical Society, 11(81):83–84, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981441.

Anonymous:1869:SMMe

[Ano69-39] Anonymous. Stated meeting, March 4, 1870. Proceedings of the American Philosophical Society, 11(81):282–284, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981476.

Anonymous:1869:SMMa

[Ano69-40] Anonymous. Stated meeting, March 5, 1869. Proceedings of the American Philosophical Society, 11(81):23–24, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981438.

Anonymous:1869:SMMh

[Ano69-41] Anonymous. Stated meeting, May 20, 1870. Proceedings of the American Philosophical Society, 11(81):385, 438–441, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981491.

Anonymous:1869:SMMd

[Ano69-42] Anonymous. Stated meeting, May 21, 1869. Proceedings of the American Philosophical Society, 11(81):114–115, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981449.

Anonymous:1869:SMMg

[Ano69-43] Anonymous. Stated meeting, May 6, 1870. Proceedings of the American Philosophical Society, 11(81):368–370, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981488.

Anonymous:1869:SMMc

[Ano69-44] Anonymous. Stated meeting, May 7, 1869. Proceedings of the American Philosophical Society, 11(81):112–114, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981448.

Anonymous:1869:SMNb

[Ano69-45] Anonymous. Stated meeting, Nov. 19th, 1869. Proceedings of the American Philosophical Society, 11(81):213–214, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981462.

Anonymous:1869:SMNd

[Ano69-46] Anonymous. Stated meeting, Nov. 21st, 1870. Proceedings of the American Philosophical Society, 11(81):523–524, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981507.

Anonymous:1869:SMNc

[Ano69-47] Anonymous. Stated meeting, Nov. 4th, 1870. Proceedings of the American Philosophical Society, 11(81):521–523, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981506.

Anonymous:1869:SMNa

[Ano69-48] Anonymous. Stated meeting, Nov. 5, 1869. Proceedings of the American Philosophical Society, 11(81):211–212, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981461.

Anonymous:1869:SMOb

[Ano69-49] Anonymous. Stated meeting, Oct. 15, 1869. Proceedings of the American Philosophical Society, 11(81):209–211, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981460.

Anonymous:1869:SMOa

[Ano69-50] Anonymous. Stated meeting, October 1, 1869. Proceedings of the American Philosophical Society, 11(81):201–202, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981457.

Anonymous:1869:SMOd

[Ano69-51] Anonymous. Stated meeting, October 21, 1870. Proceedings of the American Philosophical Society, 11(81):516–521, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981505.

Anonymous:1869:SMOc

[Ano69-52] Anonymous. Stated meeting, October 7th, 1870. Proceedings of the American Philosophical Society, 11(81):514–515, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981504.

Anonymous:1869:SMSa

[Ano69-53] Anonymous. Stated meeting, Sept. 17, 1869. Proceedings of the American Philosophical Society, 11(81):193–194, January 1869. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981455.

Anonymous:1869:SMSb

[Ano69-54] Anonymous. Stated meeting, September 16, 1870. Proceedings of the American Philosophical Society, 11(81):513, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981503.

Anonymous:1869:VCD

[Ano69-55] Anonymous. Verbal communication by E. D. Cope, at the meeting of the A. Phil. Soc., Feb. 18, 1870. Proceedings of the American Philosophical Society, 11(81):284, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981477.

Anonymous:1871:A

[Ano71a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 12(86):??, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1871:FM

[Ano71b] Anonymous. Front matter. Proceedings of the American Philosophical Society, 12(86):587–592, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981653.

Anonymous:1871:HBN

[Ano71c] Anonymous. Heller and Brightley’s new transit. Proceedings of the American Philosophical Society, 12(86):115–120, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981680.

Anonymous:1871:I

[Ano71d] Anonymous. [Illustration]. Proceedings of the American Philosophical Society, 12(86):??, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981773.

Anonymous:1871:IMP

[Ano71e] Anonymous. [Illustration]: Map of the Pittsburg & Baltimore Coal Coke & Iron Co’s Lands Ursina. Somerset Co Penn’a. Proceedings of the American Philosophical Society, 12(86):??, 1871. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981772.

Anonymous:1871:RHC

[Ano71f] Anonymous. “Remarks” on Hyrtl’s collection. Proceedings of the American Philosophical Society, 12(86):191–192, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981697.

Anonymous:1871:SMAe

[Ano71g] Anonymous. Stated meeting, April 19th, 1872. Proceedings of the American Philosophical Society, 12(86):438–440, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981735.

Anonymous:1871:SMAb

[Ano71h] Anonymous. Stated meeting, April 21st, 1871. Proceedings of the American Philosophical Society, 12(86):110–112, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981677.

Anonymous:1871:SMAd

[Ano71i] Anonymous. Stated meeting, April 5, 1872. Proceedings of the American Philosophical Society, 12(86):437–438, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981734.

Anonymous:1871:SMAa

[Ano71j] Anonymous. Stated meeting, April 7th, 1871. Proceedings of the American Philosophical Society, 12(86):109, 1871. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981676.

Anonymous:1871:SMAc

[Ano71k] Anonymous. Stated meeting, Aug. 18th, 1871. Proceedings of the American Philosophical Society, 12(86):161, 1871. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981688.

Anonymous:1871:SMAf

[Ano71l] Anonymous. Stated meeting, August 15, 1872. Proceedings of the American Philosophical Society, 12(86):458–459, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981741.

Anonymous:1871:SMDa

[Ano71m] Anonymous. Stated meeting, December 1, 1871. Proceedings of the American Philosophical Society, 12(86):216, 1871. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981702.

Anonymous:1871:SMDb

[Ano71n] Anonymous. Stated meeting, December 15, 1871. Proceedings of the American Philosophical Society, 12(86):228–229, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981705.

Anonymous:1871:SMDd

[Ano71o] Anonymous. Stated meeting, December 20th, 1872. Proceedings of the American Philosophical Society, 12(86):576–577, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981767.

Anonymous:1871:SMDc

[Ano71p] Anonymous. Stated meeting, December 6th, 1872. Proceedings of the American Philosophical Society, 12(86):573–575, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981766.

Anonymous:1871:SMFd

[Ano71q] Anonymous. Stated meeting, February 16, 1872. Proceedings of the American Philosophical Society, 12(86):357–359, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981716.

Anonymous:1871:SMFb

[Ano71r] Anonymous. Stated meeting, February 17, 1871. Proceedings of the American Philosophical Society, 12(86):16–17, 1871. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981660.

Anonymous:1871:SMFc

[Ano71s] Anonymous. Stated meeting, February 2d, 1872. Proceedings of the American Philosophical Society, 12(86):321–323, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981713.

Anonymous:1871:SMFa

[Ano71t] Anonymous. Stated meeting, February 3, 1871. Proceedings of the American Philosophical Society, 12(86):14–16, 1871. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981659.

Anonymous:1871:SMJf

[Ano71u] Anonymous. Stated meeting, January 19, 1872. Proceedings of the American Philosophical Society, 12(86):307–308, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981710.

Anonymous:1871:SMJb

[Ano71v] Anonymous. Stated meeting, January 20, 1871. Proceedings of the American Philosophical Society, 12(86):8, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981656.

Anonymous:1871:SMJe

[Ano71w] Anonymous. Stated meeting, January 5th, 1872. Proceedings of the American Philosophical Society, 12(86):287–288, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981708.

Anonymous:1871:SMJa

[Ano71x] Anonymous. Stated meeting, January 6. Proceedings of the American Philosophical Society, 12(86):1–3, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981654.

Anonymous:1871:SMJh

[Ano71y] Anonymous. Stated meeting, July 19, 1872. Proceedings of the American Philosophical Society, 12(86):457–458, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981740.

Anonymous:1871:SMJd

[Ano71z] Anonymous. Stated meeting, July 21, 1871. Proceedings of the American Philosophical Society, 12(86):158–159, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981686.

Anonymous:1871:SMJc

[Ano71-27] Anonymous. Stated meeting, June 16, 1871. Proceedings of the American Philosophical Society, 12(86):113–114, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981679.

Anonymous:1871:SMJg

[Ano71-28] Anonymous. Stated meeting, June 21, 1872. Proceedings of the American Philosophical Society, 12(86):442–444, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981738.

Anonymous:1871:SMMe

[Ano71-29] Anonymous. Stated meeting, March 1, 1872. Proceedings of the American Philosophical Society, 12(86):359–360, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981717.

Anonymous:1871:SMMf

[Ano71-30] Anonymous. Stated meeting, March 15, 1872. Proceedings of the American Philosophical Society, 12(86):434–435, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981732.

Anonymous:1871:SMMb

[Ano71-31] Anonymous. Stated meeting, March 17, 1871. Proceedings of the American Philosophical Society, 12(86):63–64, 1871. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981670.

Anonymous:1871:SMMa

[Ano71-32] Anonymous. Stated meeting, March 3, 1871. Proceedings of the American Philosophical Society, 12(86):37–38, 1871. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981663.

Anonymous:1871:SMMh

[Ano71-33] Anonymous. Stated meeting, May 17th, 1872. Proceedings of the American Philosophical Society, 12(86):441–442, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981737.

Anonymous:1871:SMMd

[Ano71-34] Anonymous. Stated meeting, May 19, 1871. Proceedings of the American Philosophical Society, 12(86):159–160, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981687.

Anonymous:1871:SMMg

[Ano71-35] Anonymous. Stated meeting, May 3d, 1872. Proceedings of the American Philosophical Society, 12(86):440–441, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981736.

Anonymous:1871:SMMc

[Ano71-36] Anonymous. Stated meeting, May 5, 1871. Proceedings of the American Philosophical Society, 12(86):112–113, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981678.

Anonymous:1871:SMNa

[Ano71-37] Anonymous. Stated meeting, Nov. 3d, 1871. Proceedings of the American Philosophical Society, 12(86):176–177, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981695.

Anonymous:1871:SMNc

[Ano71-38] Anonymous. Stated meeting, November 1, 1872. Proceedings of the American Philosophical Society, 12(86):559–561, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981763.

Anonymous:1871:SMNd

[Ano71-39] Anonymous. Stated meeting, November 15th, 1872. Proceedings of the American Philosophical Society, 12(86):570–573, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981765.

Anonymous:1871:SMNb

[Ano71-40] Anonymous. Stated meeting, November 17, 1871. Proceedings of the American Philosophical Society, 12(86):192–194, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981698.

Anonymous:1871:SMOc

[Ano71-41] Anonymous. Stated meeting, Oct. 18th, 1872. Proceedings of the American Philosophical Society, 12(86):552–554, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981758.

Anonymous:1871:SMOb

[Ano71-42] Anonymous. Stated meeting, Oct. 4th, 1872. Proceedings of the American Philosophical Society, 12(86):545–546, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981756.

Anonymous:1871:SMOa

[Ano71-43] Anonymous. Stated meeting, Oct. 6th., 1871. Proceedings of the American Philosophical Society, 12(86):167–168, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981691.

Anonymous:1871:SMSa

[Ano71-44] Anonymous. Stated meeting Sept. 15th, 1871. Proceedings of the American Philosophical Society, 12(86):161–162, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981689.

Anonymous:1871:SMSb

[Ano71-45] Anonymous. Stated meeting, Sept. 20, 1872. Proceedings of the American Philosophical Society, 12(86):514–516, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981751.

Anonymous:1873:A

[Ano73a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 13(90):??, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1873:BM

[Ano73b] Anonymous. Back matter. Proceedings of the American Philosophical Society, 13(90):509–512, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981652.

Anonymous:1873:DLM

[Ano73c] Anonymous. Description of Lesley’s micrometer for field-note plotting. Proceedings of the American Philosophical Society, 13(90):233– 235, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981628.

Anonymous:1873:FM

[Ano73d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 13(90):??, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981596.

Anonymous:1873:SMAb

[Ano73e] Anonymous. Stated meeting, April 18th, 1873. Proceedings of the American Philosophical Society, 13(90):209–212, 1873. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981626.

Anonymous:1873:SMAa

[Ano73f] Anonymous. Stated meeting, April 4th, 1873. Proceedings of the American Philosophical Society, 13(90):181–182, 1873. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981621.

Anonymous:1873:SMDb

[Ano73g] Anonymous. Stated meeting, December 19th, 1873. Proceedings of the American Philosophical Society, 13(90):505–508, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981651.

Anonymous:1873:SMDa

[Ano73h] Anonymous. Stated meeting, December 5th, 1873. Proceedings of the American Philosophical Society, 13(90):269, 1873. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981640.

Anonymous:1873:SMFb

[Ano73i] Anonymous. Stated meeting, February 21st, 1873. Proceedings of the American Philosophical Society, 13(90):36–37, 1873. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981603.

Anonymous:1873:SMFa

[Ano73j] Anonymous. Stated meeting, February 7th, 1873. Proceedings of the American Philosophical Society, 13(90):33–34, 1873. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981601.

Anonymous:1873:SMJb

[Ano73k] Anonymous. Stated meeting, January 17th, 1873. Proceedings of the American Philosophical Society, 13(90):22–23, 1873. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981599.

Anonymous:1873:SMJa

[Ano73l] Anonymous. Stated meeting, January 3d, 1873. Proceedings of the American Philosophical Society, 13(90):1–2, 1873. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981597.

Anonymous:1873:SMJd

[Ano73m] Anonymous. Stated meeting, July 18th, 1873. Proceedings of the American Philosophical Society, 13(90):258–259, 1873. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981634.

Anonymous:1873:SMJc

[Ano73n] Anonymous. Stated meeting, June 20th, 1873. Proceedings of the American Philosophical Society, 13(90):257–258, 1873. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981633.

Anonymous:1873:SMMb

[Ano73o] Anonymous. Stated meeting, March 21st, 1873. Proceedings of the American Philosophical Society, 13(90):168–169, 1873. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981618.

Anonymous:1873:SMMa

[Ano73p] Anonymous. Stated meeting, March 7th, 1873. Proceedings of the American Philosophical Society, 13(90):154–156, 1873. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981616.

Anonymous:1873:SMMd

[Ano73q] Anonymous. Stated meeting, May 16th, 1873. Proceedings of the American Philosophical Society, 13(90):253–254, 1873. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981631.

Anonymous:1873:SMMc

[Ano73r] Anonymous. Stated meeting, May 2d, 1873. Proceedings of the American Philosophical Society, 13(90):235–236, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981629.

Anonymous:1873:SMNb

[Ano73s] Anonymous. Stated meeting, November 21st, 1873. Proceedings of the American Philosophical Society, 13(90):267–268, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981639.

Anonymous:1873:SMNa

[Ano73t] Anonymous. Stated meeting, November 7th, 1873. Proceedings of the American Philosophical Society, 13(90):266–267, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981638.

Anonymous:1873:SMOb

[Ano73u] Anonymous. Stated meeting, October 17th, 1873. Proceedings of the American Philosophical Society, 13(90):263–265, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981637.

Anonymous:1873:SMOa

[Ano73v] Anonymous. Stated meeting, October 2d, 1873. Proceedings of the American Philosophical Society, 13(90):262–263, 1873. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981636.

Anonymous:1873:SMS

[Ano73w] Anonymous. Stated meeting, September 19th, 1873. Proceedings of the American Philosophical Society, 13(90):260–262, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981635.

Anonymous:1874:ARP

[Ano74a] Anonymous. Abstract of the remarks of Prof. Cope at the meeting of the American Philosophical Society, January 16, 1874. Proceedings of the American Philosophical Society, 14(92):110, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982066.

Anonymous:1874:Aa

[Ano74b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 14(92):??, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1874:Ab

[Ano74c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 14(93):??, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1874:SMAb

[Ano74d] Anonymous. Stated meeting, April 17th, 1874. Proceedings of the American Philosophical Society, 14(92):16–18, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982062.

Anonymous:1874:SMAa

[Ano74e] Anonymous. Stated meeting April, 3d, 1874. Proceedings of the American Philosophical Society, 14(92):15–16, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982061.

Anonymous:1874:SMAc

[Ano74f] Anonymous. Stated meeting, August 21st, 1874. Proceedings of the American Philosophical Society, 14(93):180–181, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982102.

Anonymous:1874:SMDb

[Ano74g] Anonymous. Stated meeting, December 18th, 1874. Proceedings of the American Philosophical Society, 14(93):298–300, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982121.

Anonymous:1874:SMDa

[Ano74h] Anonymous. Stated meeting, December 4th, 1874. Proceedings of the American Philosophical Society, 14(93):190–192, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982108.

Anonymous:1874:SMFb

[Ano74i] Anonymous. Stated meeting, February 20th, 1874. Proceedings of the American Philosophical Society, 14(92):9–11, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982058.

Anonymous:1874:SMFa

[Ano74j] Anonymous. Stated meeting, February 6th, 1874. Proceedings of the American Philosophical Society, 14(92):8–9, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982057.

Anonymous:1874:SMJb

[Ano74k] Anonymous. Stated meeting, January 16th, 1874. Proceedings of the American Philosophical Society, 14(92):5–8, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982056.

Anonymous:1874:SMJa

[Ano74l] Anonymous. Stated meeting, January 2d, 1874. Proceedings of the American Philosophical Society, 14(92):1–5, January/June 1874. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982055.

Anonymous:1874:SMJd

[Ano74m] Anonymous. Stated meeting, July 17th, 1874. Proceedings of the American Philosophical Society, 14(93):179–180, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982101.

Anonymous:1874:SMJc

[Ano74n] Anonymous. Stated meeting, June 19th, 1874. Proceedings of the American Philosophical Society, 14(93):177–179, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982100.

Anonymous:1874:SMMb

[Ano74o] Anonymous. Stated meeting, March 20th, 1874. Proceedings of the American Philosophical Society, 14(92):12–14, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982060.

Anonymous:1874:SMMa

[Ano74p] Anonymous. Stated meeting, March 6th, 1874. Proceedings of the American Philosophical Society, 14(92):11–12, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982059.

Anonymous:1874:SMMd

[Ano74q] Anonymous. Stated meeting, May 15th, 1874. Proceedings of the American Philosophical Society, 14(92):176, January/June 1874. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982080.

Anonymous:1874:SMMe

[Ano74r] Anonymous. [Stated meeting, May 15th, 1874: Conclusion]. Proceedings of the American Philosophical Society, 14(93):177, June/ December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982099.

Anonymous:1874:SMMc

[Ano74s] Anonymous. Stated meeting, May 1st, 1874. Proceedings of the American Philosophical Society, 14(92):174–176, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982079.

Anonymous:1874:SMNb

[Ano74t] Anonymous. Stated meeting, November 20th, 1874. Proceedings of the American Philosophical Society, 14(93):187–190, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982107.

Anonymous:1874:SMNa

[Ano74u] Anonymous. Stated meeting, November 6th, 1874. Proceedings of the American Philosophical Society, 14(93):186–187, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982106.

Anonymous:1874:SMOb

[Ano74v] Anonymous. Stated meeting, October 16th, 1874. Proceedings of the American Philosophical Society, 14(93):184–186, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982105.

Anonymous:1874:SMOa

[Ano74w] Anonymous. Stated meeting, October 2d, 1874. Proceedings of the American Philosophical Society, 14(93):183–184, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982104.

Anonymous:1874:SMS

[Ano74x] Anonymous. Stated meeting, September 18th, 1874. Proceedings of the American Philosophical Society, 14(93):181–183, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982103.

Anonymous:1874:VI

[Ano74y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 14(92):??, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982054.

Anonymous:1875:Aa

[Ano75a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 14(94):??, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1875:Ab

[Ano75b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 14(95):??, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1875:ELN

[Ano75c] Anonymous. Errata: List of the North American Platypterices, Attaci, Hemileucini, Ceratocampadae, Lachneides, Teredines and Hepiali, with Notes. Proceedings of the American Philosophical Society, 14(94):??, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981884.

Anonymous:1875:SMAb

[Ano75d] Anonymous. Stated meeting, April 16, 1875. Proceedings of the American Philosophical Society, 14(94):435–437, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981880.

Anonymous:1875:SMAa

[Ano75e] Anonymous. Stated meeting, April 2, 1875. Proceedings of the American Philosophical Society, 14(94):434, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981879.

Anonymous:1875:SMAc

[Ano75f] Anonymous. Stated meeting, August 20th, 1875. Proceedings of the American Philosophical Society, 14(95):639–640, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981949.

Anonymous:1875:SMDb

[Ano75g] Anonymous. Stated meeting, December 17th, 1875. Proceedings of the American Philosophical Society, 14(95):662–666, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981958.

Anonymous:1875:SMDa

[Ano75h] Anonymous. Stated meeting, December 3d, 1875. Proceedings of the American Philosophical Society, 14(95):659–662, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981957.

Anonymous:1875:SMFb

[Ano75i] Anonymous. Stated meeting, February 19, 1875. Proceedings of the American Philosophical Society, 14(94):426–428, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981875.

Anonymous:1875:SMFa

[Ano75j] Anonymous. Stated meeting, February 5, 1875. Proceedings of the American Philosophical Society, 14(94):423–426, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981874.

Anonymous:1875:SMJa

[Ano75k] Anonymous. Stated meeting, January 1, 1875. Proceedings of the American Philosophical Society, 14(94):418–420, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981872.

Anonymous:1875:SMJb

[Ano75l] Anonymous. Stated meeting, January 15, 1875. Proceedings of the American Philosophical Society, 14(94):421–422, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981873.

Anonymous:1875:SMJd

[Ano75m] Anonymous. Stated meeting, July 16th, 1875. Proceedings of the American Philosophical Society, 14(95):637–639, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981948.

Anonymous:1875:SMJc

[Ano75n] Anonymous. Stated meeting, June 18th, 1875. Proceedings of the American Philosophical Society, 14(95):635–637, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981947.

Anonymous:1875:SMMb

[Ano75o] Anonymous. Stated meeting, March 19, 1875. Proceedings of the American Philosophical Society, 14(94):432–433, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981877.

Anonymous:1875:SMMa

[Ano75p] Anonymous. Stated meeting, March 5, 1875. Proceedings of the American Philosophical Society, 14(94):428–432, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981876.

Anonymous:1875:SMMd

[Ano75q] Anonymous. Stated meeting, May 21, 1875. Proceedings of the American Philosophical Society, 14(94):439–440, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981882.

Anonymous:1875:SMMc

[Ano75r] Anonymous. Stated meeting, May 7, 1875. Proceedings of the American Philosophical Society, 14(94):437–439, January/June 1875. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981881.

Anonymous:1875:SMNb

[Ano75s] Anonymous. Stated meeting, November 19th, 1875. Proceedings of the American Philosophical Society, 14(95):648–649, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981954.

Anonymous:1875:SMNa

[Ano75t] Anonymous. Stated meeting, November 5th, 1875. Proceedings of the American Philosophical Society, 14(95):646–648, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981953.

Anonymous:1875:SMOb

[Ano75u] Anonymous. Stated meeting, October 15th, 1875. Proceedings of the American Philosophical Society, 14(95):643–645, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981952.

Anonymous:1875:SMOa

[Ano75v] Anonymous. Stated meeting, October 1st, 1875. Proceedings of the American Philosophical Society, 14(95):642–643, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981951.

Anonymous:1875:SMS

[Ano75w] Anonymous. Stated meeting, September 17th, 1875. Proceedings of the American Philosophical Society, 14(95):640–642, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981950.

Anonymous:1875:VI

[Ano75x] Anonymous. Volume information. Proceedings of the American Philosophical Society, 14(95):675–684, June/December 1875. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981934.

Anonymous:1876:Aa

[Ano76a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 15(96):??, December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1876:Ab

[Ano76b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 16(97):??, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1876:Ac

[Ano76c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 16(98):??, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1876:BM

[Ano76d] Anonymous. Back matter. Proceedings of the American Philosophical Society, 15(96):443–455, December 1876. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982413.

Anonymous:1876:FM

[Ano76e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 15(96):i–v, December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982411.

Anonymous:1876:OWR

[Ano76f] Anonymous. Oil well records, selected from the collections of Mr. J. F. Carll, Assistant Geologist in charge of the Geological Survey of the Oil District of Pennsylvania. Proceedings of the American Philosophical Society, 16(98):346–380, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982823.

Anonymous:1876:SMAb

[Ano76g] Anonymous. Stated meeting, April 21, 1876. Proceedings of the American Philosophical Society, 16(97):232–234, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982435.

Anonymous:1876:SMAa

[Ano76h] Anonymous. Stated meeting, April 7, 1876. Proceedings of the American Philosophical Society, 16(97):229–231, January/June 1876. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982434.

Anonymous:1876:SMAc

[Ano76i] Anonymous. Stated meeting, August 18, 1876. Proceedings of the American Philosophical Society, 16(98):282–283, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982809.

Anonymous:1876:SMDa

[Ano76j] Anonymous. Stated meeting, Dec. 1st, 1876. Proceedings of the American Philosophical Society, 16(98):335–336, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982820.

Anonymous:1876:SMDb

[Ano76k] Anonymous. Stated meeting, December 15th, 1876. Proceedings of the American Philosophical Society, 16(98):336–339, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982821.

Anonymous:1876:SMFb

[Ano76l] Anonymous. Stated meeting, February 18th, 1876. Proceedings of the American Philosophical Society, 16(97):197–198, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982427.

Anonymous:1876:SMFa

[Ano76m] Anonymous. Stated meeting, February 4th, 1876. Proceedings of the American Philosophical Society, 16(97):196–197, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982426.

Anonymous:1876:SMJb

[Ano76n] Anonymous. Stated meeting, January 21, 1876. Proceedings of the American Philosophical Society, 16(97):179–183, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982423.

Anonymous:1876:SMJa

[Ano76o] Anonymous. Stated meeting, January 7th, 1876. Proceedings of the American Philosophical Society, 16(97):52–54, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982420.

Anonymous:1876:SMJd

[Ano76p] Anonymous. Stated meeting, July 21, 1876. Proceedings of the American Philosophical Society, 16(98):280–281, June/December 1876. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982808.

Anonymous:1876:SMJc

[Ano76q] Anonymous. Stated meeting, June 16, 1876. Proceedings of the American Philosophical Society, 16(98):277–279, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982807.

Anonymous:1876:SMMb

[Ano76r] Anonymous. Stated meeting, March 17, 1876. Proceedings of the American Philosophical Society, 16(97):217–218, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982432.

Anonymous:1876:SMMa

[Ano76s] Anonymous. Stated meeting, March 3, 1876. Proceedings of the American Philosophical Society, 16(97):212–213, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982430.

Anonymous:1876:SMMd

[Ano76t] Anonymous. Stated meeting, May 19, 1876. Proceedings of the American Philosophical Society, 16(97):236–237, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982437.

Anonymous:1876:SMMc

[Ano76u] Anonymous. Stated meeting, May 5, 1876. Proceedings of the American Philosophical Society, 16(97):234–235, January/June 1876. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982436.

Anonymous:1876:SMNb

[Ano76v] Anonymous. Stated meeting, November 17th, 1876. Proceedings of the American Philosophical Society, 16(98):290–292, 333–335, June/ December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982814.

Anonymous:1876:SMNa

[Ano76w] Anonymous. Stated meeting, November 3rd, 1876. Proceedings of the American Philosophical Society, 16(98):289–290, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982813.

Anonymous:1876:SMOb

[Ano76x] Anonymous. Stated meeting, October 20th, 1876. Proceedings of the American Philosophical Society, 16(98):287–289, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982812.

Anonymous:1876:SMOa

[Ano76y] Anonymous. Stated meeting, October 7th, 1876. Proceedings of the American Philosophical Society, 16(98):285–286, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982811.

Anonymous:1876:SMS

[Ano76z] Anonymous. Stated meeting, September 15, 1876. Proceedings of the American Philosophical Society, 16(98):283–285, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982810.

Anonymous:1876:VI

[Ano76-27] Anonymous. Volume information. Proceedings of the American Philosophical Society, 16(97):??, January/June 1876. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982418.

Anonymous:1877:Aa

[Ano77a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 16(99):??, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1877:Ab

[Ano77b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 17(100):??, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1877:EPP

[Ano77c] Anonymous. Errata for Proceedings A. P. S. Vol. XVI. No. 99: On the Relative Ages of the Sun and Certain of the Fixed Stars. Proceedings of the American Philosophical Society, 17(100):280, June/ December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982312.

See [Kir77b].

Anonymous:1877:SMAb

[Ano77d] Anonymous. Stated meeting, April 20th, 1877. Proceedings of the American Philosophical Society, 16(99):663–665, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982495.

Anonymous:1877:SMAa

[Ano77e] Anonymous. Stated meeting, April 6, 1877. Proceedings of the American Philosophical Society, 16(99):661–663, January/May 1877. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982494.

Anonymous:1877:SMAc

[Ano77f] Anonymous. Stated meeting, August 17th, 1877. Proceedings of the American Philosophical Society, 17(100):8–10, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982270.

Anonymous:1877:SMDb

[Ano77g] Anonymous. Stated meeting, December 21, 1877. Proceedings of the American Philosophical Society, 17(100):277–280, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982311.

Anonymous:1877:SMDa

[Ano77h] Anonymous. Stated meeting, December 7, 1877. Proceedings of the American Philosophical Society, 17(100):274–277, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982310.

Anonymous:1877:SMFb

[Ano77i] Anonymous. Stated meeting February 16th, 1877. Proceedings of the American Philosophical Society, 16(99):389–392, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982470.

Anonymous:1877:SMFa

[Ano77j] Anonymous. Stated meeting, February 2nd, 1877. Proceedings of the American Philosophical Society, 16(99):385–389, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982469.

Anonymous:1877:SMJb

[Ano77k] Anonymous. Stated meeting, January 19th, 1877. Proceedings of the American Philosophical Society, 16(99):383–385, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982468.

Anonymous:1877:SMJa

[Ano77l] Anonymous. Stated meeting, January 5th, 1877. Proceedings of the American Philosophical Society, 16(99):381–382, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982467.

Anonymous:1877:SMJd

[Ano77m] Anonymous. Stated meeting, July 20th, 1877. Proceedings of the American Philosophical Society, 17(100):4–8, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982269.

Anonymous:1877:SMJc

[Ano77n] Anonymous. Stated meeting, June 15th, 1877. Proceedings of the American Philosophical Society, 17(100):2–4, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982268.

Anonymous:1877:SMMb

[Ano77o] Anonymous. Stated meeting, March 16th, 1877. Proceedings of the American Philosophical Society, 16(99):394–396, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982472.

Anonymous:1877:SMMa

[Ano77p] Anonymous. Stated meeting, March 2nd, 1877. Proceedings of the American Philosophical Society, 16(99):392–394, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982471.

Anonymous:1877:SMMd

[Ano77q] Anonymous. Stated meeting, May 20th, 1877. Proceedings of the American Philosophical Society, 17(100):1–2, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982267.

Anonymous:1877:SMMc

[Ano77r] Anonymous. Stated meeting, May 4th, 1877. Proceedings of the American Philosophical Society, 16(99):665–668, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982496.

Anonymous:1877:SMNb

[Ano77s] Anonymous. Stated meeting, November 16, 1877. Proceedings of the American Philosophical Society, 17(100):272–274, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982309.

Anonymous:1877:SMNa

[Ano77t] Anonymous. Stated meeting, November 2, 1877. Proceedings of the American Philosophical Society, 17(100):268–272, 280, June/ December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982308.

Anonymous:1877:SMOb

[Ano77u] Anonymous. Stated meeting, October 19th, 1877. Proceedings of the American Philosophical Society, 17(100):16, 267–268, June/ December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982273.

Anonymous:1877:SMOa

[Ano77v] Anonymous. Stated meeting, October 5th, 1877. Proceedings of the American Philosophical Society, 17(100):13–16, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982272.

Anonymous:1877:SMS

[Ano77w] Anonymous. Stated meeting, September 21st, 1877. Proceedings of the American Philosophical Society, 17(100):10–13, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982271.

Anonymous:1877:VIa

[Ano77x] Anonymous. Volume information. Proceedings of the American Philosophical Society, 16(99):671–677, January/May 1877. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982466.

Anonymous:1877:VIb

[Ano77y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 17(100):??, June/December 1877. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982266.

Anonymous:1878:Aa

[Ano78a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 17(101):??, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1878:Ab

[Ano78b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 18(102):??, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1878:E

[Ano78c] Anonymous. Errata. Proceedings of the American Philosophical Society, 18(102):??, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982768.

Anonymous:1878:SMAb

[Ano78d] Anonymous. Stated meeting, April 19th, 1878. Proceedings of the American Philosophical Society, 17(101):713–715, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982665.

Anonymous:1878:SMAa

[Ano78e] Anonymous. Stated meeting, April 5th, 1878. Proceedings of the American Philosophical Society, 17(101):711–713, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982664.

Anonymous:1878:SMAc

[Ano78f] Anonymous. Stated meeting, August 16, 1878. Proceedings of the American Philosophical Society, 18(102):1–2, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982747.

Anonymous:1878:SMDb

[Ano78g] Anonymous. Stated meeting, December 20, 1878. Proceedings of the American Philosophical Society, 18(102):87–88, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982763.

Anonymous:1878:SMDa

[Ano78h] Anonymous. Stated meeting, December 6, 1878. Proceedings of the American Philosophical Society, 18(102):83–87, July/December 1878.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982762.

Anonymous:1878:SMFa

[Ano78i] Anonymous. Stated meeting, Feb. 1, 1878. Proceedings of the American Philosophical Society, 17(101):308–311, January/June 1878. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982636.

Anonymous:1878:SMFb

[Ano78j] Anonymous. Stated meeting, Feb. 15, 1878. Proceedings of the American Philosophical Society, 17(101):311–312, January/June 1878. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982637.

Anonymous:1878:SMJb

[Ano78k] Anonymous. Stated meeting, Jan. 18, 1878. Proceedings of the American Philosophical Society, 17(101):290–293, January/June 1878. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982634.

Anonymous:1878:SMJa

[Ano78l] Anonymous. Stated meeting, Jan. 4, 1878. Proceedings of the American Philosophical Society, 17(101):281–283, January/June 1878. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982632.

Anonymous:1878:SMJc

[Ano78m] Anonymous. Stated meeting, June 21st, 1878. Proceedings of the American Philosophical Society, 17(101):726–728, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982668.

Anonymous:1878:SMMa

[Ano78n] Anonymous. Stated meeting, March 1, 1878. Proceedings of the American Philosophical Society, 17(101):313–314, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982638.

Anonymous:1878:SMMb

[Ano78o] Anonymous. Stated meeting, March 15th, 1878. Proceedings of the American Philosophical Society, 17(101):709–710, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982663.

Anonymous:1878:SMMd

[Ano78p] Anonymous. Stated meeting, May 17th, 1878. Proceedings of the American Philosophical Society, 17(101):720–725, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982667.

Anonymous:1878:SMMc

[Ano78q] Anonymous. Stated meeting, May 3d, 1878. Proceedings of the American Philosophical Society, 17(101):715–720, January/June 1878. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982666.

Anonymous:1878:SMNb

[Ano78r] Anonymous. Stated meeting, Nov. 15, 1878. Proceedings of the American Philosophical Society, 18(102):61–78, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982760.

Anonymous:1878:SMNa

[Ano78s] Anonymous. Stated meeting, November 1, 1878. Proceedings of the American Philosophical Society, 18(102):43–45, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982756.

Anonymous:1878:SMOb

[Ano78t] Anonymous. Stated meeting, October 18, 1878. Proceedings of the American Philosophical Society, 18(102):37–40, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982754.

Anonymous:1878:SMOa

[Ano78u] Anonymous. Stated meeting, October 4, 1878. Proceedings of the American Philosophical Society, 18(102):3–8, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982749.

Anonymous:1878:SMS

[Ano78v] Anonymous. Stated meeting, September 20, 1878. Proceedings of the American Philosophical Society, 18(102):2–3, July/December 1878.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982748.

Anonymous:1878:VIa

[Ano78w] Anonymous. Volume information. Proceedings of the American Philosophical Society, 17(101):729–734, January/June 1878. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982631.

Anonymous:1878:VIb

[Ano78x] Anonymous. Volume information. Proceedings of the American Philosophical Society, 18(102):??, July/December 1878. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982746.

Anonymous:1879:Aa

[Ano79a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 18(103):??, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1879:Ab

[Ano79b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 18(104):??, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1879:SMAb

[Ano79c] Anonymous. Stated meeting, April 18, 1879. Proceedings of the American Philosophical Society, 18(103):235–237, January/June 1879. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982331.

Anonymous:1879:SMAa

[Ano79d] Anonymous. Stated meeting, April 4, 1879. Proceedings of the American Philosophical Society, 18(103):217–221, January/June 1879. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982327.

Anonymous:1879:SMAc

[Ano79e] Anonymous. Stated meeting, August 15, 1879. Proceedings of the American Philosophical Society, 18(104):355–356, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982512.

Anonymous:1879:SMDb

[Ano79f] Anonymous. Stated meeting, December 19, 1879. Proceedings of the American Philosophical Society, 18(104):377–379, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982520.

Anonymous:1879:SMDa

[Ano79g] Anonymous. Stated meeting, December 5, 1879. Proceedings of the American Philosophical Society, 18(104):364–369, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982518.

Anonymous:1879:SMFb

[Ano79h] Anonymous. Stated meeting, February 21, 1879. Proceedings of the American Philosophical Society, 18(103):206–208, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982321.

Anonymous:1879:SMFa

[Ano79i] Anonymous. Stated meeting, February 7, 1879. Proceedings of the American Philosophical Society, 18(103):181–184, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982317.

Anonymous:1879:SMJb

[Ano79j] Anonymous. Stated meeting, January 17, 1879. Proceedings of the American Philosophical Society, 18(103):178–181, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982316.

Anonymous:1879:SMJa

[Ano79k] Anonymous. Stated meeting, January 3, 1879. Proceedings of the American Philosophical Society, 18(103):153–155, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982314.

Anonymous:1879:SMJd

[Ano79l] Anonymous. Stated meeting, July 18, 1879. Proceedings of the American Philosophical Society, 18(104):353–355, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982511.

Anonymous:1879:SMJc

[Ano79m] Anonymous. Stated meeting, June 20, 1879. Proceedings of the American Philosophical Society, 18(104):351–353, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982510.

Anonymous:1879:SMMb

[Ano79n] Anonymous. Stated meeting, March 21, 1879. Proceedings of the American Philosophical Society, 18(103):212–213, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982324.

Anonymous:1879:SMMa

[Ano79o] Anonymous. Stated meeting, March 7, 1879. Proceedings of the American Philosophical Society, 18(103):210–212, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982323.

Anonymous:1879:SMMd

[Ano79p] Anonymous. Stated meeting, May 16, 1879. Proceedings of the American Philosophical Society, 18(103):247–251, January/June 1879. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982334.

Anonymous:1879:SMMc

[Ano79q] Anonymous. Stated meeting, May 2, 1879. Proceedings of the American Philosophical Society, 18(103):237–239, January/June 1879. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982332.

Anonymous:1879:SMNb

[Ano79r] Anonymous. Stated meeting, November 21, 1879. Proceedings of the American Philosophical Society, 18(104):362–364, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982517.

Anonymous:1879:SMNa

[Ano79s] Anonymous. Stated meeting, November 7, 1879. Proceedings of the American Philosophical Society, 18(104):361–362, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982516.

Anonymous:1879:SMOb

[Ano79t] Anonymous. Stated meeting, October 17, 1879. Proceedings of the American Philosophical Society, 18(104):359–360, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982515.

Anonymous:1879:SMOa

[Ano79u] Anonymous. Stated meeting, October 3, 1879. Proceedings of the American Philosophical Society, 18(104):358–359, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982514.

Anonymous:1879:SMS

[Ano79v] Anonymous. Stated meeting, September 19, 1879. Proceedings of the American Philosophical Society, 18(104):356–358, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982513.

Anonymous:1880:Aa

[Ano80a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 18(105):??, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1880:Ab

[Ano80b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 18(106):??, March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1880:Ac

[Ano80c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 19(107):??, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1880:SMOb

[Ano80d] Anonymous. Special meeting, October 11, 1880. Proceedings of the American Philosophical Society, 19(107):117–118, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982617.

Anonymous:1880:SMAb

[Ano80e] Anonymous. Stated meeting, April 16, 1880. Proceedings of the American Philosophical Society, 19(107):16–17, March/December 1880.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982600.

Anonymous:1880:SMAa

[Ano80f] Anonymous. Stated meeting April 2, 1880. Proceedings of the American Philosophical Society, 19(107):14–15, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982598.

Anonymous:1880:SMAc

[Ano80g] Anonymous. Stated meeting, August 20, 1880. Proceedings of the American Philosophical Society, 19(107):82–83, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982611.

Anonymous:1880:SMDb

[Ano80h] Anonymous. Stated meeting, Dec. 17, 1880. Proceedings of the American Philosophical Society, 19(107):196–198, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982628.

Anonymous:1880:SMDa

[Ano80i] Anonymous. Stated meeting, December 3, 1880. Proceedings of the American Philosophical Society, 19(107):194–196, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982627.

Anonymous:1880:SMFb

[Ano80j] Anonymous. Stated meeting, February 20, 1880. Proceedings of the American Philosophical Society, 18(105):415–417, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982772.

Anonymous:1880:SMFa

[Ano80k] Anonymous. Stated meeting, February 6, 1880. Proceedings of the American Philosophical Society, 18(105):413–415, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982771.

Anonymous:1880:SMJb

[Ano80l] Anonymous. Stated meeting, January 16, 1880. Proceedings of the American Philosophical Society, 18(105):411–413, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982770.

Anonymous:1880:SMJa

[Ano80m] Anonymous. Stated meeting, January 2, 1880. Proceedings of the American Philosophical Society, 18(105):409–410, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982769.

Anonymous:1880:SMJd

[Ano80n] Anonymous. Stated meeting, July 16, 1880. Proceedings of the American Philosophical Society, 19(107):75–76, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982609.

Anonymous:1880:SMJc

[Ano80o] Anonymous. Stated meeting, June 18, 1880. Proceedings of the American Philosophical Society, 19(107):69–75, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982608.

Anonymous:1880:SMMb

[Ano80p] Anonymous. Stated meeting, March 19, 1880. Proceedings of the American Philosophical Society, 19(107):1–4, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982594.

Anonymous:1880:SMMa

[Ano80q] Anonymous. Stated meeting, March 5, 1880. Proceedings of the American Philosophical Society, 18(105):417–419, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982773.

Anonymous:1880:SMMd

[Ano80r] Anonymous. Stated meeting, May 21, 1880. Proceedings of the American Philosophical Society, 19(107):59–62, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982606.

Anonymous:1880:SMMc

[Ano80s] Anonymous. Stated meeting, May 7, 1880. Proceedings of the American Philosophical Society, 19(107):25–27, March/December 1880.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982603.

Anonymous:1880:SMNa

[Ano80t] Anonymous. Stated meeting, Nov. 5, 1880. Proceedings of the American Philosophical Society, 19(107):158–161, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982620.

Anonymous:1880:SMNb

[Ano80u] Anonymous. Stated meeting, November 19, 1880. Proceedings of the American Philosophical Society, 19(107):191–192, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982625.

Anonymous:1880:SMOa

[Ano80v] Anonymous. Stated meeting, October 1, 1880. Proceedings of the American Philosophical Society, 19(107):117, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982616.

Anonymous:1880:SMOc

[Ano80w] Anonymous. Stated meeting, October 15, 1880. Proceedings of the American Philosophical Society, 19(107):153–158, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982619.

Anonymous:1880:SMS

[Ano80x] Anonymous. Stated meeting, September 17, 1880. Proceedings of the American Philosophical Society, 19(107):107–111, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982614.

Anonymous:1880:VIa

[Ano80y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 18(106):591–598, March 15, 1880. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982824.

Anonymous:1880:VIb

[Ano80z] Anonymous. Volume information. Proceedings of the American Philosophical Society, 19(107):??, March/December 1880. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982593.

Anonymous:1881:Aa

[Ano81a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 19(108):??, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1881:Ab

[Ano81b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 19(109):??, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1881:Ac

[Ano81c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 20(110):??, 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1881:BM

[Ano81d] Anonymous. Back matter. Proceedings of the American Philosophical Society, 19(109):1–72, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982806.

Anonymous:1881:E

[Ano81e] Anonymous. Erratum. Proceedings of the American Philosophical Society, 20(110):??, 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982416.

Anonymous:1881:OIV

[Ano81f] Anonymous. Omission from the index of Vol. XIX. Proceedings of the American Philosophical Society, 20(110):??, 1881. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982417.

Anonymous:1881:SMAa

[Ano81g] Anonymous. Stated meeting, April 1, 1881. Proceedings of the American Philosophical Society, 19(108):348–352, January/June 1881. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982257.

Anonymous:1881:SMAb

[Ano81h] Anonymous. Stated meeting, April 15, 1881. Proceedings of the American Philosophical Society, 19(108):353–354, January/June 1881. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982258.

Anonymous:1881:SMDb

[Ano81i] Anonymous. Stated meeting, December 16, 1881. Proceedings of the American Philosophical Society, 19(109):649–651, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982805.

Anonymous:1881:SMDa

[Ano81j] Anonymous. Stated meeting, December 2d, 1881. Proceedings of the American Philosophical Society, 19(109):647–649, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982804.

Anonymous:1881:SMFb

[Ano81k] Anonymous. Stated meeting, February 18, 1881. Proceedings of the American Philosophical Society, 19(108):285–287, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982250.

Anonymous:1881:SMFa

[Ano81l] Anonymous. Stated meeting, February 4, 1881. Proceedings of the American Philosophical Society, 19(108):274–276, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982247.

Anonymous:1881:SMJb

[Ano81m] Anonymous. Stated meeting, January 21, 1881. Proceedings of the American Philosophical Society, 19(108):217–218, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982244.

Anonymous:1881:SMJa

[Ano81n] Anonymous. Stated meeting, January 7, 1881. Proceedings of the American Philosophical Society, 19(108):211–213, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982241.

Anonymous:1881:SMJd

[Ano81o] Anonymous. Stated meeting, July 15, 1881. Proceedings of the American Philosophical Society, 19(109):481–482, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982792.

Anonymous:1881:SMJc

[Ano81p] Anonymous. Stated meeting, June 17th, 1881. Proceedings of the American Philosophical Society, 19(109):436–438, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982787.

Anonymous:1881:SMMb

[Ano81q] Anonymous. Stated meeting, March 18, 1881. Proceedings of the American Philosophical Society, 19(108):299–300, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982254.

Anonymous:1881:SMMa

[Ano81r] Anonymous. Stated meeting, March 4, 1881. Proceedings of the American Philosophical Society, 19(108):298, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982253.

Anonymous:1881:SMMd

[Ano81s] Anonymous. Stated meeting, May 20, 1881. Proceedings of the American Philosophical Society, 19(108):408–409, January/June 1881. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982264.

Anonymous:1881:SMMc

[Ano81t] Anonymous. Stated meeting, May 6, 1881. Proceedings of the American Philosophical Society, 19(108):403–405, January/June 1881. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982262.

Anonymous:1881:SMNb

[Ano81u] Anonymous. Stated meeting, November 18, 1881. Proceedings of the American Philosophical Society, 19(109):566–567, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982801.

Anonymous:1881:SMNa

[Ano81v] Anonymous. Stated meeting, November 4, 1881. Proceedings of the American Philosophical Society, 19(109):562–566, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982800.

Anonymous:1881:SMOb

[Ano81w] Anonymous. Stated meeting, October 21, 1881. Proceedings of the American Philosophical Society, 19(109):521–524, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982798.

Anonymous:1881:SMOa

[Ano81x] Anonymous. Stated meeting, October 7, 1881. Proceedings of the American Philosophical Society, 19(109):496–498, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982795.

Anonymous:1881:SMS

[Ano81y] Anonymous. Stated meeting, September 17, 1881. Proceedings of the American Philosophical Society, 19(109):483, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982793.

Anonymous:1881:VIa

[Ano81z] Anonymous. Volume information. Proceedings of the American Philosophical Society, 19(109):653–70, June/December 1881. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982785.

Anonymous:1881:VIb

[Ano81-27] Anonymous. Volume information. Proceedings of the American Philosophical Society, 20(110):??, 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982414.

Anonymous:1882:Aa

[Ano82a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 20(111):??, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1882:Ab

[Ano82b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 20(112):??, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1882:SMAb

[Ano82c] Anonymous. Stated meeting, April 21, 1882. Proceedings of the American Philosophical Society, 20(111):233–236, January/June 1882. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982723.

Anonymous:1882:SMAa

[Ano82d] Anonymous. Stated meeting, April 7, 1882. Proceedings of the American Philosophical Society, 20(111):231–233, January/June 1882. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982722.

Anonymous:1882:SMAc

[Ano82e] Anonymous. Stated meeting, August 18, 1882. Proceedings of the American Philosophical Society, 20(112):379–380, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982687.

Anonymous:1882:SMDa

[Ano82f] Anonymous. Stated meeting, Dec. 1, 1882. Proceedings of the American Philosophical Society, 20(112):505–508, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982701.

Anonymous:1882:SMDb

[Ano82g] Anonymous. Stated meeting, December 15, 1882. Proceedings of the American Philosophical Society, 20(112):509, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982702.

Anonymous:1882:SMFb

[Ano82h] Anonymous. Stated meeting, February 17, 1882. Proceedings of the American Philosophical Society, 20(111):224–225, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982717.

Anonymous:1882:SMFa

[Ano82i] Anonymous. Stated meeting, February 3, 1882. Proceedings of the American Philosophical Society, 20(111):204–206, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982715.

Anonymous:1882:SMJb

[Ano82j] Anonymous. Stated meeting, January 20, 1882. Proceedings of the American Philosophical Society, 20(111):202–204, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982714.

Anonymous:1882:SMJa

[Ano82k] Anonymous. Stated meeting, January 6, 1882. Proceedings of the American Philosophical Society, 20(111):197–199, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982712.

Anonymous:1882:SMJd

[Ano82l] Anonymous. Stated meeting, July 21, 1882. Proceedings of the American Philosophical Society, 20(112):378–379, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982686.

Anonymous:1882:SMJc

[Ano82m] Anonymous. Stated meeting, June 16, 1882. Proceedings of the American Philosophical Society, 20(112):332–334, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982683.

Anonymous:1882:SMMb

[Ano82n] Anonymous. Stated meeting, March 17, 1882. Proceedings of the American Philosophical Society, 20(111):229–230, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982720.

Anonymous:1882:SMMa

[Ano82o] Anonymous. Stated meeting, March 3, 1882. Proceedings of the American Philosophical Society, 20(111):225–226, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982718.

Anonymous:1882:SMMd

[Ano82p] Anonymous. Stated meeting, May 19, 1882. Proceedings of the American Philosophical Society, 20(111):298, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982728.

Anonymous:1882:SMMe

[Ano82q] Anonymous. [Stated meeting, may 19, 1882]. Proceedings of the American Philosophical Society, 20(112):299, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982681.

Anonymous:1882:SMMc

[Ano82r] Anonymous. Stated meeting, May 5, 1882. Proceedings of the American Philosophical Society, 20(111):291–292, January/June 1882. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982726.

Anonymous:1882:SMNb

[Ano82s] Anonymous. Stated meeting, Nov. 17, 1882. Proceedings of the American Philosophical Society, 20(112):503–505, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982700.

Anonymous:1882:SMNa

[Ano82t] Anonymous. Stated meeting, November 3, 1882. Proceedings of the American Philosophical Society, 20(112):496–498, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982698.

Anonymous:1882:SMOb

[Ano82u] Anonymous. Stated meeting, Oct. 20th, 1882. Proceedings of the American Philosophical Society, 20(112):476–478, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982696.

Anonymous:1882:SMOa

[Ano82v] Anonymous. Stated meeting, Oct. 6th, 1882. Proceedings of the American Philosophical Society, 20(112):475–476, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982695.

Anonymous:1882:SMS

[Ano82w] Anonymous. Stated meeting, September 15, 1882. Proceedings of the American Philosophical Society, 20(112):405, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982689.

Anonymous:1883:Aa

[Ano83a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 20(113):??, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1883:Ab

[Ano83b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 21(114):??, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1883:MTM

[Ano83c] Anonymous. Map of the terminal moraine. Proceedings of the American Philosophical Society, 20(113):662–664, January/April 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982461.

Anonymous:1883:ENY

[Ano83d] Anonymous. On the equivalent of the New York Portage, in Perry County, Middle Pennsylvania. Proceedings of the American Philosophical Society, 21(114):230–235, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982380.

Anonymous:1883:SMAb

[Ano83e] Anonymous. Stated meeting, April 20, 1883. Proceedings of the American Philosophical Society, 21(114):45, 48–49, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982360.

Anonymous:1883:SMAa

[Ano83f] Anonymous. Stated meeting, April 6, 1883. Proceedings of the American Philosophical Society, 20(113):645–649, January/April 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982458.

Anonymous:1883:SMAc

[Ano83g] Anonymous. Stated meeting, August 17, 1883. Proceedings of the American Philosophical Society, 21(114):213–214, March 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982376.

Anonymous:1883:SMDb

[Ano83h] Anonymous. Stated meeting, December 21, 1883. Proceedings of the American Philosophical Society, 21(114):329–330, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982399.

Anonymous:1883:SMDa

[Ano83i] Anonymous. Stated meeting, December 7, 1883. Proceedings of the American Philosophical Society, 21(114):306–307, March 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982394.

Anonymous:1883:SMFb

[Ano83j] Anonymous. Stated meeting, Feb. 16, 1883. Proceedings of the American Philosophical Society, 20(113):641–642, January/April 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982455.

Anonymous:1883:SMFa

[Ano83k] Anonymous. Stated meeting, Feb. 2, 1883. Proceedings of the American Philosophical Society, 20(113):640–641, January/April 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982454.

Anonymous:1883:SMJb

[Ano83l] Anonymous. Stated meeting, Jan. 19, 1883. Proceedings of the American Philosophical Society, 20(113):638–640, January/April 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982453.

Anonymous:1883:SMJa

[Ano83m] Anonymous. Stated meeting, Jan. 5, 1883. Proceedings of the American Philosophical Society, 20(113):636–637, January/April 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982452.

Anonymous:1883:SMJe

[Ano83n] Anonymous. Stated meeting, January 4, 1884. Proceedings of the American Philosophical Society, 21(114):343–344, March 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982402.

Anonymous:1883:SMJd

[Ano83o] Anonymous. Stated meeting, July 20, 1883. Proceedings of the American Philosophical Society, 21(114):212–213, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982375.

Anonymous:1883:SMJc

[Ano83p] Anonymous. Stated meeting, June 16, 1883. Proceedings of the American Philosophical Society, 21(114):211–212, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982374.

Anonymous:1883:SMMb

[Ano83q] Anonymous. Stated meeting, March 16, 1883. Proceedings of the American Philosophical Society, 20(113):644–645, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982457.

Anonymous:1883:SMMa

[Ano83r] Anonymous. Stated meeting, March 2, 1883. Proceedings of the American Philosophical Society, 20(113):643, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982456.

Anonymous:1883:SMMd

[Ano83s] Anonymous. Stated meeting, May 18, 1883. Proceedings of the American Philosophical Society, 21(114):209–210, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982373.

Anonymous:1883:SMMc

[Ano83t] Anonymous. Stated meeting, May 4, 1883. Proceedings of the American Philosophical Society, 21(114):109–111, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982367.

Anonymous:1883:SMNa

[Ano83u] Anonymous. Stated meeting, Nov. 2, 1883. Proceedings of the American Philosophical Society, 21(114):282–286, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982388.

Anonymous:1883:SMNb

[Ano83v] Anonymous. Stated meeting, November 16, 1883. Proceedings of the American Philosophical Society, 21(114):290–294, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982390.

Anonymous:1883:SMOb

[Ano83w] Anonymous. Stated meeting, October 19, 1883. Proceedings of the American Philosophical Society, 21(114):280–282, March 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982387.

Anonymous:1883:SMOa

[Ano83x] Anonymous. Stated meeting, October 5, 1883. Proceedings of the American Philosophical Society, 21(114):239–240, March 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982383.

Anonymous:1883:SMS

[Ano83y] Anonymous. Stated meeting, September 21, 1883. Proceedings of the American Philosophical Society, 21(114):214–218, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982377.

Anonymous:1883:VIa

[Ano83z] Anonymous. Volume information. Proceedings of the American Philosophical Society, 20(113):679–683, January/April 1883. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982440.

Anonymous:1883:VIb

[Ano83-27] Anonymous. Volume information. Proceedings of the American Philosophical Society, 21(114):??, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982358.

Anonymous:1884:Aa

[Ano84a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 21(115):??, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1884:Ab

[Ano84b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 21(116):??, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1884:SMAb

[Ano84c] Anonymous. Stated meeting, April 18, 1884. Proceedings of the American Philosophical Society, 21(115):457–458, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982740.

Anonymous:1884:SMAa

[Ano84d] Anonymous. Stated meeting, April 5, 1884. Proceedings of the American Philosophical Society, 21(115):450–451, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982737.

Anonymous:1884:SMAc

[Ano84e] Anonymous. Stated meeting, Aug. 15, 1884. Proceedings of the American Philosophical Society, 21(116):672–673, June 1884. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982346.

Anonymous:1884:SMD

[Ano84f] Anonymous. Stated meeting, December 5, 1884. Proceedings of the American Philosophical Society, 21(116):705–706, June 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982357.

Anonymous:1884:SMFa

[Ano84g] Anonymous. Stated meeting, February 1, 1884. Proceedings of the American Philosophical Society, 21(115):420–422, April 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982732.

Anonymous:1884:SMFb

[Ano84h] Anonymous. Stated meeting, February 15, 1884. Proceedings of the American Philosophical Society, 21(115):423–424, April 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982733.

Anonymous:1884:SMJa

[Ano84i] Anonymous. Stated meeting, January 18, 1884. Proceedings of the American Philosophical Society, 21(115):418–420, April 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982731.

Anonymous:1884:SMJc

[Ano84j] Anonymous. Stated meeting, July 18, 1884. Proceedings of the American Philosophical Society, 21(116):613–615, June 1884. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982342.

Anonymous:1884:SMJb

[Ano84k] Anonymous. Stated meeting, June 20, 1884. Proceedings of the American Philosophical Society, 21(116):611–613, June 1884. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982341.

Anonymous:1884:SMMb

[Ano84l] Anonymous. Stated meeting, March 21, 1884. Proceedings of the American Philosophical Society, 21(115):447–450, April 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982736.

Anonymous:1884:SMMa

[Ano84m] Anonymous. Stated meeting, March 7, 1884. Proceedings of the American Philosophical Society, 21(115):424–426, April 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982734.

Anonymous:1884:SMMd

[Ano84n] Anonymous. Stated meeting, May 16, 1884. Proceedings of the American Philosophical Society, 21(115):486–489, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982743.

Anonymous:1884:SMMc

[Ano84o] Anonymous. Stated meeting, May 2, 1884. Proceedings of the American Philosophical Society, 21(115):484–485, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982742.

Anonymous:1884:SMNb

[Ano84p] Anonymous. Stated meeting, November 21, 1884. Proceedings of the American Philosophical Society, 21(116):702–705, June 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982356.

Anonymous:1884:SMNa

[Ano84q] Anonymous. Stated meeting, November 7, 1884. Proceedings of the American Philosophical Society, 21(116):700–702, June 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982355.

Anonymous:1884:SMOb

[Ano84r] Anonymous. Stated meeting, October 17, 1884. Proceedings of the American Philosophical Society, 21(116):689–690, June 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982351.

Anonymous:1884:SMOa

[Ano84s] Anonymous. Stated meeting, October 3, 1884. Proceedings of the American Philosophical Society, 21(116):685–686, June 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982349.

Anonymous:1884:SMS

[Ano84t] Anonymous. Stated meeting Sept. 19, 1884. Proceedings of the American Philosophical Society, 21(116):674–675, June 1884. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982347.

Anonymous:1884:VI

[Ano84u] Anonymous. Volume information. Proceedings of the American Philosophical Society, 21(116):707–711, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982337.

Anonymous:1885:X01

[Ano85a] Anonymous. [1744–1768]. Proceedings of the American Philosophical Society, 22(119):1–6, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982526.

Anonymous:1885:X02

[Ano85b] Anonymous. [1768]. Proceedings of the American Philosophical Society, 22(119):7–22, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982527.

Anonymous:1885:X04

[Ano85c] Anonymous. [1775]. Proceedings of the American Philosophical Society, 22(119):94–98, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982529.

Anonymous:1885:X05

[Ano85d] Anonymous. [1776]. Proceedings of the American Philosophical Society, 22(119):98, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982530.

Anonymous:1885:X06

[Ano85e] Anonymous. [1779]. Proceedings of the American Philosophical Society, 22(119):99–106, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982531.

Anonymous:1885:X07

[Ano85f] Anonymous. [1780]. Proceedings of the American Philosophical Society, 22(119):106–110, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982532.

Anonymous:1885:X08

[Ano85g] Anonymous. [1781]. Proceedings of the American Philosophical Society, 22(119):110–114, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982533.

Anonymous:1885:X09

[Ano85h] Anonymous. [1782]. Proceedings of the American Philosophical Society, 22(119):114–115, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982534.

Anonymous:1885:X10

[Ano85i] Anonymous. [1783]. Proceedings of the American Philosophical Society, 22(119):115–120, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982535.

Anonymous:1885:X11

[Ano85j] Anonymous. [1784]. Proceedings of the American Philosophical Society, 22(119):121–129, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982536.

Anonymous:1885:X12

[Ano85k] Anonymous. [1785]. Proceedings of the American Philosophical Society, 22(119):129–136, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982537.

Anonymous:1885:X13

[Ano85l] Anonymous. [1786]. Proceedings of the American Philosophical Society, 22(119):136–146, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982538.

Anonymous:1885:X14

[Ano85m] Anonymous. [1787]. Proceedings of the American Philosophical Society, 22(119):147–158, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982539.

Anonymous:1885:X15

[Ano85n] Anonymous. [1788]. Proceedings of the American Philosophical Society, 22(119):158–167, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982540.

Anonymous:1885:X16

[Ano85o] Anonymous. [1789]. Proceedings of the American Philosophical Society, 22(119):167–178, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982541.

Anonymous:1885:X17

[Ano85p] Anonymous. [1790]. Proceedings of the American Philosophical Society, 22(119):178–187, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982542.

Anonymous:1885:X18

[Ano85q] Anonymous. [1791]. Proceedings of the American Philosophical Society, 22(119):187–201, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982543.

Anonymous:1885:X19

[Ano85r] Anonymous. [1792]. Proceedings of the American Philosophical Society, 22(119):201–211, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982544.

Anonymous:1885:X20

[Ano85s] Anonymous. [1793]. Proceedings of the American Philosophical Society, 22(119):211–217, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982545.

Anonymous:1885:X21

[Ano85t] Anonymous. [1794]. Proceedings of the American Philosophical Society, 22(119):217–227, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982546.

Anonymous:1885:X22

[Ano85u] Anonymous. [1795]. Proceedings of the American Philosophical Society, 22(119):227–235, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982547.

Anonymous:1885:X23

[Ano85v] Anonymous. [1796]. Proceedings of the American Philosophical Society, 22(119):235–245, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982548.

Anonymous:1885:X24

[Ano85w] Anonymous. [1797]. Proceedings of the American Philosophical Society, 22(119):246–265, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982549.

Anonymous:1885:X25

[Ano85x] Anonymous. [1798]. Proceedings of the American Philosophical Society, 22(119):265–277, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982550.

Anonymous:1885:X26

[Ano85y] Anonymous. [1799]. Proceedings of the American Philosophical Society, 22(119):277–289, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982551.

Anonymous:1885:X27

[Ano85z] Anonymous. [1800]. Proceedings of the American Philosophical Society, 22(119):290–307, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982552.

Anonymous:1885:X28

[Ano85-27] Anonymous. [1801]. Proceedings of the American Philosophical Society, 22(119):307–319, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982553.

Anonymous:1885:X29

[Ano85-28] Anonymous. [1802]. Proceedings of the American Philosophical Society, 22(119):319–330, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982554.

Anonymous:1885:X30

[Ano85-29] Anonymous. [1803]. Proceedings of the American Philosophical Society, 22(119):330–345, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982555.

Anonymous:1885:X31

[Ano85-30] Anonymous. [1804]. Proceedings of the American Philosophical Society, 22(119):345–369, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982556.

Anonymous:1885:X32

[Ano85-31] Anonymous. [1805]. Proceedings of the American Philosophical Society, 22(119):369–381, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982557.

Anonymous:1885:X33

[Ano85-32] Anonymous. [1806]. Proceedings of the American Philosophical Society, 22(119):381–392, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982558.

Anonymous:1885:X34

[Ano85-33] Anonymous. [1807]. Proceedings of the American Philosophical Society, 22(119):392–402, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982559.

Anonymous:1885:X35

[Ano85-34] Anonymous. [1808]. Proceedings of the American Philosophical Society, 22(119):402–411, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982560.

Anonymous:1885:X36

[Ano85-35] Anonymous. [1809]. Proceedings of the American Philosophical Society, 22(119):412–419, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982561.

Anonymous:1885:X37

[Ano85-36] Anonymous. [1810]. Proceedings of the American Philosophical Society, 22(119):419–425, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982562.

Anonymous:1885:X38

[Ano85-37] Anonymous. [1811]. Proceedings of the American Philosophical Society, 22(119):425–431, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982563.

Anonymous:1885:X39

[Ano85-38] Anonymous. [1812]. Proceedings of the American Philosophical Society, 22(119):431–438, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982564.

Anonymous:1885:X40

[Ano85-39] Anonymous. [1813]. Proceedings of the American Philosophical Society, 22(119):438–442, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982565.

Anonymous:1885:X41

[Ano85-40] Anonymous. [1814]. Proceedings of the American Philosophical Society, 22(119):443–450, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982566.

Anonymous:1885:X42

[Ano85-41] Anonymous. [1815]. Proceedings of the American Philosophical Society, 22(119):450–459, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982567.

Anonymous:1885:X43

[Ano85-42] Anonymous. [1816]. Proceedings of the American Philosophical Society, 22(119):459–468, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982568.

Anonymous:1885:X44

[Ano85-43] Anonymous. [1817]. Proceedings of the American Philosophical Society, 22(119):468–475, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982569.

Anonymous:1885:X45

[Ano85-44] Anonymous. [1818]. Proceedings of the American Philosophical Society, 22(119):476–484, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982570.

Anonymous:1885:X46

[Ano85-45] Anonymous. [1819]. Proceedings of the American Philosophical Society, 22(119):484–491, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982571.

Anonymous:1885:X47

[Ano85-46] Anonymous. [1820]. Proceedings of the American Philosophical Society, 22(119):492–498, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982572.

Anonymous:1885:X48

[Ano85-47] Anonymous. [1821]. Proceedings of the American Philosophical Society, 22(119):498–507, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982573.

Anonymous:1885:X49

[Ano85-48] Anonymous. [1822]. Proceedings of the American Philosophical Society, 22(119):507–516, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982574.

Anonymous:1885:X50

[Ano85-49] Anonymous. [1823]. Proceedings of the American Philosophical Society, 22(119):516–524, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982575.

Anonymous:1885:X51

[Ano85-50] Anonymous. [1824]. Proceedings of the American Philosophical Society, 22(119):524–535, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982576.

Anonymous:1885:X52

[Ano85-51] Anonymous. [1825]. Proceedings of the American Philosophical Society, 22(119):535–548, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982577.

Anonymous:1885:X53

[Ano85-52] Anonymous. [1826]. Proceedings of the American Philosophical Society, 22(119):548–560, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982578.

Anonymous:1885:X54

[Ano85-53] Anonymous. [1827]. Proceedings of the American Philosophical Society, 22(119):560–573, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982579.

Anonymous:1885:X55

[Ano85-54] Anonymous. [1828]. Proceedings of the American Philosophical Society, 22(119):573–587, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982580.

Anonymous:1885:X56

[Ano85-55] Anonymous. [1829]. Proceedings of the American Philosophical Society, 22(119):587–598, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982581.

Anonymous:1885:X57

[Ano85-56] Anonymous. [1830]. Proceedings of the American Philosophical Society, 22(119):598–612, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982582.

Anonymous:1885:X58

[Ano85-57] Anonymous. [1831]. Proceedings of the American Philosophical Society, 22(119):612–621, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982583.

Anonymous:1885:X59

[Ano85-58] Anonymous. [1832]. Proceedings of the American Philosophical Society, 22(119):621–629, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982584.

Anonymous:1885:X60

[Ano85-59] Anonymous. [1833]. Proceedings of the American Philosophical Society, 22(119):629–643, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982585.

Anonymous:1885:X61

[Ano85-60] Anonymous. [1834]. Proceedings of the American Philosophical Society, 22(119):643–662, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982586.

Anonymous:1885:X62

[Ano85-61] Anonymous. [1835]. Proceedings of the American Philosophical Society, 22(119):663–679, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982587.

Anonymous:1885:X63

[Ano85-62] Anonymous. [1836]. Proceedings of the American Philosophical Society, 22(119):679–697, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982588.

Anonymous:1885:X64

[Ano85-63] Anonymous. [1837]. Proceedings of the American Philosophical Society, 22(119):697–711, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982589.

Anonymous:1885:X65

[Ano85-64] Anonymous. [1838]. Proceedings of the American Philosophical Society, 22(119):711, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982590.

Anonymous:1885:Aa

[Ano85-65] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 22(117):??, January 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1885:Ab

[Ano85-66] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 22(118):??, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1885:Ac

[Ano85-67] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 22(119):??, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1885:Ad

[Ano85-68] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 22(120):??, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1885:BM

[Ano85-69] Anonymous. Back matter. Proceedings of the American Philosophical Society, 22(119):713–875, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982592.

Anonymous:1885:E

[Ano85-70] Anonymous. Errata. Proceedings of the American Philosophical Society, 22(119):712, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982591.

Anonymous:1885:FM

[Ano85-71] Anonymous. Front matter. Proceedings of the American Philosophical Society, 22(119):??, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982523.

Anonymous:1885:IPB

[Ano85-72] Anonymous. [Illustration]: Plate I. Building erected by the American Philosophical Society in 1785. Proceedings of the American Philosophical Society, 22(119):??, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982524.

Anonymous:1885:SMA

[Ano85-73] Anonymous. Stated meeting, April 17, 1885. Proceedings of the American Philosophical Society, 22(120):356–360, October 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982994.

Anonymous:1885:SMD

[Ano85-74] Anonymous. Stated meeting, Dec. 19, 1884. Proceedings of the American Philosophical Society, 22(117):85–88, January 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982410.

Anonymous:1885:SMFb

[Ano85-75] Anonymous. Stated meeting, February 20, 1885. Proceedings of the American Philosophical Society, 22(120):276–278, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982986.

Anonymous:1885:SMFa

[Ano85-76] Anonymous. Stated meeting, February 6, 1885. Proceedings of the American Philosophical Society, 22(120):273–276, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982985.

Anonymous:1885:SMJb

[Ano85-77] Anonymous. Stated meeting, January 16, 1885. Proceedings of the American Philosophical Society, 22(118):101–104, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982671.

Anonymous:1885:SMJa

[Ano85-78] Anonymous. Stated meeting, January 2, 1885. Proceedings of the American Philosophical Society, 22(118):98–101, March 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982670.

Anonymous:1885:SMJd

[Ano85-79] Anonymous. Stated meeting, July 17, 1885. Proceedings of the American Philosophical Society, 22(120):436–438, October 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983007.

Anonymous:1885:SMJc

[Ano85-80] Anonymous. Stated meeting, June 19, 1885. Proceedings of the American Philosophical Society, 22(120):432–435, October 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983006.

Anonymous:1885:SMMb

[Ano85-81] Anonymous. Stated meeting, March 20, 1885. Proceedings of the American Philosophical Society, 22(120):354–356, October 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982993.

Anonymous:1885:SMMa

[Ano85-82] Anonymous. Stated meeting, March 6, 1885. Proceedings of the American Philosophical Society, 22(120):278–280, October 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982987.

Anonymous:1885:SMMc

[Ano85-83] Anonymous. Stated meeting, May 1, 1885. Proceedings of the American Philosophical Society, 22(120):429–430, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983004.

Anonymous:1885:SMMd

[Ano85-84] Anonymous. Stated meeting, May 15, 1885. Proceedings of the American Philosophical Society, 22(120):430–432, October 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983005.

Anonymous:1885:V

[Ano85-85] Anonymous. Volume 1769–1774. Proceedings of the American Philosophical Society, 22(119):23–94, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982528.

Anonymous:1885:VIa

[Ano85-86] Anonymous. Volume information. Proceedings of the American Philosophical Society, 22(117):??, January 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982403.

Anonymous:1885:VIb

[Ano85-87] Anonymous. Volume information. Proceedings of the American Philosophical Society, 22(120):457–460, October 1885. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982979.

Anonymous:1886:Aa

[Ano86a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 23(121):??, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1886:Ab

[Ano86b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 23(122):??, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1886:Ac

[Ano86c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 23(123):??, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1886:Ad

[Ano86d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 23(124):??, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1886:SMAc

[Ano86e] Anonymous. Stated meeting, April 19, 1886. Proceedings of the American Philosophical Society, 23(123):451–452, July 1886. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983230.

Anonymous:1886:SMAb

[Ano86f] Anonymous. Stated meeting, April 2, 1886. Proceedings of the American Philosophical Society, 23(123):448–451, July 1886. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983229.

Anonymous:1886:SMAa

[Ano86g] Anonymous. Stated meeting, August 21, 1885. Proceedings of the American Philosophical Society, 23(121):152–153, January 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982918.

Anonymous:1886:SMDb

[Ano86h] Anonymous. Stated meeting, December 18, 1885. Proceedings of the American Philosophical Society, 23(122):321–324, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983248.

Anonymous:1886:SMDc

[Ano86i] Anonymous. Stated meeting, December 3, 1886. Proceedings of the American Philosophical Society, 23(124):611–612, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983341.

Anonymous:1886:SMDa

[Ano86j] Anonymous. Stated meeting, December 4, 1885. Proceedings of the American Philosophical Society, 23(122):319–320, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983247.

Anonymous:1886:SMFb

[Ano86k] Anonymous. Stated meeting, February 19, 1886. Proceedings of the American Philosophical Society, 23(122):333–334, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983252.

Anonymous:1886:SMFa

[Ano86l] Anonymous. Stated meeting, February 5, 1886. Proceedings of the American Philosophical Society, 23(122):330–332, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983251.

Anonymous:1886:SMJa

[Ano86m] Anonymous. Stated meeting, January 1, 1886. Proceedings of the American Philosophical Society, 23(122):324–326, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983249.

Anonymous:1886:SMJb

[Ano86n] Anonymous. Stated meeting, January 15, 1886. Proceedings of the American Philosophical Society, 23(122):326–330, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983250.

Anonymous:1886:SMMb

[Ano86o] Anonymous. Stated meeting, March 19, 1886. Proceedings of the American Philosophical Society, 23(123):446–448, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983228.

Anonymous:1886:SMMa

[Ano86p] Anonymous. Stated meeting, March 5, 1886. Proceedings of the American Philosophical Society, 23(122):335–336, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983253.

Anonymous:1886:SMMd

[Ano86q] Anonymous. Stated meeting, May 21, 1886. Proceedings of the American Philosophical Society, 23(123):454–457, July 1886. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983232.

Anonymous:1886:SMMc

[Ano86r] Anonymous. Stated meeting, May 7, 1886. Proceedings of the American Philosophical Society, 23(123):452–454, July 1886. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983231.

Anonymous:1886:SMNd

[Ano86s] Anonymous. Stated meeting, November 19, 1886. Proceedings of the American Philosophical Society, 23(124):571–572, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983338.

Anonymous:1886:SMNb

[Ano86t] Anonymous. Stated meeting, November 20, 1885. Proceedings of the American Philosophical Society, 23(122):315–319, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983246.

Anonymous:1886:SMNc

[Ano86u] Anonymous. Stated meeting, November 5, 1886. Proceedings of the American Philosophical Society, 23(124):569–571, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983337.

Anonymous:1886:SMNa

[Ano86v] Anonymous. Stated meeting, November 6, 1885. Proceedings of the American Philosophical Society, 23(122):313–315, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983245.

Anonymous:1886:SMOc

[Ano86w] Anonymous. Stated meeting, October 1, 1886. Proceedings of the American Philosophical Society, 23(124):541–544, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983333.

Anonymous:1886:SMOd

[Ano86x] Anonymous. Stated meeting, October 15, 1886. Proceedings of the American Philosophical Society, 23(124):544–546, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983334.

Anonymous:1886:SMOb

[Ano86y] Anonymous. Stated meeting, October 16, 1885. Proceedings of the American Philosophical Society, 23(121):157–160, January 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982921.

Anonymous:1886:SMOa

[Ano86z] Anonymous. Stated meeting, October 2, 1885. Proceedings of the American Philosophical Society, 23(121):154–156, January 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982920.

Anonymous:1886:SMSa

[Ano86-27] Anonymous. Stated meeting, September 18, 1885. Proceedings of the American Philosophical Society, 23(121):153–154, January 1886.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982919.

Anonymous:1886:SMSb

[Ano86-28] Anonymous. Stated meeting, September 3, 1886. Proceedings of the American Philosophical Society, 23(124):535–540, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983332.

Anonymous:1886:VIa

[Ano86-29] Anonymous. Volume information. Proceedings of the American Philosophical Society, 23(121):??, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982909.

Anonymous:1886:VIb

[Ano86-30] Anonymous. Volume information. Proceedings of the American Philosophical Society, 23(124):613–616, December 1886. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983325.

Anonymous:1887:Aa

[Ano87a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 24(125):??, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1887:Ab

[Ano87b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 24(126):??, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1887:E

[Ano87c] Anonymous. Erratum. Proceedings of the American Philosophical Society, 24(126):449, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983097.

Anonymous:1887:NMW

[Ano87d] Anonymous. On a non-magnetizable watch. Proceedings of the American Philosophical Society, 24(126):419–420, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983089.

Anonymous:1887:G

[Ano87e] Anonymous. On the gramophone. Proceedings of the American Philosophical Society, 24(126):420–421, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983090.

Anonymous:1887:SMAa

[Ano87f] Anonymous. Stated meeting, April 1, 1887. Proceedings of the American Philosophical Society, 24(125):199–201, January/June 1887. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983140.

Anonymous:1887:SMAb

[Ano87g] Anonymous. Stated meeting, April 15, 1887. Proceedings of the American Philosophical Society, 24(125):201–203, January/June 1887. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983141.

Anonymous:1887:SMDc

[Ano87h] Anonymous. Stated meeting, December 16, 1887. Proceedings of the American Philosophical Society, 24(126):444–446, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983096.

Anonymous:1887:SMDa

[Ano87i] Anonymous. Stated meeting, December 17, 1886. Proceedings of the American Philosophical Society, 24(125):184–186, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983133.

Anonymous:1887:SMDb

[Ano87j] Anonymous. Stated meeting, December 2, 1887. Proceedings of the American Philosophical Society, 24(126):434–436, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983093.

Anonymous:1887:SMFb

[Ano87k] Anonymous. Stated meeting, February 18, 1887. Proceedings of the American Philosophical Society, 24(125):195–196, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983137.

Anonymous:1887:SMFa

[Ano87l] Anonymous. Stated meeting, February 4, 1887. Proceedings of the American Philosophical Society, 24(125):192–194, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983136.

Anonymous:1887:SMJb

[Ano87m] Anonymous. Stated meeting, January 21, 1887. Proceedings of the American Philosophical Society, 24(125):190–192, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983135.

Anonymous:1887:SMJa

[Ano87n] Anonymous. Stated meeting, January 7, 1887. Proceedings of the American Philosophical Society, 24(125):187–189, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983134.

Anonymous:1887:SMMb

[Ano87o] Anonymous. Stated meeting, March 18, 1887. Proceedings of the American Philosophical Society, 24(125):198–199, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983139.

Anonymous:1887:SMMa

[Ano87p] Anonymous. Stated meeting, March 4, 1887. Proceedings of the American Philosophical Society, 24(125):196–198, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983138.

Anonymous:1887:SMMd

[Ano87q] Anonymous. Stated meeting, May 20, 1887. Proceedings of the American Philosophical Society, 24(125):206–208, January/June 1887. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983143.

Anonymous:1887:SMMc

[Ano87r] Anonymous. Stated meeting, May 6th, 1887. Proceedings of the American Philosophical Society, 24(125):203–206, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983142.

Anonymous:1887:SMNb

[Ano87s] Anonymous. Stated meeting November 18, 1887. Proceedings of the American Philosophical Society, 24(126):417–419, 421, July/ December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983088.

Anonymous:1887:SMNa

[Ano87t] Anonymous. Stated meeting, November 4, 1887. Proceedings of the American Philosophical Society, 24(126):416–417, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983087.

Anonymous:1887:SMOb

[Ano87u] Anonymous. Stated meeting, October 21, 1887. Proceedings of the American Philosophical Society, 24(126):414–415, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983086.

Anonymous:1887:SMOa

[Ano87v] Anonymous. Stated meeting, October 7, 1887. Proceedings of the American Philosophical Society, 24(126):412–413, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983085.

Anonymous:1887:SMSb

[Ano87w] Anonymous. Stated meeting, September 16, 1887. Proceedings of the American Philosophical Society, 24(126):410–411, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983084.

Anonymous:1887:SMSa

[Ano87x] Anonymous. Stated meeting, September 2, 1887. Proceedings of the American Philosophical Society, 24(126):403–409, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983083.

Anonymous:1887:VIa

[Ano87y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 24(125):??, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983119.

Anonymous:1887:VIb

[Ano87z] Anonymous. Volume information. Proceedings of the American Philosophical Society, 24(126):447–449, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983069.

Anonymous:1888:Aa

[Ano88a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 25(127):??, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1888:Ab

[Ano88b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 25(128):??, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1888:SMAb

[Ano88c] Anonymous. Stated meeting, April 20, 1888. Proceedings of the American Philosophical Society, 25(127):148–153, January/June 1888. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983298.

Anonymous:1888:SMAa

[Ano88d] Anonymous. Stated meeting, April 6, 1888. Proceedings of the American Philosophical Society, 25(127):125–131, January/June 1888. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983294.

Anonymous:1888:SMDb

[Ano88e] Anonymous. Stated meeting, December 21, 1888. Proceedings of the American Philosophical Society, 25(128):319–321, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983067.

Anonymous:1888:SMDa

[Ano88f] Anonymous. Stated meeting, December 7, 1888. Proceedings of the American Philosophical Society, 25(128):307–318, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983066.

Anonymous:1888:SMFb

[Ano88g] Anonymous. Stated meeting, February 17, 1888. Proceedings of the American Philosophical Society, 25(127):117–118, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983291.

Anonymous:1888:SMFa

[Ano88h] Anonymous. Stated meeting, February 3, 1888. Proceedings of the American Philosophical Society, 25(127):113–116, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983290.

Anonymous:1888:SMJb

[Ano88i] Anonymous. Stated meeting, January 20, 1888. Proceedings of the American Philosophical Society, 25(127):111–113, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983289.

Anonymous:1888:SMJa

[Ano88j] Anonymous. Stated meeting, January 6, 1888. Proceedings of the American Philosophical Society, 25(127):1–18, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983277.

Anonymous:1888:SMMb

[Ano88k] Anonymous. Stated meeting, March 16, 1888. Proceedings of the American Philosophical Society, 25(127):122–125, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983293.

Anonymous:1888:SMMa

[Ano88l] Anonymous. Stated meeting, March 2, 1888. Proceedings of the American Philosophical Society, 25(127):119–122, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983292.

Anonymous:1888:SMMd

[Ano88m] Anonymous. Stated meeting, May 18, 1888. Proceedings of the American Philosophical Society, 25(128):195–196, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983054.

Anonymous:1888:SMMc

[Ano88n] Anonymous. Stated meeting, May 4, 1888. Proceedings of the American Philosophical Society, 25(127):153–158, January/June 1888. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983299.

Anonymous:1888:SMNb

[Ano88o] Anonymous. Stated meeting, November 16, 1888. Proceedings of the American Philosophical Society, 25(128):304–306, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983065.

Anonymous:1888:SMNa

[Ano88p] Anonymous. Stated meeting, November 2, 1888. Proceedings of the American Philosophical Society, 25(128):302–304, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983064.

Anonymous:1888:SMOb

[Ano88q] Anonymous. Stated meeting, October 19, 1888. Proceedings of the American Philosophical Society, 25(128):299–302, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983063.

Anonymous:1888:SMOa

[Ano88r] Anonymous. Stated meeting, October 5, 1888. Proceedings of the American Philosophical Society, 25(128):297–299, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983062.

Anonymous:1888:SMSb

[Ano88s] Anonymous. Stated meeting, September 21, 1888. Proceedings of the American Philosophical Society, 25(128):200–204, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983056.

Anonymous:1888:SMSa

[Ano88t] Anonymous. Stated meeting, September 7, 1888. Proceedings of the American Philosophical Society, 25(128):197–200, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983055.

Anonymous:1888:VIa

[Ano88u] Anonymous. Volume information. Proceedings of the American Philosophical Society, 25(127):??, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983276.

Anonymous:1888:VIb

[Ano88v] Anonymous. Volume information. Proceedings of the American Philosophical Society, 25(128):329–331, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983048.

Anonymous:1889:Aa

[Ano89a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 26(129):??, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1889:Ab

[Ano89b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 26(130):??, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1889:Ac

[Ano89c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 27(131):??, November 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1889:BMa

[Ano89d] Anonymous. Back matter. Proceedings of the American Philosophical Society, 26(129):369–376, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982960.

Anonymous:1889:BMb

[Ano89e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 26(130):572–575, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983201.

Anonymous:1889:BMc

[Ano89f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 27(131):107–213, November 1889. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982881.

Anonymous:1889:CCA

[Ano89g] Anonymous. Commemoration of the Centennial Anniversary of the American Philosophical Society’s Occupation of Its Present Hall, November 21, 1889. Proceedings of the American Philosophical Society, 27(131):3, November 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982874.

Anonymous:1889:FM

[Ano89h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 27(131):1–83, November 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982869.

Anonymous:1889:IEMa

[Ano89i] Anonymous. [Illustration]: East Meeting Room of the American Philosophical Society, looking east. Proceedings of the American Philosophical Society, 27(131):??, November 1889. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982870.

Anonymous:1889:IEMb

[Ano89j] Anonymous. [Illustration]: East Meeting Room of the American Philosophical Society, looking south-east. Proceedings of the American Philosophical Society, 27(131):??, November 1889. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982871.

Anonymous:1889:IWMb

[Ano89k] Anonymous. [Illustration]: West Meeting Room of the American Philosophical Society, looking north-west. Proceedings of the American Philosophical Society, 27(131):??, November 1889. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982873.

Anonymous:1889:IWMa

[Ano89l] Anonymous. [Illustration]: West Meeting Room of the American Philosophical Society, looking south-west. Proceedings of the American Philosophical Society, 27(131):??, November 1889. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982872.

Anonymous:1889:LAP

[Ano89m] Anonymous. The laws of the American Philosophical Society. Proceedings of the American Philosophical Society, 27(131):55–73, November 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982877.

Anonymous:1889:P

[Ano89n] Anonymous. Presidents. Proceedings of the American Philosophical Society, 27(131):85–105, November 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982880.

Anonymous:1889:PD

[Ano89o] Anonymous. Proceedings at the dinner. Proceedings of the American Philosophical Society, 27(131):18–51, November 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982876.

Anonymous:1889:PH

[Ano89p] Anonymous. Proceedings at the Hall. Proceedings of the American Philosophical Society, 27(131):4–18, November 1889. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982875.

Anonymous:1889:RCA

[Ano89q] Anonymous. Report of the Committee Appointed (January 6, 1888) by the American Philosophical Society to Assist the Commission on Amended Orthografy, Created by Virtue of a Resolution of the Legislature of Pennsylvania. Proceedings of the American Philosophical Society, 26(129):306–329, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982954.

Anonymous:1889:ROC

[Ano89r] Anonymous. Rules of order, &c. of the American Philosophical Society. Proceedings of the American Philosophical Society, 27 (131):74–76, November 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982878.

Anonymous:1889:SMA

[Ano89s] Anonymous. Stated meeting, April 5, 1889. Proceedings of the American Philosophical Society, 26(129):305–306, January/July 1889. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982953.

Anonymous:1889:SMDb

[Ano89t] Anonymous. Stated meeting, December 20, 1889. Proceedings of the American Philosophical Society, 26(130):569–571, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983200.

Anonymous:1889:SMDa

[Ano89u] Anonymous. Stated meeting, December 6, 1889. Proceedings of the American Philosophical Society, 26(130):566–568, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983199.

Anonymous:1889:SMFa

[Ano89v] Anonymous. Stated meeting, February 1, 1889. Proceedings of the American Philosophical Society, 26(129):294–297, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982949.

Anonymous:1889:SMFb

[Ano89w] Anonymous. Stated meeting; February 15, 1889. Proceedings of the American Philosophical Society, 26(129):297–301, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982950.

Anonymous:1889:SMJb

[Ano89x] Anonymous. Stated meeting, January 18, 1889. Proceedings of the American Philosophical Society, 26(129):291–294, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982948.

Anonymous:1889:SMJa

[Ano89y] Anonymous. Stated meeting, January 4, 1889. Proceedings of the American Philosophical Society, 26(129):174–176, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982943.

Anonymous:1889:SMMa

[Ano89z] Anonymous. Stated meeting, March 1, 1889. Proceedings of the American Philosophical Society, 26(129):301–303, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982951.

Anonymous:1889:SMMb

[Ano89-27] Anonymous. Stated meeting, March 15, 1889. Proceedings of the American Philosophical Society, 26(129):303–304, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982952.

Anonymous:1889:SMMd

[Ano89-28] Anonymous. Stated meeting, May 17, 1889. Proceedings of the American Philosophical Society, 26(129):355–358, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982957.

Anonymous:1889:SMMc

[Ano89-29] Anonymous. Stated meeting, May 3, 1889. Proceedings of the American Philosophical Society, 26(129):353–355, January/July 1889. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982956.

Anonymous:1889:SMNa

[Ano89-30] Anonymous. Stated meeting, November 1, 1889. Proceedings of the American Philosophical Society, 26(130):531–534, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983193.

Anonymous:1889:SMNb

[Ano89-31] Anonymous. Stated meeting, November 15, 1889. Proceedings of the American Philosophical Society, 26(130):564–566, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983198.

Anonymous:1889:SMOb

[Ano89-32] Anonymous. Stated meeting, October 18, 1889. Proceedings of the American Philosophical Society, 26(130):527–531, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983192.

Anonymous:1889:SMOa

[Ano89-33] Anonymous. Stated meeting, October 4, 1889. Proceedings of the American Philosophical Society, 26(130):489–491, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983188.

Anonymous:1889:SMSb

[Ano89-34] Anonymous. Stated meeting, September 20, 1889. Proceedings of the American Philosophical Society, 26(130):488–489, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983187.

Anonymous:1889:SMSa

[Ano89-35] Anonymous. Stated meeting, September 6, 1889. Proceedings of the American Philosophical Society, 26(130):482–488, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983186.

Anonymous:1889:VIa

[Ano89-36] Anonymous. Volume information. Proceedings of the American Philosophical Society, 26(129):??, January/July 1889. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982939.

Anonymous:1889:VIb

[Ano89-37] Anonymous. Volume information. Proceedings of the American Philosophical Society, 26(130):577–580, July/December 1889. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983180.

Anonymous:1890:Aa

[Ano90a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 28(132):??, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1890:Ab

[Ano90b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 28(133):??, April 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1890:Ac

[Ano90c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 28(134):??, July/December 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1890:BM

[Ano90d] Anonymous. Back matter. Proceedings of the American Philosophical Society, 28(134):1–13, July/December 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983118.

Anonymous:1890:FM

[Ano90e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 28(133):??, April 17, 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982834.

Anonymous:1890:COH

[Ano90f] Anonymous. In commemoration of the one hundredth anniversary of the decease of Benjamin Franklin. April 17, 1890. Proceedings of the American Philosophical Society, 28(133):161–226, April 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982835.

Anonymous:1890:SMAb

[Ano90g] Anonymous. Special meeting, April 25, 1890. Proceedings of the American Philosophical Society, 28(132):102, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983321.

Anonymous:1890:SMAa

[Ano90h] Anonymous. Stated meeting, April 18, 1890. Proceedings of the American Philosophical Society, 28(132):97–101, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983320.

Anonymous:1890:SMDb

[Ano90i] Anonymous. Stated meeting, December 19, 1890. Proceedings of the American Philosophical Society, 28(134):261–263, July/December 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983117.

Anonymous:1890:SMDa

[Ano90j] Anonymous. Stated meeting, December 5, 1890. Proceedings of the American Philosophical Society, 28(134):259–261, July/December 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983116.

Anonymous:1890:SMFb

[Ano90k] Anonymous. Stated meeting, February 21, 1890. Proceedings of the American Philosophical Society, 28(132):88–89, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983317.

Anonymous:1890:SMFa

[Ano90l] Anonymous. Stated meeting, February 7, 1890. Proceedings of the American Philosophical Society, 28(132):32–37, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983306.

Anonymous:1890:SMJb

[Ano90m] Anonymous. Stated meeting, January 17, 1890. Proceedings of the American Philosophical Society, 28(132):28–31, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983305.

Anonymous:1890:SMJa

[Ano90n] Anonymous. Stated meeting, January 3, 1890. Proceedings of the American Philosophical Society, 28(132):26–28, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983304.

Anonymous:1890:SMMb

[Ano90o] Anonymous. Stated meeting, March 21, 1890. Proceedings of the American Philosophical Society, 28(132):92–97, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983319.

Anonymous:1890:SMMa

[Ano90p] Anonymous. Stated meeting, March 7, 1890. Proceedings of the American Philosophical Society, 28(132):90–92, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983318.

Anonymous:1890:SMMd

[Ano90q] Anonymous. Stated meeting, May 16, 1890. Proceedings of the American Philosophical Society, 28(132):105–108, January/June 1890. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983323.

Anonymous:1890:SMMc

[Ano90r] Anonymous. Stated meeting, May 2, 1890. Proceedings of the American Philosophical Society, 28(132):103–104, January/June 1890. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983322.

Anonymous:1890:SMN

[Ano90s] Anonymous. Stated meeting, November 21, 1890. Proceedings of the American Philosophical Society, 28(134):259, July/December 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983115.

Anonymous:1890:VIa

[Ano90t] Anonymous. Volume information. Proceedings of the American Philosophical Society, 28(132):??, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983300.

Anonymous:1890:VIb

[Ano90u] Anonymous. Volume information. Proceedings of the American Philosophical Society, 28(134):265–269, July/December 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983112.

Anonymous:1891:Aa

[Ano91a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 29(135):??, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1891:Ab

[Ano91b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 29(136):??, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1891:AMW

[Ano91c] Anonymous. An adjourned meeting was held may 29, 1891. Proceedings of the American Philosophical Society, 29(135):97–98, January/ June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983032.

Anonymous:1891:SMAb

[Ano91d] Anonymous. Stated meeting, April 17, 1891. Proceedings of the American Philosophical Society, 29(135):91–92, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983029.

Anonymous:1891:SMAa

[Ano91e] Anonymous. Stated meeting, April 3, 1891. Proceedings of the American Philosophical Society, 29(135):91, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983028.

Anonymous:1891:SMDb

[Ano91f] Anonymous. Stated meeting, December 18, 1891. Proceedings of the American Philosophical Society, 29(136):219–222, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982938.

Anonymous:1891:SMDa

[Ano91g] Anonymous. Stated meeting, December 4, 1891. Proceedings of the American Philosophical Society, 29(136):162–165, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982934.

Anonymous:1891:SMFb

[Ano91h] Anonymous. Stated meeting, February 20, 1891. Proceedings of the American Philosophical Society, 29(135):84–85, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983025.

Anonymous:1891:SMFa

[Ano91i] Anonymous. Stated meeting, February 6, 1891. Proceedings of the American Philosophical Society, 29(135):84, January/June 1891. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983024.

Anonymous:1891:SMJb

[Ano91j] Anonymous. Stated meeting, January 16, 1891. Proceedings of the American Philosophical Society, 29(135):81–83, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983023.

Anonymous:1891:SMJa

[Ano91k] Anonymous. Stated meeting, January 2, 1891. Proceedings of the American Philosophical Society, 29(135):78–81, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983022.

Anonymous:1891:SMMb

[Ano91l] Anonymous. Stated meeting, March 20, 1891. Proceedings of the American Philosophical Society, 29(135):88–91, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983027.

Anonymous:1891:SMMa

[Ano91m] Anonymous. Stated meeting, March 6, 1891. Proceedings of the American Philosophical Society, 29(135):85–87, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983026.

Anonymous:1891:SMMc

[Ano91n] Anonymous. Stated meeting, May 1, 1891. Proceedings of the American Philosophical Society, 29(135):93–95, January/June 1891. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983030.

Anonymous:1891:SMMd

[Ano91o] Anonymous. Stated meeting, May 15, 1891. Proceedings of the American Philosophical Society, 29(135):95–97, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983031.

Anonymous:1891:SMNb

[Ano91p] Anonymous. Stated meeting, November 20, 1891. Proceedings of the American Philosophical Society, 29(136):149–162, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982933.

Anonymous:1891:SMNa

[Ano91q] Anonymous. Stated meeting, November 6, 1891. Proceedings of the American Philosophical Society, 29(136):133–137, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982930.

Anonymous:1891:SMOb

[Ano91r] Anonymous. Stated meeting, October 16, 1891. Proceedings of the American Philosophical Society, 29(136):129–132, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982929.

Anonymous:1891:SMOa

[Ano91s] Anonymous. Stated meeting, October 2, 1891. Proceedings of the American Philosophical Society, 29(136):127–129, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982928.

Anonymous:1891:SMSb

[Ano91t] Anonymous. Stated meeting, September 18, 1891. Proceedings of the American Philosophical Society, 29(136):123–127, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982927.

Anonymous:1891:SMSa

[Ano91u] Anonymous. Stated meeting, September 4, 1891. Proceedings of the American Philosophical Society, 29(136):121–123, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982926.

Anonymous:1891:VIa

[Ano91v] Anonymous. Volume information. Proceedings of the American Philosophical Society, 29(135):??, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983009.

Anonymous:1891:VIb

[Ano91w] Anonymous. Volume information. Proceedings of the American Philosophical Society, 29(136):223–226, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982923.

Anonymous:1892:Aa

[Ano92a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 30(137):??, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1892:Ab

[Ano92b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 30(138):??, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1892:Ac

[Ano92c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 30(139):??, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1892:FM

[Ano92d] Anonymous. Front matter. Proceedings of the American Philosophical Society, 30(137):1–12, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983203.

Anonymous:1892:SMA

[Ano92e] Anonymous. Stated meeting, April 1, 1892. Proceedings of the American Philosophical Society, 30(138):254–256, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983177.

Anonymous:1892:SMDb

[Ano92f] Anonymous. Stated meeting, December 16, 1892. Proceedings of the American Philosophical Society, 30(139):313–316, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983354.

Anonymous:1892:SMDa

[Ano92g] Anonymous. Stated meeting, December 2, 1892. Proceedings of the American Philosophical Society, 30(139):312–313, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983353.

Anonymous:1892:SMFb

[Ano92h] Anonymous. Stated meeting, February 19, 1892. Proceedings of the American Philosophical Society, 30(137):112–117, January 1892.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983212.

Anonymous:1892:SMFa

[Ano92i] Anonymous. Stated meeting, February 5, 1892. Proceedings of the American Philosophical Society, 30(137):110–112, January 1892. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983211.

Anonymous:1892:SMJa

[Ano92j] Anonymous. Stated meeting, January 1, 1892. Proceedings of the American Philosophical Society, 30(137):1–3, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983204.

Anonymous:1892:SMJb

[Ano92k] Anonymous. Stated meeting, January 15, 1892. Proceedings of the American Philosophical Society, 30(137):106–109, January 1892. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983210.

Anonymous:1892:SMMb

[Ano92l] Anonymous. Stated meeting, March 18, 1892. Proceedings of the American Philosophical Society, 30(137):120–122, January 1892. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983214.

Anonymous:1892:SMMa

[Ano92m] Anonymous. Stated meeting, March 4, 1892. Proceedings of the American Philosophical Society, 30(137):117–120, January 1892. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983213.

Anonymous:1892:SMMd

[Ano92n] Anonymous. Stated meeting, May 20, 1892. Proceedings of the American Philosophical Society, 30(138):263–268, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983179.

Anonymous:1892:SMMc

[Ano92o] Anonymous. Stated meeting, May 6, 1892. Proceedings of the American Philosophical Society, 30(138):256–262, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983178.

Anonymous:1892:SMNb

[Ano92p] Anonymous. Stated meeting, November 18, 1892. Proceedings of the American Philosophical Society, 30(139):309–312, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983352.

Anonymous:1892:SMNa

[Ano92q] Anonymous. Stated meeting, November 4, 1892. Proceedings of the American Philosophical Society, 30(139):307–309, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983351.

Anonymous:1892:SMO

[Ano92r] Anonymous. Stated meeting, October 7, 1892. Proceedings of the American Philosophical Society, 30(139):303–307, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983350.

Anonymous:1892:SMSb

[Ano92s] Anonymous. Stated meeting, September 16, 1892. Proceedings of the American Philosophical Society, 30(139):300–303, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983349.

Anonymous:1892:SMSa

[Ano92t] Anonymous. Stated meeting, September 2, 1892. Proceedings of the American Philosophical Society, 30(139):291–300, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983348.

Anonymous:1892:VIa

[Ano92u] Anonymous. Volume information. Proceedings of the American Philosophical Society, 30(137):??, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983202.

Anonymous:1892:VIb

[Ano92v] Anonymous. Volume information. Proceedings of the American Philosophical Society, 30(139):325–328, December 1892. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983342.

Anonymous:1893:Aa

[Ano93a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 31(140):??, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1893:Ab

[Ano93b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 31(141):??, April/June 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1893:Ac

[Ano93c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 31(142):??, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1893:BM

[Ano93d] Anonymous. Back matter. Proceedings of the American Philosophical Society, 31(142):365, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982978.

Anonymous:1893:FM

[Ano93e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 31(140):1–12, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983099.

Anonymous:1893:SMAb

[Ano93f] Anonymous. Stated meeting, April 21, 1893. Proceedings of the American Philosophical Society, 31(141):137–138, April/June 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982827.

Anonymous:1893:SMAa

[Ano93g] Anonymous. Stated meeting, April 7, 1893. Proceedings of the American Philosophical Society, 31(141):135–137, April/June 1893. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982826.

Anonymous:1893:SMDa

[Ano93h] Anonymous. Stated meeting, December 1, 1893. Proceedings of the American Philosophical Society, 31(142):355, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982975.

Anonymous:1893:SMDb

[Ano93i] Anonymous. Stated meeting, December 15, 1893. Proceedings of the American Philosophical Society, 31(142):356–358, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982976.

Anonymous:1893:SMFb

[Ano93j] Anonymous. Stated meeting, February 17, 1893. Proceedings of the American Philosophical Society, 31(140):14–16, January 1893. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983103.

Anonymous:1893:SMFa

[Ano93k] Anonymous. Stated meeting, February 3, 1893. Proceedings of the American Philosophical Society, 31(140):9–14, January 1893. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983102.

Anonymous:1893:SMJb

[Ano93l] Anonymous. Stated meeting, January 20, 1893. Proceedings of the American Philosophical Society, 31(140):5–9, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983101.

Anonymous:1893:SMJa

[Ano93m] Anonymous. Stated meeting, January 6, 1893. Proceedings of the American Philosophical Society, 31(140):1–4, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983100.

Anonymous:1893:SMMb

[Ano93n] Anonymous. Stated meeting, March 17, 1893. Proceedings of the American Philosophical Society, 31(140):130–133, January 1893. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983111.

Anonymous:1893:SMMa

[Ano93o] Anonymous. Stated meeting, March 3, 1893. Proceedings of the American Philosophical Society, 31(140):128–130, January 1893. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983110.

Anonymous:1893:SMMd

[Ano93p] Anonymous. Stated meeting, May 19, 1893. Proceedings of the American Philosophical Society, 31(141):228–231, April/June 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982833.

Anonymous:1893:SMMc

[Ano93q] Anonymous. Stated meeting, May 5, 1893. Proceedings of the American Philosophical Society, 31(141):226–227, April/June 1893. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982832.

Anonymous:1893:SMNb

[Ano93r] Anonymous. Stated meeting, November 17, 1893. Proceedings of the American Philosophical Society, 31(142):351–354, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982974.

Anonymous:1893:SMNa

[Ano93s] Anonymous. Stated meeting, November 3, 1893. Proceedings of the American Philosophical Society, 31(142):348–350, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982973.

Anonymous:1893:SMOb

[Ano93t] Anonymous. Stated meeting, October 20, 1893. Proceedings of the American Philosophical Society, 31(142):252–254, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982965.

Anonymous:1893:SMOa

[Ano93u] Anonymous. Stated meeting, October 6, 1893. Proceedings of the American Philosophical Society, 31(142):247–252, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982964.

Anonymous:1893:SMSa

[Ano93v] Anonymous. Stated meeting, September 1, 1893. Proceedings of the American Philosophical Society, 31(142):233–242, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982962.

Anonymous:1893:SMSb

[Ano93w] Anonymous. Stated meeting, September 15, 1893. Proceedings of the American Philosophical Society, 31(142):242–247, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982963.

Anonymous:1893:VIa

[Ano93x] Anonymous. Volume information. Proceedings of the American Philosophical Society, 31(140):??, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983098.

Anonymous:1893:VIb

[Ano93y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 31(142):367–372, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982961.

Anonymous:1894:Aa

[Ano94a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 32(143):??, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1894:Ab

[Ano94b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 33(144):??, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1894:Ac

[Ano94c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 33(145):??, June 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1894:Ad

[Ano94d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 33(146):??, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1894:CWA

[Ano94e] Anonymous. Catalogue of works on atmospheric physics and climatology. Proceedings of the American Philosophical Society, 32 (143):248–269, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983038.

Anonymous:1894:COH

[Ano94f] Anonymous. Commemoration of the one hundred and fiftieth anniversary of the foundation of the American Philosophical Society Held at Philadelphia for the Promotion of Useful Knowledge, May 22d to 26th, 1893. Proceedings of the American Philosophical Society, 32(143):5–6, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983034.

Anonymous:1894:FMa

[Ano94g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 32(143):1–161, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983033.

Anonymous:1894:FMb

[Ano94h] Anonymous. Front matter. Proceedings of the American Philosophical Society, 33(144):1–12, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983357.

Anonymous:1894:OHF

[Ano94i] Anonymous. One hundred and fiftieth anniversary of the foundation of the American Philosophical Society [meeting minutes]. Proceedings of the American Philosophical Society, 32(143):7–160, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983035.

Anonymous:1894:SMAb

[Ano94j] Anonymous. Stated meeting, April 20, 1894. Proceedings of the American Philosophical Society, 33(145):160–162, June 1894. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983263.

Anonymous:1894:SMAa

[Ano94k] Anonymous. Stated meeting, April 6, 1894. Proceedings of the American Philosophical Society, 33(145):159–160, June 1894. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983262.

Anonymous:1894:SMDb

[Ano94l] Anonymous. Stated meeting, December 21, 1894. Proceedings of the American Philosophical Society, 33(146):370–372, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982851.

Anonymous:1894:SMDa

[Ano94m] Anonymous. Stated meeting, December 7, 1894. Proceedings of the American Philosophical Society, 33(146):367–370, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982850.

Anonymous:1894:SMFb

[Ano94n] Anonymous. Stated meeting, February 16, 1894. Proceedings of the American Philosophical Society, 33(144):146–149, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983369.

Anonymous:1894:SMFa

[Ano94o] Anonymous. Stated meeting, February 2, 1894. Proceedings of the American Philosophical Society, 33(144):144–146, January 1894. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983368.

Anonymous:1894:SMJb

[Ano94p] Anonymous. Stated meeting, January 19, 1894. Proceedings of the American Philosophical Society, 33(144):141–144, January 1894. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983367.

Anonymous:1894:SMJa

[Ano94q] Anonymous. Stated meeting, January 5, 1894. Proceedings of the American Philosophical Society, 33(144):1–5, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983358.

Anonymous:1894:SMMb

[Ano94r] Anonymous. Stated meeting, March 16, 1894. Proceedings of the American Philosophical Society, 33(144):155–157, January 1894. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983371.

Anonymous:1894:SMMa

[Ano94s] Anonymous. Stated meeting, March 2, 1894. Proceedings of the American Philosophical Society, 33(144):150–155, January 1894. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983370.

Anonymous:1894:SMMd

[Ano94t] Anonymous. Stated meeting, May 18, 1894. Proceedings of the American Philosophical Society, 33(145):256–259, June 1894. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983275.

Anonymous:1894:SMMc

[Ano94u] Anonymous. Stated meeting, May 4, 1894. Proceedings of the American Philosophical Society, 33(145):253–256, June 1894. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983274.

Anonymous:1894:SMNb

[Ano94v] Anonymous. Stated meeting, November 16, 1894. Proceedings of the American Philosophical Society, 33(146):365–367, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982849.

Anonymous:1894:SMNa

[Ano94w] Anonymous. Stated meeting, November 2, 1894. Proceedings of the American Philosophical Society, 33(146):362–365, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982848.

Anonymous:1894:SMOb

[Ano94x] Anonymous. Stated meeting, October 19, 1894. Proceedings of the American Philosophical Society, 33(146):276–281, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982840.

Anonymous:1894:SMOa

[Ano94y] Anonymous. Stated meeting, October 5, 1894. Proceedings of the American Philosophical Society, 33(146):268–276, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982839.

Anonymous:1894:SMSb

[Ano94z] Anonymous. Stated meeting, September 21, 1894. Proceedings of the American Philosophical Society, 33(146):262–268, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982838.

Anonymous:1894:SMSa

[Ano94-27] Anonymous. Stated meeting, September 7, 1894. Proceedings of the American Philosophical Society, 33(146):261, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982837.

Anonymous:1894:VIa

[Ano94-28] Anonymous. Volume information. Proceedings of the American Philosophical Society, 33(144):??, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983356.

Anonymous:1894:VIb

[Ano94-29] Anonymous. Volume information. Proceedings of the American Philosophical Society, 33(146):373–377, July/December 1894. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982836.

Anonymous:1895:Aa

[Ano95a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 34(147):??, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1895:Ab

[Ano95b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 34(148):??, July 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1895:Ac

[Ano95c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 34(149):??, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1895:HMP

[Ano95d] Anonymous. Henry M. Phillips Prize essays. Proceedings of the American Philosophical Society, 34(148):173–179, July 1895. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983254.

Anonymous:1895:SMAb

[Ano95e] Anonymous. Stated meeting, April 19, 1895. Proceedings of the American Philosophical Society, 34(147):165–168, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983167.

Anonymous:1895:SMAa

[Ano95f] Anonymous. Stated meeting, April 5, 1895. Proceedings of the American Philosophical Society, 34(147):160–165, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983166.

Anonymous:1895:SMDb

[Ano95g] Anonymous. Stated meeting, December 20, 1895. Proceedings of the American Philosophical Society, 34(149):488–491, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982908.

Anonymous:1895:SMDa

[Ano95h] Anonymous. Stated meeting, December 6, 1895. Proceedings of the American Philosophical Society, 34(149):484–488, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982907.

Anonymous:1895:SMFa

[Ano95i] Anonymous. Stated meeting, February 1, 1895. Proceedings of the American Philosophical Society, 34(147):6–8, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983153.

Anonymous:1895:SMFb

[Ano95j] Anonymous. Stated meeting, February 15, 1895. Proceedings of the American Philosophical Society, 34(147):8–11, January 1895. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983154.

Anonymous:1895:SMJb

[Ano95k] Anonymous. Stated meeting, January 18, 1895. Proceedings of the American Philosophical Society, 34(147):3–6, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983152.

Anonymous:1895:SMJa

[Ano95l] Anonymous. Stated meeting, January 4, 1895. Proceedings of the American Philosophical Society, 34(147):1–3, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983151.

Anonymous:1895:SMMa

[Ano95m] Anonymous. Stated meeting, March 1, 1895. Proceedings of the American Philosophical Society, 34(147):155–156, January 1895. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983164.

Anonymous:1895:SMMb

[Ano95n] Anonymous. Stated meeting, March 15, 1895. Proceedings of the American Philosophical Society, 34(147):157–160, January 1895. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983165.

Anonymous:1895:SMMd

[Ano95o] Anonymous. Stated meeting, May 17, 1895. Proceedings of the American Philosophical Society, 34(149):335–339, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982883.

Anonymous:1895:SMMc

[Ano95p] Anonymous. Stated meeting, May 3, 1895. Proceedings of the American Philosophical Society, 34(147):168–171, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983168.

Anonymous:1895:SMNa

[Ano95q] Anonymous. Stated meeting, November 1, 1895. Proceedings of the American Philosophical Society, 34(149):477–479, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982905.

Anonymous:1895:SMNb

[Ano95r] Anonymous. Stated meeting, November 15, 1895. Proceedings of the American Philosophical Society, 34(149):479–484, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982906.

Anonymous:1895:SMOb

[Ano95s] Anonymous. Stated meeting, October 18, 1895. Proceedings of the American Philosophical Society, 34(149):353–354, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982887.

Anonymous:1895:SMOa

[Ano95t] Anonymous. Stated meeting, October 4, 1895. Proceedings of the American Philosophical Society, 34(149):351–352, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982886.

Anonymous:1895:SMSb

[Ano95u] Anonymous. Stated meeting, September 20, 1895. Proceedings of the American Philosophical Society, 34(149):348–351, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982885.

Anonymous:1895:SMSa

[Ano95v] Anonymous. Stated meeting, September 6, 1895. Proceedings of the American Philosophical Society, 34(149):340–348, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982884.

Anonymous:1895:VIa

[Ano95w] Anonymous. Volume information. Proceedings of the American Philosophical Society, 34(147):??, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983150.

Anonymous:1895:VIb

[Ano95x] Anonymous. Volume information. Proceedings of the American Philosophical Society, 34(149):493–498, December 1895. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982882.

Anonymous:1896:Aa

[Ano96a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 35(150):??, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1896:Ab

[Ano96b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 35(151):??, August 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1896:Ac

[Ano96c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 35(152):??, November 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1896:Ad

[Ano96d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 35(153):??, December 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1896:AMF

[Ano96e] Anonymous. Adjourned meeting, February 28, 1896. Proceedings of the American Philosophical Society, 35(150):13–16, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982857.

Anonymous:1896:RMD

[Ano96f] Anonymous. Remarks made at the demonstration of the Röntgen ray, at stated meeting, February 21, 1896. Proceedings of the American Philosophical Society, 35(150):17–36, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982858.

Anonymous:1896:SMAa

[Ano96g] Anonymous. Stated meeting, April 10, 1896. Proceedings of the American Philosophical Society, 35(150):68–71, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982863.

Anonymous:1896:SMAb

[Ano96h] Anonymous. Stated meeting, April 17, 1896. Proceedings of the American Philosophical Society, 35(150):74–76, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982865.

Anonymous:1896:SMDb

[Ano96i] Anonymous. Stated meeting, December 18, 1896. Proceedings of the American Philosophical Society, 35(153):307–309, December 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983591.

Anonymous:1896:SMDa

[Ano96j] Anonymous. Stated meeting, December 4, 1896. Proceedings of the American Philosophical Society, 35(153):302–304, December 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983589.

Anonymous:1896:SMFb

[Ano96k] Anonymous. Stated meeting, February 21, 1896. Proceedings of the American Philosophical Society, 35(150):12, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982856.

Anonymous:1896:SMFa

[Ano96l] Anonymous. Stated meeting, February 7, 1896. Proceedings of the American Philosophical Society, 35(150):6–11, January 1896. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982855.

Anonymous:1896:SMJb

[Ano96m] Anonymous. Stated meeting, January 17, 1896. Proceedings of the American Philosophical Society, 35(150):3–6, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982854.

Anonymous:1896:SMJa

[Ano96n] Anonymous. Stated meeting, January 3, 1896. Proceedings of the American Philosophical Society, 35(150):1–3, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982853.

Anonymous:1896:SMMb

[Ano96o] Anonymous. Stated meeting, March 20, 1896. Proceedings of the American Philosophical Society, 35(150):65–68, January 1896. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982862.

Anonymous:1896:SMMa

[Ano96p] Anonymous. Stated meeting, March 6, 1896. Proceedings of the American Philosophical Society, 35(150):36–39, January 1896. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982859.

Anonymous:1896:SMMc

[Ano96q] Anonymous. Stated meeting, May 1, 1896. Proceedings of the American Philosophical Society, 35(150):76–77, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982866.

Anonymous:1896:SMMd

[Ano96r] Anonymous. Stated meeting, May 15, 1896. Proceedings of the American Philosophical Society, 35(151):115–119, August 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983256.

Anonymous:1896:SMNb

[Ano96s] Anonymous. Stated meeting, November 20, 1896. Proceedings of the American Philosophical Society, 35(153):296–298, December 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983587.

Anonymous:1896:SMNa

[Ano96t] Anonymous. Stated meeting, November 6, 1896. Proceedings of the American Philosophical Society, 35(153):294–296, December 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983586.

Anonymous:1896:SMOb

[Ano96u] Anonymous. Stated meeting, October 16, 1896. Proceedings of the American Philosophical Society, 35(153):291–293, December 1896.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983585.

Anonymous:1896:SMOa

[Ano96v] Anonymous. Stated meeting, October 2, 1896. Proceedings of the American Philosophical Society, 35(152):286–290, November 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983149.

Anonymous:1896:SMSb

[Ano96w] Anonymous. Stated meeting, September 18, 1896. Proceedings of the American Philosophical Society, 35(152):200–202, November 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983145.

Anonymous:1896:SMSa

[Ano96x] Anonymous. Stated meeting, September 4, 1896. Proceedings of the American Philosophical Society, 35(152):193–200, November 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983144.

Anonymous:1896:VIa

[Ano96y] Anonymous. Volume information. Proceedings of the American Philosophical Society, 35(150):??, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982852.

Anonymous:1896:VIb

[Ano96z] Anonymous. Volume information. Proceedings of the American Philosophical Society, 35(153):449–453, December 1896. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983584.

Anonymous:1897:Aa

[Ano97a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 36(154):??, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1897:Ab

[Ano97b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 36(155):??, May 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1897:Ac

[Ano97c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 36(156):??, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1897:ASMa

[Ano97d] Anonymous. Adjourned stated meeting, May 13, 1897. Proceedings of the American Philosophical Society, 36(155):216, May 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983656.

Anonymous:1897:ASMb

[Ano97e] Anonymous. Adjourned stated meeting of November 5, 1897, November 10, 1897. Proceedings of the American Philosophical Society, 36(156):465, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983400.

Anonymous:1897:CAR

[Ano97f] Anonymous. [Corrections]: Australian rock carvings. Proceedings of the American Philosophical Society, 36(156):478, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983403. See [Mat97b].

Anonymous:1897:FM

[Ano97g] Anonymous. Front matter. Proceedings of the American Philosophical Society, 36(154):iii–xv, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983455.

Anonymous:1897:SMAb

[Ano97h] Anonymous. Special meeting, April 23, 1897. Proceedings of the American Philosophical Society, 36(155):178, May 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983651.

Anonymous:1897:SMOb

[Ano97i] Anonymous. Special meeting, October 11, 1897. Proceedings of the American Philosophical Society, 36(156):434, December 1897. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983393.

Anonymous:1897:SMOd

[Ano97j] Anonymous. Special meeting, October 29, 1897. Proceedings of the American Philosophical Society, 36(156):442, December 1897. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983397.

Anonymous:1897:SMAa

[Ano97k] Anonymous. Stated meeting, April 2, 1897. Proceedings of the American Philosophical Society, 36(155):175–178, May 1897. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983650.

Anonymous:1897:SMDb

[Ano97l] Anonymous. Stated meeting, December 17, 1897. Proceedings of the American Philosophical Society, 36(156):499, December 1897. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983407.

Anonymous:1897:SMDa

[Ano97m] Anonymous. Stated meeting, December 3, 1897. Proceedings of the American Philosophical Society, 36(156):478–479, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983404.

Anonymous:1897:SMFb

[Ano97n] Anonymous. Stated meeting, February 19, 1897. Proceedings of the American Philosophical Society, 36(155):141–142, May 1897. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983646.

Anonymous:1897:SMFa

[Ano97o] Anonymous. Stated meeting, February 5, 1897. Proceedings of the American Philosophical Society, 36(154):92–93, January 1897. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983461.

Anonymous:1897:SMJa

[Ano97p] Anonymous. Stated meeting, January 1, 1897. Proceedings of the American Philosophical Society, 36(154):1–2, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983456.

Anonymous:1897:SMJb

[Ano97q] Anonymous. Stated meeting, January 15, 1897. Proceedings of the American Philosophical Society, 36(154):2–5, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983457.

Anonymous:1897:SMMb

[Ano97r] Anonymous. Stated meeting, March 19, 1897. Proceedings of the American Philosophical Society, 36(155):143, May 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983648.

Anonymous:1897:SMMa

[Ano97s] Anonymous. Stated meeting, March 5, 1897. Proceedings of the American Philosophical Society, 36(155):142, May 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983647.

Anonymous:1897:SMMd

[Ano97t] Anonymous. Stated meeting, May 21, 1897. Proceedings of the American Philosophical Society, 36(155):319–320, May 1897. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983658.

Anonymous:1897:SMMc

[Ano97u] Anonymous. Stated meeting, May 7, 1897. Proceedings of the American Philosophical Society, 36(155):192–195, May 1897. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983653.

Anonymous:1897:SMNb

[Ano97v] Anonymous. Stated meeting, November 19, 1897. Proceedings of the American Philosophical Society, 36(156):465–466, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983401.

Anonymous:1897:SMNa

[Ano97w] Anonymous. Stated meeting, November 5, 1897. Proceedings of the American Philosophical Society, 36(156):464, December 1897. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983399.

Anonymous:1897:SMOa

[Ano97x] Anonymous. Stated meeting, October 1, 1897. Proceedings of the American Philosophical Society, 36(156):433–434, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983392.

Anonymous:1897:SMOc

[Ano97y] Anonymous. Stated meeting, October 15, 1897. Proceedings of the American Philosophical Society, 36(156):440–441, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983396.

Anonymous:1897:SMSb

[Ano97z] Anonymous. Stated meeting, September 17, 1897. Proceedings of the American Philosophical Society, 36(156):420, December 1897. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983389.

Anonymous:1897:SMSa

[Ano97-27] Anonymous. Stated meeting, September 3, 1897. Proceedings of the American Philosophical Society, 36(156):325–326, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983387.

Anonymous:1897:VIa

[Ano97-28] Anonymous. Volume information. Proceedings of the American Philosophical Society, 36(154):i, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983454.

Anonymous:1897:VIb

[Ano97-29] Anonymous. Volume information. Proceedings of the American Philosophical Society, 36(156):501–504, December 1897. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983386.

Anonymous:1898:Aa

[Ano98a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 37(157):??, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1898:Ab

[Ano98b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 37(158):??, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1898:AMM

[Ano98c] Anonymous. Adjourned meeting, May 13, 1898. Proceedings of the American Philosophical Society, 37(157):155–162, January 1898. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983705.

Anonymous:1898:AMN

[Ano98d] Anonymous. Adjourned meeting, November 25, 1898. Proceedings of the American Philosophical Society, 37(158):336, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983860.

Anonymous:1898:SMAa

[Ano98e] Anonymous. Stated meeting, April 1, 1898. Proceedings of the American Philosophical Society, 37(157):81–84, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983696.

Anonymous:1898:SMAb

[Ano98f] Anonymous. Stated meeting, April 15, 1898. Proceedings of the American Philosophical Society, 37(157):142–143, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983701.

Anonymous:1898:SMDb

[Ano98g] Anonymous. Stated meeting, December 16, 1898. Proceedings of the American Philosophical Society, 37(158):337–338, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983862.

Anonymous:1898:SMDa

[Ano98h] Anonymous. Stated meeting, December 2, 1898. Proceedings of the American Philosophical Society, 37(158):337, December 1898. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983861.

Anonymous:1898:SMFb

[Ano98i] Anonymous. Stated meeting, February 18, 1898. Proceedings of the American Philosophical Society, 37(157):52–53, January 1898. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983691.

Anonymous:1898:SMFa

[Ano98j] Anonymous. Stated meeting, February 4, 1898. Proceedings of the American Philosophical Society, 37(157):51–52, January 1898. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983690.

Anonymous:1898:SMJb

[Ano98k] Anonymous. Stated meeting, January 21, 1898. Proceedings of the American Philosophical Society, 37(157):7–8, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983686.

Anonymous:1898:SMJa

[Ano98l] Anonymous. Stated meeting, January 7, 1898. Proceedings of the American Philosophical Society, 37(157):3–4, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983684.

Anonymous:1898:SMMb

[Ano98m] Anonymous. Stated meeting, March 18, 1898. Proceedings of the American Philosophical Society, 37(157):54, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983693.

Anonymous:1898:SMMa

[Ano98n] Anonymous. Stated meeting, March 4, 1898. Proceedings of the American Philosophical Society, 37(157):53, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983692.

Anonymous:1898:SMMd

[Ano98o] Anonymous. Stated meeting, May 20, 1898. Proceedings of the American Philosophical Society, 37(157):162–170, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983706.

Anonymous:1898:SMMc

[Ano98p] Anonymous. Stated meeting, May 6, 1898. Proceedings of the American Philosophical Society, 37(157):154, January 1898. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983704.

Anonymous:1898:SMNb

[Ano98q] Anonymous. Stated meeting, November 18, 1898. Proceedings of the American Philosophical Society, 37(158):327, December 1898. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983858.

Anonymous:1898:SMNa

[Ano98r] Anonymous. Stated meeting, November 4, 1898. Proceedings of the American Philosophical Society, 37(158):324–326, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983857.

Anonymous:1898:SMOb

[Ano98s] Anonymous. Stated meeting, October 21, 1898. Proceedings of the American Philosophical Society, 37(158):319–320, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983855.

Anonymous:1898:SMOa

[Ano98t] Anonymous. Stated meeting, October 7, 1898. Proceedings of the American Philosophical Society, 37(158):175–178, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983851.

Anonymous:1898:VIa

[Ano98u] Anonymous. Volume information. Proceedings of the American Philosophical Society, 37(157):1, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983683.

Anonymous:1898:VIb

[Ano98v] Anonymous. Volume information. Proceedings of the American Philosophical Society, 37(158):339–340, December 1898. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983850.

Anonymous:1899:Aa

[Ano99a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 38(159):??, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1899:Ab

[Ano99b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 38(160):??, October 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:1899:CCB

[Ano99c] Anonymous. Calendar of the correspondence of Brigadier-General George Weedon, U.S.A., with celebrated characters of the American Revolution. In the Library of the American Philosophical Society. Proceedings of the American Philosophical Society, 38(159):81–114, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983627.

Anonymous:1899:CCR

[Ano99d] Anonymous. Calendar of the correspondence of Richard Henry Lee and Arthur Lee. In the library of the American Philosophical Society. Proceedings of the American Philosophical Society, 38(159):114–131, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983628.

Anonymous:1899:FM

[Ano99e] Anonymous. Front matter. Proceedings of the American Philosophical Society, 38(159):iii–xvi, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983611.

Anonymous:1899:SMAb

[Ano99f] Anonymous. Stated meeting, April 21, 1899. Proceedings of the American Philosophical Society, 38(159):71–74, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983623.

Anonymous:1899:SMAa

[Ano99g] Anonymous. Stated meeting, April 7, 1899. Proceedings of the American Philosophical Society, 38(159):24–25, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983620.

Anonymous:1899:SMDa

[Ano99h] Anonymous. Stated meeting, December 1, 1899. Proceedings of the American Philosophical Society, 38(160):189, October 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983569.

Anonymous:1899:SMDb

[Ano99i] Anonymous. Stated meeting, December 15, 1899. Proceedings of the American Philosophical Society, 38(160):190–191, October 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983570.

Anonymous:1899:SMFb

[Ano99j] Anonymous. Stated meeting, February 17, 1899. Proceedings of the American Philosophical Society, 38(159):22–23, January 1899. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983617.

Anonymous:1899:SMFa

[Ano99k] Anonymous. Stated meeting, February 3, 1899. Proceedings of the American Philosophical Society, 38(159):21–22, January 1899. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983616.

Anonymous:1899:SMJb

[Ano99l] Anonymous. Stated meeting, January 20, 1899. Proceedings of the American Philosophical Society, 38(159):7, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983614.

Anonymous:1899:SMJa

[Ano99m] Anonymous. Stated meeting, January 6, 1899. Proceedings of the American Philosophical Society, 38(159):3–4, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983612.

Anonymous:1899:SMMb

[Ano99n] Anonymous. Stated meeting, March 17, 1899. Proceedings of the American Philosophical Society, 38(159):23–24, January 1899. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983619.

Anonymous:1899:SMMa

[Ano99o] Anonymous. Stated meeting, March 3, 1899. Proceedings of the American Philosophical Society, 38(159):23, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983618.

Anonymous:1899:SMMd

[Ano99p] Anonymous. Stated meeting, May 19, 1899. Proceedings of the American Philosophical Society, 38(159):131–132, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983629.

Anonymous:1899:SMMc

[Ano99q] Anonymous. Stated meeting, May 5, 1899. Proceedings of the American Philosophical Society, 38(159):74–75, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983624.

Anonymous:1899:SMNb

[Ano99r] Anonymous. Stated meeting, November 17, 1899. Proceedings of the American Philosophical Society, 38(160):156, October 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983566.

Anonymous:1899:SMNa

[Ano99s] Anonymous. Stated meeting, November 3, 1899. Proceedings of the American Philosophical Society, 38(160):155, October 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983565.

Anonymous:1899:SMOb

[Ano99t] Anonymous. Stated meeting, October 20, 1899. Proceedings of the American Philosophical Society, 38(160):154–155, October 1899. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983564.

Anonymous:1899:SMOa

[Ano99u] Anonymous. Stated meeting, October 6, 1899. Proceedings of the American Philosophical Society, 38(160):133–134, October 1899. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983561.

Anonymous:1899:VIa

[Ano99v] Anonymous. Volume information. Proceedings of the American Philosophical Society, 38(159):i, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983610.

Anonymous:1899:VIb

[Ano99w] Anonymous. Volume information. Proceedings of the American Philosophical Society, 38(160):193–196, October 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983560.

Anthon:1876:SLF

[Ant76] Charles E. Anthon. Silver louis of fifteen sous, struck under Louis XIV, for circulation in French America. Proceedings of the American Philosophical Society, 16(98):293–298, June/December 1876. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982815.

Ashburner:1877:MSP

[Ash77a] Charles A. Ashburner. A measured section of the Paleozoic rocks of Central Pennsylvania, from the top of the Allegheny River Coal Series, down to the Trenton Limestone in the Lower, or Cambro– Silurian System. Proceedings of the American Philosophical Society, 16(99):519–560, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982479.

Ashburner:1877:DWS

[Ash77b] Chas. A. Ashburner. Description of the Wilcox spouting waterwell. Proceedings of the American Philosophical Society, 17(100): 127–135, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982288.

Ashburner:1878:OWR

[Ash78] Chas. A. Ashburner. Oil well records in McKean and Elk Counties, Pennsylvania. Proceedings of the American Philosophical Society, 18 (102):9–26, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982750.

Ashburner:1880:CBO

[Ash80] Chas. A. Ashburner. On the constitution of the Bradford oil sand. Proceedings of the American Philosophical Society, 18(105): 419–422, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982774.

Ashburner:1881:GSS

[Ash81] Charles A. Ashburner. Geological section at St. Mary’s, Elk County, Pennsylvania. Proceedings of the American Philosophical Society, 19 (108):337–348, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982256.

Ashburner:1884:NNB

[Ash84] Charles A. Ashburner. Notes on the Natural Bridge of Virginia. Proceedings of the American Philosophical Society, 21(116):699–700, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982354.

Ashley:1897:GPA

[Ash97] George H. Ashley. Geology of the Paleozoic area of Arkansas south of the Novaculite region. Proceedings of the American Philosophical Society, 36(155):217–318, May 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983657.

Bache:1891:FOU

[Bac91a] R. Meade Bache. A fragment of objectionable university-extension teaching. Proceedings of the American Philosophical Society, 29 (135):50–59, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983020.

Bache:1891:PSC

[Bac91b] R. Meade Bache. Possible sterilization of city water. Proceedings of the American Philosophical Society, 29(135):26–39, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983016.

Bache:1892:CMP

[Bac92] R. Meade Bache. Civil and military photogrammetry. Proceedings of the American Philosophical Society, 30(138):229–240, April 1892.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983172.

Bache:1893:COR

[Bac93] R. Meade Bache. The conservation of osmazome in roasting. Proceedings of the American Philosophical Society, 31(142):318–325, July/ December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982970.

Bache:1894:DB

[Bac94a] R. Meade Bache. The dynamics of boxing. Proceedings of the American Philosophical Society, 33(145):179–187, June 1894. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983266.

Bache:1894:SBM

[Bac94b] R. Meade Bache. The secret of the Brownian movements. Proceedings of the American Philosophical Society, 33(145):163–177, June 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983264.

Baird:1891:CTH

[Bai91] Henry Carey Baird. Carey and two of his recent critics, Eugen V. Böhm-Bawerk and Alfred Marshall. Proceedings of the American Philosophical Society, 29(136):166–173, July/December 1891. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982935.

Baird:1894:ADN

[Bai94] Henry Carey Baird. Association: The dominating need of man and the keynote of social science. Proceedings of the American Philosophical Society, 33(144):134–140, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983366.

Bailey:1896:FOE

[Bai96] L. H. Bailey. The factors of organic evolution from a botanical standpoint. Proceedings of the American Philosophical Society, 35(150):88–114, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982868.

Barker:1875:CPL

[Bar75] George F. Barker. Contributions from the Physical Laboratory of the University of Pennsylvania. No. I. A new vertical-lantern galvanometer. Proceedings of the American Philosophical Society, 14 (94):440–445, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981883.

Barrett:1877:LEA

[Bar77] O. Barrett, Jr. List of elevations, above mean tide, through the County of Indiana, in Pennsylvania, copied from notes of survey under Gen. H. Haupt, for the Sea Board Pipe Line. Proceedings of the American Philosophical Society, 17(100):145–148, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982290.

Barker:1878:TSE

[Bar78] George F. Barker. On the total solar eclipse of July 29th, 1878. Proceedings of the American Philosophical Society, 18(102):103– 114, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982766.

Barker:1883:HD

[Bar83a] Geo. F. Barker. Henry Draper. Proceedings of the American Philosophical Society, 20(113):656–662, January/April 1883. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982460.

Barker:1883:MEF

[Bar83b] George F. Barker. On the measurement of electromotive force. Proceedings of the American Philosophical Society, 20(113):649–655, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982459.

Barker:1887:HDM

[Bar87] George F. Barker. On the Henry Draper Memorial Photographs of stellar spectra. Proceedings of the American Philosophical Society, 24 (125):166–172, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983130.

Baur:1892:ANT

[Bau92a] G. Baur. Addition to the note on the taxonomy of the genus Emys C. Duméril. Proceedings of the American Philosophical Society, 30(138): 245, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983174.

Baur:1892:TGE

[Bau92b] G. Baur. On the taxonomy of the genus Emys, C. Duméril. Proceedings of the American Philosophical Society, 30(137):40–44, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983208.

Baur:1893:NCT

[Bau93] G. Baur. Notes on the classification and taxonomy of the testudinata. Proceedings of the American Philosophical Society, 31(141):210–225, April/June 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982831.

Byington:1869:GCL

[BB69] Cyrus Byington and Dr. Brinton. Grammar of the Choctaw language. Proceedings of the American Philosophical Society, 11(81):317–367, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981486.

Britton:1875:ARM

[BC75] J. Blodget Britton and Charles M. Cresson. Analyses of rocky mountain coal. Proceedings of the American Philosophical Society, 14 (94):358–361, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981865.

Bell:1873:ONC

[Bel73] John Bell. Obituary notice of Charles D. Meigs, M.D. Proceedings of the American Philosophical Society, 13(90):170–179, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981619.

Biddle:1889:ONC

[Bid89] Craig Biddle. Obituary notice of Caspar Wister, M.D. Proceedings of the American Philosophical Society, 26(130):492–495, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983189.

Bland:1871:NRP

[Bla71] Thomas Bland. Notes relating to the physical geography and geology of, and the distribution of terrestrial mollusca in certain of the West India Islands. Proceedings of the American Philosophical Society, 12 (86):56–63, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981669.

Blasius:1875:SRC

[Bla75] William Blasius. Some remarks on the connection of meteorology with health. Proceedings of the American Philosophical Society, 14 (95):667–673, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981959.

Blasius:1876:BDS

[Bla76] W. Blasius. A brief discussion of some opinions in meteorology. Proceedings of the American Philosophical Society, 16(97):198–206, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982428.

Blasius:1877:CHD

[Bla77] Wm. Blasius. Causes of the Huron disaster. Proceedings of the American Philosophical Society, 17(100):212–215, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982298.

Blasius:1885:RSG

[Bla85] Wm. Blasius. The remarkable sun-glows in the falls of 1883 and 1884. Proceedings of the American Philosophical Society, 22(118):213–215, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982679.

Blasius:1887:SSB

[Bla87] William Blasius. The Signal Service Bureau. Its methods and results. Proceedings of the American Philosophical Society, 24(125): 179–183, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983132.

Blasius:1889:SSD

[Bla89] William Blasius. Has the signal service degenerated? Proceedings of the American Philosophical Society, 26(129):285–288, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982946.

Blodget:1873:NPD

[Blo73] Lorin Blodget. The non-periodic distribution of heat in the atmosphere. Proceedings of the American Philosophical Society, 13(90): 138–140, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981608.

Blodget:1874:DAC

[Blo74] Lorin Blodget. A downward atmospheric circulation, as one cause of extremes of cold. Proceedings of the American Philosophical Society, 14(92):150–153, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982073.

Boas:1887:NEB

[Boa87] F. Boas. Notes on the ethnology of British Columbia. Proceedings of the American Philosophical Society, 24(126):422–428, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983091.

Boas:1891:VTH

[Boa91] Franz Boas. Vocabularies of the Tlingit, Haida and Tsimshian languages. Proceedings of the American Philosophical Society, 29(136): 173–208, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982936.

Boas:1893:VKL

[Boa93] Franz Boas. Vocabulary of the Kwakiutl language. Proceedings of the American Philosophical Society, 31(140):34–82, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983105.

Boas:1895:ST

[Boa95] Franz Boas. Salishan texts. Proceedings of the American Philosophical Society, 34(147):31–48, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983157.

Bayard:1889:A

[BP89] John Bayard and Thomas Paine. An act. Proceedings of the American Philosophical Society, 27(131):77–82, November 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982879.

Branner:1884:CGF

[Bra84] John Casper Branner. The course and growth of the fibro-vascular bundles in palms. Proceedings of the American Philosophical Society, 21(115):459–483, 502, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982741.

Branner:1886:GPW

[Bra86] John C. Branner. The glaciation of parts of the Wyoming and Lackawanna Valleys. Proceedings of the American Philosophical Society, 23(123):337–357, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983217.

Branner:1889:NBT

[Bra89] John C. Branner. Notes on the Botocúdus and their ornaments. Proceedings of the American Philosophical Society, 26(129):171–173, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982942.

Brinton:1869:CGM

[Bri69] D. G. Brinton. Contributions to a grammar of the Muskokee language. Proceedings of the American Philosophical Society, 11 (81):301–309, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981481.

Brinton:1873:LN

[Bri73] D. G. Brinton. On the language of the Natchez. Proceedings of the American Philosophical Society, 13(90):483–499, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981647.

Briggs:1876:FWT

[Bri76a] Robert Briggs. The flow of water through an opening in a pierced plate. Proceedings of the American Philosophical Society, 16(98):

310–318, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982817.

Britton:1876:IMB

[Bri76b] J. Blodget Britton. Improved mounted burettes for volumetric analysis. Proceedings of the American Philosophical Society, 16(97): 192–195, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982425.

Briggs:1877:FWT

[Bri77a] Robert Briggs. The flow of water through an opening in a pierced plate. Proceedings of the American Philosophical Society, 17(100): 124–127, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982287.

Britton:1877:ACP

[Bri77b] J. Blodget Britton. American condensed peat: Experimental tests to determine its value for blast and puddling furnace use, &c. Proceedings of the American Philosophical Society, 16(99):656–660, January/ May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982493.

Brinton:1879:ONL

[Bri79] Daniel G. Brinton. Obituary notice of the Late Isaac Hays, M.D. Proceedings of the American Philosophical Society, 18(103):259–260, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982336.

Brinton:1880:OND

[Bri80] Dr. Brinton. Obituary notice of Dr. John Neill. Proceedings of the American Philosophical Society, 19(107):161–162, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982621.

Brinton:1881:MHM

[Bri81a] D. G. Brinton. Memoir of S. S. Haldeman, A.M., Ph.D., etc. Proceedings of the American Philosophical Society, 19(108):279–285, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982249.

Brinton:1881:NGK

[Bri81b] Daniel G. Brinton. The names of the gods in the Kiche myths, Central America. Proceedings of the American Philosophical Society, 19 (109):613–647, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982803.

Brinton:1884:GCL

[Bri84] Daniel G. Brinton. A grammar of the Cakchiquel language of Guatemala. Proceedings of the American Philosophical Society, 21 (115):345–412, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982729.

Brinton:1885:LMS

[Bri85a] Daniel G. Brinton. The lineal measures of the semi-civilized nations of Mexico and Central America. Proceedings of the American Philosophical Society, 22(118):194–207, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982676.

Brinton:1885:LEP

[Bri85b] Daniel G. Brinton. On the language and ethnologic position of the Xinca Indians of Guatemala. Proceedings of the American Philosophical Society, 22(118):89–97, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982669.

Brinton:1885:PGA

[Bri85c] Daniel G. Brinton. The philosophic grammar of American languages, as set forth by Wilhelm von Humboldt, with the translation of an unpublished memoir by him on the American verb. Proceedings of the American Philosophical Society, 22(120):306–331, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982991.

Brinton:1886:CLS

[Bri86a] Daniel G. Brinton. The conception of love in some American languages. Proceedings of the American Philosophical Society, 23 (124):546–561, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983335.

Brinton:1886:NME

[Bri86b] Daniel G. Brinton. Notes on the Mangue; an extinct dialect formerly spoken in Nicaragua. Proceedings of the American Philosophical Society, 23(122):238–257, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983235.

Brinton:1886:PIC

[Bri86c] Daniel G. Brinton. On polysynthesis and incorporation as characteristics of American languages. Proceedings of the American Philosophical Society, 23(121):48–86, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982913.

Brinton:1886:IMP

[Bri86d] Daniel G. Brinton. On the ikonomatic method of phonetic writing, with special reference to American archaeology. Proceedings of the American Philosophical Society, 23(124):503–514, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983329.

Brinton:1887:AHF

[Bri87a] D. G. Brinton. On an ancient human footprint from Nicaragua. Proceedings of the American Philosophical Society, 24(126):437– 444, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983095.

Brinton:1887:CAL

[Bri87b] D. G. Brinton. On the so-called Alaguilac language of Guatemala. Proceedings of the American Philosophical Society, 24(126):366– 377, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983080.

Brinton:1887:CRE

[Bri87c] Daniel G. Brinton. Critical remarks on the editions of Diego de Landa’s writings. Proceedings of the American Philosophical Society, 24(125):1–8, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983120.

Brinton:1887:WTH

[Bri87d] Daniel G. Brinton. Were the Toltecs an historic nationality? Proceedings of the American Philosophical Society, 24(126):229–241, July/ December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983071.

Brinton:1888:LPM

[Bri88a] Daniel G. Brinton. The language of paleolithic man. Proceedings of the American Philosophical Society, 25(128):212–225, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983058.

Brinton:1888:ONP

[Bri88b] Daniel G. Brinton. Obituary notice of Philip H. Law, Esq. Proceedings of the American Philosophical Society, 25(128):225–231, July/ December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983059.

Brinton:1889:EAA

[Bri89a] Daniel G. Brinton. The ethnologic affinities of the Ancient Etruscans. Proceedings of the American Philosophical Society, 26(130): 506–527, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983191.

Brinton:1889:KSC

[Bri89b] Daniel G. Brinton. The ta ki, the svastika and the cross in America. Proceedings of the American Philosophical Society, 26(129): 177–187, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982944.

Brinton:1890:NPL

[Bri90a] Daniel G. Brinton. Note on the Puquina language of Peru. Proceedings of the American Philosophical Society, 28(134):242–259, July/ December 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983114.

Brinton:1890:ELN

[Bri90b] Daniel G. Brinton. On Etruscan and Libyan names. A comparative study. Proceedings of the American Philosophical Society, 28 (132):39–52, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983308.

Brinton:1891:VMC

[Bri91] Daniel G. Brinton. Vocabularies from the musquito coast. Proceedings of the American Philosophical Society, 29(135):1–4, January/ June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983010.

Brinton:1892:ELE

[Bri92a] D. G. Brinton. The Etrusco–Libyan elements in the song of the Arval Brethren. Proceedings of the American Philosophical Society, 30(139):317–324, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983355.

Brinton:1892:FNF

[Bri92b] D. G. Brinton. Further notes on Fuegian languages. Proceedings of the American Philosophical Society, 30(138):249–254, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983176.

Brinton:1892:FNB

[Bri92c] Daniel G. Brinton. Further notes on the Betoya dialects; from unpublished sources. Proceedings of the American Philosophical Society, 30(139):271–278, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983344.

Brinton:1892:OCL

[Bri92d] Daniel G. Brinton. Observations on the Chinantec language of Mexico. Proceedings of the American Philosophical Society, 30(137):22– 31, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983206.

Brinton:1892:MLM

[Bri92e] Daniel G. Brinton. On the Mazatec language of Mexico and its affinities. Proceedings of the American Philosophical Society, 30(137):31– 39, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983207.

Brinton:1892:SSA

[Bri92f] Daniel G. Brinton. Studies in South American native languages. Proceedings of the American Philosophical Society, 30(137):45–105, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983209.

Brinton:1893:NCC

[Bri93a] Daniel G. Brinton. The native calendar of Central America and Mexico. A study in linguistics and symbolism. Proceedings of the American Philosophical Society, 31(142):258–314, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982967.

Brinton:1893:VND

[Bri93b] Daniel G. Brinton. A vocabulary of the Nanticoke dialect. Proceedings of the American Philosophical Society, 31(142):325–333, July/ December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982971.

Brinton:1894:ONG

[Bri94a] D. G. Brinton. Obituary notice of George de Benneville Keim. Proceedings of the American Philosophical Society, 33(145):187–192, June 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983267.

Brinton:1894:NSN

[Bri94b] Daniel G. Brinton. Nagualism. A study in Native American folklore and history. Proceedings of the American Philosophical Society, 33(144):11–73, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983361.

Brinton:1895:OND

[Bri95a] D. G. Brinton. Obituary notice of Dr. William Samuel Waithman Ruschenberger. Proceedings of the American Philosophical Society, 34(149):361–364, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982890.

Brinton:1895:MLS

[Bri95b] Daniel G. Brinton. The Matagalpan linguistic stock of Central America. Proceedings of the American Philosophical Society, 34 (149):403–415, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982898.

Brinton:1895:PEW

[Bri95c] Daniel G. Brinton. The protohistoric ethnography of Western Asia. Proceedings of the American Philosophical Society, 34(147):71–102, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983161.

Brinton:1895:SWA

[Bri95d] Daniel G. Brinton. Some words from the Andagueda dialect of the Choco stock. Proceedings of the American Philosophical Society, 34 (149):401–402, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982897.

Brinton:1896:VND

[Bri96] Daniel G. Brinton. Vocabulary of the Noanama dialect of the Choco stock. Proceedings of the American Philosophical Society, 35(152):202–204, November 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983146.

Brinton:1897:EAG

[Bri97a] Daniel G. Brinton. The ethnic affinities of the Guetares of Costa Rica. Proceedings of the American Philosophical Society, 36(156):496–498, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983406.

Brinton:1897:MTF

[Bri97b] Daniel G. Brinton. The measurement of thought as function. Proceedings of the American Philosophical Society, 36(156):438–440, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983395.

Brinton:1898:LCC

[Bri98a] Daniel G. Brinton. The linguistic cartography of the Chaco Region. Proceedings of the American Philosophical Society, 37(158):178–205, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983852.

Brinton:1898:TUR

[Bri98b] Daniel G. Brinton. On two unclassified recent vocabularies from South America. Proceedings of the American Philosophical Society, 37(158):321–323, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983856.

Brockway:1884:TOQ

[Bro84] C. B. Brockway. Thermometrical observations in Quito, Ecuador. Farenheit. Proceedings of the American Philosophical Society, 21 (116):676–684, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982348.

Brown:1894:CSC

[Bro94] Amos Peaslee Brown. A comparative study of the chemical behavior of pyrite and marcasite. Proceedings of the American Philosophical Society, 33(145):225–243, June 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983271.

Buti:1899:IFB

[But99] Rudolph Buti. On an interesting fragment of the “book of the dead”. Proceedings of the American Philosophical Society, 38(159):79–81, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983626.

Cadwalader:1869:ONM

[Cad69] Judge Cadwalader. Obituary notice of Mr. Bancker. Proceedings of the American Philosophical Society, 11(81):85–90, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981442.

Carey:1871:ONS

[Car71] Henry C. Carey. Obituary notice of Stephen Colwell. Proceedings of the American Philosophical Society, 12(86):195–209, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981700.

Carson:1875:N

[Car75] Dr. Carson. Note. Proceedings of the American Philosophical Society, 14(94):433–434, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981878.

Carll:1877:RSM

[Car77] J. F. Carll. On the results of surveys in 1876–7 made for the purpose of rectifying the system of rail road and oil well levels throughout North West Pennsylvania. Proceedings of the American Philosophical Society, 17(100):17–28, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982274.

Carson:1885:AAH

[Car85a] Hampton L. Carson. Andrew Athinson Humphreys, Brigadier-General U. S. Army, Brevet Major-General U. S. Army, Chief of Engineers. Proceedings of the American Philosophical Society, 22(117):48– 71, January 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982407.

Carter:1885:DAO

[Car85b] Oscar C. S. Carter. On the detection of adulterations in oils. Proceedings of the American Philosophical Society, 22(120):296–299, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982989.

Carter:1891:FBL

[Car91] Oscar C. S. Carter. Feldspar bed in Laurentian (?) gneiss. Proceedings of the American Philosophical Society, 29(135):49–50, January/ June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983019.

Chase:1860:SEA

[Cha60] Pliny E. Chase. Sanscrit and English analogues. Proceedings of the American Philosophical Society, 7(63):177–291, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981889.

Chase:1861:CIE

[Cha61] Pliny Earle Chase. Chinese and Indo-European roots and analogues. Proceedings of the American Philosophical Society, 8(65):5–48, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982026.

Chase:1868:OMM

[Cha68a] Pliny E. Chase. Observations on the meteors of the morning of November 14. Proceedings of the American Philosophical Society, 10(80):539, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981799.

Chase:1868:SGC

[Cha68b] Pliny Earle Chase. On some general connotations of magnetism. Proceedings of the American Philosophical Society, 10(78):368–380, January 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981587.

Chase:1868:TRP

[Cha68c] Pliny Earle Chase. Tidal rainfall of Philadelphia. Proceedings of the American Philosophical Society, 10(80):523–535, 537, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981798.

Chase:1869:TR

[Cha69a] P. E. Chase. Tidal rainfall. Proceedings of the American Philosophical Society, 11(81):202–203, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981458.

Chase:1869:CME

[Cha69b] Pliny Earle Chase. Comparison of mechanical equivalents. Proceedings of the American Philosophical Society, 11(81):313, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981483.

Chase:1869:CRL

[Cha69c] Pliny Earle Chase. Cosmical relations of light to gravity. Proceedings of the American Philosophical Society, 11(81):103–107, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981445.

Chase:1869:MVR

[Cha69d] Pliny Earle Chase. Monthly variations of rainfall at Philadelphia. Proceedings of the American Philosophical Society, 11(81):314–315, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981484.

Chase:1869:PLT

[Cha69e] Pliny Earle Chase. Philadelphia life tables. Proceedings of the American Philosophical Society, 11(81):17–22, January 1869. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981437.

Chabas:1871:NFC

[Cha71a] F. Chabas. Note of F. Chabas, of the Institute of France, on the footmark in the hieroglyphic script. Proceedings of the American Philosophical Society, 12(86):193, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981699.

Chase:1871:ADP

[Cha71b] Pliny Earle Chase. Æthereal density and polarity. Proceedings of the American Philosophical Society, 12(86):407–410, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981726.

Chase:1871:AOP

[Cha71c] Pliny Earle Chase. Æthereal oscillation, the primordial material force. Proceedings of the American Philosophical Society, 12(86):411–417, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981728.

Chase:1871:AWN

[Cha71d] Pliny Earle Chase. American weather notes. Proceedings of the American Philosophical Society, 12(86):40, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981665.

Chase:1871:CCM

[Cha71e] Pliny Earle Chase. Correlations of cosmical and molecular force. Proceedings of the American Philosophical Society, 12(86):392–394, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981719.

Chase:1871:CRS

[Cha71f] Pliny Earle Chase. Cyclical rainfall at San Francisco. Proceedings of the American Philosophical Society, 12(86):523–542, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981754.

Chase:1871:CRL

[Cha71g] Pliny Earle Chase. Cyclical rainfalls at Lisbon. Proceedings of the American Philosophical Society, 12(86):178–190, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981696.

Chase:1871:DAM

[Cha71h] Pliny Earle Chase. Daily auroral and meteoric means. Proceedings of the American Philosophical Society, 12(86):516–518, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981752.

Chase:1871:EAR

[Cha71i] Pliny Earle Chase. European and American rain-fails. Proceedings of the American Philosophical Society, 12(86):38–39, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981664.

Chase:1871:FAS

[Cha71j] Pliny Earle Chase. Further approximations to the Sun’s distance. Proceedings of the American Philosophical Society, 12(86):398–400, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981721.

Chase:1871:GRA

[Cha71k] Pliny Earle Chase. General relation of auroras to rainfall. Proceedings of the American Philosophical Society, 12(86):400–401, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981722.

Chase:1871:HSP

[Cha71l] Pliny Earle Chase. The Herschell–Stephenson postulate. Proceedings of the American Philosophical Society, 12(86):395–397, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981720.

Chase:1871:IMS

[Cha71m] Pliny Earle Chase. Influence of meteoric showers on auroras. Proceedings of the American Philosophical Society, 12(86):401–403, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981723.

Chase:1871:LCR

[Cha71n] Pliny Earle Chase. Lunar-cyclical rainfall in the northern temperate zone. Proceedings of the American Philosophical Society, 12(86):558– 559, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981762.

Chase:1871:NPT

[Cha71o] Pliny Earle Chase. On the normal position of the tidal ellipsoid. Proceedings of the American Philosophical Society, 12(86):123–124, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981683.

Chase:1871:RAG

[Cha71p] Pliny Earle Chase. On the relation of the auroras to gravitating currents. Proceedings of the American Philosophical Society, 12(86): 121–122, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981681.

Chase:1871:PIE

[Cha71q] Pliny Earle Chase. Planetary illustrations of explosive oscillation. Proceedings of the American Philosophical Society, 12(86):403–405, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981724.

Chase:1871:RMR

[Cha71r] Pliny Earle Chase. Recent monthly rainfall in the United States. Proceedings of the American Philosophical Society, 12(86):555–557, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981761.

Chase:1871:RAM

[Cha71s] Pliny Earle Chase. Resemblance of atmospheric, magnetic and oceanic currents. Proceedings of the American Philosophical Society, 12(86):68–70, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981672.

Chase:1871:SPR

[Cha71t] Pliny Earle Chase. Solar and planetary rotation. Proceedings of the American Philosophical Society, 12(86):406–407, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981725.

Chase:1871:SPC

[Cha71u] Pliny Earle Chase. Stellar and planetary correlations. Proceedings of the American Philosophical Society, 12(86):518–522, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981753.

Chase:1871:SSC

[Cha71v] Pliny Earle Chase. The sun-spot cycle of 11.07 years. Proceedings of the American Philosophical Society, 12(86):410–411, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981727.

Chase:1871:WE

[Cha71w] Pliny Earle Chase. Winds of Europe. Proceedings of the American Philosophical Society, 12(86):123, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981682.

Chase:1871:WUS

[Cha71x] Pliny Earle Chase. Winds of the United States. Proceedings of the American Philosophical Society, 12(86):65–67, 1871. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981671.

Chase:1873:CMH

[Cha73a] Pliny Earle Chase. Cosmical and molecular harmonies, No. II. Proceedings of the American Philosophical Society, 13(90):237–252, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981630.

Chase:1873:ESM

[Cha73b] Pliny Earle Chase. Estimation of solar mass and distance, from the equilibrium of elastic and gravitating forces. Proceedings of the American Philosophical Society, 13(90):142–143, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981610.

Chase:1873:GSL

[Cha73c] Pliny Earle Chase. The gamuts of sound and light. Proceedings of the American Philosophical Society, 13(90):149–154, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981615.

Chase:1873:MS

[Cha73d] Pliny Earle Chase. The music of the spheres. Proceedings of the American Philosophical Society, 13(90):193–198, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981624.

Chase:1873:NPT

[Cha73e] Pliny Earle Chase. Note on planeto-taxis. Proceedings of the American Philosophical Society, 13(90):143–145, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981611.

Chase:1873:OFS

[Cha73f] Pliny Earle Chase. Oscillatory forces in the solar system. Proceedings of the American Philosophical Society, 13(90):140–141, 1873. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981609.

Chase:1873:PRS

[Cha73g] Pliny Earle Chase. Planetary relations to the sun-spot period. Proceedings of the American Philosophical Society, 13(90):147–148, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981613.

Chase:1873:RCA

[Cha73h] Pliny Earle Chase. Recent confirmation of an astronomical prediction. Proceedings of the American Philosophical Society, 13(90):470– 477, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981645.

Chase:1873:RVL

[Cha73i] Pliny Earle Chase. Relative velocities of light and gravity. Proceedings of the American Philosophical Society, 13(90):148–149, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981614.

Chase:1873:RSI

[Cha73j] Pliny Earle Chase. Rotation of the Sun and the intra-asteroidal planets. Proceedings of the American Philosophical Society, 13(90):145– 147, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981612.

Chase:1873:TCM

[Cha73k] Pliny Earle Chase. Transcript of a curious manuscript work in cypher, supposed to be astrological. Proceedings of the American Philosophical Society, 13(90):477–482, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981646.

Channing:1874:MPN

[Cha74a] William F. Channing. Meteorological peculiarities of new england. Proceedings of the American Philosophical Society, 14(92):154, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982074.

Chase:1874:CE

[Cha74b] Pliny Earle Chase. Cosmical evolution. Proceedings of the American Philosophical Society, 14(92):159–161, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982077.

Chase:1874:CT

[Cha74c] Pliny Earle Chase. Cosmical thermodynamics. Proceedings of the American Philosophical Society, 14(92):141–147, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982071.

Chase:1874:CRB

[Cha74d] Pliny Earle Chase. Cyclical rainfall at Barbadoes. Proceedings of the American Philosophical Society, 14(93):195–216, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982110.

Chase:1874:JCR

[Cha74e] Pliny Earle Chase. Jupiter-cyclical rainfall. Proceedings of the American Philosophical Society, 14(93):193–195, June/December 1874. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982109.

Chase:1874:OAF

[Cha74f] Pliny Earle Chase. Origin of attractive force. Proceedings of the American Philosophical Society, 14(92):111–113, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982067.

Chase:1874:SFL

[Cha74g] Pliny Earle Chase. Saving-fund life-insurance. Proceedings of the American Philosophical Society, 14(92):148–149, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982072.

Chase:1875:BD

[Cha75a] Pliny Earle Chase. The beginnings of development. Proceedings of the American Philosophical Society, 14(95):622–631, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981944.

Chase:1875:FDC

[Cha75b] Pliny Earle Chase. Further dynamic co-ordinations. Proceedings of the American Philosophical Society, 14(95):651–658, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981956.

Chase:1875:FRM

[Cha75c] Pliny Earle Chase. Further relations of magnetic, gravitating, and luminous force. Proceedings of the American Philosophical Society, 14 (95):607–609, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981939.

Chase:1875:GW

[Cha75d] Pliny Earle Chase. Gravitating waves. Proceedings of the American Philosophical Society, 14(94):344–346, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981862.

Chase:1875:LMR

[Cha75e] Pliny Earle Chase. Lunar-monthly rain-fall in the United States. Proceedings of the American Philosophical Society, 14(94):416–418, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981871.

Chase:1875:PIC

[Cha75f] Pliny Earle Chase. Planetary illustrations of the creative fiat. Proceedings of the American Philosophical Society, 14(95):609–612, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
981940.

Chase:1875:YRU

[Cha75g] Pliny Earle Chase. Yearly rainfall in the United States. Proceedings of the American Philosophical Society, 14(95):613–614, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981941.

Chase:1876:NAS

[Cha76a] Pliny Earle Chase. Nebular action in the solar system. Proceedings of the American Philosophical Society, 16(97):184–192, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982424.

Chase:1876:SFP

[Cha76b] Pliny Earle Chase. On some fundamental propositions of central force. Proceedings of the American Philosophical Society, 16(98): 298–310, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982816.

Chase:1877:FIC

[Cha77a] Pliny Earle Chase. Further illustrations of central force. Proceedings of the American Philosophical Society, 17(100):98–112, June/ December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982285.

Chase:1877:CAD

[Cha77b] Pliny Earle Chase. On centres of aggregation and dissociation. Proceedings of the American Philosophical Society, 16(99):496–508, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982477.

Chance:1878:HSO

[Cha78a] H. M. Chance. Hyner’s station oil well section, included in a detailed section of the rocks between the lower productive coal measures (XIII), and the dark slates of the Devonian (VIII) in the vicinity of Renovo, Clinton Co., Pennsylvania. Proceedings of the American Philosophical Society, 17(101):670–672, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982658.

Chase:1878:LCG

[Cha78b] Pliny E. Chase. The limiting constant of gravitation. Proceedings of the American Philosophical Society, 18(102):41–43, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982755.

Chase:1878:CNH

[Cha78c] Pliny Earle Chase. Criteria of the nebular hypothesis. Proceedings of the American Philosophical Society, 17(101):341–345, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982643.

Chase:1878:CH

[Cha78d] Pliny Earle Chase. Crucial harmonies. Proceedings of the American Philosophical Society, 18(102):34–36, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982753.

Chase:1878:RR

[Cha78e] Pliny Earle Chase. Radiation and rotation. Proceedings of the American Philosophical Society, 17(101):701–704, January/June 1878. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982660.

Chase:1878:RWI

[Cha78f] Pliny Earle Chase. Results of wave interferences. Proceedings of the American Philosophical Society, 17(101):294–307, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982635.

Chase:1879:AQC

[Cha79a] Pliny Earle Chase. Approximate quadrature of the circle. Proceedings of the American Philosophical Society, 18(104):281, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982501.

Chase:1879:AAA

[Cha79b] Pliny Earle Chase. Astronomical approximations. I. Apparent semidiameter of the Sun, and nebular origin of the terrestrial day. Proceedings of the American Philosophical Society, 18(104):380–381, July/ December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982521.

Chase:1879:FCP

[Cha79c] Pliny Earle Chase. Further confirmations of prediction. Proceedings of the American Philosophical Society, 18(103):209, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982322.

Chase:1879:PC

[Cha79d] Pliny Earle Chase. The philosophy of Christianity. Proceedings of the American Philosophical Society, 18(103):123–153, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982313.

Chase:1879:SR

[Cha79e] Pliny Earle Chase. Solar records. Proceedings of the American Philosophical Society, 18(103):224–232, January/June 1879. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982329.

Chase:1880:AAIa

[Cha80a] Pliny Earle Chase. Astronomical approximations. II, III. Proceedings of the American Philosophical Society, 18(105):425–435, January/ March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982776.

Chase:1880:AAIb

[Cha80b] Pliny Earle Chase. Astronomical approximations. IV. Nodal estimation of the velocity of light. Proceedings of the American Philosophical Society, 19(107):4–9, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982595.

Chase:1880:AAV

[Cha80c] Pliny Earle Chase. Astronomical approximations. V., VI. Proceedings of the American Philosophical Society, 19(107):18–21, March/ December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982601.

Chase:1880:LPC

[Cha80d] Pliny Earle Chase. List of papers communicated to the American Philosophical Society. Proceedings of the American Philosophical Society, 19(107):184–190, March/December 1880. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982624.

Chase:1880:P

[Cha80e] Pliny Earle Chase. Photodynamics. Proceedings of the American Philosophical Society, 19(107):203–209, March/December 1880. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982630.

Chase:1880:RCA

[Cha80f] Pliny Earle Chase. Relations of chemical affinity to luminous and cosmical energies. Proceedings of the American Philosophical Society, 19 (107):21–25, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982602.

Chance:1881:AFD

[Cha81a] H. M. Chance. An analysis of the fire-damp explosions in the anthracite coal mines, from 1870 to 1880. Proceedings of the American Philosophical Society, 19(108):405–407, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982263.

Chance:1881:AGN

[Cha81b] H. M. Chance. The auriferous gravels of North Carolina. Proceedings of the American Philosophical Society, 19(109):477–481, June/ December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982791.

Chase:1881:PN

[Cha81c] Pliny Earle Chase. Photodynamic notes. Proceedings of the American Philosophical Society, 19(108):262–274, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982246.

Chase:1881:PNIa

[Cha81d] Pliny Earle Chase. Photodynamic notes, II. Proceedings of the American Philosophical Society, 19(108):354–377, January/June 1881. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982259.

Chase:1881:PNIb

[Cha81e] Pliny Earle Chase. Photodynamic notes, III. Proceedings of the American Philosophical Society, 19(109):446–472, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982789.

Chase:1881:PNIc

[Cha81f] Pliny Earle Chase. Photodynamic notes, IV. Proceedings of the American Philosophical Society, 19(109):567–612, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982802.

Chase:1882:PNVa

[Cha82a] Pliny Earle Chase. Photodynamic notes, V. Proceedings of the American Philosophical Society, 20(111):237–283, January/June 1882. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982724.

Chase:1882:PNVb

[Cha82b] Pliny Earle Chase. Photodynamic notes, VI. Proceedings of the American Philosophical Society, 20(112):406–438, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982690.

Chase:1883:PNVa

[Cha83a] Pliny Earle Chase. Photodynamic notes. VII. Proceedings of the American Philosophical Society, 20(113):566–598, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982446.

Chase:1883:PNVb

[Cha83b] Pliny Earle Chase. Photodynamic notes, VIII. Proceedings of the American Philosophical Society, 21(114):120–134, March 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982369.

Chase:1884:PNI

[Cha84] Pliny Earle Chase. Photodynamic notes, IX. Proceedings of the American Philosophical Society, 21(116):590–611, June 1884. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982340.

Chase:1885:CMR

[Cha85a] Pliny Earle Chase. The Chase–Maxwell ratio. Proceedings of the American Philosophical Society, 22(120):375–376, October 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982997.

Chase:1885:EWF

[Cha85b] Pliny Earle Chase. An experiment in weather forecast. Proceedings of the American Philosophical Society, 22(118):207–210, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982677.

Chase:1885:TEP

[Cha85c] Pliny Earle Chase. Thermo-electro-photo-baric unit. Proceedings of the American Philosophical Society, 22(120):377–378, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982998.

Claypole:1883:GN

[Cla83a] E. W. Claypole. Geological notes. Proceedings of the American Philosophical Society, 20(113):529–536, January/April 1883. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982442.

Claypole:1883:NLC

[Cla83b] E. W. Claypole. Note on a large crustacean from the Catskill Group of Pennsylvania. Proceedings of the American Philosophical Society, 21(114):236–239, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982382.

Claypole:1883:NLF

[Cla83c] E. W. Claypole. Note on a large fish-plate from the Upper Chemung (?) beds of Northern Pennsylvania. Proceedings of the American Philosophical Society, 20(113):664–666, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982462.

Claypole:1883:NRN

[Cla83d] E. W. Claypole. Note on a relic of the native flora of Pennsylvania, surviving in Perry County. Proceedings of the American Philosophical Society, 21(114):226–230, March 1883. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982379.

Claypole:1883:NGR

[Cla83e] E. W. Claypole. Note on the genus Rensselaeria in the Hamilton Group in Perry Co. Proceedings of the American Philosophical Society, 21(114):235–236, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982381.

Claypole:1883:KWS

[Cla83f] E. W. Claypole. On the Kingsmill white sandstone. Proceedings of the American Philosophical Society, 20(113):666–673, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982463. See com-

ment [Les83d].

Claypole:1883:PCF

[Cla83g] E. W. Claypole. The Perry County Fault. Note on an important correction in the geological map of Pennsylvania. Proceedings of the American Philosophical Society, 21(114):218–225, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982378.

Claypole:1884:COS

[Cla84] E. W. Claypole. On the Clinton and other shales, &c., composing the fifth group of Rogers in the First Survey of Pennsylvania. Proceedings of the American Philosophical Society, 21(115):492–502, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982745.

Claypole:1887:OVI

[Cla87] E. W. Claypole. Organic variation indefinite not definite in direction — an outcome of environment. Proceedings of the American Philosophical Society, 24(125):113–135, January/June 1887. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983128.

Conklin:1896:DFO

[Con96] E. G. Conklin. Discussion of the factors of organic evolution from the embryological standpoint. Proceedings of the American Philosophical Society, 35(150):77–88, January 1896. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982867.

Cooper:1879:NSL

[Coo79] J. G. Cooper. Notes on some land-shells of the Pacific slope. Proceedings of the American Philosophical Society, 18(104):282–288, July/ December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982502.

Cope:1869:CIM

[Cop69a] E. D. Cope. Contribution to the ichthyology of the Marañon. Proceedings of the American Philosophical Society, 11(81):559–570, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981513.

Cope:1869:ECH

[Cop69b] E. D. Cope. Eighth contribution to the herpetology of Tropical America. Proceedings of the American Philosophical Society, 11(81):553–559, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981512.

Cope:1869:FCH

[Cop69c] E. D. Cope. Fourth contribution to the history of the fauna of the Miocene and Eocene periods of the United States. Proceedings of the American Philosophical Society, 11(81):285–294, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981478.

Cope:1869:NSH

[Cop69d] E. D. Cope. Note on Saurocephalus, Harlan. Proceedings of the American Philosophical Society, 11(81):608, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981521.

Cope:1869:OFT

[Cop69e] E. D. Cope. Observations on the fishes of the tertiary shales of Green River, Wyoming Territory. Proceedings of the American Philosophical Society, 11(81):380–384, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981490.

Cope:1869:AGC

[Cop69f] E. D. Cope. On Adocus, a genus of Cretaceous Emydidæ. Proceedings of the American Philosophical Society, 11(81):295–298, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981479.

Cope:1869:SEP

[Cop69g] E. D. Cope. On some etheostomine perch from Tennessee and North Carolina. Proceedings of the American Philosophical Society, 11(81):261–270, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981470.

Cope:1869:SSP

[Cop69h] E. D. Cope. On some species of Pythonomorpha from the Cretaceous beds of Kansas and New Mexico. Proceedings of the American Philosophical Society, 11(81):574–584, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981517.

Cope:1869:A

[Cop69i] E. D. Cope. On the Adocidæ. Proceedings of the American Philosophical Society, 11(81):547–553, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981511.

Cope:1869:FFW

[Cop69j] E. D. Cope. On the fishes of a fresh water tertiary in Idaho, discovered by Capt. Clarence King. Proceedings of the American Philosophical Society, 11(81):538–547, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981510.

Cope:1869:TEA

[Cop69k] E. D. Cope. On three extinct Astaci from the fresh-water territory of Idaho. Proceedings of the American Philosophical Society, 11(81):605–607, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981520.

Cope:1869:PSF

[Cop69l] Edw. D. Cope. A partial synopsis of the fishes of the fresh waters of North Carolina. Proceedings of the American Philosophical Society, 11(81):448–495, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981499.

Cope:1869:SAH

[Cop69m] Edw. D. Cope. Second addition to the history of the fishes of the Cretaceous of the United States. Proceedings of the American Philosophical Society, 11(81):240–244, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981468.

Cope:1869:S

[Cop69n] Edward D. Cope. On the Saurodontidæ. Proceedings of the American Philosophical Society, 11(81):529–538, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981509.

Cope:1869:R

[Cop69o] Edward D. Cope. Remarks. Proceedings of the American Philosophical Society, 11(81):419, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981493.

Cope:1869:SCH

[Cop69p] Edward D. Cope. Seventh contribution to the herpetology of Tropical America. Proceedings of the American Philosophical Society, 11(81):147–192, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981453.

Cope:1871:CPF

[Cop71a] E. D. Cope. Catalogue of the “Pythonomorpha” found in the Cretaceous strata of Kansas. Proceedings of the American Philosophical Society, 12(86):264–287, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981707.

Cope:1871:DNE

[Cop71b] E. D. Cope. Descriptions of new extinct reptiles from the Upper Green River Eocene Basin, Wyoming. Proceedings of the American Philosophical Society, 12(86):554–555, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981760.

Cope:1871:DSN

[Cop71c] E. D. Cope. Descriptions of some new vertebrata from the Bridger Group of the Eocene. Proceedings of the American Philosophical Society, 12(86):460–465, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981742.

Cope:1871:NTC

[Cop71d] E. D. Cope. On a new testudinate from the chalk of Kansas. Proceedings of the American Philosophical Society, 12(86):308–310, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981711.

Cope:1871:MB

[Cop71e] E. D. Cope. On Megaptera Bellicosa. Proceedings of the American Philosophical Society, 12(86):103–108, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981675.

Cope:1871:FFC

[Cop71f] E. D. Cope. On the families of fishes of the Cretaceous formation of Kansas. Proceedings of the American Philosophical Society, 12(86): 327–357, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981715.

Cope:1871:OFC

[Cop71g] E. D. Cope. On the occurrence of fossil Cobitidæ in Idaho. Proceedings of the American Philosophical Society, 12(86):55, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981668.

Cope:1871:TEF

[Cop71h] E. D. Cope. On two extinct forms of physostomi of the neotropical region. Proceedings of the American Philosophical Society, 12(86): 52–55, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981667.

Cope:1871:PRV

[Cop71i] E. D. Cope. Preliminary report on the vertebrata discovered in the Port Kennedy Bone Cave. Proceedings of the American Philosophical Society, 12(86):73–102, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981674.

Cope:1871:SSE

[Cop71j] E. D. Cope. Supplement to the “Synopsis of the Extinct Batrachia and Reptilia of North America”. Proceedings of the American Philosophical Society, 12(86):41–52, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981666.

Cope:1871:MCO

[Cop71k] Ed. D. Cope. The method of creation of organic forms. Proceedings of the American Philosophical Society, 12(86):229–263, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981706.

Cope:1871:NPE

[Cop71l] Edw. D. Cope. Notice of proboscians from the Eocene of Southern Wyoming. Proceedings of the American Philosophical Society, 12(86): 580, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981770.

Cope:1871:SMO

[Cop71m] Edw. D. Cope. Stated meeting, Oct. 20, 1871. Proceedings of the American Philosophical Society, 12(86):173–176, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981694.

Cope:1871:DGP

[Cop71n] Edward D. Cope. A description of the genus Protostega, a form of extinct Testudinata. Proceedings of the American Philosophical Society, 12(86):422–433, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981731.

Cope:1871:NSC

[Cop71o] Edward D. Cope. Note of some Cretaceous vertebrata in the State Agricultural College of Kansas, U.S.A. Proceedings of the American Philosophical Society, 12(86):168–170, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981692.

Cope:1871:NNV

[Cop71p] Edward D. Cope. Notices of new vertebrata from the upper waters of Bitter Creek, Wyoming Territory. Proceedings of the American Philosophical Society, 12(86):483–486, 1871. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981748.

Cope:1871:ODC

[Cop71q] Edward D. Cope. Observations on the distribution of certain extinct vertebrata in North Carolina. Proceedings of the American Philosophical Society, 12(86):210–216, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981701.

Cope:1871:NGP

[Cop71r] Edward D. Cope. On a new genus of Pleurodira from the Eocene of Wyoming. Proceedings of the American Philosophical Society, 12(86): 472–477, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981745.

Cope:1871:NVG

[Cop71s] Edward D. Cope. On a new vertebrate genus from the northern part of the tertiary basin of Green River. Proceedings of the American Philosophical Society, 12(86):554, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981759.

Cope:1871:BEG

[Cop71t] Edward D. Cope. On bathmodon, an extinct genus of ungulates. Proceedings of the American Philosophical Society, 12(86):417–420, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981729.

Cope:1871:DM

[Cop71u] Edward D. Cope. On the dentition of metalophodon. Proceedings of the American Philosophical Society, 12(86):542–545, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981755.

Cope:1871:EDT

[Cop71v] Edward D. Cope. On the existence of dinosauria in the transition beds of Wyoming. Proceedings of the American Philosophical Society, 12(86):481–483, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981747.

Cope:1871:TCF

[Cop71w] Edward D. Cope. On the tertiary coal and fossils of Osino, Nevada. Proceedings of the American Philosophical Society, 12(86):478–481, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981746.

Cope:1871:TNO

[Cop71x] Edward D. Cope. On two new ornithosaurians from Kansas. Proceedings of the American Philosophical Society, 12(86):420–422, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981730.

Cope:1871:SAN

[Cop71y] Edward D. Cope. Second account of new vertebrata from the Bridger Eocene. Proceedings of the American Philosophical Society, 12(86): 466–468, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981743.

Cope:1871:SNE

[Cop71z] Edward D. Cope. Second notice of extinct vertebrates from Bitter Creek, Wyoming. Proceedings of the American Philosophical Society, 12(86):487–488, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981749.

Cope:1871:TAN

[Cop71-27] Edward D. Cope. Third account of new vertebrata from the Bridger Eocene of Wyoming Territory. Proceedings of the American Philosophical Society, 12(86):469–472, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981744.

Cope:1873:OET

[Cop73a] E. D. Cope. On the osteology of the extinct tapiroid Hyrachyus. Proceedings of the American Philosophical Society, 13(90):212–224, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981627.

Cope:1873:CIA

[Cop73b] Edward D. Cope. A contribution to the ichthyology of Alaska. Proceedings of the American Philosophical Society, 13(90):24–32, 1873.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981600.

Cope:1873:FCC

[Cop73c] Edward D. Cope. On the flat-clawed carnivora of the Eocene of Wyoming. Proceedings of the American Philosophical Society, 13(90): 198–209, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981625.

Cope:1873:NPB

[Cop73d] Edward D. Cope. On the new perissodactyles from the Bridger Eocene. Proceedings of the American Philosophical Society, 13(90): 35–36, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981602.

Cope:1873:SFU

[Cop73e] Edward D. Cope. On the short footed ungulata of the Eocene of Wyoming. Proceedings of the American Philosophical Society, 13(90): 38–74, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981604.

Cope:1874:ZTP

[Cop74a] E. D. Cope. On the zoology of a temporary pool on the plains of Colorado. Proceedings of the American Philosophical Society, 14 (92):139–140, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982070.

Cope:1874:PIU

[Cop74b] Edward D. Cope. On the Plagopterinæ and the ichthyology of Utah. Proceedings of the American Philosophical Society, 14(92): 129–139, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982069.

Cope:1875:RPO

[Cop75a] E. D. Cope. On the remains of population observed on and near the Eocene Plateau of North-Western New Mexico. Proceedings of the American Philosophical Society, 14(95):475–482, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981936.

Cope:1875:SVM

[Cop75b] E. D. Cope. Synopsis of the vertebrata of the Miocene of Cumberland County, New Jersey. Proceedings of the American Philosophical Society, 14(94):361–364, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981866.

Cope:1877:CRA

[Cop77a] E. D. Cope. A continuation of researches among the batrachia of the coal measures of Ohio. Proceedings of the American Philosophical Society, 16(99):573–578, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982481.

Cope:1877:DEV

[Cop77b] E. D. Cope. Descriptions of extinct vertebrata from the Permian and Triassic formamations of the United States. Proceedings of the American Philosophical Society, 17(100):182–193, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982295.

Cope:1877:DNV

[Cop77c] E. D. Cope. Descriptions of new vertebrata from the upper tertiary formations of the west. Proceedings of the American Philosophical Society, 17(100):219–231, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982301.

Cope:1877:DTU

[Cop77d] E. D. Cope. On a dinosaurian from the trias of Utah. Proceedings of the American Philosophical Society, 16(99):579–584, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982482.

Cope:1877:NP

[Cop77e] E. D. Cope. On a new proboscidian. Proceedings of the American Philosophical Society, 16(99):584–585, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982483.

Cope:1877:NSA

[Cop77f] E. D. Cope. On a new species of Adocidæ from the Tertiary of Georgia. Proceedings of the American Philosophical Society, 17 (100):82–84, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982283.

Cope:1877:RRD

[Cop77g] E. D. Cope. On reptilian remains from the Dakota beds of Colorado. Proceedings of the American Philosophical Society, 17(100): 193–196, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982296.

Cope:1877:SNLa

[Cop77h] E. D. Cope. On some new and little known reptiles and fishes from the Austroriparian Region. Proceedings of the American Philosophical Society, 17(100):63–68, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982279.

Cope:1877:SNLb

[Cop77i] E. D. Cope. On some new or little known reptiles and fishes of the Cretaceous no. 3, of Kansas. Proceedings of the American Philosophical Society, 17(100):176–181, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982294.

Cope:1877:SSF

[Cop77j] E. D. Cope. On some saurians found in the Triassic of Pennsylvania, by C. M. Wheatley. Proceedings of the American Philosophical Society, 17(100):231–232, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982302.

Cope:1877:BPO

[Cop77k] E. D. Cope. On the brain of Procamelus Occidentalis. Proceedings of the American Philosophical Society, 17(100):49–52, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982277.

Cope:1877:BC

[Cop77l] E. D. Cope. On the brain of coryphodon. Proceedings of the American Philosophical Society, 16(99):616–620, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982487.

Cope:1877:VBB

[Cop77m] E. D. Cope. On the vertebrata of the bone bed in Eastern Illinois. Proceedings of the American Philosophical Society, 17(100): 52–63, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982278.

Cope:1877:VDE

[Cop77n] E. D. Cope. On the vertebrata of the Dakota Epoch of Colorado. Proceedings of the American Philosophical Society, 17(100): 233–247, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982303.

Cope:1877:SCB

[Cop77o] E. D. Cope. Synopsis of the cold blooded vertebrata, procured by Prof. James Orton during his exploration of Peru in 1876–77. Proceedings of the American Philosophical Society, 17(100):33–49, June/ December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982276.

Cope:1877:TCH

[Cop77p] E. D. Cope. Tenth contribution to the herpetology of Tropical America. Proceedings of the American Philosophical Society, 17 (100):85–98, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982284.

Cope:1878:DEB

[Cop78a] E. D. Cope. Descriptions of extinct Batrachia and Reptilia from the Permian Formation of Texas. Proceedings of the American Philosophical Society, 17(101):505–530, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982652.

Cope:1878:SFP

[Cop78b] E. D. Cope. Synopsis of the fishes of the Peruvian Amazon, obtained by Professor Orton during his expeditions of 1873 and 1877. Proceedings of the American Philosophical Society, 17(101):673–701, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982659.

Cope:1879:ECH

[Cop79a] E. D. Cope. Eleventh contribution to the herpetology of Tropical America. Proceedings of the American Philosophical Society, 18 (104):261–277, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982498.

Cope:1879:SCK

[Cop79b] E. D. Cope. Second contribution to a knowledge of the Miocene fauna of Oregon. Proceedings of the American Philosophical Society, 18 (104):370–376, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982519.

Cope:1880:FPP

[Cop80a] E. D. Cope. On the foramina perforating the posterior part of the squamosal bone of the mammalia. Proceedings of the American Philosophical Society, 18(105):452–461, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982780.

Cope:1880:GC

[Cop80b] E. D. Cope. On the genera of the Creodonta. Proceedings of the American Philosophical Society, 19(107):76–82, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982610.

Cope:1880:SCH

[Cop80c] E. D. Cope. Second contribution to the history of the vertebrata of the Permian Formation of Texas. Proceedings of the American Philosophical Society, 19(107):38–58, March/December 1880. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982605.

Cope:1881:NSP

[Cop81a] E. D. Cope. Note on the structure of the posterior foot of toxodon. Proceedings of the American Philosophical Society, 19(108): 402–403, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982261.

Cope:1881:SML

[Cop81b] E. D. Cope. On some mammalia of the lowest Eocene beds of New Mexico. Proceedings of the American Philosophical Society, 19(109): 484–495, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982794.

Cope:1881:SAO

[Cop81c] E. D. Cope. The systematic arrangement of the order Perissodactyla. Proceedings of the American Philosophical Society, 19(108): 377–401, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982260.

Cope:1882:CUM

[Cop82a] E. D. Cope. The classification of the ungulate mammalia. Proceedings of the American Philosophical Society, 20(112):438–447, June/ December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982691.

Cope:1882:CHV

[Cop82b] E. D. Cope. Contributions to the history of the vertebrata of the Lower Eocene of Wyoming and New Mexico, made during 1881. Proceedings of the American Philosophical Society, 20(111):139–197, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982711.

Cope:1882:SRC

[Cop82c] E. D. Cope. On the systematic relations of the carnivora fissipedia. Proceedings of the American Philosophical Society, 20(112): 471–475, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982694.

Cope:1882:SVP

[Cop82d] E. D. Cope. Synopsis of the vertebrata of the Puerco Eocene epoch. Proceedings of the American Philosophical Society, 20(112): 461–471, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982693.

Cope:1882:TCH

[Cop82e] E. D. Cope. Third contribution to the history of the vertebrata of the Permian Formation of Texas. Proceedings of the American Philosophical Society, 20(112):447–461, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982692.

Cope:1883:FAF

[Cop83a] E. D. Cope. First addition to the fauna of the Puerco Eocene. Proceedings of the American Philosophical Society, 20(113):545–563, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982444.

Cope:1883:FCH

[Cop83b] E. D. Cope. Fourth contribution to the history of the Permian Formation of Texas. Proceedings of the American Philosophical Society, 20 (113):628–636, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982451.

Cope:1883:BEM

[Cop83c] E. D. Cope. On the brains of the Eocene mammalia Phenacodus and Periptychus. Proceedings of the American Philosophical Society, 20 (113):563–565, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982445.

Cope:1883:DLF

[Cop83d] E. D. Cope. On the distribution of the Loup Fork Formation in New Mexico. Proceedings of the American Philosophical Society, 21(114):308–309, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982395.

Cope:1883:TTM

[Cop83e] E. D. Cope. On the trituberculate type of molar tooth in the mammalia. Proceedings of the American Philosophical Society, 21 (114):324–326, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982397.

Cope:1883:SAK

[Cop83f] E. D. Cope. Second addition to the knowledge of the Puerco Epoch. Proceedings of the American Philosophical Society, 21(114):309–324, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982396.

Cope:1884:SSE

[Cop84a] E. D. Cope. On the structure of the skull in the Elasmobranch genus Didymodus. Proceedings of the American Philosophical Society, 21(116):572–590, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982339.

Cope:1884:SSO

[Cop84b] E. D. Cope. Synopsis of the species of Oreodontidæ. Proceedings of the American Philosophical Society, 21(116):503–572, 615, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982338.

Cope:1885:CHM

[Cop85a] E. D. Cope. A contribution to the herpetology of Mexico. Proceedings of the American Philosophical Society, 22(120):379–404, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982999.

Cope:1885:EMV

[Cop85b] E. D. Cope. The extinct mammalia of the Valley of Mexico. Proceedings of the American Philosophical Society, 22(117):1–21, January 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982404.

Cope:1885:FCK

[Cop85c] E. D. Cope. Fifth contribution to the knowledge of the fauna of the Permian Formation of Texas and the Indian Territory. Proceedings of the American Philosophical Society, 22(117):28–47, January 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982406.

Cope:1885:SFE

[Cop85d] E. D. Cope. On the structure of the feet in the extinct Artiodactyla of North America. Proceedings of the American Philosophical Society, 22(117):21–27, January 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982405.

Cope:1885:SCR

[Cop85e] E. D. Cope. Second continuation of researches among the batrachia of the coal measures of Ohio. Proceedings of the American Philosophical Society, 22(120):405–408, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983001.

Cope:1885:TCH

[Cop85f] E. D. Cope. Twelfth contribution to the herpetology of Tropical America. Proceedings of the American Philosophical Society, 22(118):167–194, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982675.

Cope:1886:ATG

[Cop86a] E. D. Cope. An analytical table of the genera of snakes. Proceedings of the American Philosophical Society, 23(124):479–499, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983327.

Cope:1886:CSB

[Cop86b] E. D. Cope. Catalogue of the species of batrachians and reptiles contained in a collection made at Pebas, Upper Amazon, by John Hauxwell. Proceedings of the American Philosophical Society, 23(121):94–103, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982915.

Cope:1886:CVP

[Cop86c] E. D. Cope. A contribution to the vertebrate paleontology of Brazil. Proceedings of the American Philosophical Society, 23(121):1–21, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982910.

Cope:1886:SI

[Cop86d] E. D. Cope. On the species of Iguaninæ. Proceedings of the American Philosophical Society, 23(122):261–271, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983238.

Cope:1886:SAA

[Cop86e] E. D. Cope. On the structure and affinities of the Amphiumidæ. Proceedings of the American Philosophical Society, 23(123):442–445, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983227.

Cope:1886:SBA

[Cop86f] E. D. Cope. On the structure of the brain and auditory apparatus of a theromorphous reptile of the Permian Epoch. Proceedings of the American Philosophical Society, 23(122):234–238, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983234.

Cope:1886:TNS

[Cop86g] E. D. Cope. On two new species of three-toed horses from the Upper Miocene, with notes on the fauna of the Ticholeptus beds. Proceedings of the American Philosophical Society, 23(123):357–361, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983218.

Cope:1886:RCD

[Cop86h] E. D. Cope. Report on the coal deposits near Zacualtipan, in the State of Hidalgo, Mexico. Proceedings of the American Philosophical Society, 23(121):146–151, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982917.

Cope:1886:SLN

[Cop86i] E. D. Cope. Synonymic list of the North American Species of Bufo and Rana, with descriptions of some new species of Batrachia, from specimens in the National Museum. Proceedings of the American Philosophical Society, 23(124):514–526, December 1886. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983330.

Cope:1886:TCH

[Cop86j] E. D. Cope. Thirteenth contribution to the herpetology of Tropical America. Proceedings of the American Philosophical Society, 23(122):271–287, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983239.

Cope:1887:CPA

[Cop87a] E. D. Cope. The classification and phylogeny of the artiodactyla. Proceedings of the American Philosophical Society, 24(126):377– 400, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983081.

Cope:1887:CHV

[Cop87b] E. D. Cope. A contribution to the history of the vertebrata of the trias of North America. Proceedings of the American Philosophical Society, 24(126):209–228, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983070.

Cope:1887:SBR

[Cop87c] E. D. Cope. Synopsis of the Batrachia and Reptilia obtained by H. II. Smith, in the province of Mato Grosso, Brazil. Proceedings of the American Philosophical Society, 24(125):44–60, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983124.

Cope:1888:DJD

[Cop88a] E. D. Cope. On the Dicotylinæ of the John Day Miocene of North America. Proceedings of the American Philosophical Society, 25 (127):62–79, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983282.

Cope:1888:MOD

[Cop88b] E. D. Cope. On the mechanical origin of the dentition of the amblypoda. Proceedings of the American Philosophical Society, 25 (127):80–88, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983283.

Cope:1889:OPE

[Cop89a] E. D. Cope. An outline of the philosophy of evolution. Proceedings of the American Philosophical Society, 26(130):495–505, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983190.

Cope:1889:RNA

[Cop89b] E. D. Cope. A review of the North American species of Hippotherium. Proceedings of the American Philosophical Society, 26 (130):429–458, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983183.

Cope:1892:CVP

[Cop92a] E. D. Cope. A contribution to the vertebrate paleontology of Texas. Proceedings of the American Philosophical Society, 30(137):123–131, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983215.

Cope:1892:FEJ

[Cop92b] E. D. Cope. On false elbow joints. Proceedings of the American Philosophical Society, 30(139):285–291, December 1892. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983347.

Cope:1892:SNL

[Cop92c] E. D. Cope. On some new and little known Paleozoic vertebrates. Proceedings of the American Philosophical Society, 30(138):221–229, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983171.

Cope:1892:SPK

[Cop92d] E. D. Cope. On some points in the kinetogenesis of the limbs of vertebrates. Proceedings of the American Philosophical Society, 30 (139):282–285, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983346.

Cope:1892:PV

[Cop92e] E. D. Cope. On the phylogeny of the vertebrata. Proceedings of the American Philosophical Society, 30(139):278–281, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983345.

Cope:1892:SDL

[Cop92f] E. D. Cope. On the skull of the dinosaurian Lælaps incrassatus Cope. Proceedings of the American Philosophical Society, 30(138):240–245, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983173.

Cope:1892:TNG

[Cop92g] E. D. Cope. On Tiaporus, a new genus of Teiidæ. Proceedings of the American Philosophical Society, 30(137):132–133, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983216.

Cope:1892:OL

[Cop92h] E. D. Cope. The osteology of the Lacertilia. Proceedings of the American Philosophical Society, 30(138):185–221, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983170.

Cope:1893:GT

[Cop93a] E. D. Cope. On the genus Tomiopsis. Proceedings of the American Philosophical Society, 31(142):317–318, July/December 1893. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982969.

Cope:1893:SAK

[Cop93b] E. D. Cope. Second addition to the knowledge of the Batrachia and Reptilia of Costa Rica. Proceedings of the American Philosophical Society, 31(142):333–347, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982972.

Cope:1894:FON

[Cop94a] E. D. Cope. On the fishes obtained by the naturalist expedition in rio grande do sul. Proceedings of the American Philosophical Society, 33(144):84–108, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983363.

Cope:1894:LO

[Cop94b] E. D. Cope. On the lungs of the Ophidia. Proceedings of the American Philosophical Society, 33(145):217–224, June 1894. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983270.

Cope:1894:SSP

[Cop94c] E. D. Cope. On the structure of the skull in the Plesiosaurian Reptilia, and on two new species from the Upper Cretaceous. Proceedings of the American Philosophical Society, 33(144):109–113, January 1894.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983364.

Cope:1895:FCM

[Cop95a] E. D. Cope. Fourth contribution to the marine fauna of the Miocene period of the United States. Proceedings of the American Philosophical Society, 34(147):135–155, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983163.

Cope:1895:SPM

[Cop95b] E. D. Cope. On some Plistocene mammalia from Petite Anse, La. Proceedings of the American Philosophical Society, 34(149):458–468, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982901.

Cope:1895:ROC

[Cop95c] E. D. Cope. The reptilian order Cotylosauria. Proceedings of the American Philosophical Society, 34(149):436–457, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982900.

Cope:1896:SCH

[Cop96a] E. D. Cope. Second contribution to the history of the Cotylosauria. Proceedings of the American Philosophical Society, 35(151):122–139, August 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983258.

Cope:1896:SCK

[Cop96b] E. D. Cope. Sixth contribution to the knowledge of the marine Miocene fauna of North America. Proceedings of the American Philosophical Society, 35(151):139–146, August 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983259.

Cope:1897:NPV

[Cop97] E. D. Cope. On new Paleozoic vertebrata from Illinois, Ohio and Pennsylvania. Proceedings of the American Philosophical Society, 36(154):71–91, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983460.

Coues:1893:DOM

[Cou93] Elliott Coues. Description of the original manuscript journals and field notebooks of Lewis and Clark, on which was based Biddle’s history of the expedition of 1804–6, and which are now in the possession of the American Philosophical Society in Philadelphia. Proceedings of the American Philosophical Society, 31(140):17–33, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983104.

Crane:1883:MSB

[Cra83] T. F. Crane. Medieval sermon-books and stories. Proceedings of the American Philosophical Society, 21(114):49–78, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982361.

Cresson:1874:REE

[Cre74] Charles M. Cresson. Results of an examination of an exploded locomotive boiler, and of experiments to ascertain the causes of explosion. Proceedings of the American Philosophical Society, 14(93): 264–271, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982116.

Cresson:1875:EMG

[Cre75] Charles M. Cresson. The effect of magnetic and galvanic forces upon the strength of, and destruction of iron and steel structures. Proceedings of the American Philosophical Society, 14(95):603–606, June/ December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981938.

Cresson:1877:BMP

[Cre77] Charles M. Cresson. Bituminous material from Pulaski County, Virginia, U.S. Proceedings of the American Philosophical Society, 17 (100):215–216, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982299.

Crotch:1873:AFC

[Cro73] G. R. Crotch. On the arrangement of the families of Coleoptera. Proceedings of the American Philosophical Society, 13(90):75–87, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981605.

Curwen:1885:ONT

[Cur85] John Curwen. Obituary notice of Thomas S. Kirkbride, M.D. Proceedings of the American Philosophical Society, 22(120):217–227, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982980.

Cushing:1896:EAK

[Cus96] Frank Hamilton Cushing. Exploration of ancient key dwellers’ remains on the Gulf Coast of Florida. Proceedings of the American Philosophical Society, 35(153):329–448, December 1896. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983594.

DaCosta:1885:BSP

[Da 85] J. M. Da Costa. Biographical sketch of Professor Samuel D. Gross. Proceedings of the American Philosophical Society, 22(117):78–85, January 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982409.

Davidson:1866:NLM

[Dav66] George Davidson. A new level for the measurement of the inclination of the transit axis of portable transit instruments, for all altitudes beyond the reach of the ordinary striding level. Proceedings of the American Philosophical Society, 10(76):288–290, June 1866. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982219.

Davenport:1883:SCT

[Dav83a] George E. Davenport. Some comparative tables showing the distribution of ferns in the United States of North America. Proceedings of the American Philosophical Society, 20(113):605–612, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982449.

Davis:1883:CCH

[Dav83b] Wm. Morris Davis. On the conversion of chlorine into hydrochloric acid, as observed in the deposition of gold from its solutions by charcoal. Proceedings of the American Philosophical Society, 21(114):102– 109, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982366.

Dawson:1863:NML

[Daw63] J. W. Dawson. Note on Mr. Lesley’s paper on the coal-measures of Cape Breton. Proceedings of the American Philosophical Society, 9(69):165–204, January 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981410.

Day:1883:MET

[Day83] Frank M. Day. The microscopic examination of timber with regard to its strength. A contribution from the Eli K. Price Botanical Laboratory of the University of Pennsylvania. Proceedings of the American Philosophical Society, 21(114):333–342, March 1883. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982401.

Day:1897:SOP

[Day97] David T. Day. A suggestion as to the origin of Pennsylvania petroleum. Proceedings of the American Philosophical Society, 36 (154):112–115, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983464.

Day:1898:PAR

[Day98] Wm. C. Day. The production of an asphalt resembling gilsonite by the distillation of a mixture of fish and wood. Proceedings of the American Philosophical Society, 37(157):171–174, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983707.

Delmar:1874:RPS

[Del74] Alexander Delmar. On the resources, productions and social condition of Egypt. Proceedings of the American Philosophical Society, 14 (93):232–255, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982114.

Delmar:1875:RPS

[Del75] Alexander Delmar. The resources, productions and social condition of spain. Proceedings of the American Philosophical Society, 14 (94):301–343, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981861.

Derby:1879:CGL

[Der79a] Orville A. Derby. A contribution to the geology of the Lower Amazonas. Proceedings of the American Philosophical Society, 18 (103):155–178, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982315.

Derby:1879:GDR

[Der79b] Orville A. Derby. The geology of the diamantiferous region of the Province of Paraná, Brazil. Proceedings of the American Philosophical Society, 18(103):251–258, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982335.

Dolley:1893:TDP

[Dol93] Charles S. Dolley. The Thyrsos of Dionysos and the palm inflorescence of the winged figures of Assyrian monuments. Proceedings of the American Philosophical Society, 31(140):109–116, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983107.

Dolley:1894:ONJ

[Dol94] Charles S. Dolley. Obituary notice of John M. Maisch. Proceedings of the American Philosophical Society, 33(146):345–352, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982846.

Doolittle:1897:VTL

[Doo97] C. L. Doolittle. The variation of terrestrial latitude. Proceedings of the American Philosophical Society, 36(156):434–438, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983394.

Doremus:1896:ICI

[Dor96] Charles A. Doremus. The identification of colored inks by their absorption spectra. Proceedings of the American Philosophical Society, 35(150):71–74, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982864.

Draper:1877:DOS

[Dra77] Henry Draper. Discovery of oxygen in the Sun by photography, and a new theory of the solar spectrum. Proceedings of the American Philosophical Society, 17(100):74–80, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982281.

Drake:1897:GRC

[Dra97] Noah Fields Drake. A geological reconnaissance of the coal fields of the Indian Territory. Proceedings of the American Philosophical Society, 36(156):326–419, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983388.

Dubois:1871:ONM

[Dub71a] Mr. Dubois. Obituary notice of Mr. Jacob R. Eckfeldt. Proceedings of the American Philosophical Society, 12(86):547–552, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981757.

Dubois:1871:QCR

[Dub71b] Wm. E. Dubois. On a quasi coin reported found in a boring in Illinois. Proceedings of the American Philosophical Society, 12(86):224– 228, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981704.

DuBois:1888:JCB

[DuB88] Patterson DuBois. James Curtis Booth, Ph.D., LL.D. Proceedings of the American Philosophical Society, 25(128):204–211, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983057.

DuBois:1895:MP

[DuB95] Patterson DuBois. A matter of priority. Proceedings of the American Philosophical Society, 34(147):67–71, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983160.

Dudley:1886:TRT

[Dud86] Thos. H. Dudley. Is there reciprocity in trade? And the consumption of manufactured commodities. Proceedings of the American Philosophical Society, 23(124):526–535, December 1886. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983331.

Duges:1886:RAF

[Dug86] Alfredo Dugés. Sur le Rhinocheilus Antonii. (French) [On the Rhinocheilus Antonii]. Proceedings of the American Philosophical Society, 23(122):290–291, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983241.

Duges:1888:DEN

[Dug88] Alfredo Dugés. Sur deux espèces nouvelles des Ophidiens de Mexique. (French) [On two new species of Ophidians of Mexico]. Proceedings of the American Philosophical Society, 25(128):181–183, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983051.

Dutton:1871:CRE

[Dut71] C. E. Dutton. The causes of regional elevations and subsidences. Proceedings of the American Philosophical Society, 12(86):70–72, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981673.

Dwight:1893:OPP

[Dwi93] Thomas Dwight. Observations on the Psoas Parvus and Pyramidalis. A study of variation. Proceedings of the American Philosophical Society, 31(140):117–123, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983108.

Eddy:1882:RHE

[Edd82] H. T. Eddy. Radiant heat an exception to the second law of thermodynamics. Proceedings of the American Philosophical Society, 20 (112):334–343, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982684.

Edmunds:1897:IA

[Edm97] George F. Edmunds. International arbitration. Proceedings of the American Philosophical Society, 36(155):320–323, May 1897. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983659.

Emerson:1871:LIU

[Eme71] Dr. Emerson. On lunar influence upon the conditions of wet or dry weather. Proceedings of the American Philosophical Society, 12(86):

17–18, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981661.

Farr:1896:NOW

[Far96] Marcus S. Farr. Notes on the osteology of the White River horses. Proceedings of the American Philosophical Society, 35(151):147–175, August 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983260.

Fraley:1880:SMS

[FGG^{+}80] Frederick Fraley, Asa Gray, D. C. Gilman, W. A. Hammond, Wm. C. Cattell, Daniel G. Brinton, A. Loudon Snowden, John L. LeConte, W. V. McKean, and J. P. Lesley. Special meeting at the St. George Hotel [Celebration of the Hundredth Anniversary of the Incorporation. American Philosophical Society. March 15, 1880]. Proceedings of the American Philosophical Society, 18(106):511–589, March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982825.

Field:1871:ONS

[Fie71] Henry W. Field. Obituary notice of Sir John Frederick William Herschel, Bart. Proceedings of the American Philosophical Society, 12(86):217–223, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981703.

Field:1897:SL

[Fie97] Robert P. Field. The span of life. Proceedings of the American Philosophical Society, 36(156):420–424, December 1897. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983390.

Frazer:1871:SOA

[Fra71] Persifor Frazer, Jr. On a spectroscopic observation of the aurora of April 10, 1872. Proceedings of the American Philosophical Society, 12(86):579–580, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981769.

Frazer:1874:ERN

[Fra74a] P. Frazer, Jr. On exfoliation of rocks near Gettysburg. Proceedings of the American Philosophical Society, 14(93):295–297, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982119.

Frazer:1874:NCM

[Fra74b] Persifer Frazer, Jr. Note on the color of the Moon. Proceedings of the American Philosophical Society, 14(92):155–156, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982075.

Frazer:1875:TMS

[Fra75a] Persifor Frazer, Jr. On the traps of the Mesozoic sandstone in York and Adams Counties, Pennsylvania. Proceedings of the American Philosophical Society, 14(94):402–414, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981869.

Frazer:1875:OLS

[Fra75b] Persifor Frazer, Jr. Origin of the Lower Silurian limonites of York and Adams Counties. Proceedings of the American Philosophical Society, 14(94):364–369, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981867.

Frazer:1876:NLD

[Fra76] Persifor Frazer, Jr. Note on the “Lithologie du fond des mers” of M. Delesse. Proceedings of the American Philosophical Society, 16 (97):238–240, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982438.

Fraley:1877:ONJ

[Fra77a] Frederick Fraley. Obituary notice of John C. Cresson. Proceedings of the American Philosophical Society, 17(100):149–162, June/ December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982291.

Frazer:1877:ATC

[Fra77b] Persifor Frazer, Jr. The approaches to a theory of the cause of magnetic declination. Proceedings of the American Philosophical Society, 16(99):642–650, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982491.

Frazer:1877:CS

[Fra77c] Persifor Frazer, Jr. Crystallography in sculpture. Proceedings of the American Philosophical Society, 17(100):258–260, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982306.

Frazer:1877:RSM

[Fra77d] Persifor Frazer, Jr. Regarding some Mesozoic ores. Proceedings of the American Philosophical Society, 16(99):651–655, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982492.

Frazer:1877:RPP

[Fra77e] Persifor Frazer, Jr. Remarks on Professor Prime’s paper. Proceedings of the American Philosophical Society, 17(100):255–258, June/ December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982305.

Frazer:1878:CLP

[Fra78a] Persifor Frazer, Jr. Contribution to the lithology of Pennsylvania. On the physical and chemical characteristics of a trap occurring at Williamson’s Point. Proceedings of the American Philosophical Society, 18(102):96–103, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982765.

Frazer:1878:SMO

[Fra78b] Persifor Frazer, Jr. Some microscopical observations of the phonograph record. Proceedings of the American Philosophical Society, 17 (101):531–536, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982653.

Frazer:1878:STI

[Fra78c] Persifor Frazer, Jr. Some tables for the interconversion of metric and English units. Proceedings of the American Philosophical Society, 17 (101):536–539, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982654.

Frazer:1879:FFQ

[Fra79a] Persifor Frazer, Jr. Fossil (?) forms in the quartzose rocks of the Lower Susquehanna. Proceedings of the American Philosophical Society, 18(104):277–279, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982499.

Frazer:1879:ONR

[Fra79b] Persifor Frazer, Jr. Obituary notice of Robert Frazer. Proceedings of the American Philosophical Society, 18(103):233–235, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982330.

Frazer:1880:MIO

[Fra80] Persifor Frazer, Jr. A mirror for illuminating opaque objects for the projecting microscope. Proceedings of the American Philosophical Society, 18(105):503–510, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982783.

Frazer:1882:HSV

[Fra82] Persifor Frazer. The horizon of the South Valley Hill rocks in Pennsylvania. Proceedings of the American Philosophical Society, 20(112): 510–518, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982703.

Frazer:1883:ICH

[Fra83] Persifor Frazer. An improvement in the construction of the hypsometrical aneroid. Proceedings of the American Philosophical Society, 20(113):604, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982448.

Frazer:1884:TDA

[Fra84] Persifor Frazer. Trap dykes in the archæan rocks of Southeastern Pennsylvania. Proceedings of the American Philosophical Society, 21(116):691–694, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982352.

Frazer:1885:CDA

[Fra85] Persifor Frazer. A convenient device to be applied to the hand compass. Proceedings of the American Philosophical Society, 22(118):216, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982680.

Frazer:1886:CPA

[Fra86a] Persifor Frazer. Composite photography applied to handwriting. Proceedings of the American Philosophical Society, 23(123):433–441, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983226.

Frazer:1886:GNS

[Fra86b] Persifor Frazer. General notes. — Sketch on the geology of York County, Pennsylvania. Proceedings of the American Philosophical Society, 23(123):391–410, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983224.

Frazer:1886:RWI

[Fra86c] Persifor Frazer. Résumé of the work of the International Geological Congress, Held at Berlin, Sept. 28 to Oct. 3, 1885. Proceedings of the American Philosophical Society, 23(122):259–261, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983237.

Frazer:1894:RML

[Fra94a] Persifor Frazer. Remarks on Mr. Lyman’s Paper. Proceedings of the American Philosophical Society, 33(145):216–217, June 1894. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983269.

Frazer:1894:TNM

[Fra94b] Persifor Frazer. Three new methods for the detection of forgery. Proceedings of the American Philosophical Society, 33(145):251–252, June 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983273.

Frazer:1895:EAT

[Fra95a] Persifor Frazer. Evidences of the action of two hands in joint signature marks. Proceedings of the American Philosophical Society, 34(149):473–477, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982904.

Frazer:1895:MEY

[Fra95b] Persifor Frazer. In memory of Edward Yorke Macauley, U.S.N. Proceedings of the American Philosophical Society, 34(149):364–380, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982891.

Fulton:1874:NSC

[Ful74] John Fulton. Note on the Somerset County Coal Beds in Pennsylvania. Proceedings of the American Philosophical Society, 14 (92):157–158, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982076.

Furness:1896:GB

[Fur96] William Henry Furness, 3rd. Glimpses of Borneo. Proceedings of the American Philosophical Society, 35(153):309–320, December 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983592.

G:1886:EPA

[G.86] A. R. G. Errata in paper by Augustus R. Grote. Proceedings of the American Philosophical Society, 23(121):??, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982922.

Gabb:1861:SMC

[Gab61] William M. Gabb. Synopsis of the mollusca of the Cretaceous formation, including the geographical and stratigraphical range and synonymy. Proceedings of the American Philosophical Society, 8 (65):57–257, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982032.

Gabb:1875:ITL

[Gab75] Wm. M. Gabb. On the Indian tribes and languages of Costa Rica. Proceedings of the American Philosophical Society, 14(95): 483–602, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981937.

Garman:1887:RBG

[Gar87a] Samuel Garman. On the reptiles and batrachians of Grand Cayman. Proceedings of the American Philosophical Society, 24(126): 273–277, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983076.

Garman:1887:WIR

[Gar87b] Samuel Garman. On West Indian reptiles in the Museum of Comparative Zoölogy, at Cambridge, Mass. Proceedings of the American Philosophical Society, 24(126):278–286, July/December 1887. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983077.

Garrett:1887:MPE

[Gar87c] Philip C. Garrett. Memoir of Pliny Earle Chase. Proceedings of the American Philosophical Society, 24(126):287–295, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983078.

Gatschet:1876:RUT

[Gat76] Alb. S. Gatschet. Remarks upon the Tónkawa language. Proceedings of the American Philosophical Society, 16(98):318–327, June/ December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982818.

Gatschet:1877:TL

[Gat77] Albert S. Gatschet. The Timucua language. Proceedings of the American Philosophical Society, 16(99):626–642, January/May 1877. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982490.

Gatschet:1878:TL

[Gat78] Albert S. Gatschet. The Timucua language. Proceedings of the American Philosophical Society, 17(101):490–504, January/June 1878. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982651.

Gatschet:1880:TL

[Gat80] Albert S. Gatschet. The Timucua language. Proceedings of the American Philosophical Society, 18(105):465–502, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982782.

Gatschet:1885:ALP

[Gat85a] Alb. S. Gatschet. The Aruba language and the Papiamento jargon. Proceedings of the American Philosophical Society, 22(120):299–305, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982990.

Gatschet:1885:BI

[Gat85b] Albert S. Gatschet. The Beothuk Indians. Proceedings of the American Philosophical Society, 22(120):408–424, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983002.

Gatschet:1886:BI

[Gat86] Albert S. Gatschet. The Beothuk Indians. Proceedings of the American Philosophical Society, 23(123):411–432, July 1886. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983225.

Gatschet:1890:BIT

[Gat90] Albert S. Gatschet. The Beothuk Indians. Third article. Proceedings of the American Philosophical Society, 28(132):1–16, January/ June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983301.

Gatschet:1891:MTI

[Gat91] Albert S. Gatschet. A mythic tale of the Isleta Indians. Proceedings of the American Philosophical Society, 29(136):208–218, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982937.

Genth:1873:CLU

[Gen73] F. A. Genth. Contributions from the Laboratory of the University of Pennsylvania. No. I. Corundum, its alterations and associated minerals. Proceedings of the American Philosophical Society, 13(90):361– 406, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981643.

Genth:1874:CLU

[Gen74a] F. A. Genth. Contributions from the Laboratory of the University of Pennsylvania. No. III. On American tellurium and bismuth minerals. Proceedings of the American Philosophical Society, 14(93): 223–231, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982113.

Genth:1874:IIO

[Gen74b] F. A. Genth. Investigation of iron ores and limestones from Messrs. Lyon, Shorb & Co’s iron ore banks on Spruce Creek, Half Moon Run and Warrior’s Mark Run, in Centre, Blair and Huntingdon Counties, PA. Proceedings of the American Philosophical Society, 14(92):84–99, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982064.

Genth:1874:RDS

[Gen74c] F. A. Genth. Reply to Dr. T. Sterry Hunt. Proceedings of the American Philosophical Society, 14(93):216–218, June/December 1874. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982111.

Genth:1877:CLU

[Gen77] F. A. Genth. Contributions from the Laboratory of the University of Pennsylvania. No. XI. On some tellurium and vanadium minerals. Proceedings of the American Philosophical Society, 17(100): 113–123, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982286.

Genth:1879:PSC

[Gen79] F. A. Genth. On pyrophyllite from Schuylkill County, Pennsylvania. Proceedings of the American Philosophical Society, 18(104): 279–280, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982500.

Genth:1882:CLU

[Gen82] F. A. Genth. Contributions from the Laboratory of the University of Pennsylvania. No. XX. Contributions to mineralogy. Proceedings of the American Philosophical Society, 20(112):381–404, June/ December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982688.

Genth:1884:H

[Gen84] F. A. Genth. On herderite. Proceedings of the American Philosophical Society, 21(116):694–699, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982353.

Genth:1886:CLU

[Gen86] F. A. Genth. Contributions from the Laboratory of the University of Pennsylvania. No. XXIV. Contributions to mineralogy. Proceedings of the American Philosophical Society, 23(121):30–47, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982912.

Genth:1887:CCL

[Gen87] F. A. Genth. Contributions from the Chemical Laboratory of the University of Pennsylvania. No. XXIX. Contributions to mineralogy. Proceedings of the American Philosophical Society, 24(125):23–44, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983123.

Gill:1884:NS

[Gil84] Theodore Gill. Notes on the Stromateidæ. Proceedings of the American Philosophical Society, 21(116):664–672, June 1884. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982345.

Ginley:1878:NDP

[Gin78] William Ginley. Notice of the discovery of the position of the crural processes in the genus Atrypa. Proceedings of the American Philosophical Society, 17(101):337–338, January/June 1878. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982641.

Goddard:1896:SAS

[God96] Martha Freeman Goddard. On the second abdominal segment in a few Libellulidæ. Proceedings of the American Philosophical Society, 35(152):205–212, November 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983147.

Goodwin:1869:ONS

[Goo69] Daniel R. Goodwin. Obituary notice of Samuel Vaughan Merrick, Esq. Proceedings of the American Philosophical Society, 11 (81):584–597, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981518.

Graham:1858:CGNb

[Gra58a] J. D. Graham. Contributions to geography, no. 2. On the latitude and longitude of four additional positions on Lake Michigan, and of Madison, the capital of the state of Wisconsin. Proceedings of the American Philosophical Society, 6(60):352–388, July/December 1858.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981999.

Graham:1858:CGNa

[Gra58b] J. D. Graham. Contributions to geography, No. I. On the latitude and longitude of Chicago, Illinois. Proceedings of the American Philosophical Society, 6(60):347–352, July/December 1858. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981998.

Graff:1884:ONS

[Gra84] Frederic Graff. Obituary notice of Strickland Kneass. Proceedings of the American Philosophical Society, 21(115):451–455, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982738.

Greene:1879:FDA

[Gre79a] William H. Greene. On the formation of dibenzyl by the action of ethylene chloride on benzol in the presence of aluminium chloride. Proceedings of the American Philosophical Society, 18(104): 345–346, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982507.

Greene:1879:DMP

[Gre79b] Wm. H. Greene. On dioxyethyl-methylene, and the preparation of methylene chloride. Proceedings of the American Philosophical Society, 18(104):346–347, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982508.

Greene:1880:NSS

[Gre80a] Wm. H. Greene. On a new synthesis of saligenin. Proceedings of the American Philosophical Society, 18(105):451–452, January/ March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982779.

Greene:1880:AHA

[Gre80b] Wm. H. Greene. On the action of hydrochloric acid and of chlorine on acetobenzoic anhydride. Proceedings of the American Philosophical Society, 19(107):13–14, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982597.

Grote:1874:LNA

[Gro74] Aug. R. Grote. List of the North American Platypterices, Attaci, Hemileucini, Ceratocampadæ,Lachneides, Teredines and Hepiali, with notes. Proceedings of the American Philosophical Society, 14 (93):256–264, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982115.

Grote:1879:PBA

[Gro79] Aug. R. Grote. The philosophy of the biblical account of creation. Proceedings of the American Philosophical Society, 18(104): 316–323, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982504.

Grote:1883:ISN

[Gro83] A. R. Grote. Introduction to a study of the North American Noctuidæ. Proceedings of the American Philosophical Society, 21(114):134– 176, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982370.

Grote:1895:HMA

[Gro95a] A. Radcliffe Grote. The hypenoid moths and allied groups. Proceedings of the American Philosophical Society, 34(149):416–436, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982899.

Grote:1895:A

[Gro95b] A. Radcliffe Grote. On Apatela. Proceedings of the American Philosophical Society, 34(149):388–392, December 1895. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982894.

Grote:1898:SLW

[Gro98] A. Radcliffe Grote. Specializations of the lepidopterous wing; the Pieri-Nymphalidæ. Proceedings of the American Philosophical Society, 37(157):17–44, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983688.

Grote:1899:GTB

[Gro99a] A. Radcliffe Grote. Genealogical trees of butterflies. Proceedings of the American Philosophical Society, 38(160):147–154, October 1899.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983563.

Grote:1899:SLWa

[Gro99b] A. Radcliffe Grote. Specializations of the lepidopterous wing; Parnassi-Papilionidæ. Proceedings of the American Philosophical Society, 38(159):7–21, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/983615. See also Part II [Gro99c].

Grote:1899:SLWb

[Gro99c] A. Radcliffe Grote. Specializations of the lepidopterous wing: The Parnassi-Papilionidæ. II. Proceedings of the American Philosophical Society, 38(159):25–48, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983621. See also Part I [Gro99b].

Genth:1885:CLU

[GvR85] F. A. Genth and Gerhard vom Rath. Contributions from the Laboratory of the University of Pennsylvania. No. XXIII. On the vanadates and iodyrite, from Lake Valley, Sierra Co., New Mexico. Proceedings of the American Philosophical Society, 22(120):363–375, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982996.

Hagen:1882:IAT

[Hag82] John Hagen. On the inclination of the apparent to the true horizon and the errors rising thereof in transit, altitude, and azimuthobservations. Proceedings of the American Philosophical Society, 20 (111):206–223, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982716.

Hagen:1883:RSA

[Hag83] J. G. Hagen. On the reversion of series and its application to the solution of numerical equations. Proceedings of the American Philosophical Society, 21(114):93–101, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982365.

Hall:1875:NGA

[Hal75a] Charles E. Hall. Notes on glacial action visible along the Kittatinny or Blue Mountain, Carbon, Northampton, and Monroe Counties, Pennsylvania. Proceedings of the American Philosophical Society, 14 (95):620–621, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981943.

Hall:1875:GDW

[Hal75b] Charles E. Hall. On glacial deposits at West Philadelphia. Proceedings of the American Philosophical Society, 14(95):633–634, June/ December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981946.

Hall:1876:PMP

[Hal76] Charles E. Hall. On the progress of the museum and paleontological work of the Second Geological Survey of Pennsylvania, for the year 1875. Proceedings of the American Philosophical Society, 16 (97):55–60, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982421.

Hall:1877:CPM

[Hal77] Charles E. Hall. Contribution to paleontology from the Museum of the Second Geological Survey of Pennsylvania. Proceedings of the American Philosophical Society, 16(99):621–622, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982488.

Hall:1880:RCR

[Hal80] Charles E. Hall. The relations of the crystalline rocks of Eastern Pennsylvania to the Silurian limestones and the Hudson River age of the hydromice schists. Proceedings of the American Philosophical Society, 18(105):435–443, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982777.

Hale:1883:TTL

[Hal83] Horatio Hale. The Tutelo tribe and language. Proceedings of the American Philosophical Society, 21(114):1–47, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982359.

Hammond:1882:ONJ

[Ham82] William A. Hammond. An obituary notice of John W. Draper, M.D., LL.D. Proceedings of the American Philosophical Society, 20 (111):227–228, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982719.

Hancock:1888:DDM

[Han88] Joseph L. Hancock. Description of Datames Magna Hancock. Proceedings of the American Philosophical Society, 25(127):107–110, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983288.

Hartshorne:1871:ONE

[Har71a] Henry Hartshorne. Obituary notice of Edward Rhoads, M.D. Proceedings of the American Philosophical Society, 12(86):171–173, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981693.

Hartshorne:1871:OP

[Har71b] Henry Hartshorne. On organic physics. Proceedings of the American Philosophical Society, 12(86):311–321, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981712.

Hartshorne:1876:SDP

[Har76] Henry Hartshorne. On some disputed points in physiological optics. Proceedings of the American Philosophical Society, 16(97): 218–229, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982433.

Hartshorne:1880:MGB

[Har80] Henry Hartshorne. Memoir of George B. Wood, M.D., LL.D. Proceedings of the American Philosophical Society, 19(107):118–152, March/ December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982618.

Hartshorne:1895:MJR

[Har95] Henry Hartshorne. Memoir of James E. Rhoads, M.D., LL.D. Proceedings of the American Philosophical Society, 34(149):354–357, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982888.

Haupt:1871:CEG

[Hau71] Herman Haupt. Computation of effect of gredients. Proceedings of the American Philosophical Society, 12(86):9–11, 1871. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/981657.

Haupt:1877:LNC

[Hau77] H. Haupt. Level notes and compass courses of the Seaboard Oil Pipe Line, from the the mouth of Black Fox Run, in Clarion County, Pa., to Patapsco River, near Baltimore, Md. lines run by O. Barrett, Jr., C. E., Western Division; B. F. Warren, C. E., Middle Division; J. B. Haupt, C. E., Eastern Division. Proceedings of the American Philosophical Society, 17(100):136–144, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982289.

Haupt:1878:SM

[Hau78] L. M. Haupt. On the scales of maps. Proceedings of the American Philosophical Society, 18(102):47–57, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982758.

Haupt:1879:CVM

[Hau79] Lewis M. Haupt. On the coördination of the various methods of expressing thought as applied to the system of public school instruction. Proceedings of the American Philosophical Society, 18(104): 348–351, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982509.

Haupt:1888:PPH

[Hau88] Lewis M. Haupt. The physical phenomena of harbor entrances. Their causes and remedies. Defects of present methods of improvement. Proceedings of the American Philosophical Society, 25(127): 19–41, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983278.

Haupt:1889:DDA

[Hau89] Lewis M. Haupt. Discussion on the dynamic action of the ocean in building bars. Proceedings of the American Philosophical Society, 26 (129):146–171, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982941.

Haupt:1899:HRE

[Hau99] Lewis M. Haupt. Historical résumé of the efforts made to demonstrate the practicability of the reaction breakwater at Aransas Pass, Tex., with results to February, 1899. Proceedings of the American Philosophical Society, 38(160):135–147, October 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983562.

Hayden:1868:BNP

[Hay68a] F. V. Hayden. Brief notes on the Pawnee, Winnebago, and Omaha languages. Proceedings of the American Philosophical Society, 10(78):389–421, January 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981592.

Hayden:1868:NGW

[Hay68b] F. V. Hayden. Notes on the geology of Wyoming and Colorado territories. — No. 1. Proceedings of the American Philosophical Society, 10(80):463–478, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981791.

Hayden:1869:NGW

[Hay69a] F. V. Hayden. Notes on the geology of Wyoming and Colorado Territories. No. 2. Proceedings of the American Philosophical Society, 11(81):25–56, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981439.

Hayden:1869:SSB

[Hay69b] F. V. Hayden. Sections of strata belonging to the “Bear River Group,” near Bear River City, Wyoming Territory. Proceedings of the American Philosophical Society, 11(81):420–425, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981494.

Haynes:1877:AAB

[Hay77] Francis L. Haynes. On the asserted antagonism between nicotin and strychnia. Proceedings of the American Philosophical Society, 16(99):596–616, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982486.

Hays:1898:JKD

[Hay98a] I. Minis Hays. A journal kept during the siege of Fort William Henry, August, 1757. Proceedings of the American Philosophical Society, 37(157):143–150, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983702.

Hays:1898:NHJ

[Hay98b] I. Minis Hays. A note on the history of the Jefferson manuscript draught of the Declaration of Independence in the Library of the American Philosophical Society. Proceedings of the American Philosophical Society, 37(157):88–107, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983698.

Hays:1899:NWP

[Hay99] I. Minis Hays. A note on William Penn’s commission for the Government of Pennsylvania during his first visit to England in 1684. Proceedings of the American Philosophical Society, 38(159):4–6, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983613.

Heilprin:1891:OFN

[Hei91] Angelo Heilprin. Observations on the flora of Northern Yucatan. Proceedings of the American Philosophical Society, 29(136):137– 144, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982931.

Heilprin:1892:TAF

[Hei92] Angelo Heilprin. The temperate and alpine floras of the giant volcanoes of Mexico. (Being a report from the Committee on the Michaux Legacy). Proceedings of the American Philosophical Society, 30(137):4–22, January 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983205.

Hiller:1896:BRJ

[Hil96] H. M. Hiller. A brief report of a journey up the Rejang River in Borneo. Proceedings of the American Philosophical Society, 35 (153):321–328, December 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983593.

Houston:1897:IMS

[HK97] Edwin J. Houston and A. E. Kennelly. The insulating medium surrounding a conductor the real path of its current. Proceedings of the American Philosophical Society, 36(155):144–175, May 1897. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983649.

Hoffman:1874:PUC

[Hof74] W. J. Hoffman. Pah-Ute cremation. Proceedings of the American Philosophical Society, 14(93):297–298, June/December 1874. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982120.

Hoffman:1875:CAD

[Hof75] W. J. Hoffman. On cremation among the Digger Indians. Proceedings of the American Philosophical Society, 14(94):414–415, January/ June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981870.

Hoffman:1886:RIT

[Hof86a] W. J. Hoffman. Remarks on Indian tribal names. Proceedings of the American Philosophical Society, 23(122):294–303, April 1886. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983243.

Hoffman:1886:VSL

[Hof86b] W. J. Hoffman. Vocabulary of the Selish language. Proceedings of the American Philosophical Society, 23(123):361–371, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983219.

Hoffman:1886:VWD

[Hof86c] W. J. Hoffman. Vocabulary of the Waitshum’ni dialect, of the Kawi’a language. Tule Agency, Cal. Proceedings of the American Philosophical Society, 23(123):372–379, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983220.

Hoffman:1889:FMP

[Hof89a] W. J. Hoffman. Folk-medicine of the Pennsylvania Germans. Proceedings of the American Philosophical Society, 26(129):329–352, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982955.

Hoffman:1889:GNV

[Hof89b] W. J. Hoffman. Grammatic notes and vocabulary of the Pennsylvania German dialect. Proceedings of the American Philosophical Society, 26(129):187–285, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982945.

Hoffman:1894:GFA

[Hof94] W. J. Hoffman. Gshicht fun dä al’tä tsai’tä in Pensilfâni. (Pennsylvania Dutch) [Sketch of the olden times in Pennsylvania]. Proceedings of the American Philosophical Society, 32(143):325–331, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983043.

Horn:1873:CKC

[Hor73a] George H. Horn. Contributions to a knowledge of the Curculionidæ of the United States. Proceedings of the American Philosophical Society, 13(90):407–469, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981644.

Horn:1873:RGS

[Hor73b] George H. Horn. Revision of the genera and species of the tribe Hydrobiini. Proceedings of the American Philosophical Society, 13(90): 118–137, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981607.

Horn:1873:RSS

[Hor73c] George H. Horn. Revision of the species of several genera of Meloidæ of the United States. Proceedings of the American Philosophical Society, 13(90):88–117, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981606.

Horn:1873:SHU

[Hor73d] George H. Horn. Synopsis of the Histeridæ of the United States. Proceedings of the American Philosophical Society, 13(90):273–360, 1873.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981642.

Horn:1878:RSS

[Hor78a] George H. Horn. Revision of the species of the sub-family “Bostrichidæ” of the United States. Proceedings of the American Philosophical Society, 17(101):540–555, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982655.

Horn:1878:SCU

[Hor78b] George H. Horn. Synopsis of the “Colydiidæ” of the United States. Proceedings of the American Philosophical Society, 17(101): 555–592, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982656.

Horn:1879:MRS

[Hor79a] George H. Horn. A monographic revision of the species of “Cremastochilus” of the United States. Proceedings of the American Philosophical Society, 18(104):382–397, 408, July/December 1879. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982522.

Horn:1879:SEU

[Hor79b] George H. Horn. Synopsis of the “Euphoriæ” of the United States. Proceedings of the American Philosophical Society, 18(104): 397–408, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/3473021.

Horn:1880:CNS

[Hor80a] George H. Horn. Critical notes on the species of “Selenophorus” of the United States. Proceedings of the American Philosophical Society, 19(107):178–183, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982623.

Horn:1880:RSA

[Hor80b] George H. Horn. A review of the species of “Anisodactylus” inhabiting the United States. Proceedings of the American Philosophical Society, 19(107):162–178, March/December 1880. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982622.

Horn:1883:MJL

[Hor83] George H. Horn. Memoir of John L. LeConte, M.D. Proceedings of the American Philosophical Society, 21(114):294–300, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982391.

Horn:1891:NCL

[Hor91] George H. Horn. Notes on calospasta lec. Proceedings of the American Philosophical Society, 29(136):99–102, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982924.

Houston:1874:SAM

[Hou74] Edwin J. Houston. On a supposed allotropic modification of phosphorus. Proceedings of the American Philosophical Society, 14(92): 108–110, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982065.

Houston:1883:SMT

[Hou83] Houston. On the synchronous-multiplex telegraph. Proceedings of the American Philosophical Society, 21(114):326–328, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982398.

Houston:1886:PLF

[Hou86] Edwin J. Houston. Photography by a lightning flash. Proceedings of the American Philosophical Society, 23(122):257–259, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983236.

Houston:1888:SPM

[Hou88] Edwin J. Houston. On some possible methods for the preparation of gramophone and telephone records. Proceedings of the American Philosophical Society, 25(127):144–148, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983297.

Houston:1890:MCF

[Hou90] Edwin J. Houston. On muscular contractions following death by electricity. Proceedings of the American Philosophical Society, 28 (132):37–38, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983307.

Hubbard:1878:CM

[HSL78] H. G. Hubbard, E. A. Schwarz, and John L. LeConte. The Coleoptera of Michigan. Proceedings of the American Philosophical Society, 17 (101):593–669, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982657.

Houston:1878:CIE

[HT78] Edwin J. Houston and Eiihu Thomson. Circumstances influencing the efficiency of dynamo-electric machines. Proceedings of the American Philosophical Society, 18(102):58–61, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982759.

Hubbard:1894:YNC

[Hub94] Jesse W. Hubbard. The yolk nucleus in Cymatogaster Aggregatus Gibbons. Proceedings of the American Philosophical Society, 33(144):74– 83, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983362.

Hunt:1888:CNM

[Hun88] T. Sterry Hunt. The classification and nomenclature of metalline minerals. Proceedings of the American Philosophical Society, 25(128): 170–180, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983050.

Hyatt:1894:PAC

[Hya94] Alpheus Hyatt. Phylogeny of an acquired characteristic. Proceedings of the American Philosophical Society, 32(143):349–647, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983047.

James:1895:EEP

[Jam95] Edmund J. James. An early essay on proportional representation. Proceedings of the American Philosophical Society, 34(149):468–471, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982902.

Jayne:1882:RDU

[Jay82] Horace F. Jayne. Revision of the Dermestidæ of the United States. Proceedings of the American Philosophical Society, 20(112): 343–377, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982685.

Jones:1880:NCP

[Jon80] Howard Grant Jones. Notes on the Cumberland or Potomac Coal Basin. Proceedings of the American Philosophical Society, 19(107): 111–116, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982615.

Jordan:1888:APM

[Jor88] Francis Jordan, Jr. Aboriginal pottery of the Middle Atlantic States. Proceedings of the American Philosophical Society, 25(127): 104–107, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983287.

Kane:1877:C

[Kan77] Thomas L. Kane. Coahuila. Proceedings of the American Philosophical Society, 16(99):561–573, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982480.

Keating:1878:RUE

[Kea78] J. P. Keating. Railway under the English Channel. Address of the President, M. Michael Chevalier. Proceedings of the American Philosophical Society, 17(101):283–290, January/June 1878. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982633.

Kerr:1873:TAR

[Ker73] W. C. Kerr. Topography as affected by the rotation of the Earth. Proceedings of the American Philosophical Society, 13(90):190–192, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981623.

Keyes:1888:APP

[Key88] Charles R. Keyes. On the attachment of Platyceras to Palæocrinoids, and its effects in modifying the form of the shell. Proceedings of the American Philosophical Society, 25(128):231–243, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983060.

King:1888:EMV

[Kin88] C. W. King. Epitaph of M. Verrius Flaccus. Proceedings of the American Philosophical Society, 25(127):55–59, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983280.

Kirkwood:1868:SNM

[Kir68] Daniel Kirkwood. On the shower of November, meteors as observed at Bloomington, Indiana, November 13–14, 1868. Proceedings of the American Philosophical Society, 10(80):541–542, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981801.

Kirkwood:1869:CM

[Kir69a] Daniel Kirkwood. On comets and meteors. Proceedings of the American Philosophical Society, 11(81):215–220, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981463.

Kirkwood:1869:MAB

[Kir69b] Daniel Kirkwood. On the mass of asteroids between Mars and Jupiter. Proceedings of the American Philosophical Society, 11(81):498, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981501.

Kirkwood:1869:PSS

[Kir69c] Daniel Kirkwood. On the periodicity of the solar spots. Proceedings of the American Philosophical Society, 11(81):94–102, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981444.

Kirkwood:1869:PCM

[Kir69d] Daniel Kirkwood. On the periods of certain meteoric rings. Proceedings of the American Philosophical Society, 11(81):299–300, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981480.

Kirkwood:1871:SRR

[Kir71a] Daniel Kirkwood. On some remarkable relations between the mean motions of Jupiter, Saturn, Uranus, and Neptune. Proceedings of the American Philosophical Society, 12(86):435–436, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981733.

Kirkwood:1871:FPS

[Kir71b] Daniel Kirkwood. On the formation and primitive structure of the solar system. Proceedings of the American Philosophical Society, 12(86): 163–167, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981690.

Kirkwood:1873:MJ

[Kir73] Daniel Kirkwood. On the meteors of January 2nd. Proceedings of the American Philosophical Society, 13(90):501–502, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981649.

Kirkwood:1877:EMF

[Kir77a] Daniel Kirkwood. On eight meteoric fireballs seen in the United States from July, 1876, to February, 1877. Proceedings of the American Philosophical Society, 16(99):590–596, January/May 1877. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982485.

Kirkwood:1877:RAS

[Kir77b] Daniel Kirkwood. On the relative ages of the Sun and certain of the fixed stars. Proceedings of the American Philosophical Society, 16(99):622–625, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/982489. See errata [Ano77c].

Kirkwood:1878:AEN

[Kir78] Daniel Kirkwood. On the aerolitic epoch of November 12th– 13th. Proceedings of the American Philosophical Society, 17(101): 339–341, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982642.

Kirkwood:1879:CL

[Kir79a] Daniel Kirkwood. The cosmogony of Laplace. Proceedings of the American Philosophical Society, 18(104):324–326, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982505.

Kirkwood:1879:MFS

[Kir79b] Daniel Kirkwood. On meteoric fireballs seen in the United States during the year ending march 31, 1879. Proceedings of the American Philosophical Society, 18(103):239–247, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982333.

Kirkwood:1880:OP

[Kir80] Daniel Kirkwood. On the origin of planets. Proceedings of the American Philosophical Society, 19(107):15, March/December 1880. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982599.

Kirkwood:1883:ZAR

[Kir83] Daniel Kirkwood. The zone of asteroids and the Ring of Saturn. Proceedings of the American Philosophical Society, 21(114):263–266, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982385.

Kirkwood:1885:CMN

[Kir85a] Daniel Kirkwood. The comet of 1866 and the meteors of November 14th. Proceedings of the American Philosophical Society, 22 (120):424–428, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983003.

Kirkwood:1885:LSN

[Kir85b] Daniel Kirkwood. The limits of stability of nebulous planets, and the consequences resulting from their mutual relations. Proceedings of the American Philosophical Society, 22(118):104–113, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982672.

Kirkwood:1887:BCL

[Kir87a] Daniel Kirkwood. Biela’s Comet and the large meteors of November 27–30. Proceedings of the American Philosophical Society, 24 (126):242–243, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983072.

Kirkwood:1887:NPE

[Kir87b] Daniel Kirkwood. Note on the possible existence of fireballs and meteorites in the stream of bielids. Proceedings of the American Philosophical Society, 24(126):436, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983094.

Kirkwood:1887:RAS

[Kir87c] Daniel Kirkwood. The relation of aerolites to shooting stars. Proceedings of the American Philosophical Society, 24(125):111–112, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983127.

Kirkwood:1889:IA

[Kir89] Daniel Kirkwood. On the inclinations of the asteroids. Proceedings of the American Philosophical Society, 26(130):428–429, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983182.

Kirkwood:1892:MRB

[Kir92] Daniel Kirkwood. On the mutual relations between the orbits of certain asteroids. Proceedings of the American Philosophical Society, 30(139):269–270, December 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983343.

Koenig:1874:CLU

[Koe74] Geo. A. Koenig. Contributions from the Laboratory of the University of Pennsylvania. No. II. On an improvement of the burette valve. Proceedings of the American Philosophical Society, 14(93): 218–223, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982112.

Koenig:1885:NCA

[Koe85] George A. Koenig. Note on cosalite, alaskaite and beegerite. Proceedings of the American Philosophical Society, 22(118):211–213, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982678.

Konig:1877:CLU

[Kön77a] George Augustus König. Contributions from the Laboratory of the University of Pennsylvania. No. VII. On astrophyllite, arfvedsonite and zircon, from El Paso Co. Colorado, and a colorimetric estimation of titanium before the blow-pipe. Proceedings of the American Philosophical Society, 16(99):509–518, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982478.

Konig:1877:NEF

[Kön77b] Rudolph König. Note on the exactitude of the French Normal Fork; a reply to the paper of Mr. A. J. Ellis. Proceedings of the American Philosophical Society, 17(100):80–82, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982282.

Konig:1878:CUP

[Kön78] George Aug. König. Contributions from the University of Pa., No. XV. Preliminary notice on chromometry, a new branch of quantitative analysis with the blowpipe. Proceedings of the American Philosophical Society, 18(102):29–34, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982752.

Konig:1881:CUP

[Kön81] George A. König. Contributions from the University of Pennsylvania, no. 16. On alaskaite, a new member from the series of bismuth sulphosalts. Proceedings of the American Philosophical Society, 19 (109):472–477, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982790.

Krauss:1886:BH

[Kra86] Friedrich S. Krauss. Aus Bosnien und der Hercegovina. Proceedings of the American Philosophical Society, 23(121):87–94, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982914.

Krauss:1888:ING

[Kra88] Friedrich S. Krauss. Ibrâhim Nukič, Ein Guslarenlied der Mohammedanischen Slaven in der Hercegovina. (German) [Ibrâhim Nukič, a gusle song of the Mohammedan slaves in Hercegovina].

Proceedings of the American Philosophical Society, 25(128):183– 194, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983052.

Krauss:1894:WMK

[Kra94] Friedrich S. Krauss. Wie Mohammed Köprülü Vezier geworden. Ein Guslarenlied der slavischen Mohammedaner im Herzogtum. Aufgezeichnet, verdeutscht and erläutert. (German) [How Mohammed became köprülü Vezier. a guslar song of the Slavic Mohammedans in the Duchy. Recorded, Germanized and explained]. Proceedings of the American Philosophical Society, 32(143):293–325, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983042.

Law:1885:ONW

[Law85] Philip H. Law. Obituary notice of William S. Vaux. Proceedings of the American Philosophical Society, 22(120):404–405, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983000.

Law:1888:OGU

[Law88] Philip H. Law. Observations on Gildas and the uncertainties of Early English history. Proceedings of the American Philosophical Society, 25 (127):132–135, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983295.

Lesley:1891:AWL

[LC91] J. P. Lesley and Oscar C. S. Carter. Artesian well in lowest trias at Norristown. Proceedings of the American Philosophical Society, 29(135):43–49, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983018.

LeConte:1873:ONJ

[LeC73] John L. LeConte. Obituary notice of John F. Frazer. Proceedings of the American Philosophical Society, 13(90):183–190, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981622.

LeConte:1877:TSR

[LeC77] John LeConte. Tabular synopsis of the Rhynchophora of America. Proceedings of the American Philosophical Society, 16(99):

417–424, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982474.

LeConte:1878:ADN

[LeC78] John L. LeConte. Additional descriptions of new species. Proceedings of the American Philosophical Society, 17(101):373–472, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982648.

Leidy:1887:BNI

[Lei87] Joseph Leidy. Biographical notice of Isaac Lea, LL.D. Proceedings of the American Philosophical Society, 24(126):400–403, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983082.

Leonard:1896:NPP

[Leo96] Charles Lester Leonard. New physical phenomena of the X-ray. Proceedings of the American Philosophical Society, 35(153):298–301, December 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983588.

Lesley:1859:IGW

[Les59] J. P. Lesley. On the insensible graduation of words. Proceedings of the American Philosophical Society, 7(62):129–155, July/December 1859. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981822.

Lesley:1862:SCM

[Les62] J. P. Lesley. Section of coal-measures on the Cape Breton coast. Proceedings of the American Philosophical Society, 9(68):93–109, June/ December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982048.

Lesley:1868:NSH

[Les68a] J. P. Lesley. Notes on some of the historical and mythological features of the D’Orbiney Papyrus. Proceedings of the American Philosophical Society, 10(80):543–582, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981802.

Lesley:1868:ONW

[Les68b] J. P. Lesley. Obituary notice of Wm. Parker Foulke. Proceedings of the American Philosophical Society, 10(80):481–510, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981794.

Lesley:1871:GST

[Les71a] J. P. Lesley. The geological structure of Tazewell, Russell and Wise Counties, in Virginia. Proceedings of the American Philosophical Society, 12(86):489–513, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981750.

Lesley:1871:NFU

[Les71b] J. P. Lesley. Note on a fine upthrow fault at Embreeville Furnace in East Tennessee. Proceedings of the American Philosophical Society, 12(86):444–457, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981739.

Lesley:1871:NAV

[Les71c] J. P. Lesley. Note on an apparent violation of the law of regular progressive debituminisation of the American coal beds coming east. Proceedings of the American Philosophical Society, 12(86):125–138, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981684.

Lesley:1871:NTI

[Les71d] J. P. Lesley. Note on the titaniferous iron ore belt, near Greensboro, North Carolina. Proceedings of the American Philosophical Society, 12(86):139–158, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981685.

Lesley:1871:RFO

[Les71e] J. P. Lesley. A record of fourteen oil wells at Brady’s Bend, Armstrong County, Pennsylvania. Proceedings of the American Philosophical Society, 12(86):562–570, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981764.

Lesley:1873:IOS

[Les73a] J. P. Lesley. The iron ores of the South Mountain. Along the line of the Harrisburg and Potomac Railway, in Cumberland County, Pennsylvania. Proceedings of the American Philosophical Society, 13(90): 3–21, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/981598.

Lesley:1873:SCB

[Les73b] J. P. Lesley. St. Clairsville and Bedford Railroad; and Dunning’s Creek fossil iron ore. Proceedings of the American Philosophical Society, 13(90):156–168, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981617.

Lesley:1873:SSE

[Les73c] J. P. Lesley. A study of the structure and erosion of Brush Mountain, in Blair County, Pennsylvania (with a map). Proceedings of the American Philosophical Society, 13(90):503–504, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981650.

Lesley:1873:OBE

[Les73d] Joseph Lesley. The outcrop belt of the East Kentucky Coal Field. Proceedings of the American Philosophical Society, 13(90):270–272, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981641.

Lesley:1874:BHO

[Les74] J. P. Lesley. The brown hematite ore banks of Spruce Creek, Warrior’s Mark Run, and Half Moon Run, in Huntingdon and Centre Counties, Pennsylvania, along the line of the Lewisburg, Centre County and Tyrone Railroad. Proceedings of the American Philosophical Society, 14(92):19–83, 99–107, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982063.

Lesley:1876:CPG

[Les76] J. P. Lesley. Contributions to the physical geography of the United States, by Charles Allen, Assistant in Charge of the Collection and Collation of Railroad and Other Levels for the Second Geological Survey of Pennsylvania. Proceedings of the American Philosophical Society, 16(97):61–179, January/June 1876. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982422.

Lesley:1877:SCA

[Les77a] J. P. Lesley. On a series of chemical analysis of Siluro–Cambrian limestone beds in Cumberland County, Pennsylvania. Proceedings of the American Philosophical Society, 17(100):260–266, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982307.

Lesquereux:1877:PNA

[Les77b] L. Lesquereux. On the progress of the North American carboniferous flora, in preparation for the Second Geological Survey of Pennsylvania. Proceedings of the American Philosophical Society, 16 (99):397–416, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982473.

Lesquereux:1877:LPR

[Les77c] Leo Lesquereux. Land plants, recently discovered in the Silurian rocks of the United States. Proceedings of the American Philosophical Society, 17(100):163–173, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982292.

Lesquereux:1877:SFR

[Les77d] Leo Lesquereux. A species of fungus recently discovered in the shales of the Darlington coal bed (Lower Productive Coal Measures, Alleghany River Series) at Cannelton, in Beaver County, Pennsylvania. Proceedings of the American Philosophical Society, 17(100): 173–175, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982293.

Lesley:1878:NSA

[Les78a] J. P. Lesley. Notes on a series of analyses of the dolomitic limestone rocks of Cumberland County, Pa., made by Messrs. Hartshorne and Hartranft in the Laboratory of the Second Geological Survey of Pennsylvania. Proceedings of the American Philosophical Society, 18 (102):114–122, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982767.

Lesquereux:1878:CTR

[Les78b] L. Lesquereux. On the cordaites and their related generic divisions, in the carboniferous formation of the United States. Proceedings of the American Philosophical Society, 17(101):315–335, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982639.

Lesquereux:1879:BCB

[Les79] Leo Lesquereux. On a branch of cordaites, bearing fruit. Proceedings of the American Philosophical Society, 18(103):222–223, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982328.

Lesley:1880:NME

[Les80a] J. P. Lesley. Notes on the models exhibited at the meeting November 19, 1880. Proceedings of the American Philosophical Society, 19(107): 193–194, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982626.

Lesley:1880:ONL

[Les80b] J. P. Lesley. An obituary notice of the Late John W. Harden. Proceedings of the American Philosophical Society, 18(105):422–424, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982775.

Lesley:1881:NEEb

[Les81a] [J. P. Lesley]. [Notes on an Egyptian element in the names of Hebrew kings, and its bearing on the history of the Exodus: Conclusion]. Proceedings of the American Philosophical Society, 19(109): 419–435, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982786.

Lesley:1881:NEEa

[Les81b] J. P. Lesley. Notes on an Egyptian element in the names of the Hebrew kings, and its bearing on the history of the Exodus. Proceedings of the American Philosophical Society, 19(108):409–418, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982265.

Lesquereux:1881:CBF

[Les81c] Leo Lesquereux. On a cours de botanique fossile by Prof. M. B. Renault. Proceedings of the American Philosophical Society, 19 (108):287–291, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982251.

Lesley:1882:ONE

[Les82] J. P. Lesley. Obituary notice of Edouard Desor. Proceedings of the American Philosophical Society, 20(112):519–522, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982704.

Lesley:1883:ERS

[Les83a] [J. P. Lesley]. Early records of the society. Proceedings of the American Philosophical Society, 20(113):675–678, January/April 1883. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982465.

Lesley:1883:NPS

[Les83b] J. P. Lesley. Note on the progress of the Second Geological Survey of Pennsylvania. Proceedings of the American Philosophical Society, 20 (113):537–544, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982443.

Lesley:1883:ONE

[Les83c] [J. P. Lesley]. [Obituary notice of Édouard desor]. Proceedings of the American Philosophical Society, 20(113):522–528, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982441.

Lesley:1883:KWS

[Les83d] [J. P. Lesley]. [On the Kingsmill white sandstone]: Comment. Proceedings of the American Philosophical Society, 20(113):673–675, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982464.

See [Cla83f].

Lesquereux:1883:LLL

[Les83e] Leo Lesquereux. Letter of Leo Lesquereux. Proceedings of the American Philosophical Society, 21(114):286–289, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982389.

Lesley:1884:NPG

[Les84] J. P. Lesley. Note on a possible geographical meaning for the set griffin. Proceedings of the American Philosophical Society, 21(115):455– 456, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982739.

Lesley:1886:ONJ

[Les86a] J. P. Lesley. An obituary notice of James Macfarlane. Proceedings of the American Philosophical Society, 23(122):287–289, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983240.

Lesley:1886:HWS

[Les86b] J. P. Lesley. On the Hebrew word ShDI (shaddai), translated “The Almighty”. Proceedings of the American Philosophical Society, 23(122):303–312, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983244.

Lesley:1888:ONF

[Les88] J. P. Lesley. Obituary notice of Ferdinand Vandevere Hayden, M.D., Ph.D., LL.D. Proceedings of the American Philosophical Society, 25 (127):59–61, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983281.

Lesley:1890:ONC

[Les90a] J. P. Lesley. Obituary notice of Charles Albert Ashburner. Proceedings of the American Philosophical Society, 28(132):53–59, January/ June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983309.

Lesley:1890:ONL

[Les90b] J. P. Lesley. Obituary notice of Leo Lesquereux. Proceedings of the American Philosophical Society, 28(132):65–70, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983312.

Lesley:1891:NHE

[Les91a] J. P. Lesley. Notes on Hebrew etymologies from the Egyptian ANX. Enoch; Anoki; Enos. Proceedings of the American Philosophical Society, 29(135):17–20, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983014.

Lesley:1891:ONP

[Les91b] J. P. Lesley. Obituary notice of P. W. Sheafer. Proceedings of the American Philosophical Society, 29(135):39–42, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983017.

Lesley:1891:IBT

[Les91c] J. P. Lesley. On an important boring through 2000 feet of trias, in Eastern Pennsylvania. Proceedings of the American Philosophical Society, 29(135):20–25, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983015.

Lesley:1891:GGW

[Les91d] J. P. Lesley. On the Grapeville gas-wells. Proceedings of the American Philosophical Society, 29(135):11–16, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983013.

Lewis:1882:NAA

[Lew82a] H. Carvill Lewis. Note on the aurora of April 16–17, 1882. Proceedings of the American Philosophical Society, 20(111):283–291, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982725.

Lewis:1882:NSR

[Lew82b] Henry Carvill Lewis. On a new substance resembling dopplerite from a peat bog at Scranton. Proceedings of the American Philosophical Society, 20(111):112–117, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982708.

Lewis:1885:GTD

[Lew85] H. Carvill Lewis. A great trap dyke across Southeastern Pennsylvania. Proceedings of the American Philosophical Society, 22(120):438–456, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983008.

Leyburn:1860:ONJ

[Ley60] John Leyburn. Obituary notice of Joseph Addison Alexander, D.D. Proceedings of the American Philosophical Society, 7(63): 320–325, January/June 1860. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981894.

LeConte:1876:RAN

[LH76] John L. LeConte and George H. Horn. The Rhynchophora of America, north of Mexico. Proceedings of the American Philosophical Society, 15(96):vii–ix, xi–xvi, 1–442, December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982412.

Lilley:1883:DSC

[Lil83] A. T. Lilley. Detailed section of Chemung rocks exposed in the Gulf Brook Gorge at Le Roy, Bradford County, Pennsylvania. Proceedings of the American Philosophical Society, 21(114):304–305, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982393.

Lilley:1886:RSC

[Lil86] A. T. Lilley. A revision of the section of Chemung rocks exposed in the Gulf Brook Gorge at LeRoy, in Bradford County, Pennsylvania. Proceedings of the American Philosophical Society, 23(122):291–293, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983242.

Lockington:1883:RPP

[Loc83] W. N. Lockington. The role of parasitic protophytes. Are they the primary, or the secondary cause of zymotic diseases? Proceedings of the American Philosophical Society, 21(114):88–93, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982364.

Lowrie:1869:SNC

[Low69a] Walter H. Lowrie. A search for a normal cause of the recession of cosmical nodes. Proceedings of the American Philosophical Society, 11(81):220–225, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981464.

Lowrie:1869:SSM

[Low69b] Walter H. Lowrie. Some suggestions on the maintaining forces of cosmical motion. Proceedings of the American Philosophical Society, 11(81):195–201, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981456.

Lyman:1893:GMF

[Lym93] Benjamin Smith Lyman. The Great Mesozoic Fault in New Jersey. Proceedings of the American Philosophical Society, 31(142): 314–317, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982968.

Lyman:1894:ANB

[Lym94a] Benj. Smith Lyman. Age of the Newark Brownstone. Proceedings of the American Philosophical Society, 33(144):5–9, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983359.

Lyman:1894:SCM

[Lym94b] Benj. Smith Lyman. Some coal measure sections near Peytona, West Virginia. Proceedings of the American Philosophical Society, 33(146): 282–309, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982841.

Lyman:1894:FAN

[Lym94c] Benjamin Smith Lyman. Further on the age of the Newark Brownstone. Proceedings of the American Philosophical Society, 33(144):9– 10, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983360.

Lyman:1894:SNR

[Lym94d] Benjamin Smith Lyman. Some new red horizons. Proceedings of the American Philosophical Society, 33(145):192–215, June 1894. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983268.

Lyman:1895:CFR

[Lym95a] Benjamin Smith Lyman. The Chalfont fault rock, so called. Proceedings of the American Philosophical Society, 34(149):384–388, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982893.

Lyman:1895:YF

[Lym95b] Benjamin Smith Lyman. The Yardley Fault. Proceedings of the American Philosophical Society, 34(149):381–384, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982892.

Mabery:1897:CAP

[Mab97] Charles F. Mabery. On the composition of American petroleum. Proceedings of the American Philosophical Society, 36(154):126–140, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983467.

Mansfield:1877:RT

[Man77] A. K. Mansfield. On refraction tables. Proceedings of the American Philosophical Society, 16(99):425–428, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982475.

Marsh:1871:CDN

[Mar71] O. C. Marsh. Communication on the discovery of new rocky mountain fossils. Proceedings of the American Philosophical Society, 12(86): 578–579, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981768.

Marsh:1873:GMA

[Mar73] O. C. Marsh. On the gigantic mammals of the American Eocene. Proceedings of the American Philosophical Society, 13(90):255–256, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981632.

Marsh:1874:LHE

[Mar74] B. V. Marsh. On the latent heat of expansion in connection with the luminosity of meteors, etc. Proceedings of the American Philosophical Society, 14(92):114–129, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982068.

Marshall:1878:CLU

[Mar78] John Marshall. Contributions from the Laboratory of the University of Pennsylvania. No. XIII. A study of some of the derivatives of monoand dichlor-salicylic acids. Proceedings of the American Philosophical Society, 17(101):476–489, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982650.

Mathews:1897:ARC

[Mat97a] R. H. Mathews. Australian rock carvings. Proceedings of the American Philosophical Society, 36(155):195–208, May 1897. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983654.

Mathews:1897:RCP

[Mat97b] R. H. Mathews. Rock carvings and paintings of the Australian Aborigines. Proceedings of the American Philosophical Society, 36 (156):466–478, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/ stable/983402. See corrections [Ano97f].

Mathews:1898:DAT

[Mat98a] R. H. Mathews. Divisions of Australian tribes. Proceedings of the American Philosophical Society, 37(157):151–154, January 1898. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983703.

Mathews:1898:DQA

[Mat98b] R. H. Mathews. Divisions of Queensland Aborigines. Proceedings of the American Philosophical Society, 37(158):327–336, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983859.

Mathews:1898:ICA

[Mat98c] R. H. Mathews. Initiation ceremonies of Australian tribes. Proceedings of the American Philosophical Society, 37(157):54–73, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983694.

Mathews:1899:DNA

[Mat99] R. H. Mathews. Divisions of North Australian tribes. Proceedings of the American Philosophical Society, 38(159):75–79, January 1899.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983625.

Mayer:1869:ASR

[May69] Alfred Mayer. An abstract of some of the results of measurements and examinations of the photographs of the total solar eclipse of August 7, 1869. Proceedings of the American Philosophical Society, 11(81):204–208, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981459.

McCauley:1881:MUS

[McC81] Edward Yorke McCauley. A manual for the use of students in Egyptology. Proceedings of the American Philosophical Society, 20(110): 1–89, 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982415.

Meek:1869:PLF

[Mee69] F. B. Meek. A preliminary list of fossils, collected by Dr. Hayden in Colorado, New Mexico and California, with brief descriptions of a few of the new species. Proceedings of the American Philosophical Society, 11(81):425–431, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981495.

Meinert:1886:MMC

[Mei86] Fr. Meinert. Myriapoda Musei Cantabrigensis, Mass. Part I. Chilopoda. Proceedings of the American Philosophical Society, 23 (122):161–233, April 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983233.

Merrick:1890:OND

[Mer90] J. Vaughan Merrick. Obituary notice of Daniel Raynes Goodwin, D.D., LL.D. Proceedings of the American Philosophical Society, 28 (134):227–241, July/December 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983113.

Mercer:1895:JSQ

[Mer95] H. C. Mercer. Jasper and stalagmite quarried by Indians in the Wyandotte Cave. Proceedings of the American Philosophical Society, 34(149):396–400, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982896.

Mercer:1897:FRF

[Mer97a] Henry C. Mercer. The finding of the remains of the fossil sloth at Big Bone Cave, Tennessee, in 1896. Proceedings of the American Philosophical Society, 36(154):36–70, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983459.

Mercer:1897:SMA

[Mer97b] Henry C. Mercer. The survival of the medieval art of illuminative writing among Pennsylvania Germans. Proceedings of the American Philosophical Society, 36(156):424–433, December 1897. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983391.

Meyer:1888:MIV

[Mey88] Otto Meyer. On Miocene invertebrates from Virginia. Proceedings of the American Philosophical Society, 25(127):135–144, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983296.

Miller:1897:TSG

[Mil97] G. A. Miller. On the transitive substitution groups that are simply isomorphic to the symmetric or the alternating group of degree six. Proceedings of the American Philosophical Society, 36(155):208–215, May 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983655.

Miller:1898:QG

[Mil98] G. A. Miller. On the quaternion group. Proceedings of the American Philosophical Society, 37(158):312–319, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983854.

Mears:1871:CUH

[MKAP71] J. Ewing Mears, W. W. Keen, Harrison Allen, and William Pepper. A case of universal hyperostosis, associated with osteoporosis, with a description of the specimens. Proceedings of the American Philosophical Society, 12(86):19–36, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981662.

Monks:1878:CSS

[Mon78] Sarah P. Monks. The Columella and Stapes in some North American turtles. Proceedings of the American Philosophical Society, 17 (101):335–337, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982640.

Mooney:1887:MMI

[Moo87] James Mooney. The medical mythology of Ireland. Proceedings of the American Philosophical Society, 24(125):136–166, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983129.

Mooney:1888:FCI

[Moo88] James Mooney. The funeral customs of Ireland. Proceedings of the American Philosophical Society, 25(128):243–296, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983061.

Mooney:1889:HCI

[Moo89] James Mooney. The holiday customs of Ireland. Proceedings of the American Philosophical Society, 26(130):377–427, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983181.

Morlot:1862:DCA

[Mor62] A. Morlot. On the date of the copper age in the United States. Proceedings of the American Philosophical Society, 9(68):111–114, June/ December 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982050.

Morris:1891:NHP

[Mor91] J. Cheston Morris. Notes on Hebrew phonetics. Proceedings of the American Philosophical Society, 29(135):7–11, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983012.

Morris:1894:ES

[Mor94] J. Cheston Morris. The ethics of solomon. Proceedings of the American Philosophical Society, 33(146):310–332, July/December 1894.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982842.

Morris:1896:GXP

[Mor96] J. Cheston Morris. On Genesis XI.1–9 as a poetic fragment. Proceedings of the American Philosophical Society, 35(153):305–306, December 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983590.

Morris:1897:RPD

[Mor97] J. Cheston Morris. Relation of the pentagonal dodecahedron found near Marietta, Ohio, to shamanism. Proceedings of the American Philosophical Society, 36(155):179–192, May 1897. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983652.

Michaux:1889:PJA

[MS89] André Michaux and C. S. Sargent. Portions of the journal of André Michaux, botanist, written during his travels in the United States and Canada, 1785 to 1796. Proceedings of the American Philosophical Society, 26(129):1–145, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982940.

Muhlenberg:1883:ONR

[Muh83] F. A. Muhlenberg. Obituary notice of the Rev. Dr. Charles Porterfield Krauth. Proceedings of the American Philosophical Society, 20 (113):613–628, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982450.

Nansen:1897:SRN

[Nan97] Fridtjof Nansen. Some results of the Norwegian Polar Expedition, 1893–96. Proceedings of the American Philosophical Society, 36 (156):442–464, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983398.

Nettleton:1877:FSC

[Net77] E. S. Nettleton. On the first systematic collection and discussion of the Venango County oil wells of Western Pennsylvania. Proceedings of the American Philosophical Society, 16(99):429–495, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982476.

Newberry:1882:ODB

[New82] J. S. Newberry. On the origin and drainage of the basins of the great lakes. Proceedings of the American Philosophical Society, 20 (111):91–101, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982705.

Oliver:1886:SAC

[Oli86] Charles A. Oliver. Subjective after-color (complementary color). Proceedings of the American Philosophical Society, 23(124):500–502, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983328.

Oliver:1893:ONJ

[Oli93] Charles A. Oliver. Obituary notice of Joseph Zentmayer. Proceedings of the American Philosophical Society, 31(142):358–364, July/ December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982977.

Oliver:1894:SNU

[Oli94] Charles A. Oliver. A short note upon so-called “hereditary opticnerve atrophy” — as a contribution to the question of transmission of structural peculiarity. Proceedings of the American Philosophical Society, 32(143):269–271, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983039.

Ortmann:1896:NSS

[Ort96] Arnold E. Ortmann. On natural selection and separation. Proceedings of the American Philosophical Society, 35(151):175–192, August 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983261.

Osborn:1887:TMD

[Osb87] Henry F. Osborn. The Triassic mammals, Dromatherium and Microconodon. Proceedings of the American Philosophical Society, 24 (125):109–111, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983126.

Outerbridge:1874:ESM

[Out74] Alex E. Outerbridge, Jr. Electrical spectra of metals. Results of an examination as to the practicability of assaying metals used in coinage, by means of spectrum analysis, made in and for the Assay Department of the U. S. Mint at Philadelphia. Proceedings of the American Philosophical Society, 14(92):162–173, January/June 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982078.

P:1873:APR

[P.73] P. An acoustic phenomenon in the rocky mountains. Proceedings of the American Philosophical Society, 13(90):499–500, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981648.

Packard:1883:MM

[Pac83a] A. S. Packard, Jr. On the morphology of the myriopoda. Proceedings of the American Philosophical Society, 21(114):197–209, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982372.

Packard:1883:RLF

[Pac83b] A. S. Packard, Jr. A revision of the Lysiopetalidæ, a family of Chilognath myriopoda, with a notice of the genus Cambala. Proceedings of the American Philosophical Society, 21(114):177–197, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982371.

Packard:1885:ELP

[Pac85] A. S. Packard. On the embryology of limulus polyphemus. III. Proceedings of the American Philosophical Society, 22(120):268–272, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982984.

Packard:1886:DTF

[Pac86] A. S. Packard. Discovery of the thoracic feet in a carboniferous phyllocaridan. Proceedings of the American Philosophical Society, 23(123):380–383, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983221.

Packard:1887:SPM

[Pac87] A. S. Packard. On the systematic position of the Mallophaga. Proceedings of the American Philosophical Society, 24(126):264–272, July/ December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983075.

Packard:1893:LHCa

[Pac93a] A. S. Packard. The life history of certain moths of the family Cochliopodidæ, with notes on their spines and tubercles. Proceedings of the American Philosophical Society, 31(140):83–108, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983106.

Packard:1893:LHCb

[Pac93b] Alpheus S. Packard. The life histories of certain moths of the families Ceratocampidæ, Hemileucidæ, etc., with notes on the armature of the larvæ. Proceedings of the American Philosophical Society, 31 (141):139–192, April/June 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982828.

Packard:1894:STA

[Pac94] Alpheus S. Packard. A study of the transformations and anatomy of Lagoa crispata, a bombycine moth. Proceedings of the American Philosophical Society, 32(143):275–292, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983041.

Patterson:1875:LMR

[Pat75] Robert Patterson. The letter of Mr. Robert Patterson referred to above. Proceedings of the American Philosophical Society, 14 (95):650, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981955.

Patterson:1882:ONW

[Pat82] Robert Patterson. An obituary notice of William E. DuBois. Proceedings of the American Philosophical Society, 20(111):102–107, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982706.

Peale:1861:SII

[Pea61] Franklin Peale. On the stone implements of the Indians of North America, with a classification and description of the methods of making them. Proceedings of the American Philosophical Society, 8(65):265–272, January/June 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982038.

Peale:1863:CAS

[Pea63] F. Peale. Communication on an antique stone hand-hammer. Proceedings of the American Philosophical Society, 9(70):401–403, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982160.

Peckham:1868:NOB

[Pec68] S. F. Peckham. Notes of the origin of bitumens, together with experiments upon the formation of asphaltum. Proceedings of the American Philosophical Society, 10(80):445–462, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981790.

Peckham:1897:NOP

[Pec97] S. F. Peckham. On the nature and origin of petroleum. Proceedings of the American Philosophical Society, 36(154):103–112, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983463.

Peckham:1898:GBR

[Pec98] S. F. Peckham. The genesis of bitumens, as related to chemical geology. Proceedings of the American Philosophical Society, 37(157):108–139, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983699.

Pepper:1883:OJF

[Pep83] William Pepper. Obituary of John Forsyth Meigs, M.D. Proceedings of the American Philosophical Society, 21(114):266–279, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982386.

Pepper:1896:EAT

[Pep96] Edward Pepper. Eucalyptus in Algeria and Tunisia, from an hygienic and climatological point of view. Proceedings of the American Philosophical Society, 35(150):39–56, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982860.

Phillips:1878:DLM

[Phi78] Henry Phillips, Jr. Descriptive list of medals struck to commemorate the Battle of Waterloo. Proceedings of the American Philosophical Society, 18(102):78–83, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982761.

Phillips:1879:ANU

[Phi79a] Henry Phillips, Jr. Additional notes upon the collection of coins and medals now upon exhibition at the Pennsylvania Museum and School of Industrial Art, Memorial Hall, Fairmount Park, Philadelphia. Proceedings of the American Philosophical Society, 18(104): 327–345, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982506.

Phillips:1879:BAE

[Phi79b] Henry Phillips, Jr. A brief account of the earthquake at Aix la Chapelle (Aachen) on Monday, August 26, 1878. Proceedings of the American Philosophical Society, 18(103):216–217, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982326.

Phillips:1879:NUC

[Phi79c] Henry Phillips, Jr. Notes upon the collection of coins and medals now upon exhibition at the Pennsylvania Museum and School of Industrial Art, Memorial Hall, Fairmount Park, Philadelphia. Proceedings of the American Philosophical Society, 18(103):191–206, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982320.

Phillips:1880:AOW

[Phi80a] Henry Phillips, Jr. An account of an old work on cosmography. Proceedings of the American Philosophical Society, 18(105):443–451, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982778.

Phillips:1880:ATM

[Phi80b] Henry Phillips, Jr. An account of two maps of America published respectively in the years 1550 and 1555. Proceedings of the American Philosophical Society, 19(107):10–12, March/December 1880. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982596.

Phillips:1880:SRD

[Phi80c] Henry Phillips, Jr. Some recent discoveries of stone implements in Africa and Asia. Proceedings of the American Philosophical Society, 19(107):63–68, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982607.

Phillips:1881:CAP

[Phi81a] Henry Phillips, Jr. Certain almanacs published in Philadelphia between 1705 and 1744. Proceedings of the American Philosophical Society, 19(108):291–297, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982252.

Phillips:1881:ONP

[Phi81b] Henry Phillips, Jr. Obituary notice of Peter McCall. Proceedings of the American Philosophical Society, 19(108):213–215, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982242.

Phillips:1883:BAM

[Phi83a] Henry Phillips, Jr. A brief account of the more important public collections of American archaeology in the United States. Proceedings of the American Philosophical Society, 21(114):111–119, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982368.

Phillips:1883:NRC

[Phi83b] Henry Phillips, Jr. A note respecting the correct name of the last letter of the English alphabet. Proceedings of the American Philosophical Society, 21(114):330–333, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982400.

Phillips:1884:NUC

[Phi84a] Henry Phillips, Jr. Notes upon the Codex Ramirez, with a translation of the same. Proceedings of the American Philosophical Society, 21(116):616–651, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982343.

Phillips:1884:SRI

[Phi84b] Henry Phillips, Jr. On a supposed Runic inscription at Yarmouth, Nova Scotia. Proceedings of the American Philosophical Society, 21(115):490–492, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982744.

Phillips:1885:P

[Phi85] Henry Phillips, Jr. Preface. Proceedings of the American Philosophical Society, 22(119):i–iii, July 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982525.

Phillips:1888:FCF

[Phi88] Henry Phillips, Jr. First contribution to the folk-lore of Philadelphia and its vicinity. Proceedings of the American Philosophical Society, 25 (128):159–170, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983049.

Phillips:1889:ACI

[Phi89a] Henry Phillips, Jr. An account of the Congo Independent State. Proceedings of the American Philosophical Society, 26(130):459– 476, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983184.

Phillips:1889:ALO

[Phi89b] Henry Phillips, Jr. Alphabetical list of obituary notices published in the Transactions and Proceedings of the American Philosophical Society. Proceedings of the American Philosophical Society, 26 (129):289–291, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982947.

Phillips:1892:SCS

[Phi92] Henry Phillips, Jr. Second contribution to the study of folk-lore in Philadelphia and vicinity. Proceedings of the American Philosophical Society, 30(138):246–249, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983175.

Phillips:1897:GNG

[Phi97a] Francis C. Phillips. On the genesis of natural gas and petroleum. Proceedings of the American Philosophical Society, 36(154):116–121, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983465.

Phillips:1897:OPC

[Phi97b] Francis C. Phillips. On the occurrence of petroleum in the cavities of fossils. Proceedings of the American Philosophical Society, 36(154):121–126, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983466.

Platt:1879:CSS

[Pla79] Franklin Platt. Character of some Sullivan County coals. Proceedings of the American Philosophical Society, 18(103):186–191, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982319.

Potts:1895:BSH

[Pot95] William John Potts. Biographical sketch of the Hon. Thomas H. Dudley, of Camden, N.J., who died April 15, 1893. Proceedings of the American Philosophical Society, 34(147):102–134, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983162.

Price:1871:APM

[Pri71a] E. K. Price. Another phase of modern philosophy. Proceedings of the American Philosophical Society, 12(86):361–391, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981718.

Price:1871:SPM

[Pri71b] Eli K. Price. Some phases of modern philosophy. Proceedings of the American Philosophical Society, 12(86):289–307, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/981709.

Price:1874:ONC

[Pri74] Eli K. Price. An obituary notice of Chief Justice John Meredith Read. Proceedings of the American Philosophical Society, 14(93): 271–283, June/December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982117.

Price:1876:GE

[Pri76a] Eli K. Price. The glacial epochs. Proceedings of the American Philosophical Society, 16(97):241–276, January/June 1876. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982439.

Price:1876:MT

[Pri76b] Eli K. Price. The Michaux trees. Proceedings of the American Philosophical Society, 16(98):340–345, June/December 1876. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982822.

Price:1877:S

[Pri77a] Eli K. Price. Sylviculture. Proceedings of the American Philosophical Society, 17(100):197–211, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982297.

Prime:1877:PRL

[Pri77b] Frederick Prime, Jr. On the Paleozoic rocks of Lehigh and Northampton Counties, Pennsylvania. Proceedings of the American Philosophical Society, 17(100):248–254, June/December 1877. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982304.

Price:1878:NR

[Pri78] Eli K. Price. Nature’s reforesting. Proceedings of the American Philosophical Society, 18(102):26–29, July/December 1878. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/982751.

Price:1882:RUP

[Pri82] Eli K. Price. Rockery at the University of Pennsylvania, built in 1881. Proceedings of the American Philosophical Society, 20(111): 118–122, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982709.

Prime:1894:ONT

[Pri94] Frederick Prime. Obituary notice of Thomas Mutter Cleeman. Proceedings of the American Philosophical Society, 33(145):177–179, June 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983265.

Prime:1895:VRB

[Pri95] Frederick Prime. The varying ratio between gold and silver. Proceedings of the American Philosophical Society, 34(147):49–50, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983158.

Prince:1897:PWR

[Pri97] J. Dyneley Prince. The Passamaquoddy wampum records. Proceedings of the American Philosophical Society, 36(156):479–495, December 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983405.

Prince:1899:SPW

[Pri99] J. Dyneley Prince. Some Passamaquoddy witchcraft tales. Proceedings of the American Philosophical Society, 38(160):181–189, October 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983568.

Platt:1878:SPR

[PS78] Franklin Platt and R. H. Sanders. Section of the Paleozoic rocks in Blair County. Proceedings of the American Philosophical Society, 17 (101):349–352, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982646.

Pennington:1894:AMT

[PS94] Mary E. Pennington and Edgar F. Smith. The atomic mass of tungsten. Proceedings of the American Philosophical Society, 33(146):

332–336, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982843.

Rand:1881:NPO

[Ran81] B. Howard Rand. Note on the protection of oil tanks from lightning stroke. Proceedings of the American Philosophical Society, 19 (108):216, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982243.

Rathbun:1880:E

[Rat80] R. Rathbun. Erratum. Proceedings of the American Philosophical Society, 18(105):510, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982784.

Reichenbach:1863:SS

[Rei63] O. Reichenbach. Solar spots. Proceedings of the American Philosophical Society, 9(70):234–236, June 1863. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982137.

Reid:1871:SDC

[Rei71] Joseph D. Reid. Sketch and description of a carved rock on the bank of the Monongahela River, Pa. Proceedings of the American Philosophical Society, 12(86):11–13, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981658.

Richardson:1898:DFN

[Ric98] Harriet Richardson. Description of four new species of “Rocinela”, with a synopsis of the genus. Proceedings of the American Philosophical Society, 37(157):8–17, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983687.

Rink:1885:RDE

[Rin85] H. Rink. The recent Danish explorations in Greenland and their significance as to Arctic science in general. Proceedings of the American Philosophical Society, 22(120):280–296, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982988.

Ritter:1893:NMD

[Rit93] W. F. McK. Ritter. On a new method of determining the general perturbations of the minor planets. Proceedings of the American Philosophical Society, 31(140):124–127, January 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983109.

Roberts:1871:ONE

[Rob71] Solomon W. Roberts. Obituary notice of Edward Miller, Civil Engineer. Proceedings of the American Philosophical Society, 12(86): 581–586, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981771.

Roberts:1875:ONC

[Rob75] Solomon W. Roberts. An obituary notice of Charles B. Trego. Proceedings of the American Philosophical Society, 14(94):356–358, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981864.

Robinson:1880:ONM

[Rob80] Moncure Robinson. Obituary notice of Michel Chevalier. Proceedings of the American Philosophical Society, 19(107):28–37, March/ December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982604.

Roberts:1882:ONW

[Rob82] W. Milnor Roberts. An obituary notice of William Milnor Roberts. Proceedings of the American Philosophical Society, 20(111):199– 202, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982713.

Robinson:1883:ONH

[Rob83] Moncure Robinson. Obituary notice of Henry Seybert. Proceedings of the American Philosophical Society, 21(114):241–263, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982384.

Rogers:1880:BNP

[Rog80] Fairman Rogers. Biographical notice of Professor Joseph Henry. Proceedings of the American Philosophical Society, 18(105):461–464, January/March 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982781.

Rollett:1894:FGR

[Rol94] Hermann Rollett. Den Forschern. (German) [Researchers]. Proceedings of the American Philosophical Society, 32(143):345–349, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983046.

Rosengarten:1894:CR

[Ros94] Joseph G. Rosengarten. The Château de Rochambeau. Proceedings of the American Philosophical Society, 33(146):353–361, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982847.

Rosengarten:1895:ONH

[Ros95a] J. G. Rosengarten. Obituary notice of Henry Coppée, LL.D. Proceedings of the American Philosophical Society, 34(149):357–361, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982889.

Rosengarten:1895:PBE

[Ros95b] J. G. Rosengarten. The Paris Book Exhibition of 1894. Proceedings of the American Philosophical Society, 34(147):12–21, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983155.

Rosengarten:1895:MPF

[Ros95c] Joseph G. Rosengarten. Memoir of P. F. Rothermel. Proceedings of the American Philosophical Society, 34(149):393–396, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982895.

Rothrock:1877:SLS

[Rot77] J. T. Rothrock. Syllabus of lectures on silvyculture. Proceedings of the American Philosophical Society, 16(99):668–669, January/May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982497.

Rothrock:1882:BST

[Rot82] J. T. Rothrock. Biographical sketch of Thomas Potts James. Proceedings of the American Philosophical Society, 20(111):293–297, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982727.

Rothrock:1883:SMD

[Rot83] J. T. Rothrock. Some microscopic distinctions between good and bad timber of the same species. Proceedings of the American Philosophical Society, 20(113):599–604, January/April 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982447.

Rothrock:1886:BML

[Rot86] J. T. Rothrock. Biographical memoir of the Late Honorable Eli K. Price, LL.D. Proceedings of the American Philosophical Society, 23 (124):572–605, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983339.

Rothrock:1889:BSL

[Rot89] J. T. Rothrock. Biographical sketch of the Late Nathaniel Archer Randolph, M.D. Proceedings of the American Philosophical Society, 26(129):359–365, January/July 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982958.

Rothrock:1891:SOB

[Rot91] J. T. Rothrock. Some observations on the Bahamas and Jamaica. Proceedings of the American Philosophical Society, 29(136): 145–148, July/December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982932.

Rothrock:1894:FP

[Rot94a] J. T. Rothrock. Forests of Pennsylvania. Proceedings of the American Philosophical Society, 33(144):114–133, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983365.

Rothrock:1894:GFI

[Rot94b] J. T. Rothrock. On the growth of the forestry idea in Pennsylvania. Proceedings of the American Philosophical Society, 32(143):332–342, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983044.

Ruschenberger:1884:OND

[Rus84] W. S. W. Ruschenberger. Obituary notice of Dr. Robert Bridges. Proceedings of the American Philosophical Society, 21(115):427–447, April 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982735.

Ruschenberger:1886:SLR

[Rus86] W. S. W. Ruschenberger. A sketch of the life of Robert E. Rogers, M.D., LL.D., with biographical notices of his father and brothers. Proceedings of the American Philosophical Society, 23(121):104–146, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982916.

Ruschenberger:1891:SLD

[Rus91] W. S. W. Ruschenberger. A sketch of the life of Dr. Gouverneur Emerson. Proceedings of the American Philosophical Society, 29 (135):60–78, January/June 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983021.

Ruschenberger:1892:SLJ

[Rus92] W. S. W. Ruschenberger. A sketch of the life of Joseph Leidy, M.D., LL.D. Proceedings of the American Philosophical Society, 30(138):135–184, April 1892. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983169.

Ruschenberger:1893:NWB

[Rus93] W. S. W. Ruschenberger. A notice of William Barton Rogers. Proceedings of the American Philosophical Society, 31(142):254– 257, July/December 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982966.

Ryder:1889:PSG

[Ryd89a] John A. Ryder. The phylogeny of the sweat-glands. Proceedings of the American Philosophical Society, 26(130):534–540, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983194.

Ryder:1889:PTC

[Ryd89b] John A. Ryder. A physiological theory of the calcification of the skeleton. Proceedings of the American Philosophical Society, 26(130): 550–558, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983196.

Ryder:1889:PEH

[Ryd89c] John A. Ryder. Proofs of the effects of habitual use in the modification of animal organisms. Proceedings of the American Philosophical Society, 26(130):541–549, July/December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983195.

Ryder:1890:EOM

[Ryd90a] John A. Ryder. The eye, ocular muscles and lachrymal glands of the shrew-mole (Blarina talpoides Gray). Proceedings of the American Philosophical Society, 28(132):16–18, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983302.

Ryder:1890:OST

[Ryd90b] John A. Ryder. The origin of sex through cumulative integration, and the relation of sexuality to the genesis of species. Proceedings of the American Philosophical Society, 28(132):109–159, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983324.

Ryder:1893:EFO

[Ryd93a] John A. Ryder. Energy as a factor in organic evolution. Proceedings of the American Philosophical Society, 31(141):192–203, April/ June 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982829.

Ryder:1893:MGF

[Ryd93b] John A. Ryder. The mechanical genesis of the form of the fowl’s egg. Proceedings of the American Philosophical Society, 31(141):203–209, April/June 1893. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982830.

Ryder:1894:AFV

[Ryd94] John A. Ryder. The adaptive forms and vortex-motion of the substance of the red blood-corpuscles of vertebrates. Proceedings of the American Philosophical Society, 32(143):272–275, January 1894. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983040.

Sachse:1895:HAC

[Sac95] Julius F. Sachse. Horologium Achaz (Christophorus Schissler, Artifex). Proceedings of the American Philosophical Society, 34(147):21– 30, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983156.

Sachse:1896:JPC

[Sac96] Julius F. Sachse. The joly process of color photography. Proceedings of the American Philosophical Society, 35(151):119–122, August 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983257.

Sachse:1898:OBR

[Sac98] Julius F. Sachse. An old broadside, with a reference to the throne of congress. Proceedings of the American Philosophical Society, 37(157):45–51, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983689.

Sadtler:1875:CLU

[Sad75] Samuel P. Sadtler. Contributions from the Laboratory of the University of Pennsylvania. No. V. On a new occurrence of tartronic acid, with some remarks on the molecular structure of glyceric acid. Proceedings of the American Philosophical Society, 14(95): 615–619, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981942.

Sadtler:1876:CLU

[Sad76] Samuel P. Sadtler. Contributions from the Laboratory of the University of Pennsylvania. No. VI. On the composition of the natural gas from certain wells in Western Pennsylvania. Proceedings of the American Philosophical Society, 16(97):206–212, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982429.

Sadtler:1877:CLU

[Sad77] Samuel P. Sadtler. Contributions from the Laboratory of the University of Pennsylvania. No. VIII. On the composition of the natural gas from certain wells in Western Pennsylvania and New York. Proceedings of the American Philosophical Society, 16(99):585–589, January/ May 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982484.

Sadtler:1878:CRG

[Sad78] Samuel P. Sadtler. On the calculation of results in gas-analyses. Proceedings of the American Philosophical Society, 17(101):473–475, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982649.

Sadtler:1897:GCR

[Sad97] Samuel P. Sadtler. The genesis and chemical relations of petroleum and natural gas. Proceedings of the American Philosophical Society, 36(154):93–102, January 1897. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983462.

Schwarz:1878:CF

[Sch78] E. A. Schwarz. The Coleoptera of Florida. Proceedings of the American Philosophical Society, 17(101):353–372, January/June 1878. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982647.

Scott:1894:NOA

[Sco94] W. B. Scott. Notes on the osteology of Agriochærus Leidy (Artionyx O. & W.). Proceedings of the American Philosophical Society, 33(145):243–251, June 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983272.

Scott:1898:PNS

[Sco98] W. B. Scott. Preliminary note on the selenodont artiodactyls of the Uinta Formation. Proceedings of the American Philosophical Society, 37(157):73–81, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983695.

Scudder:1894:TTS

[Scu94] Samuel H. Scudder. Tertiary Tipulidæ, with special reference to those of Florissant, Colorado. Proceedings of the American Philosophical Society, 32(143):163–245, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983036.

Scudder:1897:SGM

[Scu97] Samuel H. Scudder. The species of the genus Melanoplus. Proceedings of the American Philosophical Society, 36(154):5–35, January 1897.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983458.

Smith:1894:AMT

[SD94] Edgar F. Smith and En. D. Desi. The atomic mass of tungsten. Proceedings of the American Philosophical Society, 33(146): 337–338, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982844.

Sellers:1873:DHD

[Sel73] W. Coleman Sellers. D. S. Holman’s device for the microscopic examination of fluids. Proceedings of the American Philosophical Society, 13(90):180–181, 1873. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981620.

Sellers:1875:ONM

[Sel75] Coleman Sellers. An obituary notice of Mr. Joseph Harrison, Jr. Proceedings of the American Philosophical Society, 14(94):347–355, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981863.

Sellers:1886:ONL

[Sel86] William Sellers. An obituary notice of the Late George Whitney. Proceedings of the American Philosophical Society, 23(123):388–390, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983223.

Sellers:1899:CAD

[Sel99] Coleman Sellers. Chairman’s annual discourse for 1899, on the transmission of energy by electricity. Proceedings of the American Philosophical Society, 38(159):49–71, January 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983622.

Sharswood:1868:ONJ

[Sha68] George Sharswood. Obituary notice of Joseph Reed Ingersoll. Proceedings of the American Philosophical Society, 10(80):513–522, July 1868. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/981797.

Sharpless:1883:LHC

[Sha83] Isaac Sharpless. The latitude of Haverford College Observatory. Proceedings of the American Philosophical Society, 21(114):78–81, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982362.

Sharples:1895:NPT

[Sha95] S. P. Sharples. Notes on photographic testing of inks. Proceedings of the American Philosophical Society, 34(149):471–472, December 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982903.

Sheafer:1869:BRA

[She69] P. W. Sheafer. Boring records from the anthracite basin. Proceedings of the American Philosophical Society, 11(81):107–110, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981446.

Sherwood:1878:SDR

[She78] Andrew Sherwood. Section of Devonian rocks made in the Catskill Mountain at Palenville; Kauterskill Creek, New York. Proceedings of the American Philosophical Society, 17(101):346–347, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982644.

Smith:1887:ELS

[SKHE87] Edgar F. Smith, E. B. Knerr, W. S. Hoskinson, and Charles H. Ehrenfeld. Electrolysis of lead solutions. Determination of boric acid. dihalogen derivatives of salicylic acid. Barite. Proceedings of the American Philosophical Society, 24(126):428–434, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983092.

Slade:1895:SJA

[Sla95] Daniel Denison Slade. The significance of the Jugal Arch. Proceedings of the American Philosophical Society, 34(147):50–67, January 1895. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983159.

Sadtler:1879:PNI

[SM79] Samuel P. Sadtler and H. G. McCarter. Preliminary notice of an investigation on “petrocene”, a product of the destructive distillation of petroleum. Proceedings of the American Philosophical Society, 18 (103):185–186, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982318.

Smith:1877:CLUa

[Smi77a] Edgar F. Smith. Contributions from the Laboratory of the University of Pennsylvania. No. IX. Upon some new chlorine derivatives from toluol. Proceedings of the American Philosophical Society, 17 (100):29–32, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982275.

Smith:1877:CLUb

[Smi77b] Edgar F. Smith. Contributions from the Laboratory of the University of Pennsylvania. No. X. Dichlorsalicylic acid. Proceedings of the American Philosophical Society, 17(100):68–73, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982280.

Smith:1877:CLUc

[Smi77c] Edgar F. Smith. Contributions from the Laboratory of the University of Pennsylvania. No. XII. A new method for the decomposition of chromic iron. Proceedings of the American Philosophical Society, 17 (100):216–218, June/December 1877. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982300.

Smith:1878:CLUa

[Smi78a] Edgar F. Smith. Contributions from the Laboratory of the University of Pennsylvania. No. XIV. I. A new monochlordinitrophenol and an aniline derivative of a-monochlordinitrophenol. II. Beryllium borate. Proceedings of the American Philosophical Society, 17 (101):706–708, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982662.

Smith:1878:CLUb

[Smi78b] Edgar F. Smith. Contributions from the Laboratory of the University of Pennsylvania, No. XVI. On the electrolytic estimation of cadmium. Proceedings of the American Philosophical Society, 18(102):46, July/ December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982757.

Smith:1879:ACF

[Smi79] Edgar F. Smith. Analysis of a calculus found in a deer. Proceedings of the American Philosophical Society, 18(103):213–215, January/June 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982325.

Smith:1888:AGH

[Smi88] Edgar F. Smith. Action of the gas from As_{2}O_{3} and HNO_{3} upon m-oxybenzoic acid. Proceedings of the American Philosophical Society, 25(128):194–195, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/983053.

Smith:1895:TS

[Smi95] George H. Smith. The theory of the state. Proceedings of the American Philosophical Society, 34(148):181, 183, 185–334, July 1895. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983255.

Smith:1896:MFC

[Smi96] James Perrin Smith. Marine fossils from the coal measures of Arkansas. Proceedings of the American Philosophical Society, 35 (152):213–285, November 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983148.

Smyth:1898:SPA

[Smy98] Albert H. Smyth. Shakespeare’s Pericles and Apollonius of Tyre. Proceedings of the American Philosophical Society, 37(158):206–312, December 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983853.

Scott:1887:PRV

[SO87] W. B. Scott and Henry F. Osborn. Preliminary report on the vertebrate fossils of the Uinta Formation, collected by the Princeton Expedition of 1886. Proceedings of the American Philosophical Society, 24 (126):255–264, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983074.

Spencer:1881:DPO

[Spe81] J. W. Spencer. Discovery of the preglacial outlet of the basin of Lake Erie into that of Lake Ontario; with notes on the origin of our Lower Great Lakes. Proceedings of the American Philosophical Society, 19 (108):300–337, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982255.

Sharswood:1878:SAS

[SS78] Andrew Sharswood and Clark Sharswood. Section along Schoharie Creek in Schoharie County, N.Y., between Gilboa and Middleburg, from the Catskill down to the Upper Helderberg. Proceedings of the American Philosophical Society, 17(101):347–349, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982645.

StJohn:1869:DFF

[St.69] Orestes St. John. Descriptions of fossil fishes, from the Upper Coal Measures of Nebraska. Proceedings of the American Philosophical Society, 11(81):431–437, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981496.

Smith:1882:CW

[ST82] Edgar F. Smith and N. Wiley Thomas. Corundum and wavellite. Proceedings of the American Philosophical Society, 20(111):230–231, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982721.

Stevenson:1874:APC

[Ste74] Jno. J. Stevenson. On the alleged parallelism of coal beds. Proceedings of the American Philosophical Society, 14(93):283–295, June/ December 1874. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982118.

Stevenson:1875:NGW

[Ste75a] Jno. J. Stevenson. Notes on the geology of West Virginia. No. II. Proceedings of the American Philosophical Society, 14(94):370–401, January/June 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/981868.

Stevenson:1875:GRL

[Ste75b] John J. Stevenson. The geological relations of the lignitic groups. Proceedings of the American Philosophical Society, 14(95):447–475, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.jstor.org/stable/
981935.

Stevenson:1879:SGS

[Ste79] John J. Stevenson. Surface geology of South-West Pennsylvania and adjacent portions of West Virginia and Maryland. Proceedings of the American Philosophical Society, 18(104):289–316, July/December 1879. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982503.

Stevenson:1880:NGW

[Ste80a] John J. Stevenson. Notes on the geology of Wise, Lee and Scott Counties, Virginia. Proceedings of the American Philosophical Society, 19(107):88–107, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982613.

Stevenson:1880:NRR

[Ste80b] John J. Stevenson. Notes respecting a re-eroded channel-way. Proceedings of the American Philosophical Society, 19(107):84–87, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982612.

Stevenson:1881:GRP

[Ste81a] John J. Stevenson. A geological reconnaissance of parts of Lee, Wise, Scott and Washington Counties, Virginia. Proceedings of the American Philosophical Society, 19(108):219–262, January/June 1881. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982245.

Stevenson:1881:NCF

[Ste81b] John J. Stevenson. Notes on the coal-field near Cañon City, Colorado. Proceedings of the American Philosophical Society, 19(109): 505–521, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982797.

Stevenson:1881:NQC

[Ste81c] John J. Stevenson. Notes on the Quinnimont Coal Group in Mercer Co. of West Virginia and Tazewell Co. of Virginia. Proceedings of the American Philosophical Society, 19(109):498–505, June/ December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982796.

Stevenson:1881:UFC

[Ste81d] John J. Stevenson. The Upper Freeport Coal Bed along Laurel Ridge in Preston County of West Virginia. Proceedings of the American Philosophical Society, 19(108):276–279, January/June 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982248.

Stevenson:1882:NLG

[Ste82] John J. Stevenson. Note on the Laramie Group in the vicinity of Raton, New Mexico. Proceedings of the American Philosophical Society, 20(111):107–111, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982707.

Stevenson:1885:NGS

[Ste85a] John J. Stevenson. Notes on the geological structure of Tazewell, Russell, Wise, Smyth and Washington Counties of Virginia. Proceedings of the American Philosophical Society, 22(118):114–161, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982673.

Stevenson:1885:SNR

[Ste85b] John J. Stevenson. Some notes respecting metamorphism. Proceedings of the American Philosophical Society, 22(118):161–166, March 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/982674.

Stevenson:1887:GRB

[Ste87a] John J. Stevenson. A geological reconnaissance of Bland, Giles, Wythe and portions of Pulaski and Montgomery Counties of Virginia. Proceedings of the American Philosophical Society, 24(125): 61–108, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983125.

Stevenson:1887:NSG

[Ste87b] John J. Stevenson. Notes on the surface geology of South-West Virginia. Proceedings of the American Philosophical Society, 24 (125):172–178, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983131.

Stevenson:1896:RFD

[Ste96] Cornelius Stevenson. On the remains of the foreigners discovered in Egypt by Mr. Flinders-Petrie, 1895, now in the Museum of the University of Pennsylvania. Proceedings of the American Philosophical Society, 35(150):56–64, January 1896. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982861.

Stille:1869:ONH

[Sti69] Charles J. Stille. Obituary notice of Horace Binney, Jr. Proceedings of the American Philosophical Society, 11(81):371–380, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/981489.

Stowell:1882:VND

[Sto82] T. B. Stowell. The vagus nerve in the domestic cat (Felis Domestica). Proceedings of the American Philosophical Society, 20 (111):123–138, January/June 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982710.

Stoll:1885:SRG

[Sto85] Otto Stoll. Supplementary remarks to the grammar of the Cakchiquel language of Guatemala, edited by D. G. Brinton, M.D. Proceedings of the American Philosophical Society, 22(120):255–268, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982983.

Stokes:1886:NNF

[Sto86a] Alfred C. Stokes. Notices of new fresh-water infusoria. Proceedings of the American Philosophical Society, 23(124):562–568, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983336.

Stokes:1886:SNH

[Sto86b] Alfred C. Stokes. Some new hypotrichous infusoria. Proceedings of the American Philosophical Society, 23(121):21–30, January 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982911.

Stowell:1886:TND

[Sto86c] T. B. Stowell. The trigeminus nerve in the domestic cat (Felis domestica). Proceedings of the American Philosophical Society, 23 (124):459–478, December 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983326.

Stokes:1887:NNF

[Sto87a] Alfred C. Stokes. Notices of new fresh-water infusoria. Proceedings of the American Philosophical Society, 24(126):244–255, July/December 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983073.

Stowell:1887:FND

[Sto87b] T. B. Stowell. The facial nerve in the domestic cat. Proceedings of the American Philosophical Society, 24(125):8–18, January/June 1887. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/983121.

Stowell:1888:AND

[Sto88a] T. B. Stowell. The accessory nerve in the domestic cat. Proceedings of the American Philosophical Society, 25(127):94–99, January/ June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983285.

Stowell:1888:GPN

[Sto88b] T. B. Stowell. The glosso-pharyngeal nerve in the domestic cat. Proceedings of the American Philosophical Society, 25(127):89–94, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983284.

Stowell:1888:HND

[Sto88c] T. B. Stowell. The hypoglossal nerve in the domestic cat. Proceedings of the American Philosophical Society, 25(127):99–104, January/ June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983286.

Stokes:1890:NNF

[Sto90] Alfred C. Stokes. Notices of new fresh-water infusoria. Proceedings of the American Philosophical Society, 28(132):74–80, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983315.

Stokes:1894:NPU

[Sto94] Alfred C. Stokes. Notices of presumably undescribed infusoria. Proceedings of the American Philosophical Society, 33(146):338– 345, July/December 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982845.

Strong:1876:DIL

[Str76] William Strong. A discourse, illustrative of the life and character of Horace Binney, delivered before the Bar of Philadelphia, the Law Association, and the American Philosophical Society, in the Musical Fund Hall, on the evening of Jan. 5, 1876. Proceedings of the American Philosophical Society, 16(97):1–51, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982419.

Tafel:1861:ILE

[Taf61] R. L. Tafel. Investigations into the laws of English orthography and pronunciation. Proceedings of the American Philosophical Society, 8 (66):285–376, July/December 1861. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981780.

Tafel:1862:ILE

[Taf62] R. L. Tafel. [Investigations into the laws of English orthography and pronunciation: Conclusion]. Proceedings of the American Philosophical Society, 9(67):39–54, January 1862. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981542.

Taylor:1885:CP

[Tay85] W. Curtis Taylor. On composite photography. Proceedings of the American Philosophical Society, 22(120):360–362, October 1885. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982995.

Taylor:1887:ONA

[Tay87] Alfred B. Taylor. Octonary numeration, and its application to a system of weights and measures. Proceedings of the American Philosophical Society, 24(126):296–366, July/December 1887. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983079.

Thayer:1878:MTC

[Tha78] Russell Thayer. Movements of troops in cities in cases of riot or insurrection. Proceedings of the American Philosophical Society, 18 (102):89–95, July/December 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982764.

Thayer:1883:AS

[Tha83] Russell Thayer. Aerial ships. Proceedings of the American Philosophical Society, 21(114):301–303, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982392.

Trowbridge:1876:ASP

[Tro76] David Trowbridge. On the atmospheres of the Sun and planets. Proceedings of the American Philosophical Society, 16(98): 327–333, June/December 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982819.

Uhler:1888:AFN

[Uhl88] P. R. Uhler. The Albirupean formation and its nearest relatives in Maryland. Proceedings of the American Philosophical Society, 25 (127):42–54, January/June 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983279.

Vaux:1884:PPS

[Vau84] Richard Vaux. The Pennsylvania prison system. Proceedings of the American Philosophical Society, 21(116):651–664, June 1884. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982344.

Vaux:1885:BNH

[Vau85] Richard Vaux. Biographical notice of Henry M. Phillips. Proceedings of the American Philosophical Society, 22(117):72–78, January 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982408.

Vaux:1887:JRL

[Vau87] Richard Vaux. James R. Ludlow. Proceedings of the American Philosophical Society, 24(125):19–23, January/June 1887. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/983122.

Vaux:1889:STS

[Vau89] Richard Vaux. Some thoughts on the Sun and cross symbols. Proceedings of the American Philosophical Society, 26(130):476–481, July/ December 1889. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983185.

Vaux:1890:ONF

[Vau90] Richard Vaux. Obituary notice of Franklin B. Gowen. Proceedings of the American Philosophical Society, 28(132):61–65, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983311.

vanDenburgh:1898:HN

[vD98] John van Denburgh. Herpetological notes. Proceedings of the American Philosophical Society, 37(157):139–141, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983700.

vanDenburgh:1899:NSB

[vD99] John van Denburgh. Notes on some birds of Santa Clara County, California. Proceedings of the American Philosophical Society, 38 (160):157–180, October 1899. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983567.

vonHumboldt:1885:VAL

[vHB85] Wilhelm von Humboldt and D. G. Brinton. On the verb in American languages. Proceedings of the American Philosophical Society, 22(120):332–354, October 1885. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982992.

Wake:1890:AAM

[Wak90] C. Staniland Wake. The Asiatic affinities of the Malay language. Proceedings of the American Philosophical Society, 28(132):81–87, January/June 1890. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/983316.

Wall:1884:SIP

[Wal84] J. Sutton Wall. On some Indian picture rocks in Fayette County, Pa. Proceedings of the American Philosophical Society, 21(116):687–689, June 1884. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/982350.

Warwick:1891:EMF

[War91] Hill Sloane Warwick. The electrolysis of metallic formates. Proceedings of the American Philosophical Society, 29(136):103–120, July/ December 1891. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982925.

Wharton:1894:DKE

[Wha94] Joseph Wharton. Dust from the Krakatoa eruption of 1883. Proceedings of the American Philosophical Society, 32(143):343–345, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/983045.

White:1880:NPS

[Whi80] I. C. White. Notes on the place of the Sharon conglomerate in the Paleozoic series. Proceedings of the American Philosophical Society, 19(107):198–203, March/December 1880. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/982629.

White:1881:NGW

[Whi81] I. C. White. Notes on the geology of West Virginia. Proceedings of the American Philosophical Society, 19(109):438–446, June/ December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982788.

White:1882:NGW

[Whi82] I. C. White. Notes on the geology of West Virginia. Proceedings of the American Philosophical Society, 20(112):479–496, June/ December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/982697.

Williamson:1875:MOT

[Wil75] R. S. Williamson. Meteorological observations taken on the Nile between Cairo and the First Cataract, during January and February, 1873. Proceedings of the American Philosophical Society, 14 (95):632, June/December 1875. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981945.

Wilcocks:1878:LDA

[Wil78] Alexander Wilcocks. Letter of Dr. Alexander Wilcocks on shadows without penumbra, read February 1, 1878. Proceedings of the American Philosophical Society, 17(101):705, January/June 1878. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982661.

Wilder:1881:BCF

[Wil81] Burt G. Wilder. The brain of the cat, Felis Domestica. 1. Preliminary account of the gross anatomy. Proceedings of the American Philosophical Society, 19(109):524–562, June/December 1881. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/982799.

Williston:1882:CMN

[Wil82] W. Williston. Contribution to a monograph of the North American Syrphidæ. Proceedings of the American Philosophical Society, 20 (112):299–332, June/December 1882. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982682.

Williams:1883:CMS

[Wil83] Henry S. Williams. On a crinoid with movable spines. Proceedings of the American Philosophical Society, 21(114):81–88, March 1883. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/982363.

Wilson:1888:BNT

[Wil88] Joseph M. Wilson. Biographical notice of Thomas Ustick Walter, A.M., Ph.D., LL.D., Late Member of the American Philosophical Society. Proceedings of the American Philosophical Society, 25(128): 322–327, July/December 1888. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983068.

Willcox:1894:SWW

[Wil94] Joseph Willcox. Some of the work of the Wagner Free Institute of Science of Philadelphia. Proceedings of the American Philosophical Society, 32(143):245–247, January 1894. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/983037.

Williams:1898:NKD

[Wil98] Edward H. Williams, Jr. Notes on Kansan drift in Pennsylvania. Proceedings of the American Philosophical Society, 37(157):84–87, January 1898. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/983697.

Winchell:1869:GAEa

[Win69a] A. Winchell. On the geological age and equivalents of the Marshall Group. Proceedings of the American Philosophical Society, 11 (81):57–82, 418, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981440.

Winchell:1869:GAEb

[Win69b] A. Winchell. On the geological age and equivalents of the Marshall Group. Proceedings of the American Philosophical Society, 11(81):385–418, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981492.

Winchell:1869:NDF

[Win69c] Alexander Winchell. Notices and descriptions of fossils, from the Marshall Group of the Western states, with notes on fossils from other formations. Proceedings of the American Philosophical Society, 11(81):245–260, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981469.

Wood:1869:MAH

[Woo69a] Horatio C. Wood, Jr. On the medical activity of the hemp plant, as grown in North America. Proceedings of the American Philosophical Society, 11(81):226–232, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981465.

Wood:1869:PSF

[Woo69b] Horatio C. Wood, Jr. Prodromus of a study of the fresh water algae of Eastern North America. Proceedings of the American Philosophical Society, 11(81):119–145, January 1869. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/981451.

Wood:1871:RFT

[Woo71a] Geo. B. Wood. Revival of fruit trees prematurely ceasing to bear fruit, or prematurely decaying. Proceedings of the American Philosophical Society, 12(86):3–7, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981655.

Wood:1871:IFW

[Woo71b] George B. Wood. Influence of fresh wood-ashes on the growth of wheat, potatoes, &c. Proceedings of the American Philosophical Society, 12(86):323–327, 1871. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/981714.

Wootten:1876:CAW

[Woo76] John E. Wootten. A combination of apparatus by which ordinary anthracite coal-waste, from the dirt-banks at the mines, can be successfully and profitably burned in the furnaces of stationary and locomotive boilers. Proceedings of the American Philosophical Society, 16(97):214–217, January/June 1876. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/982431.

Wyckoff:1886:UOS

[Wyc86] A. B. Wyckoff. The use of oil in storms at sea. Proceedings of the American Philosophical Society, 23(123):383–388, July 1886. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/983222.